A Bibliography of Papers in Lecture Notes in Computer Science (1993)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

11 March 2017 Version 1.24

Title word cross-reference

1, 579^{n} [828]. 16 [2097]. 2 [1076, 1017, 942, 1282]. 2/3 [2561]. 3 [1753, 1464, 1875, 570, 1873, 1446, 942, 1441, 938, 932, 1918, 568, 1420, 588]. AC^{O} [253]. mod2^{32} [91]. C^{∗∗} [2436]. D [2559, 1691]. E^{3} [2353]. f [2551]. F_{ω} [199]. GeO_{2} [1529]. k [970, 2551, 975, 775, 2172, 1000, 1711]. l [2104]. m [1360]. AC^{0} [1008]. TLA^{+} [1980]. N [633, 536, 747, 1671, 2203, 2086]. N × M [1933]. O(1) [2556]. O(log log n) [735]. O(lgn) [1040]. O(\sqrt{n}) [213]. O^{∗} [2148]. q [1415]. Q_{+} [1181]. s^{2} [2099]. w [2582]. x^{m+n} = x^{n} [39]. xxx [2200]. Z_{n} [2067].

&-early [178].

* [805].

-Approximation [2561]. -Approximations [2200]. -ary [1360, 633]. -body [747]. -Branching [2582]. -Cube [2559]. -D [1753, 1464, 570, 942, 1420, 588]. -Dependent [2551]. -DES [2099]. -Dimensional [1691, 2086, 932, 1918, 536]. -Enclosing [975]. -Objects [1850]. -party [1671]. -Point [970]. -Power [775]. -Round [2097]. -Span [2104]. -Term [2172]. -th [1415]. -Trees [1000, 1711]. -worst-case-time [213].

0-1 [677, 1343].

1 [1792]. 1-safe [2531]. 1.0.1 [1837]. 1/D [2131]. 10 [1895]. 1st [269].

2 [2078, 1634]. 2-Prover [2078]. 2-stage [1418]. 2-structures [1051].

3 [960, 1943, 828]. 3-edge-components [71]. 3-edge-cuts [71]. 3-Planar [960]. 3-Satisfiability [828]. 3-valued [1747].

4 [2059]. 4-connected [72]. 4.2 [1145].

5 [353].

8-bit [850].

92 [1614].

= [81, 2247].

A- [335]. A-Sufficiency [170]. Abduction [378, 209]. abductive [1647]. Abelian [87]. Abnormal [1471]. Abstract [1748, 1642, 2323, 1331, 105, 2006, 129, 1218, 1202, 350, 1954, 1221, 1182, 1632, 87, 2327, 2079, 1165, 2065, 2400, 132, 1151, 1188, 1958, 1167, 1168, 331, 2077, 2138, 612, 2080, 1170, 2083, 2064, 1046, 1266, 1230, 947, 179, 1631, 1069, 186, 1629, 14, 1253, 1628, 949, 163]. Abstraction [1407, 162, 923, 290, 2171, 432, 1233, 1047, 1095]. Abstraction-Based [432]. Abstractions [294, 2141, 2299]. AC [250]. AC-like [250]. AC1 [784]. AC1-Unification [784]. Academia [1302]. Accelerated [468]. Accelerating [495]. Acceptability [2244]. acceptance [246]. Access [1173, 2453, 2251, 2511, 805, 1569, 1137, 238, 1659]. Accessing [1893]. Accident [1516]. Accommodation [1993]. Accounting [2401]. Accumulation [392]. Acquiring [727]. Acquisition [1625, 1627, 1607, 286, 581, 1620, 2221, 1626, 399, 1621, 1619, 1612, 1618, 2396]. Across [1111, 896]. ACSP [2534]. ACSR [1258]. Action [1048, 555, 1007, 1245]. Active [417, 357, 1329, 1436, 566, 2152, 2497, 1515, 904, 660, 940, 2226]. Activities [2141, 2031].

Activity [2375, 1884, 643]. Actor [2416]. Actor-Based [2416]. ActorStudio [1899]. Acute [311]. Acyclic [1704]. AD [156]. adaptability [1635]. Adaptable [1911]. Adapted [767, 1882]. Adaption [1904]. Adaptive [766, 552, 553, 2369, 2372, 1392, 562, 2559, 1111, 593, 1413, 2149, 479, 559, 541, 558, 451, 1478, 1828, 359, 504]. Adaptively [2083]. Adding [2153]. Addition [1181, 1182, 2054, 595, 1349, 101]. Additional [522, 2177]. Address [875]. Addressable [2106]. adequacy [1055]. adic [1934]. Aditi [2484]. Adjusting [2587]. Administration [1921]. Administrative [1925]. Admissible [757, 605]. Admission [1151, 1117]. Advanced [1317, 2232, 2418, 2146, 1535]. Advances [2573]. Adventure [862]. adversaries [113]. Advisor [311]. AERIE [2391]. Aero [1487]. Aero-Space [1487]. Aerotriangulation [1480]. Affine [1453]. After [563, 453]. Against [2108, 113, 2105, 114, 2083]. Agency [794]. Agent [143, 925, 1989, 1762, 1756]. Agents [1551, 1540, 2224, 1741, 310, 650]. Aggregates [433, 2485]. Aggregating [2189]. Aggregation [2474, 2140, 326]. Agora [914]. Agreement [2095, 106, 1658]. Agricultural [1485, 1484]. AI [1552, 1997, 2277]. Aid [419]. Aided [1460, 442, 2054, 634, 1922, 845, 1349, 99, 1486]. Aim [1549]. Aims [928]. AKL [1213, 1651]. AL [2131, 1959]. AL-1 [2131]. AL-1/D [2131]. ALADIN [1624]. ALEXSYS [1531]. AlgBench [1587, 1601]. Algebra [1584, 1586, 803, 918, 1588, 309, 1582, 2008, 1596, 1251, 1258, 1078, 1094]. Algebraic [1, 1575, 771, 1951, 2208, 1585, 2005, 2040, 17, 1600, 1973, 1944, 2000, 949, 706, 1953, 1645, 7, 8, 10, 273, 3, 2410, 2411]. Algebraic-Code [2040]. Algebraically [1970]. Algebras [21, 392, 1716, 819, 1784, 1960, 693, 384, 762, 6, 1266, 1075, 1244, 1070, 1263, 1095].

Algorithm [341, 2592, 951, 139, 956, 1675, 1782, 1679, 1440, 1826, 1812, 974, 1019, 2016, 2395, 1184, 977, 1533, 487, 464, 486, 2196, 1151, 1578, 1447, 300, 490, 1933, 1579, 989, 496, 1410, 1204, 344, 2557, 1700, 2538, 1702, 605, 2202, 1936, 999, 1742, 1708, 500, 588, 495, 1711, 1405, 67, 1364, 76, 229, 1669, 1358, 258, 1260, 1367, 1339]. Algorithmic [1985, 810, 744, 2009, 2333, 2319, 2328, 2157, 725]. Algorithms [2537, 935, 485, 1674, 1721, 954, 2573, 1676, 2587, 2585, 968, 1682, 681, 1683, 970, 971, 2544, 764, 2540, 2598, 739, 2559, 822, 2590, 2432, 2589, 1939, 979, 1693, 1816, 1432, 2588, 1408, 2162, 2434, 1934, 2594, 478, 1973, 492, 2394, 2519, 1803, 2578, 998, 2451, 1000, 503, 1117, 2596, 1001, 272, 68, 2405, 63, 81, 1631, 70, 1769, 269, 1083, 1659, 1656, 231, 1768, 1629, 243, 217, 278, 2408, 271, 1091, 2601, 2414]. Aligned [1021]. Alignment [854]. ALISA [1423]. All-Nearest-Neighbors [977]. All-Pairs [945, 985]. all-to-all [63]. Allocation [1228, 304, 1579, 2466]. Alps [1684]. Alternating [2582, 95]. Alternation [202, 201]. amalgamation [9]. Amalgame [2467]. Ambiguity [408, 2222]. Ambiguous [2565]. ami [1315]. Among [2224, 736, 2359, 992, 1818, 2581, 1366]. amorphous [1248]. amount [1064]. Anaesthesia [1549]. Analog [2041, 578]. Analogic [1490]. Analogical [1823, 440, 2170, 1577]. Analysing [497]. Analysis [916, 1994, 474, 1781, 589, 1782, 2343, 430, 1220, 1318, 1277, 785, 1620, 497, 1954, 2195, 866, 1836, 1484, 1141, 925, 685, 1415, 83, 1022, 1784, 2449, 1293, 1787, 1278, 1203, 1940, 593, 1789, 1462, 1133, 1562, 1956, 691, 2443, 2348, 1208, 571, 943, 565, 521, 2269, 1983, 2420, 391, 2015, 1959, 2399, 1798, 2461, 1813, 1470, 1802, 2147, 652, 1471, 1516, 1804, 1116, 1805, 1489, 1320, 2422, 1633, 1647, 1644, 1651, 1636, 1649]. Analytic [1109]. Analytical [585, 1382, 1144]. Analyzing [363, 863, 1354]. Anatomical [575, 576]. Anatomy [574]. AND/OR [252]. Andre [1043]. Andrew [1049]. Angiograms [1464]. Angiography [570, 1470, 594]. Angle [1710]. Angles [1395]. Animation [443, 1499, 1573, 269]. Animators [2326]. Anisotropic [1388]. ANN [513]. Annealing [584, 1768]. Anomaly [1318, 915]. anonymous [1670]. ANSAware [1834]. Answering [1506]. Answers [1720, 529, 1772]. Anticipatory [2230]. antimonotonic [192]. Any [2239, 2049, 287]. Apertos [2133]. API [1129]. Application [289, 1547, 284, 1485, 1476, 794, 516, 681, 1875, 318, 1593, 669, 533, 1900, 279, 1159, 412, 2332, 2042, 293, 1560, 1152, 1161, 1870, 1837, 2174, 1554, 634, 1515, 846, 1479, 1289, 1823, 1000, 594, 476, 762, 535, 560, 106, 66, 179, 1350, 101, 1101, 1081]. Applications [374, 1675, 2573, 2515, 1891, 1135, 2350, 2493, 972, 2101, 1280, 739, 1852, 976, 1497, 2309, 759, 2334, 744, 1570, 1548, 1839, 2004, 1126, 2048, 1522, 520, 1602, 2596, 2026, 91, 230, 70, 1349, 1356, 1367]. Applicative [1959, 173]. Applied [570]. Applying [2360, 896, 2458]. Appreciations [456]. Approach [411, 513, 1985, 2234, 589, 734, 426, 2379, 543, 1906, 2477, 966, 2375, 1728, 54, 1995, 937, 416, 2246, 1969, 2253, 1418, 1557, 1379, 604, 577, 1585, 666, 1982, 1714, 922, 2010, 2158, 298, 1315, 1999, 1413, 1401, 1200, 2136, 2338, 2381, 2180, 565, 521, 1114, 1968, 1983, 1913, 2271, 459, 425, 1463, 1952, 1517, 544, 407, 297, 2000, 432, 1450, 440, 1509, 1612, 1319, 1618, 439, 652, 446, 31, 2402, 2133, 1986, 716, 1044, 1359, 719, 262, 2485]. Approaches [2453, 1301, 1465, 2262, 381, 2153, 1398, 2109, 1474, 2083]. Approximability [2580]. Approximate [1417, 1068, 979, 738, 2341, 992, 402]. Approximated [409]. Approximating [980, 983, 740]. Approximation [811, 968, 1686, 2598, 2561, 739, 455, 1389, 1383, 985, 737, 629, 2596, 96].

Approximations [1223, 392, 1021, 2200, 1726, 1031, 2560, 151]. Apt [1042]. Aquarius [1594]. Arabic [1908]. Arbeitsplatz [1912]. Arbitrary [2324]. arbitration [1334]. Arc [1675, 966]. Architectural [524, 1635]. Architecture [2134, 2350, 426, 2305, 926, 653, 489, 840, 1864, 805, 2494, 167, 355, 1841, 1989, 517, 525, 2298, 557, 1928, 1149, 1112, 848, 1157, 2118, 560, 1489, 1761, 1752]. Architectures [1845, 523, 2308, 1806, 1803, 349, 2455, 1869, 625, 226]. Area [1783, 2120, 2231]. Argumentation [388]. Arguments [2191, 2258, 2077, 2244]. Arithmetic [1934, 1189, 276]. Arithmetical [1165]. Arnold [1043, 1178]. Arrangement [572]. Array [835, 2433, 517]. Array-Based [835]. Arrays [676, 1676, 995, 1932, 836, 132, 843, 2443, 521, 2538, 837, 1828, 848, 1336, 1373, 1058]. Art [667, 1918]. Artifact [859]. Artificial [2007, 1609, 1996, 514, 533, 507, 487, 473, 559, 413, 1493, 2272, 2600, 476, 1754, 1756]. Artists [673]. Artmap [489, 560]. Arto [1052]. ary [633, 1360]. ASIC [857]. ASICs [524]. asking [1355]. ASN.CG [2400]. Aspects [441, 771, 2253, 1925, 2119]. Assembly [2397, 72]. Assertion [1331]. Assertion-Based [1331]. assertional [1092]. Assessment [859, 357, 303, 1518, 650, 883]. Assignment [2505, 175, 173]. Assist [2418, 2374]. Assistance [1916, 1883, 1342]. Assistant [2002]. Assisted [1483, 1466, 1624, 2373]. Associated [384]. Associative [555, 29, 338, 345, 473, 1922]. Associative-Commutative [29, 345]. Assumption [1740, 379, 2043]. Assumption-Based [379]. Assumptions [1175, 2077, 2581, 2314]. Assurance [1531]. Assuring [887]. Asteroidal [958]. ASTRAL [1330]. Asymptotic [811, 832]. Asymptotics [745]. Asynchronous [1438, 839, 2301, 1660, 260]. Asynchrony [2300]. Atherosclerotic [591]. ATM [1133]. ATMS [1747, 2265, 2269]. atomic [1665]. atomicity [1656]. Attaching [911]. Attack [1883, 2059, 126, 131, 93, 1376, 1335]. Attacks [2060, 99, 2083]. Attained [1309]. Attaining [2083]. Attention [1885]. Attenuation [585]. Attitude [658]. Attraction [2163]. Attractor [475]. Attribute [2129, 1730, 1205, 1281]. Attribute-Specific [1730]. Audio [1160, 125, 1111]. Auditory [2214]. Augmenting [969]. Australian [1361]. Authentication [2113, 1335, 2030, 1333, 2112, 1334, 134, 130]. Author [135, 929, 2283]. Authoring [1878]. Authors [1902]. Auto [2210]. Auto-Organisation [2210]. Autoassociative [2163]. Autocorrelation [1294]. Autoepistemic [1166, 1185, 1726, 1760]. Automata [1985, 788, 680, 1017, 763, 2203, 541, 1990, 1036, 1059, 1248, 1072, 260, 1076, 240, 249, 1250, 235, 1085, 250, 233, 261]. Automated [1729, 2311, 575, 1421, 437, 1478, 1894, 2312, 652, 216]. Automatic [513, 842, 637, 506, 2498, 2215, 2196, 1941, 1459, 1870, 2333, 1327, 1543, 1480, 524, 1471]. Automatical [1974]. automatically [1771]. Automating [2414]. Automation [2375]. Automaton [2023]. Autonomous [1551, 1992, 316, 941, 650, 1776]. Auxiliary [825]. Available [1940]. avalanche [1343]. average [67, 263]. Avoiding [1198, 701, 1639]. AWB [2148]. AX [335]. AX-Unification [335]. Axiom [2247, 708, 1195, 1266]. Axiomatic [880]. axiomatization [1080]. Axiomatizations [244]. Axiomatizing [703, 208]. Axioms [1175]. Axis [1021]. Axis-Aligned [1021]. AXYTRANS [2123].

B [771, 490, 454]. B-Cell [454]. B-Learning [490]. B-Regular [771]. Back [1224]. Background [2186]. Backprojection [1451, 588]. Backtrack [633]. Backtrack-Bounded [633]. Backward [385]. bag [275]. Balanced [952, 2556, 1343]. Balancing [957, 312, 616, 1828]. Balconies [1914]. Band [1387]. Band-Limited [1387]. Bandwidth [1140]. Banerjee [2445]. Banerjee-Wolfe [2445]. Bank [560]. Bargaining [1741]. Bars [1901]. Base [1526, 410, 1169, 1617, 2455, 1319, 2022, 2501]. Based [1890, 598, 373, 371, 2528, 2362, 954, 1388, 835, 2237, 393, 398, 1331, 1220, 1584, 404, 379, 600, 2438, 2480, 2316, 784, 1954, 562, 2023, 115, 54, 2047, 1221, 2481, 468, 1825, 1863, 1184, 1876, 290, 2152, 2441, 679, 666, 2535, 2145, 1152, 2290, 2184, 2059, 693, 847, 2416, 2021, 1455, 2301, 403, 307, 1491, 853, 90, 2054, 1897, 934, 2440, 887, 781, 608, 636, 619, 1958, 1441, 2329, 869, 2176, 690, 1492, 2095, 149, 2267, 2132, 2077, 943, 565, 521, 1616, 2027, 2067, 2171, 1478, 1923, 1479, 283, 1895, 1463, 1718, 2330, 1898, 576, 2231, 297, 2399]. Based [2000, 432, 654, 700, 1157, 481, 625, 510, 1612, 1742, 2209, 1878, 1394, 607, 1281, 652, 1531, 709, 2396, 31, 2390, 1537, 2207, 729, 1755, 1633, 116, 209, 12, 1767, 1341, 1370, 730, 1083, 1241, 1647, 1087, 1357, 1358, 1359, 1569, 1363, 1373, 1509, 1763, 16, 1215, 838, 1571, 2148, 845, 852, 1841, 1363]. Based-DSS [1571]. Bases [1825, 2254, 1541, 2492, 1536, 2004, 476, 656]. Basic [809, 1121, 1228, 2500, 718, 1240, 1078]. basic-level [718]. basing [1766]. Basis [912, 286, 468, 469, 1772, 2410]. Batch [106]. Bayesian [1379, 584, 467, 97, 590]. BDB [1526]. Be [2117, 2077, 40, 448, 120, 635]. Beam [1452, 588]. Beauty [561]. Bed [1466]. Been [1309]. Begriffsverzeichnis [659]. Behavior [2407, 924, 1787, 1488]. Behavior-oriented [2407]. Behavioral [2135, 2010, 2338, 556]. behaviour [1756]. behavioural [4, 1100]. Behaviours [2532, 1259, 1080]. Being [1693]. Belief [2530, 627, 2241, 1728, 2276, 2280, 2282, 2240]. Benes [1363]. Benes-based [1363]. Benign [1657]. Bent [2082]. Bernstein [2174]. Best [1417, 1111]. Best-Effort [1111]. Best-First [1417]. Better [946, 2117, 2111, 349]. Between [1012, 1163, 1900, 2188, 1027, 922, 876, 2265, 1033, 413, 1106, 1710, 5, 713, 1063, 1658, 350, 1503, 1003, 2080, 1420]. Beyond [1833, 1943, 1113, 1006, 2494, 1901, 2215]. Bias [2359]. Bibliography [2472]. Bichromatic [971]. Bildvorlagen [1910]. Bimodal [2252, 488]. Bin [1693]. Binary [485, 1653, 948, 1029, 2088, 1384, 526, 2583, 1050, 223, 206]. Binding [1818, 462, 178]. Binomial [2555]. Biological [561]. Biologically [459]. Biology [1709]. Biophysics [449]. Bioreactor [554]. bipartite [210]. Birth [2392]. Bisection [1697, 1106]. Bisimulation [1240, 779, 1238, 1239, 1080]. bisimulations [1046]. Bit [1689, 114, 850, 264, 1366]. Bites [1685]. bits [84]. Blind [542]. blinde [1910, 1912]. Block [466, 2432, 1448, 1697, 90, 1390, 1338, 1375]. Blocks [319, 1228]. Blood [532]. Blue [993]. BMAP [1792]. BMAP/G/1 [1792]. Bodies [2255]. body [747]. Boolean [1667, 83, 256, 2021, 1032, 2172, 2197, 1343, 626]. Boolean-valued [2197]. borrow [1067]. BOS [426]. BOS-Method [426]. Bottleneck [1695, 1798]. Bottom [2488, 1202, 2183]. Bottom-Up [2488, 1202, 2183]. Bound [954, 1391, 1697, 2525]. Boundaries [1525, 567]. boundary [1090]. Bounded [1217, 676, 633, 930, 8, 314, 2524, 1186, 738, 1188, 1402, 2582, 2166, 2536, 831, 750, 67, 221, 235, 266, 194]. Bounds [753, 962, 1817, 129, 2065, 746, 1008, 2075, 2173, 2064, 1666, 57, 240, 84, 130, 197]. Box [1170, 2322, 215]. Boxes [2021, 1345]. Brain [1526, 578]. Branching [1248, 2582, 256, 1080, 1235]. break [111].

Bridging [791, 876, 2401]. Brightness [1395]. Brightness-Contrast [1395]. Broadcast [2572, 764, 1665, 1243]. Broadcasting [765]. broadcasts [1242]. Brokers [1501]. Browsers [1909]. Browsing [1879]. Bruijn [1360]. Buchi [760]. Bufferless [1678]. Build [1840]. Building [373, 1746, 319, 816, 606, 1715, 1170, 1610]. Built [280, 1191]. Built-in [280, 1191]. Burstall [1044]. Bus [1932, 1140]. Buses [1701, 1703]. Business [2350, 885, 1302, 2401]. Business-Oriented [2350]. Butterfly [2570].

C [921, 2424, 1851, 506, 823, 666, 1818, 2439, 2423, 2130, 1850, 1610]. C-KAT [1610]. C.I.M [428]. Cable [455]. CAC [1600]. CAD [651]. CAD/CAM [651]. CADI [1483]. Calculating [272]. Calculation [1966, 1384, 1973, 1513]. Calculi [1177, 2277, 1964, 1048, 1047, 184, 1096, 198, 208]. Calculus [2506, 280, 812, 813, 1603, 1978, 2011, 327, 786, 683, 1593, 1688, 43, 2183, 2437, 778, 2265, 26, 2329, 164, 1007, 2271, 1191, 919, 2151, 171, 1238, 1050, 176, 215, 1230, 178, 207, 1073, 180, 1231, 1241, 189, 2409, 1093, 1239, 1103, 1265]. Call [1957, 196]. Call-By-Name [1957]. Call-by-value [196]. CAM [651]. Camera [1433]. Can [1607, 2117, 2187, 2077, 120]. Cancellation [222]. CANDID [1550]. Candidate [122, 1537]. Canonical [55, 2574, 725]. Capabilities [2191, 245, 1565]. Capability [859]. Capacities [1781, 1795]. Capacity [2572]. Capsule [1326]. Captured [40]. Cards [2037]. Cartographic [1274]. cascade [1367]. Case [2094, 393, 2370, 338, 2556, 2184, 608, 149, 408, 773, 347, 1563, 2406, 213, 263, 448, 1559, 1563]. Case-Based [393, 2184, 608, 149]. Cases [286]. Cash [2117, 121]. Castings [1531]. Categorical [812, 41, 1281, 1044]. Categories [780, 708, 10]. categorization [718]. Causal [200, 2257, 2262, 2497, 2270]. Causality [1863, 720]. Causation [2382]. cause [1087]. caused [1237]. Cayley [1036]. ccc [171]. ccc-calculus [171]. CCS [322, 2152]. Cell [1837, 454]. Cells [1844]. Cellular [2023, 1017, 843, 541, 1494, 1036, 1076]. Center [653]. Centered [2136, 625, 2227]. Centering [451]. Central [2540]. Cerebral [1464, 570, 456, 1470]. CERISE [1618]. Certain [959, 2087, 63]. Certifying [2094]. CFP [1767]. CG [718]. CGMA [2395]. Chain [2594, 1410, 701, 229]. Chain-Coded [1410]. Chain-like [2594]. Chaining [385]. Chains [945, 101]. Challenge [1495, 1580, 2367, 868, 903]. Challenges [1829, 1636]. Chameleon [851]. Change [2143, 2129, 746]. Change-Making [746]. Changeling [1971]. Changes [1307, 2166, 266]. Channel [2590, 1132, 1346]. Channels [542, 1350]. Chaos [359]. Chaotic [1955, 1630, 1219]. Chapman [1734]. Character [1453, 547, 2163, 1456]. Characteristic [767, 1357]. Characteristics [575, 2099, 2098]. Characterizable [2194]. Characterization [158, 1378, 725, 1061, 70, 187]. Characterizations [763, 2193, 2012, 831, 189]. Characterizing [1749, 573]. Characters [1455]. Charts [1018]. Cheater [2028]. Checked [1025]. Checking [734, 51, 678, 680, 782, 2338, 645, 1232, 1656, 1260]. Checklist [403]. checksum [1358]. Chemical [400, 520, 1416]. Chervonenkis [2199]. Chest [533]. Chip [515, 1490]. Chips [519]. chocs [1238]. Choice [2476, 1196, 1748]. Choosing [1919]. chordal [210]. Chosen [2083]. Chromosomes [1471]. Chunks [1561]. CIM [1554, 728]. Cipher [2033, 92, 93, 1375]. Ciphers [90, 1358]. Ciphertext [2020, 2083]. Circle [975, 2553, 1404, 1374]. Circuit [115, 2589, 1008, 1702, 116, 96]. Circuits [254, 1014, 839, 1870, 2333, 846, 2197, 276, 253, 252]. Circular [1675, 966, 775]. Circular-Arc [1675, 966]. Circumscription [2501]. Clash [369]. Class [372, 1172, 815, 2498, 1497, 1423, 2145, 1984, 2149, 1722, 329, 2150, 274, 1105]. Class-Based [2145]. Class-less [1497]. Classes [955, 2541, 2568, 917, 315, 1527, 824, 2500, 2492, 2582, 2192, 2490, 831, 503, 66, 63, 1052, 16, 1745]. Classical [1173, 1299, 413, 1168, 1169, 1190, 174]. Classification [1436, 1454, 533, 87, 1423, 611, 296, 532, 594, 503, 595, 1471, 535, 1610, 1767, 1236]. Classifier [468, 467, 529]. Classifiers [2162, 537, 1492, 502]. Classifying [464, 762]. Clausal [697]. Clause [1594, 206]. Clause-Diffusion [1594]. Clauses [677, 33]. Claw [2025]. Client [1840, 1119, 1842, 1794]. Client-Server [1794]. Client/Server [1840, 1119, 1842]. Clinical [1906]. Cliques [1028]. clock [1359]. clock-controlled [1359]. Clocks [1981, 919]. CLOG [2145]. Closed [763, 1798, 708, 1234]. closest [227]. Closure [2036, 1704, 247]. CLP [156, 1592, 2518]. Clustering [970, 1744, 1825, 1690]. Clusters [2207]. CM [353, 1634]. CM-2 [1634]. CM-5 [353]. CML [1263]. CNF [603]. CNN [1490]. Co [757, 1763]. Co-Finite [757]. co-operating [1763]. Coal [1475]. Coarse [971, 1424]. Coarse-to-Fine [1424]. CoBase [1506]. Cobham [760]. Cocomparability [2595]. Code [1229, 506, 1945, 2040]. Coded [1410, 1813]. Codes [1005, 2563, 2030, 362, 1530, 1338, 1335, 1333]. Coding [1391, 1406, 1428, 1390]. coefficients [234]. Coercion [2008]. Cognitive [2389, 648, 1278, 1423, 2210, 1267, 1268]. Coherence [1440, 1582]. Coherent [2368, 2366]. cohesion [1771].

Collaborative [1322, 1914]. Collages [1268]. Collapsed [23]. Collective [2433, 356]. Collision [88, 2093, 2107, 1340]. Collision-Free [88, 2093, 2107]. Collisions [2057]. Color [1397, 1435]. Colored [318, 2204]. Coloring [2545, 2550, 1711, 62]. Colour [508, 741, 851, 1439]. Coloured [315]. Combatting [2073]. Combination [343, 335, 1422, 634, 561, 402, 534]. Combinations [787, 22, 2201, 24, 2538, 2513, 1375]. Combinatorial [339, 972, 2563, 997, 2580, 74]. combinators [1104]. Combine [1682]. Combined [26, 1972, 1428, 588]. combiners [96]. Combining [175, 1591, 290, 188, 1635, 611, 1950, 1719, 346, 1272]. commands [1058]. Comments [2069]. Commitment [1504, 114]. Committed [1196]. Committed-Choice [1196]. Commodity [984]. Common [2095]. Communicating [819, 1258, 1058]. Communication [753, 961, 2224, 2470, 1153, 443, 2225, 1145, 540, 1921, 2523, 1521, 1113, 2450, 1152, 354, 1960, 2301, 1698, 846, 356, 1661, 2078, 1234, 1086, 1243, 1082, 1084, 84]. Communications [2119, 106]. Community [1829]. Commutations [1002, 770]. Commutative [29, 345, 225]. Compact [736, 1598, 2522]. Compaction [2589, 998]. Comparability [328]. Comparative [817, 1400, 499, 1058]. Comparing [1834, 572]. Comparison [954, 1465, 490, 413, 1835, 701, 1416, 1709, 713, 714, 1374]. Comparison-Based [954]. Comparisons [1679, 572]. Compatibility [770]. Compatible [627]. Compensation [1823]. Competition [483]. Competitive [1674, 2587, 742, 2165, 478, 530]. Compilation [2450, 2416]. Compilations [1970]. Compile [1203]. Compile-Time [1203]. Compiled [2142]. Compiler [2419, 1220, 2430, 2418, 2441, 922, 2446, 2422]. Compilers [1229, 2449]. Compiling [291, 602, 1634, 2446]. Complement [345, 250]. Complete [974, 1720, 708, 830, 219, 1080, 822, 740]. Completely [448]. Completeness [370, 815, 1077, 820, 22, 1191, 23, 2513, 46, 1234]. Completion [2007, 689, 47, 470, 41, 233]. Complex [1045, 2486, 835, 2477, 2355, 2524, 1712, 2136, 548, 1874]. complex-like [1045]. Complex-Object [2486]. Complexities [1026]. Complexity [137, 753, 954, 959, 2160, 2568, 967, 2531, 491, 1221, 1689, 1023, 1767, 2523, 749, 2044, 1383, 1698, 740, 1167, 1723, 2582, 2077, 98, 1008, 747, 2197, 645, 1705, 2199, 997, 2173, 2198, 2583, 949, 880, 1062, 230, 1076, 209, 77, 1079, 1082, 64, 242, 1084, 245, 1052, 264, 263, 262, 1760]. Component [950, 2372, 380, 1538, 1159, 1433, 1325, 2127, 709]. Component-Based [709]. Components [1329, 2467, 906, 2347, 71]. Components. [318]. COMPOSE [2452]. Composite [1477, 1476, 1831, 2452]. Composition [1555, 1586]. Compositional [2424, 1982, 1950, 330]. Compositionality [162]. Compound [1419]. Compounds [1416]. Compression [598, 551, 1389, 234, 1072]. Computability [236, 1193, 1074]. Computation [2007, 1173, 424, 2001, 2085, 1587, 2309, 2566, 1585, 2308, 752, 1447, 2510, 1188, 2054, 449, 1209, 361, 1933, 1581, 767, 1944, 48, 1602, 1420, 288, 1583, 1084, 1349, 1355, 99, 262]. Computational [1135, 2006, 1379, 1576, 597, 756, 2473, 751, 2451, 1546, 64]. Computations [1996, 1932, 1857, 2567, 85, 245]. Computatrix [450]. compute [239]. computed [120]. Computer [1460, 2369, 2213, 442, 855, 1574, 1163, 2108, 2111, 1484, 1995, 1900, 1586, 1378, 2504, 1483, 2365, 1462, 903, 2347, 856, 2346, 634, 1466, 1624, 1922, 845, 1894, 544, 1588, 864, 2109, 1582, 2008, 1474, 2373, 1596]. Computer-Aided [1460, 442, 634].

Computer-Assisted [1466]. Computer-Related [2108]. computergestutzte [1920]. Computers [2017, 1003, 360, 1873, 1300, 1874, 1701, 1925, 362, 520, 346]. Computes [2476]. Computing [783, 2304, 675, 945, 2305, 1584, 2576, 1851, 975, 1943, 2284, 37, 2427, 354, 2285, 1696, 1747, 903, 1600, 351, 697, 363, 1424, 1928, 2199, 2303, 1842, 512, 252, 901, 1283]. Concentrate [878]. Concept [955, 289, 319, 659, 1861, 1864, 1889, 1720, 1893, 292, 645, 1909, 2207, 1346, 1758]. Conception [2368]. Concepts [396, 2389, 1121, 1180, 1576, 2177, 431, 625, 2357, 9, 12]. Conceptual [438, 2389, 717, 2385, 2379, 1277, 2391, 1744, 2398, 2392, 2460, 2376, 2395, 1499, 2382, 423, 2397, 437, 2381, 2386, 1498, 2401, 434, 2399, 2335, 440, 728, 2396, 2402, 2390, 1610, 729, 713, 715, 716, 731, 730, 721, 714, 725, 720, 722, 732, 726, 733]. Concerning [759]. Concurrency [788, 776, 1320, 1, 7, 1659, 1089, 1236, 1100, 241]. Concurrent [2509, 1539, 885, 2426, 785, 2317, 1022, 2325, 915, 1972, 163, 2295, 2417, 654, 2307, 777, 1196, 1658, 200, 1652, 1092]. Condition [51, 1655]. Conditional [32, 54, 2247, 43, 50, 25, 22, 24, 35, 155, 2277, 668, 386, 31, 45, 198]. Conditionals [386, 2279]. Conditions [404, 323, 25, 1434, 2445, 988, 2093, 2032, 1644]. Conductance [592]. Cone [1452, 588]. Cone-Beam [588]. Conference [126, 2038, 1368, 1369]. Conferences [2096]. Configurable [1126, 1926]. Configuration [2196, 1998, 1777]. Confirmation [1545]. Conflict [1883, 238]. Conflicting [2343]. Conflicts [1440, 2544, 307, 1511, 82]. Confluence [2506, 787, 23, 46, 45]. Confluent [786, 191]. Congruence [979, 1262, 1648, 1080]. Congruences [1956, 257]. Conjecture [1681]. Conjugacy [334]. Connected [1507, 950, 958, 470, 1701, 72]. Connection [1177, 1954, 353, 1149]. Connectionism [518]. Connectionist [648, 484, 544, 472]. Connectionist-Symbolic [648]. Connections [1163, 464]. Connectivity [2546, 2586, 1694, 456]. Conquer [1682, 2562, 745, 364, 2549, 217]. Consecutive [2591]. Consensus [640, 612]. Consequences [622, 2550]. conservation [182]. Conservative [2515, 1200, 181]. Considerations [585, 837]. Considering [1929]. Consistency [21, 2135, 680, 33, 2254, 1183, 2129, 2338, 634, 1775]. Consistent [415, 2198]. Consisting [1016]. Conspiracy [126]. Constant [340, 2044, 1093]. Constrained [47, 631, 2273, 1776, 131]. Constraint [633, 632, 785, 1202, 2246, 2317, 1730, 768, 307, 1999, 1200, 1732, 696, 2519, 672, 1549, 1742, 578, 1775, 1757, 1777]. Constraint-Based [307]. Constraints [674, 1731, 884, 1136, 1716, 327, 1519, 2339, 799, 661, 1514, 980, 1201, 447, 153, 48, 2147, 2487, 249, 1105]. Construct [279, 1325, 2024, 1619, 109]. Constructed [2563, 1374]. Constructing [1608, 2554, 2543, 1345, 742, 1765, 2549]. Construction [140, 963, 2238, 2062, 1559, 2033, 1159, 2574, 1365, 2021, 894, 1600, 1616, 94, 110, 2032, 1193, 1360, 1338, 273, 1247, 1098]. Construction-Defined [1600]. Constructions [2065, 2064]. Constructive [642, 468, 1966, 502, 155, 2278, 174]. Constructor [1600, 22, 31]. Constructor-Based [31]. Constructs [1321, 1226, 2527]. Contained [457]. Content [1157, 2106, 1056, 174]. Content-Addressable [2106]. Content-Based [1157]. Contents [2223, 1543]. Context [1496, 2389, 1015, 866, 649, 1500, 782, 18, 1006, 869, 1031, 2211, 2335, 773, 222]. Context-Free [1015, 782, 1031, 773, 222]. Context-Freeness [1006]. Contexts [168, 1590, 1589]. Contextsensitive [1881].

Contextual [44, 1211]. Contextualizing [2343]. Contingency [497]. Continuation [2290]. Continuations [2150]. Continuous [1845, 1153, 316, 1118, 1806, 1146, 1156, 594, 186]. Contour [1425, 1399]. Contour-Detection [1399]. Contours [1462, 1410]. Contraction [2433, 1699]. Contractions [1738]. Contradiction [1725, 1545]. Contradictory [288]. Contrast [1558, 1395]. Contribution [1529]. Control [1992, 2324, 1993, 1473, 1477, 554, 1615, 1440, 1153, 866, 2374, 1884, 2435, 1552, 412, 1880, 1521, 1475, 1151, 1857, 2286, 1988, 1124, 1554, 397, 1989, 557, 405, 607, 1117, 1804, 592, 174, 1375]. control-like [174]. Control-Parallel [2435]. Controllability [1990]. Controlled [623, 454, 1359]. Controller [553, 540, 850, 556, 848]. Controllers [552]. Controlling [287, 1659, 2205]. Controls [2453, 1476]. Conventional [2507, 361]. Convergence [485, 1031, 97, 495]. Converting [1839, 1842]. Convex [974, 1020, 977, 2574, 1408, 2268, 999]. Convolution [1393]. COO [1323]. cooperating [1082, 1323]. Cooperation [1551, 1892, 2127, 305]. Cooperative [1506, 1321, 2010, 1823, 2364, 1772]. Cooperativeness [1528]. Coordination [920, 2298, 1323]. Coping [1773]. Coq [193]. Coreference [2376]. Corners [1409]. Corona [1481]. Corpora [1535]. Corpus [1291]. Correct [153]. Correction [1401, 97]. Correctness [1575, 1220, 2468, 1952, 1640]. Correlated [595]. Correlation [96, 1450, 131]. Correlations [2039, 1366]. Correspondence [497, 487, 1448, 544, 1052]. Corresponding [2066, 81]. Corridors [1914]. Corrupted [582]. Cortex [456]. Cosine [2041, 1389]. Cost [352, 1505]. Could [40]. Counter [759, 235]. Counting [223, 2540, 821, 978, 221, 993, 1668]. Coupled [455, 1441]. Coupling [2345, 1672]. Couplings [1205]. Courseware [1923]. Cover [739]. Coverability [321]. Coverage [772, 1281]. Covering [968, 1028, 986]. CPWM [515]. Crash [1661, 2231]. Crater [1487]. Create [1290, 503, 1049]. Creating [2467, 1902]. Creation [2289]. Creative [642]. Creativity [1922, 1924, 1752]. Crest [1398]. CREW [961, 1683, 1088, 1935]. CREW-PRAM [1683]. Criteria [308, 2115, 2312, 658, 1710]. Criterion [1734, 1917, 2205, 1343]. Critical [1166, 1456, 1522, 41, 1294]. Critical-Pair [41]. Critical-Pair/Completion [41]. Crochemore [1679]. Crochemore-Perrin [1679]. Cross [1450]. Cross-Correlation [1450]. Crossing [2552]. Cryptanalysis [2097, 2019, 2023, 1348, 2056, 1347, 2018, 2060, 2020, 2105, 2038, 91, 92]. Crypto’91 [2057]. Cryptographic [113, 2094, 2566, 2061, 2581, 2116, 89, 2051, 1358, 1354]. cryptographically [1345]. Cryptography [2121, 2110, 2092, 2052, 1367]. Cryptology [2031]. Cryptosystem [2056, 133, 1370]. Cryptosystems [2068, 100, 2088, 2035, 2029, 2072, 2055, 1332, 1371]. CSP [2534]. CT [549, 1468]. CTL [1260]. Cube [2559, 2086]. Cultural [1280]. Culture [894]. Cumulative [1336]. Curricula [901]. Curvature [930]. Curve [102, 2055, 2067, 133]. Curves [2510, 2087, 1371]. Customisable [1572]. Customizable [2237, 2306]. Customized [1502]. Cut [810, 685, 2126, 1106]. Cut-Free [810]. cuts [71]. Cutset [1842]. Cutting [1418]. Cycles [798, 1107]. Cyclic [19, 1618]. Cylindrical [2005].

D [1873, 1753, 1464, 784, 973, 1076, 1017, 1875, 570, 1028, 1446, 942, 1441, 938, 2131, 1282, 568, 1420, 588, 2422]. D-Graphics [1875]. D-Unification [784]. Damgard [2023]. Data [1526, 1388, 508, 2389, 2323, 1123, 2425, 2431, 1277, 2480, 1071, 1500, 669, 352, 570, 976, 1443, 1521, 874, 1807, 1560, 2167, 2296, 1468, 354, 1816, 1818, 2294, 1524, 1384, 2178, 2511, 2436, 1888, 2434, 2353, 1597, 2423, 1124, 1821, 1971, 1806, 2494, 1548, 1492, 1566, 2443, 1937, 433, 941, 331, 2444, 2557, 1169, 1871, 420, 2420, 2428, 590, 2461, 582, 1530, 1707, 1470, 1281, 1564, 588, 2022, 657, 1512, 1275, 234, 81, 271, 1778, 14, 2415]. Data-Flow [2443]. Data-Independences [2511]. Data-Parallel [2436]. Data/Knowledge [433]. Database [1507, 2337, 1568, 2469, 912, 2134, 417, 2453, 156, 1817, 618, 2470, 1506, 2221, 1503, 2391, 1739, 616, 1519, 1559, 669, 799, 2471, 1829, 666, 1570, 1573, 2478, 2491, 801, 1310, 805, 2494, 2497, 1505, 2496, 1569, 1515, 415, 1926, 1572, 2484, 672, 1517, 1823, 670, 1522, 2342, 667, 446, 1564, 1516, 1512, 1097]. Databases [1810, 2486, 610, 1495, 2137, 665, 662, 791, 2070, 2459, 2493, 2498, 1309, 2339, 2462, 1527, 2474, 1557, 1829, 1528, 2145, 2341, 1523, 2502, 1525, 1508, 2338, 1879, 671, 1913, 2340, 2483, 663, 1509, 668, 2465, 418, 2454, 2390, 2482, 660, 2207, 1513, 717, 1660]. Dataflow [1888, 2446, 1255]. Datalog [2476]. datatypes [2404, 1645]. DBMS [1529, 1518, 1824]. DCE [1834, 1833, 1836, 1851, 1852, 1848, 1838, 1837, 1850, 1841, 1844, 1835, 1849, 1843, 1842]. DCE-based [1841]. DCG [682]. DCG-Terms [682]. DCGs [1204]. DDH [1812]. Deal [53, 2287]. Dealing [427]. Debugger [2292]. Debugging [2323, 1331, 1539, 2311, 2317, 2327, 2332, 1199, 2321, 2333, 2329, 2319, 2314, 1725, 2330, 2312, 2328, 2315]. Decidability [1011, 207, 759, 36, 781, 49, 1009]. Decidable [1984, 1033, 1240, 1745]. Deciding [372, 688, 776, 1102]. Decision [369, 955, 1172, 2243, 684, 635, 2206, 300, 2260, 344, 1589, 2517, 2274, 626, 658, 2583, 2280, 1666, 83, 1108]. Decision-Making [2260, 2280]. Declarations [699].

Declarative [2323, 154, 2313, 2317, 1226, 2496]. Decline [1299]. decomposing [1765]. Decomposition [1538, 2005, 326, 2131, 2454, 586, 2207, 1245, 1083, 73, 232]. Decompositions [86]. Deductibility [2276]. Deduction [1175, 1594, 142, 409, 2013, 165, 283, 701, 216]. Deductive [2479, 2486, 2495, 156, 141, 2477, 791, 1739, 2493, 2503, 669, 2474, 679, 2502, 1562, 434, 2483, 2484, 697, 668, 2390, 2482, 660]. Deep [1884]. Default [675, 1727, 141, 380, 1728, 2245, 641, 604, 1166, 2256, 2263, 148, 2267, 381, 696, 2273, 2278, 2279, 2281]. Defaults [2233, 2234, 2488, 142]. Defeasible [2241, 382]. Defence [1883]. Deferred [966]. Deferred-Query [966]. Definability [761, 821, 825, 248, 247, 187]. definable [246]. Defined [874, 1600]. Definiens [170]. Defining [2091, 604, 1069, 208]. Definition [910, 2132, 714]. Definitions [21, 193]. Deformations [564, 943]. Defying [220]. Degeneracies [748]. Degree [2546, 2543, 1189, 1768]. Degree-3 [2543]. Degrees [2530]. Delaunay [748]. Delays [1198, 1698]. Delegated [304]. Delete [2547]. Deletion [1507]. Delinquency [2108]. Deliverables [1044]. Delivery [1111, 1127, 1598]. delta [1174]. delta-Rule [1174]. Demand [1209, 1803, 1117]. Demandedness [1208]. Demands [1706]. Demonstration [1651]. Dempster [2281]. Dempster-Shafer [2281]. Demytko [1364]. Demytko-Miller-Trbovich [1364]. Dendritic [455]. Dense [1004, 1448, 1258, 1102]. density [205]. Dependability [887, 1802, 1635]. Dependence [2443, 2420, 2417, 1751]. Dependencies [2175, 2242, 2341, 2497, 657]. Dependency [1555, 2149, 1630]. Dependent [2551, 1807, 2217, 185]. Depth [756, 982, 2197, 833, 2599]. Depth-First [982]. Derivation [1739, 1971, 2408, 331]. derivations [79, 1723]. Derivative [1397]. Derivatives [564]. Derived [2492, 1512]. Deriving [1397, 1889, 425]. DES-based [1363]. DES-Like [2021, 2035, 2106]. Description [703, 2238, 704, 570, 2345, 1915, 1983, 1470, 1830, 706, 1405, 71, 1363]. Descriptions [2532, 1267]. Design [513, 2521, 857, 2305, 2379, 1906, 653, 2221, 2361, 2023, 1867, 1321, 1812, 1555, 1141, 923, 1712, 1852, 1159, 1784, 1497, 850, 1595, 1982, 470, 1518, 1138, 436, 1888, 358, 1562, 1118, 2348, 2130, 1624, 444, 2356, 834, 1914, 502, 2601, 415, 447, 2295, 1898, 526, 837, 1944, 2081, 1137, 2342, 2458, 625, 1623, 1319, 1564, 500, 1116, 1368, 1241, 243, 226, 1340, 2414, 1369, 293]. Designed [1612]. Designing [1965, 2375, 926, 2068, 2441, 276, 984, 1901, 2351]. Designs [466, 687, 896]. Desk [2011]. Desk-Top [2011]. Detecting [2068, 1409, 988, 1971, 1091]. Detection [1487, 2575, 1177, 1427, 1558, 1318, 2391, 2452, 1401, 2028, 1410, 2444, 1398, 1400, 596, 1396, 1399]. Detector [2231]. Determination [999]. Determined [875]. Determining [2429, 2191]. determinism [7]. Deterministic [2476, 260, 584, 2532, 752, 1697, 853, 408, 1100, 1039, 329, 2454, 1068, 1073]. Developed [906]. Developer [651]. Developers [1323]. Developing [1522, 1763, 1636]. Development [452, 1547, 640, 1460, 426, 448, 884, 644, 1314, 427, 669, 2129, 376, 1571, 609, 2009, 388, 2155, 1222, 862, 870, 1961, 445, 1274, 400, 432, 883, 2352, 1563, 133, 268, 1044, 2409, 2411]. Developments [1560, 2580]. Device [1889]. Devnet [317]. DFA [2160]. DFN [1835]. DFN-RPC [1835]. Diagnosing [1896, 2314]. Diagnosis [1460, 2316, 2313, 2333, 608, 1725, 1479, 1537]. Diagram [1685, 2520]. Diagrams [2528, 2583, 710]. Dialect [2439]. Dialectica [1164]. Dialog [1880]. Dialogue [1883, 1888, 1897, 1907]. dialogues [1772].

Diameter [366]. Dictionaries [2494]. Dictionary [2539, 1535]. Diekert [1045]. Difference [2059, 2383]. Differences [1907]. Different [284, 1591, 494, 1003, 1180, 1927, 851, 1465, 1446, 2219, 1398, 2277, 1375]. Differential [91, 2097, 2019, 455, 2475, 2060, 2020, 2105, 569]. Differentiating [1175]. Difficult [1687]. Difficulties [2393]. Diffie [106, 1372]. Diffuse [1433]. Diffusion [1594, 1395, 1390, 573, 592]. Digit [469]. Digital [1129, 2575, 1460, 1485, 1148, 755, 1483, 1488, 1111, 2333, 1119, 2027, 846, 2067, 2113, 1803, 1481, 2069, 1283, 1341, 2058]. Digitized [2510]. Digitizing [1293]. Digraph [1704]. Digraphs [1677, 1707, 2204, 75]. Dimension [2199, 2197]. Dimensional [973, 536, 976, 1444, 1691, 1933, 2356, 2086, 932, 582, 998, 1918, 68, 236, 1250]. Dimensionality [2218]. Dimensions [429, 2203]. Diophantine [1014, 1733]. Direct [552, 158, 1681, 1683, 1913]. Directed [1694, 153]. Direction [1432]. Directional [1397, 2542, 166, 1295]. directionality [1641]. Directory [1837]. Discipline [861]. Discontinuities [1431]. Discontinuous [295]. Discourse [2381, 711]. discourses [719]. Discovering [2335, 2207, 1513]. Discovery [1736, 2156]. Discrepancy [2080]. Discrete [2047, 317, 2041, 2068, 2085, 754, 127, 2363, 1389, 1791, 477, 751, 1341, 1370, 123]. Discrete-Time [477]. Discriminant [2195]. Discriminator [481]. Discriminator-Based [481]. Discs [997]. Discussion [865, 2002, 872, 882, 2095, 897, 902, 890]. Diseases [1462]. Disenrollment [2102]. Disentailment [674]. Disjoint [343, 1691, 2578, 1705, 1708]. Disjunction [617]. Disjunctions [1201]. Disjunctive [1004, 599]. Disk [1828, 1144]. Disparity [1428]. Displacements [1430]. Display [536, 1289]. Distance [1457, 2384, 464, 942, 1295, 1272, 63, 131, 1359]. Distillations [2288]. Distributed [909, 1845, 1551, 2318, 921, 1317, 348, 1128, 312, 1501, 640, 1834, 1216, 1833, 1135, 2229, 1594, 426, 2470, 2431, 1318, 926, 1662, 2438, 1856, 1503, 1867, 1812, 1552, 1851, 764, 2447, 1313, 1713, 2464, 1872, 1689, 1863, 322, 1784, 2152, 1858, 1227, 1714, 2296, 1981, 2439, 1971, 1839, 1505, 1931, 2272, 1871, 2295, 1146, 663, 919, 1724, 1830, 1849, 2300, 1846, 512, 216, 1672, 200, 70, 243, 1670, 1671]. Distributing [274, 1850]. Distribution [766, 2096, 2239, 1865, 1486, 354, 126, 2038, 1704, 488, 495, 1065, 1368, 128, 1374, 1350, 1372, 1369]. Distributions [536, 754, 1785, 1452]. Distributivity [337, 698]. Disunification [48]. Divergent [53, 1922]. Divertibility [114]. Divertible [2049]. Divide [2562, 745, 217, 364, 2549]. Divide-and-Conquer [2562, 745, 364, 2549, 217]. division [245]. DNA [1687, 2202]. Do [2393]. Doany [2442]. Document [1680, 2375, 1915, 1808, 1543, 2466]. Document-Driven [2375]. Documentation [1995, 2332]. Documents [1136, 1291, 1569, 1902]. Does [1023, 2383]. Dog [1685]. Doing [1574]. Dokumentation [1920]. DOL [775]. Domain [905, 1603, 764, 1306, 634, 1581, 472, 1773, 723, 1074]. domain-independent [723]. Domains [788, 2006, 1202, 1947, 1250, 1600, 2163, 1535]. Dominance [1683]. Dominant [1410]. Domination [958, 2595, 1000]. Domination-Like [1000]. Doors [1914]. Dot [1494]. Double [1181, 2555, 79]. Double-Ended [2555]. Down [1011, 1204, 1039, 1766]. downward [220]. Draft [2072]. Drawing [2574]. Drawings [960, 1459, 1458, 1416, 80]. Driven [2375, 1599, 303, 1788, 1209, 444, 607]. Drops [1815]. DSC [1580]. DSM [502]. DSPNexpress [1790]. DSS [1571]. Dual [145, 739, 1381]. Duality [2188]. Duplicate [1507]. Duration [1978, 1964, 207]. During [427, 643, 447, 2135, 2589]. dyadic [1264]. Dynamic [916, 2591, 735, 474, 956, 2096, 1607, 1198, 678, 1153, 1682, 970, 1812, 2544, 1593, 1712, 1819, 1848, 1787, 1488, 1692, 1693, 693, 985, 2259, 490, 2142, 2288, 2557, 2420, 2428, 590, 2421, 1707, 2451, 113, 2405, 1661, 1663, 1051, 1630, 188, 14, 1099]. Dynamical [1987, 412, 236]. Dynamically [1126, 1049]. Dynamically-Configurable [1126]. Dynamics [1179, 498, 1930]. Dynamism [1904]. Dynamization [978].

E-Matching [1729]. E-Unification [54, 1184]. ear [1083]. Early [2154, 508, 178]. Ease [838]. Easily [1025]. Ecological [1906, 2211]. Edge [1030, 1400, 1708, 1711, 1396, 71]. Edge-Coloring [1711]. Edge-Disjoint [1708]. Edge-Replacement [1030]. EDI [2122]. edit [131]. Editing [1591, 1289]. Editor [1899, 1808, 834, 1923]. Editors [2349, 1909]. education [1750]. Educational [1483]. EEC [2121]. EEC-RACE [2121]. EEG [563]. Effect [2427]. Effective [801, 1933, 773, 901]. Effectively [2367]. Effectiveness [352]. Effects [824, 2294, 1907, 2287, 1386, 2217, 595]. Efficiency [969, 1892, 1861, 597, 1346, 1629]. Efficient [1810, 140, 2182, 1676, 1060, 267, 1955, 966, 2426, 765, 785, 1202, 234, 2047, 973, 974, 1018, 724, 755, 118, 1083, 2167, 1652, 1228, 1656, 300, 2054, 2087, 992, 355, 1126, 1208, 943, 2067, 2538, 1702, 89, 481, 2307, 2465, 288, 2487, 119, 2406, 76, 1669, 1659, 1647, 226, 271]. Effort [1111]. EFH [1750]. EFH-Soar [1750]. Ego [1426]. Ego-Motion [1426]. Ehrenfeucht [814]. Ehrenfeucht-Fraisse [814]. Eigenvectors [1397]. Elasticity [2192]. elections [1351]. Electrical [579]. Electronic [1916, 2017, 2071, 1353]. Elementary [138]. Elements [2404, 2350, 2540, 1530]. ELF [1486]. Elicitating [428]. Elicitation [729]. Elimination [280, 692, 1259]. elitist [1768]. Ellipse [1401]. Elliptic [102, 2088, 2087, 1371, 2055, 2067, 133]. Embedded [1692, 1808, 447]. Embedding [159, 1226]. Embeddings [365, 967, 74, 572]. Emerging [1622]. Emission [584]. Empirical [884, 2344, 143, 1924]. Empirically [875]. Emulations [349]. Enabling [1623]. Enactment [1328]. Encapsulating [907]. encapsulation [1071]. Enclosing [975]. Encoded [1439]. Encoding [842]. Encryption [2060]. Encryptions [2040]. End [1224, 1886, 1623]. End-User [1886]. End-Users [1623]. Ended [2555]. Endomorphic [2413]. Energy [516]. Enforcement [2468, 672]. engine [730]. Engineer [406]. Engineering [289, 858, 905, 286, 885, 1302, 1309, 861, 1326, 1969, 1497, 643, 886, 1306, 875, 876, 877, 868, 2148, 285, 619, 895, 878, 1624, 1616, 1968, 429, 1316, 297, 1799, 654, 879, 883, 891, 1794, 888]. Engines [1623, 2106]. English [2398]. Enhanced [1603]. Enhancement [556]. Enhancing [1129, 1909, 1272]. Enriching [662]. enrichment [222]. Ensembles [2423]. Entailment [674, 2519]. Entities [2499]. Entity [666, 1284]. Entity-Relationship [666]. Entrenchment [1165]. Entropies [627]. Entropy [382, 2165, 1789]. Enumeration [1869]. Envelopes [2584]. Environment [1737, 1965, 398, 1311, 2372, 644, 1292, 1586, 1927, 850, 1571, 1328, 2148, 1562, 2130, 1119, 845, 1946, 2455, 1137, 654, 1830, 510, 1563, 512, 2297, 1671]. Environments [1834, 2368, 1896, 1309, 1559, 1945, 2129, 1554, 445, 1122, 1853, 1312, 2326, 505, 1878, 936]. Epistemic [1165, 2267]. epoch [1660]. EPOS [1327]. Equality [1172, 1079]. Equalization [542, 593]. Equation [1010, 455]. Equational [1223, 21, 1013, 1593, 1032, 31, 1266, 10]. Equations [340, 336, 1733, 1945, 455, 2059, 333, 2525, 342, 1602]. Equivalence [1011, 313, 1038, 1240, 235, 4, 1102, 258]. Equivalences [1043, 776, 1087]. Equivalent [2512, 815, 2079]. Erasing [1033]. EREW [961]. Ergonomics [1905]. Erroneous [395]. Error [2316, 1519, 2313, 1971, 1390, 2198, 97]. error-correction [97]. Error-Detecting [1971]. Errors [2320]. Essential [2505]. Essentially [1182]. Establishing [886]. Estimate [2020]. estimates [1346]. Estimating [1430, 597]. Estimation [1819, 583, 1403, 1425, 1815, 1384, 447]. ETOL [1038]. Euclidean [1578, 991]. Evaluated [291]. Evaluating [877, 1736]. Evaluation [1485, 633, 1817, 424, 618, 2488, 1783, 800, 1859, 291, 2286, 1428, 2356, 2420, 1895, 1894, 1520, 1517, 2115, 561, 1469, 1294]. Even [1174, 2062, 744, 2577]. Even-Mansour [2062]. Event [317, 324, 1988, 1791, 14, 1962, 1249, 1261]. Events [1786, 2382]. Every [2383]. Evidence [1533, 2255, 2258, 1736]. Evolution [417, 2135, 649, 416, 2478, 614, 1327, 492, 878, 1700]. Evolutionary [2211, 1770, 2457]. Evolutive [484]. Evolving [819, 1303]. Exact [745, 231, 2201, 1068]. Example [1307, 1277, 2183, 615, 2231]. Examples [2182, 399, 635, 395, 2270, 1108]. Exceptional [1903]. Exceptions [1721]. Exchange [2042, 2112]. exciting [201]. Exclusion [681, 380]. Executable [421, 693, 1197, 721]. Executing [1217, 1746]. Execution [2479, 2426, 1521, 2445, 2428, 2417, 1724, 2406, 2447]. Executions [2293]. Exhaust [553]. Exhaustion [990]. Existence [2103, 829]. Existential [645, 731]. Existing [1523]. Expandable [1855]. Expansion [2282]. Expansions [2231]. EXPECT [1617]. Expected [811, 2545]. Experience [1304, 1460, 905, 885, 2473, 2288, 2444].

Experiences [2154, 1311]. Experiment [1324]. Experimental [858, 1473, 360, 1876, 895, 878, 1914, 896, 864]. Experimentation [873, 881, 901, 888]. Experiments [860, 1485]. Expert [411, 284, 2344, 1495, 1532, 1553, 609, 606, 1554, 1544, 2270, 408, 1534, 405, 400, 1543, 1530, 402, 607, 1531, 2387, 1535]. Expertise [1606]. explain [1756]. Explaining [2479]. Explanation [2181, 1774, 302, 623]. Explanations [421, 1882]. Explanatory [282]. Explicit [19]. Exploitation [1862]. Exploration [2384, 1913]. Exploring [1541]. Exponential [1181, 1602]. Exponentially [2565]. Exponentials [1179]. Exponentiation [105, 132, 192, 1373, 101, 1367]. Expressibility [811]. Expressing [621]. Expression [1966]. Expressions [2182, 1680, 2527, 2489]. Expressive [601, 647, 1078]. expressiveness [84]. Extended [677, 105, 1680, 1978, 129, 673, 87, 2079, 2065, 779, 132, 1151, 1188, 2040, 599, 2077, 246, 795, 612, 2080, 702, 2083, 2064, 1230, 204]. Extendible [834]. Extending [1592, 1606, 1848, 2445, 1601, 2381, 2287, 1565]. Extensibility [2142]. Extensible [926, 2130]. Extension [2515, 2346, 2048, 2489, 9]. Extensional [786]. Extensions [675, 289, 327, 1160, 2256, 2260, 2263, 692, 167, 647]. External [475]. Extra [1175]. Extra-Logical [1175]. Extract [1446]. Extracting [174, 700, 1753]. Extraction [1385, 1404, 1412, 177]. Extremal [2597]. Eye [2448].

F [172]. F.F.T. [88]. F.O.B. [2365]. Face [2336, 459, 2578]. Faces [639, 1482]. Facets [1006]. Facilitating [1532]. Facilitatory [1558]. Facilities [1881]. Facility [1987, 1506, 443]. Facing [407]. Factor [589]. Factorial [497]. Factoring [102, 124, 2377]. Factors [876, 1898].

Factory [1304, 885]. Facts [668]. Fahlman [487, 486]. Fail [2065, 118, 2064, 119, 1658]. Fail-Stop [2065, 118, 2064, 119, 1658]. Failure [284, 154, 2012, 560, 1657]. failures [1664, 1666]. Fair [2072, 1254]. Fairness [1981, 1671, 216, 1252]. False [1815]. Families [315, 763, 2025, 2026, 1365]. Family [318, 2441, 376, 2024, 1759]. Fast [485, 857, 105, 1454, 1387, 1631, 1733, 976, 1443, 1472, 1786, 1389, 768, 1447, 1410, 344, 579, 767, 998, 936, 133, 1358, 1367]. Fast-Convergence [485]. Fast-Turnaround [857]. Faster [2598]. Fat [986]. Fault [1993, 1856, 749, 1870, 1659]. Fault-Tolerance [749]. fault-tolerant [1659]. Faults [947, 1303]. Faulty [2331]. FEAL [2059, 93, 2036]. FEAL-4 [2059]. Feasibility [2521, 2421]. Feasible [1193, 567, 256]. Feature [674, 157, 564, 581, 498, 1385, 300, 545, 526, 1105, 495, 1767]. Feature-Guided [581]. Features [2450, 651, 332, 595, 1105]. Federated [2469, 1311]. feebly [1365]. feebly-one-way [1365]. Feed [554, 547, 493]. Feed-Forward [554, 493]. Feedback [1993]. Feeling [1877]. Feistel [2030]. FEMUS [2469]. FFT [2057, 89, 2107]. FFT-Hash [2057, 89]. FFT-Hash-II [2107]. FGCS [2302]. Fiat [134, 2027]. fidl [1838]. Field [1397, 2344, 836, 2210, 837, 848]. Field-Programmable [837]. Fields [284, 1414, 1441, 1428, 767, 1602, 567, 1357]. Fight [2108]. File [1155, 2539]. Files [1815, 1813]. Fill [1696]. Fill-In [1696]. Filling [965]. Filter [1073, 1404]. Filtered [1811, 588]. Filtered-Backprojection [588]. Filtering [1388, 2599]. Filters [1424]. FIN [2191]. FIN-Type [2191]. Finally [359]. Financial [905]. findall [1212]. Finding [951, 139, 959, 2548, 622, 600, 318, 1020, 144, 989, 2552, 1482, 1079, 1669]. Findings [1924]. Fine [1862, 2426, 1424]. Fine-Grained [1862, 2426]. Finetely [603]. Finetely-Valued [603]. Finite [373, 139, 757, 1781, 392, 772, 1202, 159, 87, 821, 763, 2203, 776, 634, 2192, 1795, 767, 1602, 595, 1059, 1072, 123, 1080, 1107, 1109]. Finite-Generated [392]. finite-state [1080]. Finiteness [1644, 49]. Firing [1027]. First [2515, 34, 1417, 683, 33, 1966, 982, 611, 35, 283, 1952, 670, 2150, 573, 218, 175, 1105]. first-class [1105]. First-Order [2515, 34, 33, 1966, 35, 283, 1952, 670, 218]. Fitting [1461, 1176]. Five [1006]. Fixation [1436]. Fixed [237, 496, 1701]. Fixed-parameter [237]. Fixing [1383]. Fixpoint [814, 2475, 1630, 1629]. Flang [1207]. Flexible [2464, 1611, 927, 510]. FLINS [625]. FLINS-Fuzzy [625]. Flow [1430, 2573, 2324, 1318, 352, 739, 1475, 984, 2434, 2443, 548, 1424, 1081]. Flowchart [1012]. Flowcharts [150]. Flowers [561]. Flows [1991, 1874, 1706, 2577]. Flsim [181]. Fluid [1874]. Focus [1867, 1885]. Foot [2579]. Foot-Prints [2579]. Forgetting [481, 1085]. Fork [778, 1058]. Form [813, 2117, 1525, 36, 148, 1437]. Formal [1994, 138, 286, 1307, 1330, 1867, 1016, 1995, 144, 1969, 1301, 2332, 1308, 1325, 2155, 1971, 2268, 1968, 2410, 852, 2277, 2051, 1746, 268, 243]. Formalising [428]. Formalism [2233, 1965, 714, 1777]. Formalisms [2272]. Formalization [1732, 172]. formalized [190]. Formally [642]. Format [701]. Formation [1469]. Formats [1138]. Forms [1915, 1950, 238]. Formulae [2172, 1041, 13]. Formulas [2432, 624, 1745]. Formulation [41]. FORTRAN [2446, 424, 2431, 1838, 1839, 1937, 2422]. Forward [554, 291, 547, 493]. Foundation [166]. Foundations [589, 1751, 2474, 2478, 1619, 1938]. Founded [1726, 2516]. Four [741, 1538, 2197]. Fourier [1402, 1933, 1452]. FPGA [844, 839, 854, 847, 853, 845, 852]. FPGA-Based [845, 852]. FPGAs [842, 855, 838, 841, 840, 856, 849, 846]. Fractal [1477, 2562, 1415, 546, 234]. fractal-like [234]. Fraisse [814]. Frame [2247, 641]. Framework [2541, 1591, 2554, 613, 1330, 2361, 2023, 1739, 1875, 802, 2481, 2248, 920, 1159, 160, 583, 2440, 1284, 1808, 2340, 363, 524, 1269, 282]. Frameworks [1890, 905, 873, 5]. Fraud [2109]. frauds [130]. Free [810, 958, 88, 1174, 334, 1015, 782, 1525, 1031, 333, 775, 941, 344, 2025, 1035, 2093, 773, 2107, 1437, 1655, 1042, 223, 1652, 222, 253]. Free-Form [1437]. Free-Space [941]. Freeness [1640, 1006]. Freezing [1684]. Frequency [1188]. Frobenius [1014]. Frontier [1033, 762]. Frontiers [256, 351]. Fsub [1371]. Fulfilling [1113]. Full [513, 2097, 162, 772, 1022, 1047]. Fully [470, 1692, 1693, 985, 1707, 186, 1253]. Function [598, 1392, 665, 2057, 2023, 1180, 468, 1486, 469, 2201, 114, 585, 110, 574, 2026, 255, 186, 109]. Functional [1625, 2175, 2134, 1176, 1198, 680, 2242, 1210, 820, 1784, 1199, 2320, 824, 1889, 1961, 672, 2125, 1516, 657, 1513, 1641, 2406, 242]. Functionality [2213, 2345]. Functionals [744]. Functions [157, 2236, 1592, 2082, 627, 2585, 2023, 1002, 818, 2523, 2021, 90, 295, 2264, 1031, 1581, 122, 2113, 2197, 2276, 2280, 1638, 83, 1377, 1049, 1376, 1340, 1366, 1343, 95, 252]. Fundamentals [886]. Funds [2123]. Furio [1047]. Further [978, 1931]. Fusion [1440, 2433, 2388, 1466]. Future [448]. Futures [928]. fuzziness [1508]. Fuzzy [411, 618, 2242, 1454, 412, 395, 1508, 405, 848, 625, 386, 733, 401, 560, 625]. Fuzzy-Centered [625]. FX [1634].

G [1792]. Gabor [598]. Galois [1954]. Game [814, 862]. Games [972]. Gamma [1648, 2437, 1950]. GANNet [500]. Gap [1302, 247, 876]. Gap-definability [247]. Gate [836, 843, 837, 848, 252]. gates [254]. Gauss [1581]. Gaussian [1391, 542, 1461, 595]. GCLA [170]. GCV [1486]. GCV-Aided [1486]. GemCode [1530]. General [1010, 954, 335, 215, 55, 1682, 54, 974, 2103, 2035, 333, 364, 2077, 153, 2075, 1605, 9, 1629, 123]. Generalisations [2533]. Generalised [1418]. Generalising [386]. Generalization [2164, 141, 2183, 293, 948, 1768]. Generalized [1658, 978, 1024, 2104, 979, 2321, 62, 2029, 2028, 342, 48, 1001, 2186, 131, 1249, 277, 108, 1363]. Generate [287, 95]. Generated [392, 2326, 1771]. Generating [2530, 2369, 318, 1945, 1031, 1530]. Generation [1229, 637, 2237, 2089, 2053, 308, 506, 2498, 1688, 307, 1870, 1009, 626, 330, 524, 2364, 1362, 1364]. Generative [315, 1777]. Generator [1224, 927, 1348]. generators [108, 1363]. Generic [1613, 694, 643, 2126, 30, 524, 16]. Genetic [935, 513, 2016, 2158, 1768, 1421, 477, 478, 500, 1769]. Genome [2179]. Genomic [1548]. Gentzen [2012]. Gentzen-style [2012]. genus [67]. Geocoded [1492]. Geographic [1528, 1284, 1286, 2458, 1297]. Geographical [1529, 1527]. Geometric [955, 2177, 1525, 938, 1336, 1357, 231]. Geometrical [939, 1445]. geometries [1374]. Geometry [637, 2562, 2451, 573]. Geometry-Limited [573]. Gestalt [2210]. Getting [2060, 1684]. GGD [651]. GIS [1288, 1281]. Gist [2407]. Given [1603, 1842]. Glasses [1901]. GLEFATINF [1591]. Global [969, 1076, 1921, 456, 1103]. Gluons [2127]. Go [1607]. Goal [2542, 1309, 303, 153]. Goal-Directed [153]. Goal-Driven [303]. Godel [1171, 2323, 1164]. Golfkrieg [1920]. Good [310]. Goodness [595]. Goods [1553]. Gossip [2536]. Gossiping [2571, 1691, 1697, 2569]. Grade [269]. Gradient [1388, 1403, 567]. Gradual [1626]. Grain [2436]. Grained [1862, 2426, 971]. Grammar [1205, 651, 480, 1082]. Grammars [471, 1680, 482, 1015, 2129, 1026, 1205, 1038, 78]. Granularity [2242, 213, 1628]. Graph [365, 20, 2507, 86, 967, 2480, 1015, 81, 321, 2395, 1859, 2550, 983, 651, 326, 1973, 2381, 834, 1106, 2402, 57, 63, 71, 269, 1083, 714, 278, 2408, 78, 1099]. Graph-Based [2480]. Graphic [2528, 1591]. Graphical [2546, 1852, 1160, 2353, 1909, 1918, 2328]. Graphics [1875, 2367]. Graphs [1675, 958, 1012, 743, 2385, 2379, 966, 2595, 1218, 741, 2391, 960, 1686, 2392, 2376, 2586, 2545, 2574, 1691, 1692, 1694, 982, 1984, 985, 2553, 2382, 324, 2397, 991, 2386, 2401, 283, 1036, 2417, 1705, 2399, 1708, 2390, 729, 713, 66, 717, 715, 72, 67, 76, 716, 731, 59, 83, 730, 58, 73, 62, 210, 85, 725, 720, 722, 79, 732, 1107, 726, 733, 1081]. Gray [1414, 2563]. Gray-Scale [1414]. Grazing [1277]. Greedy [504, 1001, 2405]. Grey [1398]. grid [80]. Grids [1691, 1697, 68]. Gröbner [2004]. Ground [1203, 36]. Group [1890, 2100, 2571, 1342, 652, 38, 123, 95]. Group-oriented [1342]. Grouped [1116]. Grouping [2219, 1395, 940, 2226]. Groups [87, 773, 949]. Growing [2190]. Grows [121]. Growth [811, 678]. Grundy [66]. Grzegorz [1051]. Guarding [2550]. Guided [581, 600, 1630]. Guidelines [1919]. Guides [1926]. GUIs [1881].

Halting [206, 1033]. Hamming [464]. Hand [1463]. Handbooks [883]. Handicapped [2365]. Handle [1592, 2445]. Handling [1887, 1495, 2431, 1305, 1586, 2250, 1533, 1201, 2593, 2386, 2272]. Handprinted [1455]. Handwritten [469, 1915, 1458]. Hard [738, 205]. Hardware [516, 514, 2435, 1858, 1816, 356, 526]. Harmonization [1917]. Harray [2446]. Hash [2057, 2023, 90, 2113, 89, 2107, 1376, 1340, 2057]. Hashing [1811, 348, 88, 1812, 2093, 89, 2601, 1339]. Hausdorff [1056]. HAVAL [1339]. Having [813]. HCI [2343, 2344]. Health [2061].

Heap [2554, 973]. Heap-like [2554]. Heidelberg [1161]. HeiMAT [1161]. Hellman [106, 1372]. Help [1023, 1881]. Help-Facilities [1881]. Helps [2575, 1693]. Hennessy [1046]. Herstellung [1910]. Heterogeneity [1853]. Heterogeneous [1501, 712, 1123, 2467, 1818, 1383, 2478, 1094, 1823, 1509, 1510, 1489, 1761]. Heuristic [1615, 293, 2005, 690, 605, 2170]. Heuristics [969]. Hexagonal [80]. HFS [688]. HFS-Theory [688]. Hidden [2196, 538, 994, 833]. Hidden-Surface-Removal [994]. Hiding [2046]. Hierarchical [2419, 319, 1533, 2347, 740, 1732, 2513, 2366, 401]. Hierarchically [2136]. Hierarchies [1721, 2486, 632, 700, 1279, 724, 1082]. Hierarchy [2318, 1717, 1165, 1005, 1039, 201, 275, 204, 203, 241]. High [465, 857, 516, 1148, 2101, 840, 107, 528, 887, 457, 519, 845, 1534, 1928, 1140, 1531, 1474, 1056, 1937]. High-Level [1534]. High-Order [528, 457]. High-Performance [840, 1928]. High-Speed [2101, 107]. High-Tech [1474]. Higher [466, 2514, 818, 2291, 789, 628, 569, 1219, 56, 706, 175, 1250, 185, 720, 191, 1049, 1103]. higher-dimensional [1250]. Higher-Order [466, 2514, 2291, 789, 1219, 56, 191]. Highly [2350, 2416, 856, 94, 1572, 1343, 852, 1750]. Highly-Customisable [1572]. Highly-Interactive [2350]. Highways [1821]. Hilbert [1178]. Histogram [1388]. Historical [2343, 2010, 671]. History [2544]. HOL [705]. holey [1779]. holographic [251]. holonomic [932]. homogeneous [73]. Homomorphic [87, 2103]. homomorphisms [1057]. Honeycomb [1928]. Honsell [1047]. Hopfield [539, 550, 751]. Horn [2584, 206]. Hot [842]. Hough [1429, 1403, 1446, 1402, 1401]. Hough-Like [1401]. Hour [1901]. Hull [1408]. Human [2369, 2374, 1884, 597, 2347, 2346, 1922, 1894, 1898, 574, 282, 1482].

Human-Computer [2369, 2347, 2346, 1894]. Humanities [1280]. Humans [310, 2387]. Hybrid [1985, 1992, 1987, 1986, 398, 1978, 1991, 1086, 1979, 1982, 1984, 298, 1981, 1980, 2434, 1988, 1977, 1989, 1990, 1983, 484, 505]. Hypercube [364, 1699, 217]. Hypercubes [2571]. HyperDEBU [2292]. Hyperedge [1015, 1006]. Hypergraph [2584]. Hypergraphs [1187, 82]. HyperLecture [2371]. Hypermedia [1568, 2372, 1566, 1879, 1823]. HyperPATH [1568]. HyperPATH/ [1568]. Hyperplanes [953]. Hyperset [1589]. Hypersimple [757]. Hypertext [1567, 2368, 2367, 1832, 1569, 1902, 446, 1882]. Hypertexts [2366]. Hypotheses [145]. Hypothesis [2187]. Hypothetical [154].

I/O [954, 1128, 1141]. I/O-Complexity [954]. IC [2037]. IC-Cards [2037]. Iconic [461, 1422, 2352, 1104]. ICSP [2534]. IDAGs [2239]. Ideal [2076, 2540, 87]. Ideas [2319]. Identification [531, 2084, 555, 129, 412, 2169, 1988, 2066, 590, 2218]. Identifying [1940, 2331, 2178]. If [775, 401]. If-Then [401]. Ignoring [1092]. II [237, 1147, 1113, 612, 89, 831, 1623, 1235, 2107]. III [1150]. ill [1237]. ill-timed [1237]. Image [485, 598, 2336, 1388, 1460, 589, 1485, 1391, 1476, 1392, 549, 1406, 1072, 2223, 1484, 533, 1868, 551, 1557, 2220, 1472, 577, 1486, 1483, 1429, 575, 1475, 2334, 1433, 593, 1389, 1425, 1492, 1493, 1915, 550, 1380, 545, 1451, 1479, 1463, 332, 1424, 591, 572, 1480, 1394, 568, 1471, 586, 1489]. Imager [575]. Imagery [1487, 1422]. Images [1397, 2575, 581, 566, 1385, 1387, 546, 1414, 583, 1446, 1384, 1492, 1466, 571, 565, 1431, 1478, 576, 569, 1398, 1481, 1412, 578, 1753, 1091]. Imaging [582, 594]. Imai [1376]. imitation [1756]. Immutable [2299]. Impact [837, 1149, 1628]. Impacts [308]. Impedance [579]. Imperative [1331, 1226, 1953]. impersonation [1335]. Implement [856, 285]. Implementation [341, 909, 1212, 2154, 2456, 1225, 516, 463, 1826, 1812, 1215, 673, 2101, 1019, 2289, 511, 322, 1866, 854, 2290, 107, 1205, 1197, 1518, 801, 1118, 1207, 364, 1213, 2319, 4, 2484, 1289, 2458, 2307, 1849, 1550, 2170, 1512, 8, 1086, 269, 1652, 1252, 1671]. Implementational [170]. Implementations [365, 1575, 514, 1227]. Implemented [2435]. Implementing [715, 168, 44, 169, 671, 870, 557]. Implicants [697, 1749]. Implicates [697]. Implication [1514, 1169, 387, 2487, 1765]. Implications [589, 1837]. Implicit [2153, 1247, 1355]. Improve [421, 870]. Improved [2592, 68, 426, 982, 1439, 1827, 2173, 57]. Improvement [1304, 1493, 871, 503]. Improvements [1203]. Improving [2019, 1886, 1533, 1578, 306, 492]. In-Place [2554]. Inaccessible [2252]. Inchoative [2382]. Incidence [2265, 2264]. Including [508, 2023, 2170]. Inclusion [758, 1054]. incompatible [1062, 61]. Incomplete [1737, 1621, 1842]. Inconsistency [1305, 2250, 1549]. Inconsistent [2251]. Incorporated [448]. Incorporating [575, 583]. Incorrect [529]. Increased [491]. Increasing [1694, 1771, 690]. Incremental [396, 2089, 1267, 1732, 492, 941, 2180, 1616, 871, 2519, 262]. Independence [2595, 1897, 2298, 2277]. Independences [2511]. Independent [2255, 534, 723]. Independently [906]. Independently-Developed [906]. Index [135, 929, 2283, 659, 976]. Indexes [1282]. Indexical [2459]. Indexing [1438]. Indicators [1901]. Individual [398, 1907]. individuals [1758]. Induce [2527]. Induce-Expressions [2527]. Induce-Statements [2527]. Inducing [718]. Induction [144, 1186, 480, 30, 1967, 2209, 2207]. Inductive [140, 2182, 639, 1593, 399, 642, 43, 145, 33, 821, 293, 2167, 2341, 2166, 2187, 2527, 193, 2185, 2157]. Industrial [1994, 284, 1314, 630, 535]. Industrial-Scale [1994]. Industry [1475, 1474]. Inequalities [677]. Inference [916, 20, 1539, 392, 2302, 2391, 145, 2254, 2167, 2194, 2166, 2187, 700, 402, 151, 2279, 2157, 1761, 730]. Inferred [1823]. Inferring [1761]. Infinite [761, 260, 261, 259, 1109]. Influence [1674, 463, 2242]. Influences [1892]. Informatics [2179]. Information [1317, 2530, 2336, 1501, 2074, 426, 1495, 2379, 1906, 2355, 2418, 427, 562, 443, 1422, 1555, 428, 587, 2464, 617, 2251, 1809, 661, 2427, 583, 2177, 1403, 742, 1446, 2039, 2158, 2095, 2228, 384, 1922, 1913, 1924, 2113, 1297, 1435, 432, 130, 2274, 2466, 2081, 1361, 2075, 2458, 1157, 439, 377, 1842, 1772, 1753, 1056, 265, 1641, 1346, 719, 84]. information-seeking [1772]. Information-theoretic [130]. Information-Weighted [587]. Infrastructure [873]. Inheritance [916, 162, 1717, 2503, 1586, 1511, 383, 915, 611, 700, 647, 1583, 1060, 1071]. Inhibition [453]. Inhibitory [522]. initial [1359]. Inlining [1214]. Innermost [52]. Input [2182, 526, 607]. Input-Driven [607]. Input/Output [2182]. Inputs [962, 475, 522]. Inseparability [808, 1181]. Insertion [998]. Insight [1844]. Inspection [1322, 1478, 1479]. Inspired [456]. Instance [600, 2046, 645]. Instance-Based [600]. Instance-Hiding [2046]. institutions [15]. Instruction [1929]. Instrumentation [1797]. Integer [962, 2013, 688, 1579, 989, 1706, 245]. Integral [739]. Integrated [1129, 1135, 450, 1306, 2130, 1923, 1513]. Integrating [1568, 1755, 1875, 1527, 906, 2367, 2010, 636, 1569, 2529, 2272, 1914, 299, 387, 1846]. Integration [860, 2213, 1784, 419, 1984, 2478, 2148, 2228, 2378, 420, 2138, 654, 1509, 418, 656, 1320, 716, 1259, 1761]. Integrator [522]. Integrity [2339, 799, 1514, 672, 153]. Intelligence [2007, 1996, 2272, 723]. Intelligent [640, 1539, 2372, 610, 2002, 644, 2232, 2001, 1713, 642, 1988, 2329, 310, 2494, 2156, 1617, 296, 2458, 2373, 1512]. INTELTEXT [2366]. Intensive [669]. Interactice [1899]. Interaction [1887, 1919, 1292, 1414, 2345, 2346, 434, 1894, 1917, 1281, 1751, 1264]. Interactive [2336, 2350, 1142, 1875, 1915, 1126, 1127, 1936, 150, 1362, 1750, 115, 116, 120, 2079, 128]. Interchange [1138]. Interchangeability [1730]. Interconnected [473]. Interconnection [540, 2400, 358]. Interconnections [1854]. Interdependencies [2468]. Interface [1906, 313, 1292, 1876, 1160, 1858, 666, 1571, 1899, 1897, 2348, 1893, 1808, 1914, 1898, 1909, 2358, 2352, 2373]. Interfaces [2369, 1911, 2001, 1880, 1570, 2365, 444, 2351, 1918, 2328]. Interference [2247, 2363]. Interframe [1430]. interleaving [1251]. interleavings [1092]. Interlocks [1929]. Intermedia [1124]. Intermediate [2441]. Interoperability [1503, 908]. Interoperable [662]. Interoperating [2467]. Interplace [1292]. Interpolation [735]. Interpretation [1731, 618, 1331, 1202, 1422, 1018, 1164, 1958, 1492, 548, 331, 28, 940, 2226, 1631, 1632, 1069, 1629, 1778, 1628]. Interpretations [1954]. Interpreter [813, 291, 2290, 606]. Interpreting [1243, 1189]. interrelations [1052]. Intersection [950, 971, 974, 978, 194, 2505, 173]. Intersections [993]. Intertask [2287]. Interval [2591, 1675, 966, 403, 1999, 1956, 2517, 288, 58, 60, 212]. Interval-Based [403]. Intervals [2591, 736, 391, 69]. Interworking [1718]. Intestine [1462]. Intimate [903]. Intra [1330]. Intra-Level [1330]. intractability [237]. Intractable [2026, 123]. Introducing [797, 2232]. Introduction [367, 1180, 1976, 1787, 1161, 2403, 1051]. Introspection [1185]. intuitionistic [176]. Invariance [2210, 1396]. Invariant [2007, 486, 528, 453, 332, 2079]. Invariants [315, 1453, 1951, 1444]. Invent [2194]. Inverse [580, 1176, 334, 1766]. Inversion [579, 1393]. Inverting [722]. Investigation [591]. Invited [1251, 1266, 1230, 514, 2090, 1241, 527, 449, 1257, 17, 49, 30, 35, 2069, 505, 1246]. Invocation [2153]. Involvement [1918]. Involving [385, 1966, 614]. IP [2049]. Irradiance [1433]. Irregular [1381, 2422]. Irrigation [407]. ISALOG [2486]. ISAR [1575]. ISBES [1975]. Isn’t [2126]. ISO [1895, 2118]. ISO-OSI [2118]. Isolated [534]. Isomorphism [708, 38, 218, 67]. Isomorphisms [1063]. Isothetic [992]. Issues [170, 2453, 2470, 1788, 887, 1167, 1824, 363, 1294]. Items [1538, 1530]. Iterated [1451, 253]. Iteration [1955, 774, 1219, 81, 1630]. Iterative [2430, 1387, 2099, 2098, 2162, 326, 97]. ITEX [298]. IV [1154]. IZIC [1599].

Japan [2284]. Japanese [2302]. JAPROC [850]. Jigsaw [1417]. Job [1552, 1662]. jobs [1062, 61]. Jog [998]. Joint [1122]. Joint-Viewing [1122]. JPEG [1389]. Jukeboxes [1821]. Junction [1436]. Junk [2073]. Just [2442].

k-D [973]. Kalman [1404, 495]. kappa [2200]. kappa-Label [2200]. Karatsuba [1579]. KAT [1610]. Kayles [66]. KBMS [2457]. KBMS-Supported [2457]. Keeping [2536]. Kerberos [2124]. Kernel [2084, 849, 2132, 1289, 1944, 2128, 2133]. Kernels [1595, 1393]. KEW [1613]. Key [2096, 2110, 2103, 2114, 100, 2492, 126, 2035, 2095, 2040, 2072, 2038, 2112, 1372, 106, 1368, 1370, 128, 1374, 1375, 1369]. Keynote [858, 866, 898, 873, 883, 891]. Keystroke [1293]. Kid [2090]. Kinds [2193]. Kinodynamic [931]. Kit [1892]. KIV [268]. KL [713, 380]. KL-one [713]. KL-ONE-Like [380]. Kleene [829]. Knapsack [2056]. Know [146]. Knowing [613, 1166]. Knowledge [1625, 646, 1737, 1501, 417, 2091, 117, 1540, 1607, 286, 409, 1615, 1620, 115, 2047, 1626, 2078, 802, 2046, 410, 399, 2388, 2254, 2079, 643, 1876, 1270, 575, 1541, 2044, 395, 2010, 1571, 1536, 1614, 285, 636, 619, 2397, 1621, 2378, 433, 2381, 2377, 1624, 1569, 2077, 1616, 2048, 615, 1617, 2356, 1479, 292, 629, 1718, 400, 2043, 2080, 297, 2399, 574, 2000, 1619, 654, 2358, 2457, 2380, 1612, 1618, 652, 1531, 2396, 1535, 656, 2186, 2501, 1772, 727, 116, 711, 719, 1763]. Knowledge-Based [1876, 636, 619, 1479, 297, 2000, 654, 652, 1569]. Known [2059, 93]. Known-Plaintext [2059]. Kohonen [844, 497, 511, 577, 496]. Kolmogorov [527]. KONTERM [1543]. Korea [2031]. KorSo [439]. Kripke [1547]. Kripke-Type [1547]. Krypto [2090]. Krzysztof [1042].

Label [2506, 6, 2200]. Label-Selective [2506]. labeled [1051]. labeling [1770]. Labelled [679, 1029]. Laboratory [879]. Lambda [813, 1075, 189, 1193, 2506, 787, 812, 1230, 178, 1073, 786, 180, 2183, 1047, 184, 187, 169, 26, 1093, 1095, 197, 198, 208]. lambda-calculi [198, 208]. lambda-Calculus [2506, 812, 786, 2183, 26, 1230, 1073, 1093]. lambda-I [184]. lambda-pi-calculus [180]. lambda-typed [208]. LAMBDALG [706]. Lambertian [1445, 1435]. Land [1285, 1761]. Landmark [934]. Landmark-Based [934]. Landscape [1291, 2354]. Landscapes [1297]. Language [461, 2528, 2486, 793, 2495, 265, 156, 308, 1225, 509, 1584, 911, 2508, 926, 2438, 1321, 159, 1215, 2460, 1181, 1865, 1599, 920, 2462, 641, 2144, 823, 1499, 376, 824, 2145, 2049, 2169, 789, 2168, 2193, 1197, 2416, 169, 266, 147, 2436, 606, 1907, 2142, 2397, 775, 2268, 2192, 2132, 2496, 1972, 1206, 2295, 1274, 672, 1959, 309, 297, 2222, 2373, 1516, 706, 1743, 305, 1762, 12, 1635, 1771, 720, 16]. Language-Driven [1599]. Languages [139, 289, 757, 928, 2541, 2508, 1210, 301, 1016, 2518, 1865, 906, 1004, 2184, 758, 2194, 36, 1030, 801, 1031, 1720, 1722, 2153, 1035, 1009, 645, 769, 1938, 2307, 2352, 762, 2159, 1953, 1583, 1066, 248, 1633, 235, 222, 205, 257, 1758, 258, 259]. Large [1430, 1580, 1541, 1228, 1462, 2436, 2353, 1824, 1871, 1303, 407, 432, 2366, 1345, 1351]. Large-Grain [2436]. Large-Scale [432]. Latest [2536]. Lattice [1218, 2269, 2560, 724]. Lattice-Theoretic [2269]. Lattices [621]. Law [1544, 1546]. laws [275]. Layer [2590, 521]. Layered [536, 2180, 493, 1702]. Layers [1276, 1234]. Layout [1869, 1458]. Lazy [2289, 1209, 1208, 2319, 1645]. LDL [2485]. Leading [1470]. Leaky [522]. Learn [1737, 2188]. Learnability [2175, 757, 2541, 1041]. Learned [635]. Learner [1535]. Learners [2188]. Learning [137, 485, 2164, 414, 396, 138, 2160, 2236, 1766, 294, 2541, 398, 555, 2372, 549, 523, 450, 143, 2191, 1744, 1735, 1900, 2208, 1927, 1769, 638, 489, 487, 2257, 486, 2161, 2189, 2184, 1108, 395, 2177, 2165, 2168, 2193, 2206, 1491, 147, 490, 2262, 1421, 295, 636, 2176, 496, 478, 457, 2179, 492, 2180, 451, 2201, 2156, 2211, 2052, 2015, 2172, 2202, 658, 394, 2205, 2173, 2198, 481, 1612, 460, 2190, 2157, 2373, 656, 2186, 2159, 2204, 2240, 265, 266, 1770]. Least [1817, 2162, 1404]. Lebiere [487, 486]. Lecture [2371]. LEDA [271]. Left [337, 1204]. Left-Recursive [1204]. Leftmost [52]. Legacy [1839]. Legal [1609, 1543, 1763]. Legality [383]. LEGO [172, 1761]. Lego-land [1761]. Lemma [1030, 255]. Length [1678, 2238, 2589, 2510, 1030, 1339]. Lengths [100]. Less [828, 1497]. Letter [1011, 1752]. Level [485, 1731, 2499, 926, 1330, 2242, 1142, 2250, 1379, 687, 1293, 519, 845, 1534, 1398, 1623, 718, 1755]. Leveled [1678]. Levels [1892, 203]. Leverrier [767]. Liberalized [1174]. Library [1325, 922, 271]. License [547]. Life [1896]. Liftable [624]. Light [1439, 1576, 1563]. Lighting [1432]. Like [1401, 387, 1000, 1045, 174, 2554, 234, 62, 2594, 134, 380, 2021, 2462, 2035, 250, 2106]. limit [224]. Limitation [518]. Limitations [2572, 2062, 1937, 85]. Limited [1387, 573]. Limiting [1611]. Limits [238]. Linda [2425, 2438, 1633, 1939]. Linda-Based [2438, 1633]. Line [2575, 1681, 971, 1883, 755, 2588, 556, 561, 1416, 1001]. Linear [808, 531, 1011, 2181, 677, 812, 2541, 2572, 1407, 1392, 1344, 2587, 1179, 1210, 995, 1580, 1733, 2249, 822, 1024, 1486, 688, 1026, 1654, 2059, 2301, 92, 2228, 164, 492, 2201, 521, 579, 1192, 1967, 2538, 2578, 2366, 701, 1963, 503, 1708, 1711, 1196, 1053, 176, 1338, 67, 1077, 96, 1096, 1235, 1260, 1058]. Linear-Time [1708, 1654]. Linearisation [1567]. linearity [1640]. Linearizations [37]. Linearly [815]. Lines [953, 1398, 69]. Lingual [625]. Linguistic [1286, 719]. Linking [310]. Lisp [2296]. List [2310, 2297, 230, 275]. Lists [2587, 2206, 1108]. literate [2409]. Live [331]. Liveness [1014]. LL [1026]. LOA [803]. Load [957, 312, 616, 1860, 1828]. Local [140, 1013, 2045, 2120, 483, 1261, 782, 742, 2510, 457, 233, 122, 1925, 1929, 85, 1049, 1260]. Locality [363]. Localization [596]. Locally [259, 246]. Locate [1484]. located [1089]. Locating [566, 204, 2035, 1885]. Location [1685, 2320, 2522]. Locking [1831]. Löf [1170]. Log [2068]. Logan [1533]. Logarithm [2047, 1370, 123]. Logarithmic [962]. Logarithms [2085, 127, 1341]. Logic [808, 411, 2512, 157, 1223, 675, 1727, 2181, 809, 1171, 2515, 2235, 141, 2237, 51, 814, 1198, 632, 780, 162, 601, 1013, 613, 2438, 2480, 785, 2316, 382, 1179, 1210, 2243, 639, 2223, 168, 159, 1215, 154, 2245, 802, 1454, 2247, 2313, 617, 279, 50, 2252, 604, 160, 820, 800, 1949, 687, 1166, 1199, 689, 291, 679, 2256, 2145, 2161, 395, 1726, 847, 165, 2301, 1187, 300, 388, 169, 2261, 614, 801, 620, 843, 628, 1200, 781, 1958, 1167, 148, 691, 2266, 1722, 2267, 164, 599, 381, 331, 624, 1211, 695, 1169, 1725, 152, 1192, 1967, 2517]. Logic [612, 166, 153, 1760, 2273, 1719, 2275, 699, 830, 1963, 2278, 2292, 702, 377, 281, 56, 2281, 2186, 2315, 1196, 1053, 174, 176, 1248, 1641, 1067, 209, 1077, 1078, 11, 711, 730, 185, 1247, 1096, 195, 14, 710, 1636]. logic-based [209]. Logical [1175, 632, 791, 622, 1210, 761, 2248, 1714, 2532, 2478, 825, 1509, 831, 1269, 5, 248, 2244, 1765]. Logicism [1607]. Logics [703, 1591, 301, 2251, 686, 603, 1168, 624, 1719, 1748, 1759, 244, 1760, 732, 1109]. LOGIDATA [794, 797, 806, 793, 792, 790, 798, 796, 803, 804]. Login [2111]. LOKI [2019, 2018]. LOKI91 [1347]. Long [1414]. Long-Range [1414]. Longest [945, 2589, 983]. Lookahead [1674]. Lookup [841]. Lookup-Table [841]. Loop [2429, 1177, 1951, 2433, 1257, 2440, 2442]. Loops [2430]. LOP [341]. Low [2078, 1379, 1534, 236, 204]. low-dimensional [236]. Low-Level [1379, 1534]. Lower [753, 954, 962, 627, 2065, 2525, 1008, 197, 2075, 2064, 203]. LR [1026]. Luby [108]. Luby-Rackoff [108]. Lucas [1377]. Łukasiewicz [387]. LVQ [502, 503]. Lyndon [225].

M [1049]. MacDraw [2358]. Machine [807, 461, 1129, 857, 2011, 353, 636, 2052, 852, 1724, 1917, 1612, 2227, 723, 2327]. Machine-Learning [636]. Machines [348, 1173, 2419, 2430, 772, 350, 352, 2450, 759, 2189, 2511, 2439, 1033, 2187, 831, 750, 2446, 238, 221, 235, 202]. Macro [2370, 1019]. Macroscopic [1930]. Made [952, 1308]. Magnetic [582, 578]. Magnetospheric [1486]. Mail [2073]. Maintaining [2135, 1775, 754, 1270]. Maintenance [2248, 2129, 1714, 1514, 2331, 1536, 1742, 1776]. Majority [2040, 252]. Makanin [341]. Make [118, 2126, 285, 2383, 119]. Making [1776, 734, 1307, 2250, 1611, 2071, 746, 2260, 658, 2280]. Mammography [593]. Man [1544, 1917]. Man-Machine [1917]. Managed [913]. Managed-Object [913]. Management [1551, 2336, 618, 379, 1440, 1502, 1559, 802, 2114, 1809, 2129, 1876, 1227, 2146, 876, 311, 1864, 1315, 1860, 1548, 1808, 615, 795, 2306, 407, 1830, 1312, 1510, 1140, 1116, 1660]. Manager [1554]. Managing [2143, 884, 2294, 1843, 1542, 1512]. Manipulable [1286]. Manipulating [2136]. Manipulation [1913, 83, 256]. MANs [1783]. Mansour [2062]. Manual [1293]. Manufacturer [2365]. Manufacturing [653, 652]. Many [703, 639, 687, 538, 267, 1366]. Many-Valued [703, 687]. Map [2239, 1293, 1893, 1273, 1289, 1276, 1762, 586]. MapA [517]. Maple [705]. Mapping [392, 409, 2579, 498, 1687, 437, 2557, 495, 1295]. Mappings [2171]. Maps [844, 497, 511, 942, 496, 545, 526, 534, 1268, 1461, 1283]. Marginals [2260]. Marital [1544]. Marked [326]. Market [559, 1756]. Marking [324]. Markov [1414, 2196, 1441]. Markovian [2219]. Marriage [705]. Martin [1170]. Martin-Löf [1170]. Massively [2430, 102, 2085, 2308, 354, 1791, 1961, 520, 2303]. MAT [2161]. Match [1697]. Matching [964, 1679, 1417, 1440, 653, 1729, 642, 2561, 739, 2144, 2395, 976, 1587, 1941, 1447, 1448, 942, 991, 1439, 2394, 1450, 1480, 1437, 1639]. Material [1477, 1475, 386]. Materialization [665]. Materialized [2492]. Materials [1476]. Mathematica [1574, 1602]. Mathematical [1996, 1599, 144, 2016, 1597, 2003]. Mathematics [1162, 1574]. Matrices [976, 1695, 2440, 1933]. Matrim [1544]. Matrix [510, 229]. Matroid [2561]. Matsumoto [1376]. Matthew [1046]. Maurer [1356]. Maurer-Yacobi [1356]. Max [1449, 1206]. Max-Min [1449]. Maximum [382, 2551, 1020, 1028, 1789, 1030, 737, 1669]. May [1067]. McCAT [2441]. McCulloch [458]. MD5 [91]. Meaning [1182]. Means [471, 928, 2397, 331, 1604, 1202, 1635]. Measure [205, 1416, 1420, 1359, 264]. Measurement [866, 1131, 887, 871, 873, 328, 1469, 1800, 1537, 881, 901]. Measurement-Based [887]. Measurements [860, 869]. Measures [393, 595, 1082]. Measuring [870, 863, 596]. Mechanical [289, 1872, 612]. Mechanism [2152, 453, 1830, 346, 1673]. Mechanisms [2017, 873, 701]. Mechanistic [2188]. Media [1845, 1153, 1118, 1806, 1146, 1156]. Medial [571]. medians [70]. Mediator [1921]. Medical [598, 589, 566, 661, 1468, 583, 608, 1466, 571, 565, 576, 297, 572]. Medium [1292]. Meeting [1916]. Meets [2316]. Mega [425]. Mega-Project [425]. MEGAPHONE [1152]. Megascope [1844]. Mellor [1319]. Membership [2236, 45]. Memories [357, 473]. Memory [348, 1217, 1216, 1856, 2586, 1863, 1861, 2450, 1864, 903, 2439, 990, 2163, 1871, 1724, 1935, 1657, 1255, 96, 1659, 1340]. Memory-Management [1864]. MEMSY [1858, 1855]. Mental [2225, 1268]. Merge [1491]. Mesh [1701]. Meshes [2558, 947, 1703]. Message [2289, 1846, 1244, 1673]. message-passing [1244]. Messages [1678, 2150]. Meta [417, 1731, 2508, 926, 291, 1200, 624, 1572, 1546, 1564]. Meta-Interpreter [291]. Meta-Language [2508]. Meta-Level [926].

Meta-Logic [624]. Meta-Model [417].

Meta-Programming [1200]. Meta-Reasoning [1546]. Meta-Visualisation [1572]. Metaclass [927]. metacomputation [1637]. Metadata [1500]. MetaFlex [927]. Metafor [707]. Metal [2250]. Metal-Level [2250]. Metalevel [2131]. Metalogic [161]. Metaphor [1291, 1914, 1878]. Metaphor-Based [1878]. Metaphors [1290]. metatheory [1746]. Method [1212, 373, 640, 1831, 815, 1867, 748, 2498, 841, 536, 1521, 293, 132, 2088, 1315, 2148, 2266, 1168, 2020, 1451, 493, 2420, 292, 767, 1456, 1823, 2446, 1767, 92, 93, 1375, 426]. MethoDex [609]. Methodological [1517]. Methodologies [619, 2342]. Methodology [1317, 662, 642, 376, 877, 609, 1793, 1624, 667]. Methods [1608, 414, 580, 857, 633, 1862, 1477, 448, 911, 2216, 1733, 1969, 468, 597, 107, 1325, 1403, 877, 2194, 413, 1814, 870, 629, 561, 650, 914, 394, 2204, 305, 1766, 1375, 1246]. Metric [476]. Metrication [2353]. Metrics [1411, 896]. Metropolitan [1783]. MHC [484]. Micro [850]. microdomain [1752]. Microscope [1884]. Microscopic [2374, 1930]. microworlds [1750]. Migration [2585, 1570]. MIKE [1616]. Miller [1364]. Milner [1048]. MIMD [2419, 2439, 1874]. Min [1449, 1106]. Mind [2166, 266]. Minima [962]. Minimal [1646, 308, 1255, 321, 1696]. Minimization [2321, 2165, 738, 843]. Minimizing [2589, 2149]. Minimum [962, 2238, 69, 1696, 1408, 991]. Minimum-Space [1408]. Mining [1475]. Miniscoped [1967]. Minor [763]. Mints [171]. Miprac [2448]. Mirror [1010, 1827]. Mismatch [1823]. Misprints [1479]. Missing [1395]. Mixed [641, 2271]. Mixin [914]. Mixin-Methods [914]. Mixing [707]. MLP [537, 518]. Mnemonic [1530]. Mobile [933, 941, 932, 2119, 936, 1232].

Mobius [1002]. Modal [694, 1734, 604, 160, 679, 695, 2275, 732, 1247]. Modality [1466]. Mode [1691, 691]. Moded [50]. Model [1496, 465, 452, 1727, 913, 792, 1128, 280, 417, 1216, 2456, 648, 1135, 398, 442, 1681, 463, 655, 1277, 600, 2316, 1232, 2374, 649, 2435, 2503, 1500, 673, 435, 1280, 2460, 892, 2245, 816, 2571, 2462, 1527, 1323, 1576, 2139, 1199, 1866, 782, 1247, 2301, 2500, 307, 1524, 933, 1133, 1327, 1310, 1441, 1267, 480, 2174, 1566, 692, 451, 2219, 1515, 943, 521, 1931, 1616, 2287, 1478, 456, 420, 425, 484, 576, 590, 2461, 1724, 2457, 1830, 2173, 1612, 1546, 2322, 1394, 460, 347, 1564, 2364, 1144, 1537, 880, 1295, 2207, 1513, 1283, 186, 1257, 1070, 1260]. Model-Based [2316, 1616, 1478, 576, 1612, 1394]. Modeling [414, 1992, 2499, 884, 1750, 1783, 2391, 799, 2016, 1288, 1787, 1790, 285, 862, 1793, 530, 2003, 2000, 1997, 1602, 1461, 709, 1805, 728]. Modelling [1609, 2375, 1884, 1606, 1414, 1499, 1888, 887, 630, 1621, 433, 941, 1354, 431, 1498, 440, 1953, 1762, 719]. Models [531, 373, 1781, 812, 2389, 2480, 566, 2214, 350, 1744, 1003, 893, 2215, 1497, 421, 1785, 1859, 1786, 2196, 1818, 307, 1789, 1981, 479, 423, 1614, 869, 457, 1990, 889, 434, 472, 1193, 1935, 883, 1268, 568, 891, 1275, 1657, 1055, 1073, 1747, 1236, 1100, 1253]. Modes [403]. Modifiable [1886]. Modifications [2331]. Modified [2234, 183, 1888, 1356, 550]. Modifier [1886]. modifiers [1773]. Modifying [1519]. Modular [1234, 25, 1855, 132, 537, 1373]. Modularity [787, 23, 1071]. Modulation [515]. module [1098]. Modules [523, 802, 169, 1761]. Moduli [104]. modulo [101]. MoHA [398]. Molds [965]. Molecular [1709]. Moment [1453, 1444]. Monad [1032]. Monadic [1172, 825, 832]. Money [755, 1353]. Monge [1695]. Monitor [1858]. Monitoring [1946, 1778]. Monitors [1800]. Mono [1379].

Monoid [344]. monoidal [254]. Monoids [334, 1034, 1059]. Monolithic [1842]. Monotone [2201, 1702, 2172]. Monotonic [385, 2168, 147, 2266, 192]. Monotonically [1029]. MONTAGE [839]. Montgomery [132, 134]. Montgomery-suitable [134]. MOODsup [805]. morphic [1091]. Morphisms [774, 1061, 258]. Morphological [1406, 1418, 1468]. MOSIX [312]. Mostly [1693]. Motion [935, 1427, 1429, 1434, 1426, 1403, 1425, 1432, 934, 933, 565, 1428, 932, 582]. Motion-Corrupted [582]. Motions [1935]. Motivated [1736, 459]. Motor [2361, 2231]. Move [2043, 1573]. Movement [591]. moves [1235]. Moving [1685, 2230]. MR [581, 570, 1468, 1470, 594]. MRI [1388]. MSM [2361]. Muldimedia [1121]. Muller [1249, 843]. Multi [485, 1891, 1123, 2050, 2361, 143, 230, 649, 973, 1305, 920, 1423, 984, 1466, 1126, 521, 695, 425, 2275, 658, 568, 1762, 1352]. Multi-Agent [143, 1762]. Multi-Class [1423]. Multi-Commodity [984]. Multi-Context [649]. Multi-Criteria [658]. Multi-Dimensional [973]. Multi-Layer [521]. Multi-Level [485]. Multi-list [230]. Multi-Modal [2275]. Multi-Modality [1466]. Multi-Model [425]. Multi-Object [920]. Multi-Party [1126, 1352]. Multi-Perspective [1305]. Multi-Purpose [2050]. Multi-Resolution [568]. Multi-Sensori-Motor [2361]. Multi-Sourced [1123]. Multi-User [1891]. Multicast [1133, 1132]. Multicommodity [2573, 1706, 2577]. Multicomputers [971, 2307]. Multicounter [240]. Multicut [739]. Multidatabase [1502, 2468, 1504]. Multidimensional [1819, 1029]. Multilayer [534]. Multilevel [2499, 1510]. Multilinear [2161]. Multilingual [2469]. Multilisp [2287]. Multimedia [1128, 1135, 2470, 1136, 1123, 1142, 1141, 1145, 2471, 1809, 1159, 1880, 1113, 1807, 1152, 1161, 1138, 1124, 1126, 1114, 1127, 1808, 1132, 1155, 1146, 1803, 1149, 1112, 2351, 1157, 1878, 1117, 1144, 1116]. multimediale [1920]. Multimodal [1267, 2266, 2360]. Multimodality [575]. Multimodel [397]. Multiple [916, 1540, 2503, 1532, 2397, 1287, 624, 1989, 889, 2138, 1289, 896, 1480, 1510, 1276, 1638]. Multiple-Valued [624]. Multiplication [1579, 2087, 253]. Multiplicative [698]. multiplicities [1230]. Multiprecision [1578]. Multiprocessor [1784, 1855, 1860, 2466]. Multiprocessors [1217, 1856, 1871]. Multiprogramming [1857]. Multiresolution [1405]. MULTIS [1623]. Multiscale [577, 571]. multiset [1231]. Multisignature [2027, 1341]. MultiSpectral [1492]. Multistage [540]. Multithreaded [2304]. Multivariate [595]. Multiversion [1831]. MultiView [1877]. Multiwindow [2292]. Multlog [703]. MulTVision [2293]. music [2600]. Mutual [646, 681]. mutually [1342]. Myopic [1693].

NaDSet [1183]. Name [2146, 1957]. Names [826, 1049]. Naming [337]. Narrowing [1223, 55, 820, 1209, 1208, 155]. Natural [1908, 1175, 308, 1900, 641, 165, 2397, 309, 535, 1771, 720, 1649]. Nature [2504]. Navigating [2366, 2354]. Navigation [2368, 2489, 1755]. NDB [2476]. NDB-PTIME [2476]. Nearby [1947]. Nearest [977, 1399]. Necessary [1434, 2093, 2032, 1655]. Need [1696]. Needed [538]. Needs [884, 1911]. Negation [2486, 154, 1764, 617, 801, 2012, 155, 299, 830, 2485]. Negation-Complete [830]. Negations [601, 2261]. Negative [32, 800, 1185, 265]. Negative-Conditional [31]. Negotiation [1890, 304]. Neighborhood [991]. Neighbors [977, 1399]. Neighbourhood [496]. Ness [316]. Nested [914]. Net [317, 320, 525, 529, 1249]. Nets [780, 2531, 314, 315, 318, 319, 321, 536, 323, 1027, 776, 326, 325, 1515, 751, 444, 329, 82, 7, 214, 1077, 1261, 1354]. Network [531, 1551, 1781, 1627, 2238, 509, 516, 539, 515, 523, 563, 501, 506, 2101, 1868, 546, 507, 2120, 540, 1943, 468, 1873, 487, 1785, 464, 2570, 1789, 606, 469, 547, 480, 477, 478, 548, 1893, 550, 1127, 519, 2211, 1155, 456, 556, 1494, 561, 1112, 520, 500, 560]. Networked [1853]. Networks [2164, 471, 474, 552, 633, 1678, 466, 555, 554, 475, 482, 463, 1148, 765, 947, 1783, 494, 562, 2111, 514, 533, 2215, 551, 540, 412, 577, 366, 470, 2570, 1111, 528, 2258, 2162, 2262, 358, 353, 479, 1133, 1327, 473, 462, 383, 496, 413, 1492, 364, 1699, 1493, 521, 2163, 2557, 1795, 517, 557, 493, 292, 532, 1798, 2081, 647, 1801, 2217, 499, 288, 476, 2577, 512, 578, 535, 1256, 2240, 1775, 1255, 74, 1661, 1754, 1668, 64, 211, 1757, 1670]. Neural [531, 2164, 461, 465, 471, 474, 552, 553, 554, 509, 475, 543, 516, 539, 515, 523, 563, 562, 2195, 514, 501, 506, 533, 2215, 546, 507, 540, 536, 468, 487, 470, 528, 467, 2162, 449, 479, 473, 462, 480, 477, 413, 1492, 478, 1493, 548, 1757, 519, 517, 525, 557, 2211, 456, 493, 556, 292, 1494, 532, 561, 529, 520, 2217, 499, 476, 500, 512, 578, 535, 560, 1754]. Neuro [412]. Neuro-Fuzzy [412]. Neurocomputer [511]. Neurocomputers [510]. NeuroGraph [512]. Neuron [522, 524]. Neurons [485, 452, 457, 538, 458]. Neutralization [702]. Next [1449, 1537]. NFA [2565]. NIST [2058, 2069]. NMFS [1155]. NN [1399]. NN-Set [1399]. No [2043, 299]. Node [464]. Nodes [1817]. Noise [1391, 543, 593, 1386]. Noisy [522, 1400, 595, 1108, 97]. Noisy-Leaky [522]. Non [531, 1251, 1407, 1362, 2476, 115, 116, 1387, 385, 2103, 2079, 1966, 2532, 2510, 2440, 1033, 2266, 2228, 1168, 565, 579, 697, 1435, 932, 1935, 2366, 2552, 2454, 504, 476, 2240, 7, 1073, 255, 361, 128]. Non-Classical [1168]. Non-Clausal [697].

Non-Constructive [1966]. Non-Crossing [2552]. Non-Deterministic [2476, 2532, 2454]. Non-Existence [2103]. Non-Greedy [504]. Non-Hierarchical [2366]. Non-holonomic [932]. Non-interactive [1362, 115, 116, 2079, 128]. Non-Lambertian [1435]. Non-Linear [531, 1407, 2228, 579]. Non-Local [2510]. Non-Monotonic [385]. Non-Orthogonal [476]. non-probabilistic [255]. Non-Rigid [565]. Non-Singular [2440]. Non-Synchronized [1935]. Non-Uniform [1387]. Nonasymptotic [1346]. nonconvex [65]. nondeterminism [1064, 196]. Nondeterministic [750]. Nonequivalence [10, 1344]. nonessential [1092]. Nonlinear [412, 94, 530, 1343]. nonlinearity [1344]. Nonmonotonic [136, 613, 147, 1167, 611, 1194, 2275, 1748]. Nonmonotonicity [2235]. Nonparametric [1487, 2215]. Nonperfect [1337]. Nonsupersingular [2087]. Noodle [2496]. Normal [2234, 813, 36, 148, 1471]. Normalization [438, 52, 172, 177, 183]. Normalized [1451]. Notation [2400, 1241]. Note [2160, 2046, 1734, 127, 384]. Notes [1668, 2177, 827]. Nothing [1507]. Notion [169, 179, 1637, 1252]. Notions [2313]. Novel [1212, 857, 2395, 2420, 1359]. Novice [2358]. Nozzle [553]. NP} [2077, 815, 2078, 822, 825]. NP-Complete [822]. NP-Completeness [815]. NP-Optimization [825]. Nucleus [456]. Number [2575, 2053, 464, 1934, 2020, 266, 1074]. Numbers [2117, 2524, 2047]. Numeric [2001]. Numerical [1874].

O [1128, 1141, 925, 1572]. O-Complexity [954]. O-O [925, 1572]. ObjChart [924]. Object [1496, 2337, 1568, 438, 913, 2143, 2486, 2495, 912, 928, 2134, 2154, 417, 2456, 430, 2137, 665, 911, 2477, 2508, 791, 1529, 786, 2493, 1851, 2503, 435, 1875, 2498, 1559, 1865, 669, 2230, 2339, 799, 920, 507, 923, 924, 2144, 1825, 1159, 2129, 1880, 1587, 1497, 2139, 1866, 906, 2145, 1152, 486, 1296, 2146, 1523, 1818, 1815, 1455, 1864, 2492, 908, 436, 2141, 2155, 2436, 1511, 2136, 1562, 805, 2142, 2338, 2497, 915, 2496, 803, 1467, 1877, 2340, 415, 434, 1572, 1520, 2483, 918, 400, 2125, 2455, 2342, 1509, 2457, 667, 1830, 2307, 1849, 418, 1312, 1550, 2147]. Object [2489, 446, 1565, 2390, 2128, 2133, 2482, 1081, 12, 1241, 16]. Object- [1864]. Object-based [1509, 12, 1241, 16]. Object-Centered [2136]. Object-Orientation [2128]. Object-Oriented [1496, 2337, 1568, 438, 2495, 912, 928, 2154, 417, 430, 665, 911, 2477, 2508, 791, 1529, 2493, 1851, 2503, 435, 1875, 2498, 1559, 1865, 669, 2339, 799, 507, 923, 2144, 1159, 2129, 1880, 1587, 2139, 906, 2145, 1152, 1523, 1455, 1864, 436, 2141, 2155, 2436, 1511, 1562, 805, 2142, 2338, 2497, 915, 2496, 1467, 434, 1572, 2483, 918, 2125, 2342, 2457, 667, 1830, 2307, 1849, 418, 1550, 2147, 2489, 446, 1565, 2390, 2133, 2482]. Objectively [1542]. Objectives [866]. ObjectMath [1945, 1586]. Objects [916, 1965, 664, 2499, 1919, 509, 1831, 791, 939, 2, 977, 1409, 2152, 1848, 2502, 1525, 2136, 2473, 1850, 1211, 444, 1877, 2138, 332, 576, 919, 2151, 1420, 724]. Obligations [1330]. Oblivious [2559, 2042, 111]. Observability [1990]. Observable [1049, 13]. Observation [2374, 1778]. observational [13]. Observations [2504, 1312]. Observers [2070, 597]. Observing [1065, 1089, 1962]. Obstacles [2520, 992]. Obtain [496]. Obtained [570, 120]. Occam [1637]. Occlusions [1442]. Occur [2393]. OCR [1915]. odd [1357]. Odor [414]. Off [755]. Off-Line [755]. Offering [1927]. Office [2375, 1830]. Offices [1914, 1925]. offs [1099, 750]. Offsetprinting [1479]. OLE [1923]. omega [787, 257, 197, 258]. omega-languages [257]. omega-regular [258]. Omnipotent [1300]. On-Demand [1803, 1117]. On-Line [1883, 2588, 556, 561, 1001]. ONC [1835]. once [1640]. ONE [380, 753, 842, 2023, 2046, 988, 122, 1036, 2024, 698, 1398, 998, 2198, 206, 1365, 109, 1366, 95, 1339, 713]. One-Dimensional [998]. One-Hot [842]. One-Prover [2046]. one-round [95]. One-Sided [698, 2198]. One-Way [753, 2023, 122, 1036, 2024, 1366, 1339]. Online [957, 2547]. Only [1173, 613, 1166]. onto [1754]. ontolingua [714]. Ontologies [1623]. ontology [726]. OO [2453]. OOAD [2458]. OODB [2138]. Open [1845, 2234, 1135, 313, 2400, 1840, 1864, 1191, 829, 1563, 1083]. Operating [312, 1864, 306, 2306, 510, 2133, 1763]. Operation [170]. Operational [820, 1960, 277, 1262, 1266, 241]. Operations [1817, 731, 795, 1534, 356, 1776, 724, 1656]. Operator [1397, 2475, 1764]. Operators [385, 1735, 1421, 406, 729, 215, 714, 1650, 1765]. Optical [1430, 580, 548]. Optimal [2537, 951, 1675, 811, 1677, 2572, 55, 2597, 1277, 542, 63, 1683, 960, 1663, 2571, 1857, 1408, 746, 933, 364, 1699, 2180, 2034, 999, 2583, 229]. Optimally [2576, 754, 204]. Optimisations [1220]. Optimised [475]. Optimization [1223, 811, 339, 1782, 466, 2418, 2480, 2518, 2449, 1521, 324, 1629, 478, 2482, 1643, 825]. Optimizations [2430]. Optimized [335, 1970, 1847, 492, 695, 848]. Optimizers [2481]. Optimizing [500]. Opto [2231]. Opto-motor [2231]. Optoelectronic [1854]. Or-Parallel [1724, 2286]. Oracle [124, 831]. Oracles [2568, 239]. ORB [1849]. ORCA [840, 2423]. Order [674, 157, 465, 471, 2515, 466, 34, 911, 1729, 301, 2514, 683, 1415, 33, 818, 1966, 2291, 1585, 2533, 528, 789, 628, 781, 462, 457, 35, 283, 1952, 569, 1219, 670, 699, 832, 56, 573, 46, 706, 218, 175, 77, 185, 720, 191, 1049, 1103, 1246].

Order-Sorted [674, 157, 301, 1585, 699, 46]. Ordered [2488, 684, 1029, 2516, 212]. Ordering [370, 372, 2574, 2259, 2583]. Orderings [29, 2256]. Orders [981, 328, 40, 833, 59, 60]. Ordinary [2055]. Organisation [2210]. Organisational [868]. Organising [481, 495]. Organization [2500, 2218]. Organizations [1712, 1807]. Organized [498]. Organizing [844, 549, 2371, 511, 551, 480, 496, 397, 545, 526, 534, 499]. Orientation [2217, 2128, 1272, 1670]. Orientation-Dependent [2217]. Oriented [1496, 2337, 1568, 438, 2495, 912, 928, 2154, 417, 2350, 618, 161, 430, 665, 911, 2477, 2508, 791, 2221, 1529, 2493, 1851, 2503, 435, 1875, 2498, 1559, 1865, 669, 2339, 799, 507, 1326, 923, 2144, 1159, 2129, 1880, 1587, 2139, 906, 2145, 1152, 1523, 1455, 1864, 908, 436, 2141, 2155, 2436, 1511, 1562, 805, 2142, 2338, 2497, 2348, 915, 2496, 1604, 1467, 434, 1572, 2483, 918, 2125, 2342, 2457, 667, 1830, 2307, 1849, 418, 1550, 2147, 2489, 446, 1565, 2390, 2133, 2482, 2407, 1342, 1569, 1081]. Orthogonal [960, 1865, 191, 476]. Oscillator [454]. Oscillators [474]. OSF [1833, 1851, 1850, 1841, 1843]. OSI [913, 2118]. Osub [1568]. Other [1304, 947, 2543, 1963, 184]. Output [2182, 1339]. overlap [223]. overlap-free [223]. overloading [178]. Overview [790, 2530, 835, 320, 2284, 1793, 1572, 2484, 1964]. own [1904]. Oxford [1535]. Oz [2360].

p [1371, 2290, 204, 1934, 197]. P-adic [1934]. P-Continuation [2290]. P/poly [204]. PAC [137, 2175, 2177, 2174, 2180, 2173]. Package [1483]. Packaging [893, 895]. Packet [946, 1111, 1701, 1801]. Packet-Switched [1111]. Packing [1693, 231]. PACLIB [1595, 1944]. Page [1820, 2585, 2588]. Paging [1674, 2586]. PAGODE [1224]. PaiLisp [2291, 2290]. Pair [1447]. Pair/Completion [41]. Pairs [945, 741, 786, 985, 2060, 2020, 227]. Pancake [366]. Pancreatic [454]. PANDA [921]. Panel [103, 928, 104]. Paper [1143, 1147, 1150, 1154, 1158, 2002, 2117, 514, 867, 527, 898, 449, 505]. Papers [2284]. Paradigm [1304, 858, 2305, 403, 2346]. Paradigms [137, 857, 2229, 361]. Paradox [876]. Paragon [1946]. Parallel [1507, 2304, 1212, 1229, 365, 1780, 1965, 1216, 963, 2305, 2424, 2430, 129, 2302, 350, 229, 616, 960, 1686, 2435, 1003, 102, 360, 511, 1945, 540, 1809, 2085, 2590, 1873, 2432, 2309, 2293, 2284, 1521, 1859, 1939, 2574, 2308, 1857, 354, 2285, 982, 2416, 768, 1447, 987, 2511, 2436, 2434, 1934, 856, 988, 1791, 361, 691, 1937, 226, 355, 1213, 2567, 1871, 1961, 65, 351, 1702, 363, 1797, 349, 2080, 2417, 767, 1944, 852, 1724, 362, 1938, 1869, 1935, 520, 2292, 2465, 346, 2303, 1420, 2442, 588, 2446, 2463, 2297, 1068, 1240, 63, 1073, 238, 1086, 64]. parallel [1084, 1083, 69, 1659, 212, 2286]. Parallel- [588]. Parallelisation [922]. Parallelism [1862, 2425, 1868, 1940, 1870]. Parallelization [2429, 972, 1941, 1942, 1257]. Parallelizing [2419, 2449, 2439]. Parameter [1952, 237]. Parameterisation [1445]. Parameterization [1402]. Parameterized [325, 1581, 1952, 567]. Parameters [1429, 1638]. Parametric [779, 195]. Paramodulation [684]. Parsing [1975, 682, 1204]. Part [2250, 1895, 572]. Partial [424, 784, 2447, 818, 981, 2039, 781, 2171, 48, 1000, 1246, 1711, 173, 1772, 77]. Partial-order [1246]. Partially [2488, 2082, 291, 470, 225]. Partially-Bent [2082]. Participants [2310]. Participatory [1624]. Partition [1825]. Partition-Based [1825]. Partitioning [2288, 2421, 986, 1767]. Parts [433]. Party [1126, 1342, 1352, 1671]. Passing [2289, 1952, 1846, 1245, 1244]. Passive [1329]. Patches [573]. Patching [841]. Path [29, 2586, 983, 1030, 992, 2489, 1295, 272, 228].

Paths [1675, 930, 1691, 985, 2578, 1705, 2552, 1708, 1107, 211, 278]. Pathwidth [743]. Pattern [139, 964, 563, 2212, 1454, 2144, 976, 470, 2184, 758, 528, 1941, 1870, 413, 2163, 538, 1037, 1009, 2218, 1450, 1386, 595, 592, 1639]. Pattern-Matching [1941]. Patterns [1419, 923, 853, 2178, 503, 1374]. PCF [186]. PCTE [2148, 1565]. PCTE-Based [2148]. PDEs [1873]. pebbling [1099]. PEIRCE [2392, 731, 730]. Pen [1455]. Pen-Based [1455]. Penguin [1818]. People [1911]. Per-Process [2427]. Perceiving [2224]. Perception [398, 939, 2225, 1488, 2600]. Perceptive [2229]. Perceptrons [534]. Perceptual [2219, 2205, 940, 2226]. Perfect [2239, 2595, 115, 2077, 2034, 2043, 116, 1637, 1333]. Perfectly [2096]. Perfectly-Secure [2096]. Performability [1802]. Performance [857, 633, 357, 1782, 1783, 1836, 1784, 1378, 1859, 840, 2286, 1518, 1790, 1837, 1937, 1901, 521, 1119, 556, 447, 1520, 1797, 1517, 1946, 1928, 1400, 1799, 1813, 596, 1800, 1794]. Period [2039]. Periodic [774, 1107]. Perishable [1553]. permanent [1664]. Permutation [743, 2558, 2033, 2559, 57, 108, 1363]. Permutations [2094, 964, 94, 110, 2024, 2032, 2022, 1365, 109]. Permuted [2084, 1695, 843]. Perpetual [2569]. Perrin [1679]. Persistence [972]. Persistent [2143, 509, 2467, 420, 2299]. Personal [1619]. Personen [1910]. Perspective [925, 1305, 399, 1361]. Perspectives [2343, 1927, 445, 2357]. Petri [82, 215, 780, 314, 316, 214, 1077, 317, 318, 319, 320, 321, 323, 1261, 1027, 326, 1515, 1354, 444, 329]. Petri-Nets [1515]. PFL [672]. Phase [2121, 1969]. Phases [286]. Philip [1053]. Philosophies [875]. Phoneme [1743]. Photometric [1445, 1435]. Photon [585]. Physical [1687, 2123, 1809, 1807, 2581, 1997].

Physically [943, 565]. Physics [516, 2015, 520]. Physiological [590]. pi [1238, 180, 1231, 1241, 164, 1007, 1239, 1265]. pi-Calculus [164, 1007, 1238, 1231, 1241, 1239]. Picture [1482]. Pieces [1922]. Piecewise [492]. Pignistic [2280]. Pilgrim [1841, 1844]. pipe [1058]. Pipeline [1929, 1936]. Pipelining [1932]. Pitts [1049]. PL1 [2237]. Place [2554]. Placement [844]. Plaintext [2059, 93]. Plan [294, 2329, 1776]. Planar [741, 1686, 977, 2574, 1691, 1692, 982, 985, 1702, 332, 1705, 1707, 1708, 2577, 960]. Planarity [1677, 1692]. Plane [945, 637, 930, 977, 2520, 1690, 1434, 2552, 2596]. Plane-Sweep [977]. Planlite [558]. Planning [935, 1735, 931, 1553, 1285, 298, 934, 933, 608, 1449, 1554, 634, 558, 2171, 2401, 932, 1775]. Plans [1605, 2402]. Plaque [591, 1469]. Platform [1135]. Plausible [403, 2386, 2279]. Player [1156]. PLDs [835]. Plurality [907]. Point [955, 2597, 970, 1685, 979, 1410, 2081, 1437]. Pointer [2421, 2408]. Pointers [1953]. Points [1690, 1446, 1282, 1456]. Policy [2306]. Polling [1804]. poly [204, 189]. poly-time [189]. polyadic [1265]. Polychromatic [1448]. Polygon [950, 2548, 2576, 228]. Polygonal [1409, 2596]. Polygons [951, 1020, 999, 986, 1710]. Polyhedral [965, 1409, 994, 1442]. Polyhedron [974]. Polymorphic [51, 1957, 1650]. polymorphism [194, 195]. Polynomial [139, 112, 2545, 2161, 1186, 2394, 2525, 831, 258]. Polynomial-Time [112, 2161, 831]. Polynomially [738, 2565]. Polynomials [2174, 767]. Polytope [968, 1257]. polytopes [65]. Polytrees [2257]. Pool [1191]. Popperian [2191]. Portability [363]. Portable [1599, 106]. Portuguese [1743]. Poset [1023]. Posets [2540]. Position [867, 898]. Positive [32, 315, 2167, 2359, 31]. Positive/Negative [32, 31]. Positive/Negative-Conditional [31]. Possession [2044]. Possibilistic [2235, 2237, 378, 2253, 628]. Post [1052, 1109]. Posters [1298]. Postman [1685]. Power [961, 1782, 2045, 2009, 774, 775, 2567, 751, 767, 238, 1064, 85, 1340]. Power-Series [1782]. powerdomain [1097]. Practical [589, 2406, 2241, 2370, 1334, 1308, 125, 469, 2319, 2066, 359, 1938, 1000, 2083, 1338, 1233, 1351, 1341, 1629, 1636]. Practice [430, 1825, 876, 2529, 362]. Pragmatic [2311, 1319]. Pragmatics [299]. PRAM [1683, 1088, 347]. PRAMs [961]. Pratt [1050]. Pre [286]. Pre-Formal [286]. Precedence [1438]. Precise [263]. Precomputation [105]. Preconditions [144, 277]. Predecessors [989]. Predefined [21]. Predicate [1171, 917, 144, 325, 695, 2171, 1041, 1063, 273]. Predicate/Transition [325]. Predicates [1162, 2518, 825, 1262]. Predicative [1948]. Predict [1505]. Prediction [531, 543, 1401, 2176]. Prediction/Correction [1401]. Predictions [560]. Predictive [1151]. Predictor [1392]. Preemption [2509, 702]. Preferences [379]. Preferential [2234, 2275, 658]. Preferred [149]. Prefix [366]. preimages [1366]. Preliminary [675, 2076, 2229, 1131, 614, 1732, 164]. preorders [1043]. Preprocessing [115, 1478, 116]. Prerequisites [1288]. Presburger [1182]. presence [1666, 1060]. Presentation [2371, 2353, 1124]. Presented [2057]. Presenting [730]. preserved [1254]. preserving [1254]. Pressure [532, 1531]. Presynaptic [453]. Preventative [1316]. Prices [559]. Pricing [2073]. Primal [739]. Primal-Dual [739]. Primality [1377]. Primary [1885]. Prime [2053, 1364, 697, 76, 1749]. Primes [2089, 104]. Primitive [1419, 1016, 2167].

Primitives [2219]. Primitivity [1052]. Principal [1265]. Principle [2238, 1449, 696, 1192]. Principles [671]. Printed [1457]. Prints [2579]. Prior [583]. priorities [1258]. Prioritization [2285]. Priority [973, 2286, 1652]. Privacy [2061, 1352]. Private [2450, 2040]. Private-Key [2040]. Private-Memory [2450]. Probabilistic [137, 753, 2238, 382, 1744, 2246, 2254, 2570, 1524, 78, 2270, 2173, 2051, 2209, 151, 656, 2240, 256, 131, 255, 1768, 232]. Probabilistics [2262]. Probabilities [2268, 832, 2280]. Probability [627, 754, 2276, 1372]. Probable [2089]. Probe [955, 264]. Probing [953]. Problem [1608, 2477, 655, 1014, 741, 337, 2551, 1017, 2598, 2561, 1418, 977, 2126, 2588, 345, 746, 1614, 887, 1033, 1114, 1604, 2269, 1924, 296, 1702, 544, 645, 1623, 2170, 2539, 2008, 1054, 63, 214, 206, 1370, 1047, 231, 1052, 720, 2410, 213, 2592]. Problem-Oriented [1604]. Problem-Solving [1614, 1623]. Problems [485, 953, 2546, 811, 962, 966, 815, 741, 1178, 2547, 1683, 970, 1685, 971, 1686, 975, 2246, 2586, 1712, 822, 2520, 978, 2523, 1730, 2550, 738, 1999, 825, 2210, 862, 737, 740, 938, 1169, 1902, 631, 2522, 1037, 1009, 1705, 2583, 1000, 2596, 1001, 2580, 2422, 1666, 57, 229, 60, 250, 1777]. Procedural [416, 2358]. Procedure [369, 1172, 684, 2183, 2517, 153, 97]. Procedures [1740, 335, 2431, 1589, 41, 387, 1103]. Process [1727, 1012, 1391, 555, 1329, 442, 1558, 1580, 385, 11, 1784, 2535, 1306, 875, 2427, 1960, 1328, 1327, 862, 1267, 869, 634, 2228, 870, 1922, 871, 447, 405, 406, 1561, 150, 1312, 2354, 2170, 1251, 1266, 1233, 1078, 1637, 1244, 1070, 1263]. Processes [2389, 1277, 290, 1027, 782, 1793, 863, 329, 1658, 200, 1065, 1240, 1232, 1770, 1102, 1256, 1058]. Processing [1820, 485, 1128, 508, 1460, 1485, 1476, 616, 2073, 617, 1472, 1483, 1423, 1475, 2334, 2010, 2158, 1525, 1505, 557, 612, 297, 1481, 2446, 1805, 2463, 2297]. Processor [2304, 1932, 527, 517]. Processors [352]. prod [12]. Producing [2366]. Product [2432, 874, 2557, 769, 229]. Production [1772, 2237, 2385, 1555, 1553, 1554]. Production-Rule [2237]. Products [772, 1031, 561]. Professionals [2359, 1903]. Profile [859, 2418]. Program [2479, 177, 734, 2449, 2293, 2448, 2321, 376, 1958, 1604, 2411, 621, 2417, 153, 880, 268, 1643, 1647, 270, 1044, 2412]. Programmable [357, 836, 843, 846, 837, 848]. Programme [2121]. Programmer [347]. Programmers [1837]. Programming [1625, 2591, 157, 1727, 921, 2235, 1672, 1329, 161, 632, 162, 1299, 644, 2424, 1210, 639, 168, 2435, 159, 154, 802, 1945, 704, 820, 1948, 823, 2013, 688, 789, 165, 436, 2141, 169, 2261, 490, 2436, 1200, 1167, 1937, 915, 2130, 1211, 2527, 1169, 152, 1206, 163, 1242, 292, 166, 904, 2125, 1938, 1849, 2326, 505, 2357, 1846, 1196, 2405, 81, 1086, 275, 1760]. Programs [2512, 2318, 1223, 1547, 2324, 1974, 424, 734, 51, 1331, 1198, 1318, 2426, 785, 2316, 2447, 2313, 2317, 1872, 617, 50, 160, 800, 2432, 2437, 1199, 1859, 2320, 689, 2161, 2325, 620, 2149, 988, 2473, 691, 1722, 2582, 599, 331, 1190, 1961, 1725, 612, 1959, 1950, 700, 830, 2292, 702, 1842, 2186, 2315, 1042, 1641, 2406, 1648, 256, 274, 242, 1092, 1636]. Progress [2002, 1548, 1901]. Project [806, 2302, 1604, 425, 1624, 1822]. Projecting [1754]. Projection [501, 2394]. Projections [942]. projective [1749]. Projectivity [528]. Projects [1894, 896]. Prolog [1042, 676, 1216, 1162, 1202, 1221, 685, 1203, 2286, 169, 1629, 167, 1724, 700, 1628]. Prologs [2167]. Promises [1937]. Proof [1171, 2050, 1330, 1017, 2046, 1183, 1863, 1244, 679, 165, 1087, 602, 612, 153, 2278, 1605, 172, 1746]. Proofs [809, 810, 2091, 1591, 115, 1179, 2078, 704, 50, 2079, 2048, 30, 1190, 2526, 251, 177, 116, 183]. Propagation [1742]. Properties [1781, 811, 442, 2543, 1951, 1023, 1004, 36, 1188, 2192, 1972, 49, 1959, 2277, 2051, 832, 1962, 1748, 1645, 236, 1076, 77, 193, 1049, 1098, 1254]. Property [2454, 247]. Propose [1537]. Proposed [2058, 2069]. Proposition [1398]. Propositional [694, 2271]. ProSet [1939]. ProSet-Linda [1939]. PROSPECTRA [376]. Prosthesis [903]. protection [1346, 1333]. Protein [2176, 1461]. Protocol [2042, 1155, 111, 120]. Protocols [129, 1783, 2111, 2054, 2095, 2080, 2051, 113, 1654, 1349, 1355, 1354, 1353, 99, 1352]. Prototype [855, 1840, 2132, 804, 1531, 2446]. Prototype-Based [2132]. Prototypes [806]. Prototyping [1939, 849, 423, 2356, 845, 1206, 2457]. provability [1748]. Provable [1970, 2105]. Provably [2063, 2066, 113]. proven [1334]. Provence [1328]. Prover [2046, 152, 730, 1594, 2078]. Provers [1591, 690]. Provide [1402, 1549]. Providing [1801]. Province [1925]. Proving [2512, 2515, 1575, 339, 117, 1220, 600, 815, 1729, 642, 43, 33, 2183, 35, 282, 38, 1745]. Proving-by-Example [2183]. Provision [1922]. Proximity [2520, 1416]. Proxy [2473]. Prudence [2169]. Pruning [678, 493]. Pseudonyms [2061]. Pseudorandom [2033, 110, 2032, 1362, 108, 109]. PSPACE [740]. PSPACE-Complete [740]. Psychobiological [460]. Psychologies [1287]. Psychology [2210, 1619]. PTIME [2476]. Public [2110, 1921, 1370, 100, 2092, 2095, 2072, 128]. Public-Key [100, 2072, 1370, 128]. Pumping [1030]. Pure [946, 190]. Purpose [2050, 2035, 1605, 1629]. pushdown [235]. pushout [79]. Puzzle [1417]. PVM [1943, 1835]. Pyramid [1381, 1413]. Pyramidal [1744]. Pyramids [1383].

Q [295]. Q-Functions [295]. QOS [1153]. Quadratic [2086, 2560]. Quadtree [1384, 1382]. Quadtrees [963, 1383]. Qualitative [2016, 1270, 1296, 1271, 397, 406, 2015, 407, 863, 1997, 1272]. Quality [963, 1473, 1476, 1896, 1801, 2190, 1531]. Quantification [1476, 645, 1461]. Quantifications [1217, 676]. Quantified [618, 301]. Quantifiers [1966]. Quantitative [884, 1315, 869, 1775]. Quantization [501, 483, 1394]. Quantizer [1491]. Quantizers [504]. Quantum [1872]. Queries [2337, 618, 301, 1519, 1819, 1521, 1188, 1832, 992, 2201, 2489, 1275, 228]. Querrying [1523]. Query [2591, 1820, 2160, 2488, 966, 1506, 2480, 2460, 2481, 2462, 1527, 2139, 2321, 2010, 1525, 801, 1505, 2490, 831, 2482, 2463, 221]. query-bounded [221]. Querying [1541, 2496, 1276]. Questions [299]. Queue [973, 1652, 1792]. Queueing [1781, 1785, 1786, 1789, 1795]. Queues [2555, 1798]. Quick [1665]. Quiescence [1252].

R [1042, 2357]. R-Technology [2357]. Rabin [1059]. Race [988, 2444, 201, 2121]. Rackoff [108]. Radial [468, 469]. Radial-Basis-Function [468]. Radiation [311]. Radiological [2362]. Radiology [1460, 2399]. Rana [450]. Random [2568, 1688, 2045, 2545, 1414, 470, 2511, 1441, 1494, 1704, 1137, 238]. Random-Dot [1494]. Randomised [2540]. Randomization [972]. Randomized [2587, 1429, 1861, 1403, 752, 2588, 1703, 2451]. Randomizers [2034]. Randomness [2092, 987, 358, 122, 1756]. Range [962, 1414, 1431]. Rank [344]. ranking [230]. Rapid [140, 849]. Rare [1786]. Raster [1295]. Rate [2074, 2075]. Ratio [1391]. Rational [1716, 1738, 1934, 2519, 769, 248]. Rattlesnake [1112]. Ray [1463, 1481, 833]. razor [1637]. RDBMS [715, 671]. Re [842, 2447, 304]. Re-Allocation [304]. Re-Encoding [842]. Re-execution [2447]. Reachability [325, 1707]. Reaching [2542]. Reaction [400]. Reactive [924, 1672, 1632]. read [1656]. Reading [2567]. Real [411, 1473, 1427, 1592, 1978, 1145, 1301, 489, 1472, 1905, 1151, 849, 1114, 1972, 431, 1796, 1303, 447, 2517, 1101, 1112, 2300, 1319, 1144, 1074, 1320, 1070, 1102, 1256]. Real-Time [1427, 1978, 1145, 1301, 489, 1905, 1151, 849, 1114, 1972, 1303, 447, 2517, 1112, 2300, 1319, 1144, 1320, 1101, 1070, 1102, 1256]. Real-World [431]. REAL92 [1972]. Realised [521]. Realistic [508, 457]. Reality [2232, 1904]. realizability [183]. Realizable [355]. Realization [1727, 487, 469]. Really [1114]. Reals [1173, 705]. Reason [2248, 1756]. Reasoning [2233, 797, 798, 796, 1603, 393, 1609, 136, 34, 379, 613, 2241, 1728, 2243, 1590, 1593, 410, 818, 2253, 2014, 290, 705, 1185, 1296, 403, 628, 1271, 149, 2386, 611, 696, 2015, 650, 1297, 1758, 1194, 626, 402, 1546, 2170, 1779, 1269, 709, 2387, 2487, 1272, 712, 1647, 1092]. Recognition [471, 414, 2362, 563, 2212, 1415, 1453, 2215, 546, 486, 528, 1695, 1462, 1459, 469, 547, 413, 2163, 459, 576, 472, 1456, 1386, 1458, 534, 1394, 592, 657, 725, 2600]. recognizable [1066, 224]. Recognize [538, 76]. Recognizer [2363, 1455]. Recognizing [332]. Reconfigurability [995]. Reconfigurable [1676, 2558, 1932, 856, 1701, 1489]. Reconstructing [996]. Reconstruction [2385, 632, 1392, 581, 1464, 1387, 587, 584, 1486, 981, 1441, 1451, 585, 1450, 582, 1442, 1452, 1710, 588, 586, 1060, 1359, 199]. Recording [2370]. Recovering [1429, 2264]. Recovery [1426, 2465]. Rectangles [1021].

Rectangular [2520, 992]. Rectilinear [2597, 2598, 2520, 228, 2552, 69]. Recurrence [457]. Recurrences [2562, 745]. Recurrent [531, 471, 482, 462, 477]. Recursion [786, 1188, 215, 1638]. Recursions [1974]. Recursive [808, 1820, 181, 638, 2432, 1966, 1204, 261, 2485, 74, 188]. Recursiveness [2524]. Red [993]. Red/Blue [993]. Redesign [2019]. Reduce [1214, 1673, 1598]. Reduced [1847, 1706]. Reduces [464]. Reducibility [1025, 756]. Reducing [1820]. Reduction [34, 784, 2191, 786, 1385, 2220, 2535, 171, 1238, 222, 1649, 1745]. Reduction-Based [2535]. Reductions [2045, 1056, 174, 1093, 197]. Redundancy [410]. Reed [843]. Reengineering [1570]. Reexamination [1166]. Refactoring [2140]. Reference [1907]. References [375]. Refined [1929]. Refinement [372, 1245, 1949, 1617, 621, 215, 276, 2409, 1765, 1101]. Refinements [370, 692]. Refining [1979, 2444]. Reflectance [1469]. Reflections [1643, 876]. Reflective [655, 2481, 2494, 2132]. Reflex [2231, 1515]. Reformulation [2181]. Refunds [2071]. Refute [2187]. Regime [2015]. Region [463]. Regional [1547]. Regions [1285, 2552]. Register [1228, 1359, 97]. Registration [575, 591, 1437]. Regression [1310]. Regular [471, 2541, 1680, 482, 771, 2194, 1031, 1722, 770, 1035, 762, 1259, 258]. Regularities [140]. Regularity [36]. Regularization [2164, 1392, 494, 1486]. Reinforcement [490, 295, 460]. Rejection [529]. Related [953, 962, 2108, 1685, 975, 2523, 36, 364, 1699, 2582, 1193]. Relating [710]. Relation [2456, 1027, 2259, 910, 309, 572, 1748, 13, 720, 208]. Relational [673, 2462, 1966, 275, 1523, 805, 437, 670, 2404, 717]. Relations [350, 1527, 2265, 1282, 391, 700, 1513, 1050, 1766, 1769, 1760]. Relationship [1304, 666, 2534, 2581].

Relationships [5, 885, 1555]. Relative [2233, 991]. Relative-Strength [2233]. Relativism [1498]. Relativistic [391]. Relativizing [2568]. Relators [827]. Relaxation [632, 1791, 594, 1770]. Relaxed [1931]. Release [2593]. RELFUN [158]. Reliability [1786]. Reliable [734, 765, 439]. RelieF [1910]. remark [128]. Remarks [1991, 1369]. Remembering [2544]. Removal [994, 1725, 488, 27, 833]. Reorganization [616]. repeated [1639]. Repetition [701]. Repetitive [994]. Replacement [313, 1015, 1006, 1030]. Replaces [1967]. Replanning [1327]. Replicas [1809, 1079]. Replicated [1826, 2465, 1660]. Replication [2588, 1822, 1231]. Report [103, 675, 905, 104, 614, 1548]. Reporting [2362, 978, 993]. Reports [2399]. Repository [2154]. Represent [1291]. Representation [461, 471, 939, 2223, 2339, 711, 2184, 1306, 1597, 453, 629, 400, 309, 574, 1543, 1619, 2396, 401, 718, 719, 1095]. Representations [2344, 960, 2441, 1560, 1384, 1918, 2227, 1754]. Representative [1282]. Represented [396]. Representing [390, 686, 643, 1814, 2381, 720, 2380]. Requests [2547]. Requirement [441, 643]. Requirements [1994, 2546, 2154, 1307, 427, 428, 925, 1995, 1308, 1113, 1306, 1889, 1127, 1968, 425, 429, 621, 1382, 101]. RES [2233]. Research [2302, 875, 876, 903, 1824, 2483, 2031, 1997]. Resequencer [1133]. Reservation [1504]. resilient [1661]. Resistance [2019]. resistant [1340]. Resolution [373, 369, 371, 1716, 694, 683, 816, 684, 385, 2266, 1168, 696, 1549, 2185, 387, 568, 1140, 1766, 238, 722, 689]. Resolution-Like [387]. Resolving [748]. Resonance [582, 578]. Resource [304, 1844, 101, 750]. Resource-Bounded [750]. resources [1258]. Respectable [2250]. Response [1785, 1847, 585, 2069, 2364]. Restoration [1392]. Restricted [2251, 2152, 2402, 1041, 1773, 1047, 60]. Restrictions [340]. Result [734, 2550]. Results [112, 1440, 2531, 1836, 1872, 978, 759, 2532, 990, 148, 1723, 164, 2036, 1931, 1382, 207, 1076, 1077, 1047, 1760]. Rethinking [871]. Reticular [456]. Retina [1488]. Retrieval [1324, 2336, 562, 2223, 2464, 1557, 1815, 149, 2466, 670, 1157, 724]. Retrieval-by-Contents [2223]. Reusable [1561]. Reuse [509, 441, 892, 893, 1326, 923, 922, 894, 895, 2526, 904, 440, 2402]. Reusing [891]. Revealing [2039]. Reversal [750, 240]. Reversal-Space [750]. Reversals [366]. Reverse [1326, 1969]. Reversing [2563]. Review [1322, 1614, 1490]. Revision [2371, 2241, 1728, 2259, 2271, 2282, 2501]. Revisited [812, 2491, 458, 2322, 264, 182]. Revolution [1821]. Rewrite [787, 2507, 32, 50, 53, 26, 175, 191]. Rewriting [2512, 52, 20, 54, 29, 1002, 1593, 25, 44, 47, 18, 1222, 17, 24, 35, 23, 2513, 40, 46, 42, 27, 28, 1642, 173]. RHT [1403]. Rigid [1184, 1444, 565]. Ring [1507, 2067]. rings [1669]. RISC [1592, 525]. RISC-CLP [1592]. RMTP [1131]. RMTP/RTIP [1131]. Road [1516]. Robin [1048]. Robinson [1181]. Robot [935, 2579, 1484, 934, 933, 941, 932, 1776]. Robotics [644, 937, 2003]. Robots [2232, 936]. Robust [1426]. robustness [219]. Rod [1044]. ROI [463]. Role [1611, 307, 358, 481, 883, 888]. rollback [1673]. Roof [1400]. room [1893]. Rotation [1396]. Rough [2072, 658, 2209]. Roughly [396]. Round [2097, 2044, 95]. Roundness [999]. Route [2548, 1267]. Router [359]. Routes [2549]. Routing [1678, 2572, 946, 539, 2558, 360, 736, 2571, 540, 2559, 2590, 1697, 541, 1701, 1703, 68, 1667]. Rozenberg [1051]. RPC [1848, 1847, 1835, 1846]. RSA [120, 107, 1355, 99]. RTIP [1131]. Ruby [276]. Rule [1174, 2237, 404, 409, 2480, 616, 2481, 1184, 290, 606, 2176, 1463, 432, 607, 656]. Rule-Based [404, 2480, 2481, 1184, 290, 1463, 607]. Rules [791, 635, 1421, 2412, 626, 2209, 401, 2485, 193, 198]. run [1656]. Running [2291]. Runtime [2438, 1227].

S [1345, 2021]. S-Boxes [2021, 1345]. S02 [2098]. SAC [2021]. SACLIB [1595]. SACLIB/PACLIB [1595]. Safe [776, 2531, 1249]. Safer [2071]. Safety [2243, 1022]. Salomaa [1052]. Salzburg [1925]. Sample [137, 1384, 2173, 2198, 595]. Sampling [1387, 2180]. Satellite [1422, 1492]. Satisfaction [2246, 1730, 768, 631, 578, 13, 1757]. Satisfiability [815, 631, 828]. Satisfying [39, 2021, 1343]. saturated [1096]. Saturation [323, 1192, 1967]. Saving [1679]. Scalable [971, 1861, 362]. Scale [1994, 1414, 1287, 1289, 432, 1351]. Scaling [1149]. Scanned [1459]. Scattered [1034]. SCCS [1243]. Scene [940, 2226]. Scenes [994, 1442]. Scheduler [1971]. schedules [1255, 727]. Scheduling [61, 1136, 2477, 2418, 1580, 1145, 2593, 2035, 1698, 2594, 1114, 1929, 212, 1144, 1062, 1775, 1662, 60]. Schema [1551, 2135, 416, 419, 2478, 1572, 2335, 1823, 1509, 418, 716]. Schemas [912, 799, 437, 1572]. Schemata [1567]. Scheme [2084, 2305, 2047, 2104, 1861, 107, 2296, 1185, 985, 2028, 2027, 2067, 163, 2307, 1116, 1351, 1341, 1356, 134]. Schemes [438, 2076, 1012, 2102, 86, 2074, 1686, 87, 2103, 126, 2066, 2038, 2075, 1338, 1342, 1336, 1337, 1350, 1265]. Schnorr [2057]. Science [1163, 2504, 864]. Sciences [1280]. Scientific [1829, 2334]. Scintigraphic [590]. Sconing [827]. Scope [2376]. Scrambler [2041]. Script [641]. Scrutiny [1322]. Seam [1472]. Search [1810, 952, 735, 633, 2089, 678, 1417, 1277, 2587, 2556, 679, 2005, 982, 948, 602, 605, 2516, 2106]. Searching [950, 2540, 736, 978]. Seasonal [1926]. Second [471, 1397, 911, 1729, 832, 573, 1050]. Second-Order [471, 911]. Secondary [2176]. Secret [2076, 86, 2074, 2104, 2054, 2028, 2095, 2075, 2112, 1338, 1351, 1336, 1349, 1337]. Secret-Key [2112]. Secrets [2042, 267]. Secure [2096, 755, 2041, 125, 126, 1349, 114, 2066, 113, 111, 1334]. Security [112, 2084, 129, 427, 2124, 2120, 2566, 2105, 2115, 1843, 2119, 2118, 2083, 2122, 1361]. See [1494]. seeking [1772]. Seen [760]. Segment [971, 993]. Segmentation [1457, 549, 577, 1413, 550, 545, 530, 1467, 1463, 568, 578, 2599]. Segmented [2220]. sehbehinderte [1910, 1912]. Selected [2022]. Selecting [870, 1778, 221]. Selection [2570, 987, 2228, 1847, 1701, 561, 1068]. Selective [2506, 452]. Selectivity [1819]. Self [2537, 1676, 813, 2369, 844, 855, 549, 314, 2587, 523, 2371, 498, 511, 551, 1025, 853, 480, 496, 457, 397, 545, 526, 2218, 481, 534, 499, 1195, 495, 1663, 243, 916]. Self-Adaptive [2369]. Self-Adjusting [2587]. Self-Interpreter [813]. Self-Learning [523]. Self-Organising [481, 495]. Self-organization [2218]. Self-Organized [498]. Self-Organizing [844, 549, 2371, 511, 551, 480, 496, 397, 545, 526, 534, 499]. Self-Reducibility [1025]. Self-Stabilizing [2537, 314, 243]. Self-Test [853]. Self-Timed [855]. Self-Verifying [1195]. SelfLog [1225]. Semantic [369, 371, 1174, 1627, 158, 2384, 2514, 1892, 1868, 799, 617, 2253, 661, 1956, 302, 408, 1498, 602, 1959, 1550, 2482]. Semantically [662, 600]. Semantics [21, 1740, 793, 2234, 32, 162, 1633, 382, 154, 2518, 817, 1738, 820, 823, 2535, 184, 1960, 2382, 2261, 17, 1957, 433, 1282, 1273, 405, 406, 1718, 1952, 1950, 2417, 1101, 299, 2275, 2358, 1262, 1266, 178, 200, 1231, 11, 12, 1241, 1673, 277, 1097, 241, 1235, 726, 1058, 1081]. Semaphore [988]. Semi [2547, 1002, 770]. Semi-Commutations [1002, 770]. Semi-Online [2547]. Semigroup [333]. Semigroups [39]. Sense [285]. Sensor [411]. Sensori [2361]. Sensory [561]. Sentences [1181]. Separating [961, 1015, 1433, 2565, 203, 2582]. Separation [951, 492, 220]. Separator [956]. Separators [1696]. Sequence [1810, 2546, 1022, 854, 2054, 1428, 1709, 1349, 97]. Sequences [2429, 589, 1429, 1278, 1027, 1219, 1362, 1357, 1360]. Sequent [1603, 2011]. Sequential [1870, 2178, 329, 1631, 96, 274, 1235]. Sequentialization [691]. Serial [462, 2080]. Series [1782, 960, 771, 1787, 455, 2009, 530, 532, 2461]. Series-Parallel [960]. Server [1840, 2054, 1118, 1119, 1822, 1794, 1349, 99, 1842]. Server-Aided [2054, 1349, 99]. Servers [1847, 1803, 1117]. Service [1151, 1837, 1801]. Services [1893, 2037, 1841]. Servoing [937]. Session [1143, 1147, 1150, 1154, 1158, 1110, 1115, 1120, 1125, 1130, 1134, 1139]. Set [953, 1740, 958, 1162, 618, 1163, 2551, 739, 977, 979, 1815, 1589, 658, 582, 2209, 56, 1399, 249, 275, 224, 1399]. Set-Of-Support [56]. Set-Oriented [618]. Set-valued [1815]. Set/Hyperset [1589]. Sets [912, 2597, 2503, 159, 1215, 37, 2353, 989, 2268, 2274, 1056, 219]. Setting [141, 2411]. Several [2038]. SGML [1680]. Shading [2216, 1432]. Shadowing [1443]. Shafer [1533, 2281]. Shafer-Logan [1533]. Shamir [134, 2027]. Shamir-like [134]. Shannon [283]. Shape [1407, 2216, 566, 1432, 2219, 568, 1399, 1405, 1753]. Shapes [939, 1411, 1296]. Share [1892]. Share-Kit [1892]. Shareable [874]. Shared [1507, 348, 1217, 1856, 1863, 1861, 2294, 990, 1724, 1935, 2583, 2299, 1657, 1258, 1362]. Shared-Nothing [1507]. ShareME [1878]. Sharing [1501, 2076, 86, 2074, 1540, 2103, 2104, 1818, 2028, 2378, 2075, 1338, 267, 1640, 1336, 1337]. Sharply [1186]. Shells [1611]. shift [1359, 97]. Shooting [833]. Shop [1552]. Short [1143, 1147, 1150, 1154, 1158, 1414, 603]. Short-and [1414]. shortcutting [75]. Shorter [278]. Shortest [1675, 930, 2548, 2576, 980, 992, 1935, 2552, 2549]. Shortest-Path [992]. Should [635, 878]. Shuffle [1031]. Sibling [2026]. Side [2427, 2294, 2287]. Side-Effect [2427]. Sided [698, 2198]. Sieve [2086]. Signal [1391, 527, 557, 1986, 1635]. Signal-to-Noise [1391]. Signature [1815, 1581, 2067, 2066, 2058, 2069, 1813, 1342, 1350]. Signatures [2063, 2079, 2065, 118, 2113, 2064, 119, 120]. Signed [2088, 624, 997]. Significance [1709]. Signs [2352]. silence [84]. silent [1235]. SIMD [2435, 2297]. Similarity [1810, 393, 303, 440, 1420]. Simple [950, 1487, 951, 952, 20, 2548, 482, 2576, 748, 2556, 638, 1445, 2563, 1849, 2539, 1710, 1669, 58, 228]. Simpler [1194]. Simplification [1172]. Simplified [1017, 489, 1412, 108]. Simplify [2489]. Simplifying [24]. Simulate [443]. simulated [1768]. Simulating [348, 1488, 863]. Simulation [1780, 1987, 508, 1676, 482, 506, 317, 507, 1786, 1788, 1791, 1931, 525, 1874, 747, 2455, 512, 888]. Simulations [1003, 520]. Simultaneous [2060, 750]. Simultaneously [2567]. Single [2499, 1677, 2033, 2590, 79, 1490, 110, 109, 252]. Single-Chip [1490]. Single-Layer [2590]. Single-Source [1677]. Singular [2440]. Siphons [318]. Sipser [255]. Sisyphus [1614]. Situatedness [1607]. situating [711]. Situational [2348]. Situations [1881]. Size [452, 121, 986, 595]. Skeletal [2402]. Skeleton [1406]. Sketching [1576, 1889]. Skew [1124]. SKIL [704]. Skylab [1481]. Slash [829]. SLD [1723]. SLD-Derivations [1723]. SLDNF [689]. SLDNF-Resolution [689]. Slicing [2324, 2447, 2325, 1869]. SLWV [152]. Small [961, 494, 100, 455, 1393]. Smallest [975]. Smart [511]. SMIsC [2368]. Smooth [2047, 1966, 1468, 942, 234]. Smoothing [1427, 587]. snack [1097]. SNAKE [944]. Snakes [944]. snapshot [1654]. snapshots [1662, 1653]. Soar [1750]. Social [1280]. soft [1069, 15]. Software [1324, 1887, 858, 884, 442, 885, 1314, 563, 866, 1302, 1892, 1313, 1309, 861, 1326, 1323, 2129, 2332, 1905, 886, 875, 1325, 2146, 876, 877, 868, 2331, 1315, 2148, 2359, 894, 619, 862, 895, 1554, 870, 878, 2529, 889, 1303, 2127, 445, 1895, 1316, 2526, 900, 863, 873, 1799, 1903, 440, 2457, 879, 883, 891, 888]. Software-Evaluation [1895]. Solar [1481]. Solution [1010, 485, 784, 1873, 125, 455, 1434, 345, 333, 1602, 213]. Solutions [811, 884, 938, 288]. Solvability [2525]. Solve [48]. Solver [1732]. Solving [1608, 340, 1417, 655, 1202, 2562, 1733, 249, 1614, 915, 2269, 1924, 296, 2519, 828, 1623, 2170]. Some [485, 589, 2453, 1163, 1836, 972, 1076, 2103, 2208, 1991, 2523, 2504, 2550, 1047, 1082, 2532, 1723, 1959, 837, 2173, 1962, 265, 248, 938]. Sometimes [1194]. sonar [1755]. sonar-based [1755]. Sort [2538]. Sorted [674, 157, 301, 1585, 699, 281, 46]. sortedness [1069]. Sorting [2540, 1689, 1024, 366, 1697, 1973, 1871, 1701, 1668, 221]. sorts [3]. Source [1677]. Sourced [1123]. Sources [2010]. SP [2081]. Space [1487, 1721, 508, 564, 2564, 678, 2361, 1280, 1285, 1278, 1214, 1408, 1402, 1287, 941, 530, 1282, 1286, 1877, 391, 1830, 330, 1269, 1279, 750, 495, 1537, 201, 203, 1099, 202]. Space-Time [391]. Spaces [2103, 2200, 2187, 1998, 1749]. SPADE [2154]. Span [2104]. Spanners [2543]. Spanning [2537]. Sparcle [2304]. Sparse [765, 1580, 1056]. Sparsely [473]. Sparseness [2543]. Spatial [2430, 2223, 390, 1292, 2462, 1285, 1270, 1296, 1524, 1525, 1271, 1893, 1282, 1297, 574, 572, 1268, 1779, 1294, 1272, 1275]. Spatio [1277, 2334]. Spatio-Temporal [1277, 2334]. Spatiotemporal [1424]. Speaker [534]. Special [1911, 127, 2365, 1903, 1442, 66]. Specialists [2397]. Specialization [2393]. specializing [1764]. Specific [1730, 883, 120]. Specification [1985, 1887, 1584, 924, 2452, 1301, 1784, 376, 1205, 2345, 2155, 1970, 1604, 1972, 1206, 1319, 706, 305, 1320, 1, 2, 2407, 11, 12, 2410]. Specifications [1307, 680, 441, 1305, 144, 2332, 1966, 1979, 693, 1197, 13, 740, 425, 1952, 31, 2315, 8, 9, 10, 1247, 3]. specify [7]. Specifying [1965, 2481, 1713, 14]. Specimen [2374]. Speckle [1385]. Speckles [1465]. SPECT [585, 588]. spectrum [1235, 1604]. Specular [1433]. Speech [2221, 2214, 2215, 2041]. Speed [1148, 2101, 107]. Speeding [2053, 2088]. Speedup [1066]. Spirality [960]. Spirit [1752]. Spline [587]. Splines [944, 944]. Split [1491]. Split-Merge [1491]. Splitting [464]. Splittings [219]. SPMD [2428]. Spoken [1907]. Sponsor [2287]. Spontaneous [2221]. Sprague [66]. Sprague-Grundy [66]. SQL [2462, 1521]. SQL-like [2462]. Squarers [847]. Squares [2162, 1404]. SRAM [838]. SRAM-Based [838]. ST [318, 1113]. ST-Components. [318]. ST-II [1113]. stabilization [1663]. Stabilizing [2537, 314, 243]. Stable [313, 1747]. Stack [1227, 1214]. STACS’93 [201]. Staffing [876]. Stage [1924, 1418]. stairway [15]. Stallings [773]. Stamp [1831]. Stamps [2536]. Standard [1840, 2069, 347, 2058]. Standardization [2116]. standards [1361]. Star [1035]. Star-Free [1035]. Starting [2378]. State [461, 161, 313, 772, 1018, 2061, 1211, 530, 667, 330, 1063, 1359, 1080]. Statecharts [1320]. Stategies [600]. Statements [2229, 2527]. States [2225, 2203, 2267].

Static [51, 970, 685, 1940, 1651, 1636]. Statically [2142]. Statically-Compiled [2142]. Statistical [1487, 2530, 1491, 413, 561, 394, 1386, 1709, 592]. Statistics [1384]. Status [1844]. Steiner [958, 2597, 2598, 69, 1001]. Step [611]. Steps [1608, 828]. Stereo [1379, 1445, 1447, 1448, 1441, 1435]. Stereograms [1494]. Stereolithography [2521]. Stereoscopic [1440, 1428]. Still [861]. Stimuli [451]. Stimulus [452]. Sting [2306]. Stochastic [482, 1784, 323, 483, 2158, 326, 990, 2176, 594, 568]. Stochastic-Rule [2176]. stochasticity [205]. Stock [1418, 559, 2307, 1756]. Stock-Cutting [1418]. Stop [2065, 118, 2064, 119, 1658]. Storage [509, 1807, 1118, 1579, 1806, 670, 1382, 1116]. Storage-Requirements [1382]. Stored [1127]. Stores [1057, 1520]. Straight [1681]. Strain [1925]. Strategic [655, 2401]. Strategies [137, 348, 1955, 1277, 969, 1739, 1872, 2005, 2206, 1222, 1168, 2004, 904, 481, 1746, 1774, 2412]. Strategy [55, 1502, 575, 53, 1209, 1424, 56]. Stratification [1137]. Stratified [814]. stream [1358]. Strength [2233]. Stress [1925]. Stretch [1164]. strict [1343]. Strictness [1645, 1220, 1954, 1644]. String [340, 1679, 1417, 98, 663]. Strings [996]. Stripe [1439]. Strong [2024, 172, 1672, 1345, 183]. Strongly [766]. Structural [583, 2010, 415, 1038, 1470, 71]. Structure [2213, 1477, 1179, 2208, 976, 127, 1278, 1434, 1816, 2262, 2176, 1007, 1380, 1715, 569, 1297, 1543, 1998, 1412, 1707, 2539, 1537, 219, 81, 1249]. Structured [1266, 426, 931, 160, 2441, 1897, 1879, 1439, 1909, 329, 1319, 241, 1320, 1262]. Structures [783, 1547, 2389, 2006, 2554, 566, 2398, 570, 821, 643, 1227, 2296, 2533, 2149, 1597, 1821, 1832, 2377, 331, 1189, 2557, 2356, 2428, 2421, 1962, 377, 2396, 1048, 1051, 731, 81, 1261, 721, 2415, 728]. Structuring [810, 2344, 2133]. Studenten [1912]. Studies [143, 1886, 1119, 1398]. Studio [1148]. Study [2344, 2370, 360, 1837, 2360, 1400, 2358, 1563, 499, 1512, 2406, 1058]. Studying [186]. Sturmian [1061]. style [81, 2012]. Styles [1914]. Sub [315, 1755, 1754]. Sub-Classes [315]. sub-symbolic [1755, 1754]. Subband [1391]. Subclass [2202]. Subdeductions [701]. Subedge [2576]. Subgraph [737]. Subliminal [1350]. Sublists [2563]. sublogarithmic [201, 240, 203, 202]. Subnets [313]. Subpixel [1386]. Subprograms [1226]. Subsets [344]. Substances [2380]. substitution [1075, 1090, 224, 1104]. Substitutions [19]. Substransactions [1510]. Substrate [1596]. Substrings [996]. Substructural [679]. substructure [184]. Subsumption [2475]. Subsystem [1141, 1181]. Subtheories [149]. Subtype [910]. Subtyping [1583]. Success [284]. Successful [2587]. Suffice [2536]. Sufficiency [170]. Sufficient [25, 1434, 2093, 2032, 1655]. Suffix [98]. suitable [1371, 134]. Sum [1681, 767]. Summary [865, 872, 1560, 2427, 882, 897, 902, 595, 890]. summation [1348]. Sums [786, 1602]. Super [2492, 110, 2032, 109]. Super-Key [2492]. Supercomputer [1490]. Superimposing [1813]. Superimposing-Coded [1813]. Supernormal [142]. Superposition [1172]. Superstrings [980]. Supervised [1394]. Support [1994, 1129, 1153, 1142, 2243, 1145, 1852, 1323, 1499, 1581, 1126, 1598, 525, 1617, 2127, 871, 292, 1549, 356, 1312, 56, 881]. Supported [509, 2457]. Supporting [1845, 1890, 921, 765, 2139, 2148, 1824, 1922, 2428, 446]. Supports [2258]. Suppress [1465]. Surface [1431, 994, 1469, 1710, 833]. Surfaces [1599, 1445, 1383, 942, 1435, 1437]. Surjective [487]. Surjectivity [1017]. Survey [2573, 2311, 1301, 1948, 1824, 1008, 1798, 2342, 4]. Sweep [977, 2596]. Sweeping [1116]. Switch [687]. Switch-Level [687]. Switched [1111, 1801]. Switching [2036, 255]. Switching-Closure-Test [2036]. Sybase [1822]. Syllabic [472]. Symbiosis [285]. Symbolic [1046, 2007, 648, 424, 1996, 1584, 1592, 2001, 1202, 1422, 570, 1587, 2449, 2309, 2284, 1423, 2308, 1296, 2285, 2258, 1999, 292, 2303, 1583, 1755, 1754]. Symbols [1035]. symmetric [252]. Symmetrical [538, 631]. Symmetries [683]. Symposium [1178]. Symptom [2313]. synchronised [1086]. Synchronization [1229, 2470, 1123, 1121, 2287, 1146, 1122, 1655, 1671]. Synchronized [1142, 1935]. Synchronous [839, 1007, 1672, 84]. Syndrome [311]. Syntactic [1183, 1034, 1048, 257]. Syntax [2506, 2400, 12]. Synthesis [140, 2182, 1974, 1641, 144, 2432, 847, 477, 2348, 519, 845, 400, 150, 1040, 214]. Synthesis-Oriented [2348]. System [1507, 2469, 1992, 703, 2336, 1128, 312, 1501, 1575, 618, 2385, 1574, 2050, 2302, 1153, 653, 2438, 2371, 2241, 562, 443, 1503, 2374, 616, 1552, 1555, 2398, 1313, 838, 817, 704, 641, 2452, 1322, 1886, 661, 1863, 1553, 1784, 1587, 1876, 679, 376, 1855, 1475, 2286, 2059, 1818, 306, 1573, 606, 1124, 2594, 2329, 1806, 1554, 1505, 1601, 1581, 1126, 1544, 804, 1922, 1808, 1904, 585, 1155, 1479, 1303, 2306, 296, 1206, 1572, 1520, 2484, 400, 837, 1588, 1543, 1823, 1137, 654, 700, 708, 1522, 1530, 2458, 1157, 625, 1800, 1480, 1550, 2151, 1563, 1531, 2446, 706]. System [940, 2226, 172, 1368, 268, 731, 1774, 269, 1647, 128, 193, 1369]. System-Adaption [1904]. Systematic [900, 650, 1610, 6]. Systems [1845, 411, 2512, 2337, 1985, 1568, 1992, 138, 2479, 2143, 1317, 284, 1987, 20, 2134, 787, 2507, 640, 32, 860, 1216, 1986, 2509, 2453, 1833, 2213, 2237, 1993, 55, 393, 1607, 1539, 426, 2470, 1123, 1495, 2379, 1906, 404, 1121, 1978, 2375, 1506, 427, 2361, 2243, 327, 694, 1867, 54, 1321, 649, 1580, 2503, 428, 1932, 1559, 1995, 2046, 1532, 1713, 1872, 2208, 50, 2471, 1809, 1301, 2254, 1022, 412, 25, 1991, 1905, 1979, 2400, 679, 1982, 47, 779, 53, 1984, 1864, 609, 2478, 903, 1980, 2347, 1511, 285, 630, 26, 636, 1907, 1860, 310, 1977, 2494]. Systems [22, 384, 558, 1504, 1972, 1990, 1983, 2270, 24, 408, 2156, 1796, 1286, 1534, 447, 623, 1146, 2513, 612, 629, 1797, 1718, 2360, 407, 432, 2312, 2466, 2202, 329, 402, 1997, 2457, 701, 481, 1312, 2300, 2357, 607, 439, 777, 40, 652, 2106, 2387, 1195, 1794, 42, 1804, 2133, 1805, 1610, 1512, 1583, 1762, 1753, 1043, 175, 106, 1266, 1672, 1245, 236, 1663, 1632, 249, 1244, 1082, 1654, 1087, 232, 190, 1090, 224, 191, 244, 1254, 1235, 1763, 173, 1297]. Systolic [854, 132, 521, 2159, 1373]. Systolic-Arrays [132]. Szpilrajn [2533].

Table [497, 841]. Tableaux

[1174, 602, 1109]. Tables [1826, 736, 1824]. Tactus [1142]. Talk [1251, 1266, 1230, 1241, 1257, 17, 49, 30, 35, 1246]. Tamil [1286]. Taming [1853]. Tangible [924]. TAO [2128]. tape [1346]. Target [1123]. Targets [1484]. Task [2221, 2289, 2363, 1327, 2594, 1569, 883, 1623, 1319]. Task-Oriented [2221, 1569]. Task-Specific [883]. Tasks [1613, 957, 212]. tastbaren [1910]. taste [1053]. Tatzelwurm [270]. Taxonomic [797, 798, 796, 2249]. taxonomies [729]. taxonomy [1085]. Teams [2189]. Tech [1474]. Technical [1459, 2109]. Technique [1407, 1427, 1682, 318, 2035, 606, 2360, 1412, 1394, 1233]. Techniques [34, 2426, 168, 2353, 636, 634, 2444, 2116, 697, 1797, 863, 592, 1629]. Technologies [836]. Technology [2154, 1460, 2214, 673, 1313, 892, 1816, 2146, 899, 2529, 1923, 1924, 900, 2357, 709, 711, 1361]. Telecommunication [1313, 2037]. Teleconferencing [125]. Telepresence [2232]. Temporal [1438, 1540, 780, 1136, 1427, 1440, 1277, 522, 441, 649, 390, 435, 2460, 2245, 1385, 1979, 1278, 2334, 1508, 781, 2381, 2340, 1192, 1967, 1715, 391, 918, 650, 1248, 1759, 1775, 727, 11, 719, 244, 1096, 1246]. Temporal-fuzziness [1508]. temporally [1776]. Temporary [957]. Tense [1719]. Tensor [2432]. Tensors [476]. Term [2512, 20, 678, 1593, 25, 47, 24, 35, 2513, 2172, 699, 40, 46, 27, 28, 176, 7, 714, 173]. Term-Rewriting [35]. Termal [384]. Terminal [786]. Terminating [45]. Termination [2512, 787, 29, 338, 50, 25, 26, 28]. Terminological [380, 290, 614, 1719, 309, 1759]. Terminology [368]. Terms [157, 494, 17, 2201, 670, 829, 6, 83, 682]. Terrain [933, 1283]. Terrains [932]. Tes [1793]. Test [287, 850, 37, 2445, 1870, 853, 2359, 1466, 2036, 58]. Test-Bed [1466]. Test-Sets [37]. Testability [762]. Testable [1040, 259]. Testbed [1311]. Testing [2318, 1677, 1692, 2359, 1310, 1958, 777, 269, 1377, 1252, 1102]. TestIt [1894]. Tests [1487, 494, 1491, 1079, 1639]. Text [1457, 1908, 1627, 298, 1814, 1715, 625, 1542, 1743]. Text-Based [625]. Text-to-Phoneme [1743]. Textbook [395]. Texts [2366, 1771]. Textural [1412]. Texture [1415, 1413, 550]. Textured [1414]. Textures [535]. th [1415]. Thalamic [456]. Their [618, 308, 972, 744, 634, 612, 837, 656, 2026, 1057, 106, 2, 942, 1031, 1350, 883, 1081]. them [2512, 896]. Theorem [637, 2515, 339, 1591, 600, 1729, 642, 43, 33, 2533, 690, 760, 152, 35, 773, 282, 218, 1262, 1259, 730, 232, 1095, 182]. Theorems [1182, 144, 38]. Theoretic [1188, 2269, 2081, 262, 130]. Theoretical [143, 1284]. Theories [343, 280, 1717, 2249, 638, 345, 1290, 2263, 826, 250]. Theory [1890, 2007, 2575, 1163, 2245, 1947, 1182, 641, 686, 1825, 824, 2334, 744, 165, 2210, 1032, 164, 2179, 2381, 2529, 1169, 1191, 1170, 2274, 829, 362, 2278, 2209, 594, 2157, 2281, 1577, 881, 2404, 66, 1751, 9, 185, 1044, 1098, 1239, 1074, 688]. Theory-Based [1890]. Thinking [1922]. Thrashing [1820]. Three [601, 1071, 1687, 1949, 1444, 1285, 1690, 875, 361, 2153, 2356, 429, 582]. Three-Clustering [1690]. Three-Dimensional [1444, 2356, 582]. Three-Valued [601, 1949]. Threshold [2102, 1332, 87, 2523, 2566, 2029, 253, 2197, 259]. Thue [2202]. thunkification [1646]. Thwart [2109]. Tied [663]. Tight [84]. tilings [1670]. Time [2537, 411, 735, 139, 1666, 112, 508, 1473, 2564, 1427, 1978, 287, 2524, 1145, 2556, 931, 2545, 1301, 2014, 822, 489, 1285, 1024, 1472, 1785, 1905, 1787, 1807, 2161, 1151, 1203, 1408, 987, 614, 849, 477, 1621, 1847, 2536, 1114, 1972, 530, 1796, 1303, 447, 2517, 2578, 391, 532, 2231, 472, 2461, 1112, 1522, 831, 2300, 1319, 1708, 288, 1279, 1144, 1320, 1248, 57, 1258, 67, 1663, 1654, 189, 1673, 1070, 1093, 213, 1778, 1099, 1101, 1102, 1235, 258, 1256]. Time-Bounded [2524]. Time-Critical [1522]. Time-Dependent [1807]. Time-Optimal [1408]. Time-Randomness [987]. Time-Series [2461]. Time-Space [2564, 1099]. Time-Stamps [2536]. Time-to-Crash [2231]. Time-Varying [931, 1621, 1778]. Timed [855, 316, 324, 919, 777, 1320, 1237]. Times [2593, 1062]. Timing [2528, 1931, 1666]. Tired [1684]. Together [1607]. Token [766]. Tokens [2579]. Tolerance [1856, 59, 749]. tolerant [1659]. Tolerating [1664, 947]. Tomograms [1465]. Tomographic [587]. Tomography [580, 584, 579]. TOOA [918]. Tool [508, 2237, 1473, 1729, 649, 320, 1599, 2467, 507, 2293, 1518, 2500, 1838, 2148, 1837, 1624, 1894, 1170, 1612, 1878, 268]. Tool-Box [1170]. Toolkit [1142, 2396, 1161]. Toolkit-Level [1142]. Toolkits [1160]. Tools [2094, 117, 448, 1840, 877, 1548, 863, 402]. Top [1011, 2385, 2011, 1204, 1039, 1849, 1766]. Top-Down [1011, 1204, 1039, 1766]. Topographic [1380]. Topological [1024, 1998, 1276]. Topologies [1990]. Topology [64, 500]. toposes [183]. Tori [1697]. torus [1670]. Toupie [681]. Touring [1129]. Tournaments [959]. Tours [742]. Trace [1788, 2330, 2322, 42, 1066, 248, 244]. Trace-Based [2330]. Trace-Driven [1788]. Traces [761, 2126, 1045, 260]. Track [2536]. Tracking [2230, 1472, 1425, 2497, 943]. Tractable [1318]. Trade [750, 1099]. Trade-offs [750, 1099]. Tradeoff [987]. Tradeoffs [2564]. Traditional [1827]. Traffic [543]. Trained [487]. Training [491, 1900, 484, 1925]. Transaction [1502, 1323, 1510, 1805]. Transactions [1556]. Transducers [1011, 1019]. Transfer [892, 1900, 2042, 899, 900, 111, 1350]. Transferred [121]. Transform [1457, 563, 2041, 1429, 1403, 1389, 1446, 1933, 1452]. Transformation [2429, 313, 1406, 1444, 376, 1468, 2440, 342, 2406, 2412]. Transformational [423]. Transformations [2430, 419, 1254, 1039, 1057]. transformers [1063, 273]. Transforming [2512]. transient [1664]. Transition [2456, 2017, 817, 779, 325, 1843, 1043, 232, 1254]. Transitive [1704]. Transitivity [1023]. Translate [2389]. Translates [2398]. Translating [2252, 185, 1433]. Translation [2013, 695]. Translational [1429]. Translations [1012]. Translator [1743]. Transparency [1476]. Transparent [251, 922]. Transport [1129, 2123]. Transversals [2584]. Transverse [2141]. Trapdoor [2094, 104]. Trapdoors [2068]. Traps [318]. Travel [794]. Traveler [2592]. Trbovich [1364]. Treat [2547]. Treating [2473]. treatment [108]. Tree [2537, 1011, 371, 139, 2597, 1019, 2556, 2598, 468, 570, 763, 981, 2593, 1699, 23, 1035, 1037, 1039, 1470, 1001, 762, 1054, 1059, 1248, 70, 249, 1637, 275, 62, 250]. tree-like [62]. Trees [952, 955, 959, 741, 1716, 739, 635, 1029, 984, 1698, 843, 325, 98, 2519, 1000, 1040, 1711, 69]. Treewidth [743, 1696, 210, 2553]. Trekking [1684]. Trellis [909]. Trends [2108, 2212]. Triangulated [933]. Triangulation [1271, 229, 65]. Triangulations [963, 748]. Triconnected [2574]. Trie [1814]. Triple [958]. Triple-Free [958]. Triply [962]. Triply-Logarithmic [962]. TROLL [1576, 1563]. TRS [45]. True [776, 1189, 7]. Truncation [1390]. trusted [1342]. Truth [1734, 1714, 711]. TS [2296]. TS/ [2296]. TSP [969]. TUPLE [2297]. Tupling [1638]. Turbofan [553]. Turing [807, 1033, 750]. Turnaround [857]. Tutorial [136, 1792, 1796]. Tutoring [2156, 296, 1774]. TV [1148]. Tweety [146]. Two [601, 336, 1164, 2045, 1900, 1969, 976, 759, 2054, 1933, 344, 1286, 2578, 582, 1623, 1253, 1834, 68, 235, 69, 1376, 258, 202]. Two-Dimensional [976, 1933, 582, 68]. Two-Face [2578]. Two-Level [1623]. Two-Local [2045]. Two-Phase [1969]. Two-Way [1164, 759, 235, 202]. typability [180]. Type [916, 1547, 20, 51, 841, 686, 2152, 1601, 2030, 1170, 1588, 829, 700, 199, 2008, 1577, 27, 2505, 175, 1060, 185, 190, 1044, 1098, 173, 2191]. Type-Coercion [2008]. Typed [2495, 912, 2508, 786, 2183, 689, 166, 1193, 208]. Types [2323, 51, 911, 145, 1264, 1284, 826, 17, 708, 181, 188, 1650, 271, 246, 194, 14, 197, 1649]. Typing [2507, 1848, 1957, 2151, 2413, 1265, 208].

U.S. [2284]. U.S./Japan [2284]. UIMS [1897]. Ultra [1454, 887]. Ultra-High [887]. Ultrasonic [1465]. Ultrasonography [597]. Ultrasound [1385]. Ultrix [1145]. Ultrix-4.2 [1145]. Unambiguity [1680]. Unary [1035]. unbalanced [1654].

Uncertain [410, 1621, 941, 2387]. Uncertainties [656]. Uncertainty [2542, 390, 2246, 628, 2386, 2272, 406, 407, 1666, 2244]. Uncovering [1179]. Undecidability [1017, 1169, 207, 180]. Undecidable [758, 206, 199]. undeniable [1342]. Understanding [1908, 424, 946, 1868, 641, 2327, 2358]. Understanding/Debugging [2327]. Undirected [982, 76]. Undo [1891]. unfolding [1249]. Unforgeable [2063]. Unification [343, 2137, 784, 54, 2514, 338, 1019, 1184, 1587, 2291, 1960, 1037, 698, 342, 699, 2186, 660, 1042, 335, 784]. Unified [589, 1584, 2006, 613, 1728, 1094]. Unified-Algebra-Based [1584]. Uniform [112, 815, 1739, 1865, 1387, 1866, 2193, 1402, 1597, 1088]. Unifying [1500, 2248, 381, 2204]. Union [986]. Unions [139]. Unique [1965, 1434]. Uniqueness [2507]. Unit [698, 1062, 1373]. unit-times [1062]. Uniting [2125]. Unity [2387, 1253]. Universal [807, 1019, 789, 1088]. Universality [1033]. Universally [2076, 301]. Universes [826]. UNIX [312, 1058, 2124]. Unprovable [1182]. Unreliable [404]. Unsatisfiable [1181]. Unskolemization [622]. Unstructured [936]. Unsuccessful [2587]. Unsupervised [532]. Untrained [1902]. Update [2556, 795]. Updates [156, 415]. Updating [2492, 2491, 620]. upon [1388, 1895]. Upper [962, 1817, 627, 197]. Upward [1677, 220]. URANO [507]. Urban [543]. Usability [2312]. Usage [306, 1649]. Use [1731, 2237, 566, 2214, 2370, 2191, 536, 875, 988, 1504, 532, 650, 572, 2364, 901, 729, 6, 226, 3]. Used [643, 719]. useful [1636]. User [1891, 1911, 2001, 1886, 1880, 1876, 1160, 1570, 1571, 1899, 307, 310, 2348, 1893, 1808, 444, 1885, 1898, 1918, 2352, 2328, 1882, 1375]. User-Adapted [1882]. User-Driven [444]. Users [1532, 1290, 1623]. Using [1457, 935, 1487, 1397, 1173, 2389, 2499, 2238, 855, 1136, 1417, 622, 1291, 627, 2418, 515, 2585, 382, 2391, 2562, 673, 587, 144, 546, 617, 1969, 2041, 2129, 1784, 421, 577, 1429, 1840, 37, 486, 132, 1325, 1205, 395, 1940, 847, 2206, 2500, 1448, 2088, 2333, 1384, 1573, 1838, 2178, 1673, 1421, 856, 1790, 462, 1971, 480, 496, 1449, 149, 541, 1208, 550, 2377, 558, 444, 1282, 1924, 1206, 2401, 1494, 332, 1424, 1435, 2399, 670, 1800, 2227, 891, 573, 1471, 512, 578, 586, 1516, 592, 2501, 2315, 1537, 1399, 2154, 1338, 1662, 1072, 724]. using [239, 1631, 1079, 249, 1769, 1445, 1747, 2155, 942, 1441, 253, 1850, 1354, 84, 1375, 621, 48]. Utilisation [836]. Utility [883]. Utilizing [2450]. Utterance [2364].

V [1158]. Vacillatory [2169]. Valiant [2541]. Validating [423, 896]. Validation [411, 1627, 421, 422]. Valuation [2258]. Value [451, 2201, 1702, 1742, 1245, 196]. value-passing [1245]. Valued [703, 601, 1949, 603, 687, 624, 1815, 1747, 2197]. Values [1387, 829, 239]. valves [211]. Vapnik [2199]. Variable [1174, 1678, 462, 1104, 2583, 592, 1339]. Variable-Length [1678]. Variables [336, 1210, 1103]. Variant [1586, 1491]. Variants [371, 604]. Variations [2273]. Various [137, 2193, 381]. Varying [931, 1621, 1778]. Vaughan [1050]. VC [2203, 2197]. VC-Dimension [2197]. VC-Dimensions [2203]. VDM [2155]. Vector [1397, 501, 352, 1022, 483, 1873, 1491, 1428, 503, 504, 567]. Vectors [526]. Vehicle [2593, 547, 1556]. verbs [718]. verifiable [1355]. Verification [1985, 442, 681, 838, 1301, 687, 636, 1604, 2526, 852, 2051, 1632, 1246]. verifier [270]. Verifying [1713, 1977, 1098, 1195, 242, 1064]. Versatility [1201, 690]. Version [1721, 2076, 673, 773]. Versioning [665, 2502]. Versions [1496, 2137, 129, 1900, 1935]. Versus [752, 2510, 147, 537, 433, 1344, 192]. Vertex [1015, 1691, 1694]. Vertex-Connectivity [1694]. Vertex-Disjoint [1691]. vertices [63]. Very [100, 1824]. Vessel [1464, 570, 1470]. VHDL [847, 849]. Via [50, 335, 2234, 174, 2073, 584, 96, 688, 403, 1404, 349, 1623, 1710]. Video [1160, 1111, 1119, 1137, 1140]. Vienna [2431]. View [1210, 2448, 1306, 875, 419, 1446, 1449, 302, 2138, 2081, 1830]. Viewed [403]. Viewer [2227]. Viewer-Centered [2227]. Viewing [664, 285, 1122]. Viewpoints [1532, 889]. Views [2137, 2139, 1442, 1071]. VIP [1912]. Virtual [2232, 2415]. Visibility [1283]. Visible [2576]. Vision [2213, 1473, 1379, 1378, 938, 544, 2227, 1474, 940, 2226]. Visual [1625, 452, 508, 1438, 939, 937, 1488, 1425, 2228, 451, 1478, 1490, 1918, 2357]. Visualisation [2355, 1328, 422, 2356, 1572]. Visualization [1599, 2327, 1913, 574, 1936]. Visualizing [2293]. Visually [2365, 2205]. VITAL [1622]. VLF [1486]. VLF/ELF [1486]. VLIW [2419]. VLIW/MIMD [2419]. VLSI [2304, 995, 293, 2333, 1869]. VLSI-Design [293]. Vocabularies [2378]. Voice [2362]. Volker [1045]. Voltage [454]. Voltage-Controlled [454]. Volume [570, 1443, 559, 1467, 582, 1936, 1470]. Volumes [1871]. Voronoi [1685, 1280, 2520]. Voting [2040, 1351]. Voxel [594]. vs [2579, 79]. vulnerability [1357].

Wadler [1053]. wait [1655, 1652]. wait-free [1655, 1652]. Walkthrough [1573]. Wallet [2070]. WAM [1221, 1215, 291]. WAM-Based [1215]. Wang [1670]. Ward [1319]. Ward/Mellor [1319]. warp [1673]. Warren [2327]. Watchman [2548, 2549]. Wave [1486]. Wavelet [563, 586, 234]. Waves [1486]. Way [753, 2023, 1164, 861, 759, 122, 1036, 2024, 1365, 235, 1366, 1339, 202]. Ways [1687]. Weak [1022, 2263, 1672, 238]. Weakened [1185]. weakest [277]. Weakly [2576]. Weakness [1186]. Weaknesses [875, 2035]. Weight [991]. Weighted [1675, 2595, 1014, 587, 227, 2162, 991, 2539, 1072]. Weighting [2255]. Weightless [558]. Weights [500]. Well [50, 874, 1726, 2516, 1237]. well-caused [1237]. Well-Defined [874]. Well-Founded [1726, 2516]. Well-Moded [50]. Weyl [1596]. Wheels [1900]. Where [2393]. Which [120, 738]. While [1735, 2366]. Wide [1302]. widening [1650]. Widenings [1955]. Width [961]. Will [448]. Window [1921, 2088]. Windows [2396]. Wire [2589, 1346]. wire-tape [1346]. Within [2121, 186, 1442]. Without [1620, 1684, 1342, 1035]. Wizard [2360]. Wolfe [2445]. Word [336, 2363, 342, 534, 1091]. Words [1016, 223, 225, 261, 259]. Work [1890, 2585, 1823]. Workbench [1622, 2392]. Working [1896, 1892]. Workshop [2310, 2284]. Workspace [931]. Workstation [1141, 851, 1140]. Workstations [1873]. World [2377, 431, 1779]. Worlds [2252]. Wormhole [360, 2571]. Worst [2556, 213]. Worst-Case [2556]. write [238, 1656]. Writing [2567]. WYSIWYG [2349].

X [1463, 1481, 2396]. X-Ray [1463, 1481]. X-Windows [2396]. XII. [1298].

Yacobi [1356]. Yes [299]. Yes/No [299]. Yields [2544]. YPREL [537].

Zero [117, 115, 2047, 2078, 2046, 2079, 2077, 2048, 2080, 116]. Zero-Knowledge [115, 2047, 2078, 2046, 2079, 2077, 2048, 2080, 116]. Zheng [1376]. Zheng-Matsumoto-Imai [1376]. ZKIP [2044, 2049, 2043]. ZM4 [1858].

References

Astesiano:1993:ASC

[1] E. Astesiano and G. Reggio. Algebraic specification of concurrency. Lecture Notes in Computer Science, 655:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrich:1993:OTS

[2] H.-D. Ehrich, M. Gogolla, and A. Sernadas. Objects and their specification. Lecture Notes in Computer Science, 655:40–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mosses:1993:USA

[3] P. D. Mosses. The use of sorts in algebraic specifications. Lecture Notes in Computer Science, 655:66–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orejas:1993:IBE

[4] F. Orejas, M. Navarro, and A. Sanchez. Implementation and behavioural equivalence: a survey. Lecture Notes in Computer Science, 655:93–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astesiano:1993:RBL

[5] E. Astesiano and M. Cerioli. Relationships between logical frameworks. Lecture Notes in Computer Science, 655: 126–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernot:1993:LAS

[6] G. Bernot and P. Le Gall. Label algebras: a systematic use of terms. Lecture Notes in Computer Science, 655:144–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettaz:1993:HSN

[7] M. Bettaz and M. Maouche. How to specify non determinism and true concurrency with algebraic term nets. Lecture Notes in Computer Science, 655: 164–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breu:1993:BIA

[8] M. Breu. Bounded implementation of algebraic specifications. Lecture Notes in Computer Science, 655:181– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1993:NCA

[9] H. Ehrig, M. Baldamus, and F. Orejas. New concepts of amalgamation and extension for a general theory of specifications. Lecture Notes in Computer Science, 655:199–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1993:NCE

[10] H. Ehrig and F. Parisi-Presicce. Nonequivalence of categories for equational algebraic specifications. Lecture Notes in Computer Science, 655:222– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiadeiro:1993:PST

[11] J. F. Fiadeiro, J. L. Costa, A. Sernadas, and T. S. E. Maibaum. Process semantics of temporal logic specification. Lecture Notes in Computer Science, 655:236–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabriel:1993:OBS

[12] P. Gabriel. The object-based specification language prod: concepts, syntax, and semantics. Lecture Notes in Computer Science, 655:254–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knapik:1993:SOF

[13] T. Knapik. Specifications with observable formulae and observational satisfaction relation. Lecture Notes in Computer Science, 655:271–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reggio:1993:ELS

[14] G. Reggio. Event logic for specifying abstract dynamic data types. Lecture Notes in Computer Science, 655:292– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salibra:1993:SSI

[15] A. Salibra and G. Scollo. A soft stairway to institutions. Lecture Notes in Computer Science, 655:310–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1993:GCO

[16] E. G. Wagner. Generic classes in an object-based language. Lecture Notes in Computer Science, 655:330– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinke:1993:ITA

[17] Karl Meinke. Invited talk: Algebraic semantics of rewriting terms and types. Lecture Notes in Computer Science, 656:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahrs:1993:CR

[18] Stefan Kahrs. Context rewriting. Lecture Notes in Computer Science, 656: 21–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rose:1993:ECS

[19] Kristoffer Hogsbro Rose. Explicit cyclic substitutions. Lecture Notes in Computer Science, 656:36–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banach:1993:STI

[20] Richard Banach. Simple type inference for term graph rewriting systems. Lecture Notes in Computer Science, 656:51–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antimirov:1993:CSE

[21] Valentin M. Antimirov and Anatoly I. Degtyarev. Consistency and semantics of equational definitions over predefined algebras. Lecture Notes in Computer Science, 656:67–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Middeldorp:1993:CCC

[22] Aart Middeldorp. Completeness of combinations of conditional constructor systems. Lecture Notes in Computer Science, 656:82–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plump:1993:CTR

[23] Detlef Plump. Collapsed tree rewriting: Completeness, confluence, and modularity. Lecture Notes in Computer Science, 656:97–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohlebusch:1993:CSC

[24] Enno Ohlebusch. Combinations of simplifying conditional term rewriting systems. Lecture Notes in Computer Science, 656:113–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gramlich:1993:SCM

[25] Bernhard Gramlich. Sufficient conditions for modular termination of conditional term rewriting systems. Lecture Notes in Computer Science, 656: 128–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loria-Saenz:1993:TCR

[26] Carlos Loria-Saenz and Joachim Steinbach. Termination of combined (rewrite and lambda-calculus) systems. Lecture Notes in Computer Science, 656:143–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zantema:1993:TRT

[27] Hans Zantema. Type removal in term rewriting. Lecture Notes in Computer Science, 656:148–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zantema:1993:TTR

[28] Hans Zantema. Termination of term rewriting by interpretation. Lecture Notes in Computer Science, 656:155– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1993:POT

[29] Nachum Dershowitz and Subrata Mitra. Path orderings for termination of associative-commutative rewriting. Lecture Notes in Computer Science, 656:168–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padawitz:1993:ITG

[30] Peter Padawitz. Invited talk: Generic induction proofs. Lecture Notes in Computer Science, 656:175–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wirth:1993:CBA

[31] Claus-Peter Wirth and Bernhard Gramlich. A constructor-based approach for positive/negative-conditional equational specifications. Lecture Notes in Computer Science, 656:198– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1993:SPN

[32] Klaus Becker. Semantics for positive/negative conditional rewrite systems. Lecture Notes in Computer Science, 656:213–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganzinger:1993:ITP

[33] Harald Ganzinger and Jurgen Stuber. Inductive theorem proving by consistency for first-order clauses. Lecture Notes in Computer Science, 656:226– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bronsard:1993:RTF

[34] Francois Bronsard and Uday S. Reddy. Reduction techniques for first-order reasoning. Lecture Notes in Computer Science, 656:242–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaisted:1993:ITC

[35] David Plaisted, Geoffrey Alexander, Chu Heng, and Lee Shie-Jue. Invited talk: Conditional term-rewriting and first-order theorem proving. Lecture Notes in Computer Science, 656:257– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kucherov:1993:DRR

[36] Gregory Kucherov and Mohamed Tajine. Decidability of regularity and related properties of ground normal form languages. Lecture Notes in Computer Science, 656:272–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofbauer:1993:CLU

[37] Dieter Hofbauer and Maria Huber. Computing linearizations using testsets. Lecture Notes in Computer Science, 656:287–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1993:PGI

[38] Hantao Zhang. Proving group isomorphism theorems. Lecture Notes in Computer Science, 656:302–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1993:SSX

[39] Nachum Dershowitz. Semigroups satisfying x^{m+n} = x^{n}. Lecture Notes in Computer Science, 656:307–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vorobyov:1993:COC

[40] Sergei Vorobyov. Could orders be captured by term rewriting systems? Lecture Notes in Computer Science, 656: 315–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stokkermans:1993:CFC

[41] Karel Stokkermans. A categorical formulation for critical-pair/completion procedures. Lecture Notes in Computer Science, 656:328–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yabo:1993:TRS

[42] Wang Yabo and David Lorge Parnas. Trace rewriting systems. Lecture Notes in Computer Science, 656:343– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraus:1993:CCI

[43] Ulrich Fraus. A calculus for conditional inductive theorem proving. Lecture Notes in Computer Science, 656: 357–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hantao:1993:ICR

[44] Zhang Hantao. Implementing contextual rewriting. Lecture Notes in Computer Science, 656:363–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamada:1993:CTM

[45] Junnosuke Yamada. Confluence of terminating membership conditional TRS. Lecture Notes in Computer Science, 656:378–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

With:1993:CCO

[46] Lars With. Completeness and confluence of order-sorted term rewriting. Lecture Notes in Computer Science, 656:393–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoot:1993:CCT

[47] Charles Hoot. Completion for constrained term rewriting systems. Lecture Notes in Computer Science, 656: 408–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takano:1993:GPC

[48] Akihiko Takano. Generalized partial computation using disunification to solve constraints. Lecture Notes in Computer Science, 656:424–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacholski:1993:ITD

[49] Leszek Pacholski. Invited talk: Decidability of finiteness properties. Lecture Notes in Computer Science, 656:429– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganzinger:1993:TPW

[50] Harald Ganzinger and Uwe Waldmann. Termination proofs of well-moded logic programs via conditional rewrite systems. Lecture Notes in Computer Science, 656:430–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonnier:1993:LPP

[51] Staffan Bonnier and Jonas Wallgren. Logic programs with polymorphic types: A condition for static type checking. Lecture Notes in Computer Science, 656:438–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoy:1993:NLI

[52] Sergio Antoy. Normalization by leftmost innermost rewriting. Lecture Notes in Computer Science, 656:448– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inverardi:1993:SDD

[53] Paola Inverardi and Monica Nesi. A strategy to deal with divergent rewrite systems. Lecture Notes in Computer Science, 656:458–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delsart:1993:NAG

[54] Bertrand Delsart. A new approach to general E-unification based on conditional rewriting systems. Lecture Notes in Computer Science, 656:468– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bockmayr:1993:ONS

[55] Alexander Bockmayr, Stefan Krischer, and Andreas Werner. An optimal narrowing strategy for general canonical systems. Lecture Notes in Computer Science, 656:483–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wenchang:1993:SSS

[56] Fang Wenchang and Kao Jung-Hong. Set-of-support strategy for higherorder logic. Lecture Notes in Computer Science, 656:498–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandstadt:1993:ITB

[57] A. Brandstadt. On improved time bounds for permutation graph problems. Lecture Notes in Computer Science, 657:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsu:1993:STI

[58] W.-L. Hsu. A simple test for interval graphs. Lecture Notes in Computer Science, 657:11–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felsner:1993:TGO

[59] S. Felsner. Tolerance graphs and orders. Lecture Notes in Computer Science, 657:17–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1993:SPR

[60] K. Jansen. On scheduling problems restricted to interval orders. Lecture Notes in Computer Science, 657:27–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1993:SIJ

[61] H. L. Bodlaender, K. Jansen, and G. J. Woeginger. Scheduling with incompatible jobs. Lecture Notes in Computer Science, 657:37–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1993:GCT

[62] K. Jansen and P. Scheffler. Generalized coloring for tree-like graphs. Lecture Notes in Computer Science, 657:50–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahlhaus:1993:OPA

[63] E. Dahlhaus. Optimal (parallel) algorithms for the all-to-all vertices distance problem for certain graph classes. Lecture Notes in Computer Science, 657:60–??, 1993. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1993:TPN

[64] J. Hromkovic. Topology of parallel networks and computational complexity. Lecture Notes in Computer Science, 657:70–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preilowski:1993:PTN

[65] W. Preilowski. Parallel triangulation of nonconvex polytopes. Lecture Notes in Computer Science, 657:78–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1993:KSC

[66] H. L. Bodlaender. Kayles on special classes of graphs — an application of Sprague-Grundy theory. Lecture Notes in Computer Science, 657:90–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1993:LTA

[67] J. Chen. A linear time algorithm for isomorphism of graphs of bounded average genus. Lecture Notes in Computer Science, 657:103–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhatia:1993:IAR

[68] D. Bhatia, T. Leighton, F. Makedon, and C. H. Norton. Improved algorithms for routing on two-dimensional grids. Lecture Notes in Computer Science, 657:114–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ihler:1993:MRS

[69] E. Ihler. Minimum rectilinear Steiner trees for intervals on two parallel lines. Lecture Notes in Computer Science, 657:123–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerstel:1993:NCT

[70] O. Gerstel and S. Zaks. A new characterization of tree medians with applications to distributed algorithms. Lecture Notes in Computer Science, 657: 135–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinitz:1993:ECS

[71] E. Dinitz. The 3-edge-components and a structural description of all 3-edge-cuts in a graph. Lecture Notes in Computer Science, 657:145–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1993:ACG

[72] J. Chen and A. Kanevsky. On assembly of 4-connected graphs. Lecture Notes in Computer Science, 657:158–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jamison:1993:HDG

[73] B. Jamison and S. Olariu. On the homogeneous decomposition of graphs. Lecture Notes in Computer Science, 657:170–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1993:ERC

[74] S. K. Das and A. Mao. Embeddings in recursive combinatorial networks. Lecture Notes in Computer Science, 657:184–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thorup:1993:SD

[75] M. Thorup. On shortcutting digraphs. Lecture Notes in Computer Science, 657:205–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cournier:1993:EAR

[76] A. Cournier and M. Habib. An efficient algorithm to recognize prime undirected graphs. Lecture Notes in Computer Science, 657:212–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felsner:1993:CPO

[77] S. Felsner and D. Wagner. On the complexity of partial order properties. Lecture Notes in Computer Science, 657:225–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mosbah:1993:PGG

[78] M. Mosbah. Probabilistic graph grammars. Lecture Notes in Computer Science, 657:236–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parisi-Presicce:1993:SVD

[79] F. Parisi-Presicce. Single vs. double pushout derivations of graphs. Lecture Notes in Computer Science, 657:248– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kant:1993:HGD

[80] G. Kant. Hexagonal grid drawings. Lecture Notes in Computer Science, 657:263–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erwig:1993:GAI

[81] M. Erwig. Graph algorithms = iteration + data structures? the structure of graph algorithms and a corresponding style of programming. Lecture Notes in Computer Science, 657:277–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alimonti:1993:PNH

[82] P. Alimonti and E. Feuerstein. Petri nets, hypergraphs and conflicts. Lecture Notes in Computer Science, 657: 293–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gergov:1993:AMB

[83] J. Gergov and C. Meinel. Analysis and manipulation of Boolean functions in terms of decision graphs. Lecture Notes in Computer Science, 657:310– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OReilly:1993:EST

[84] U.-M. O’Reilly and N. Santoro. The expressiveness of silence: Tight bounds for synchronous communication of information using bits and silence. Lecture Notes in Computer Science, 657: 321–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Litovsky:1993:PLL

[85] I. Litovsky, Y. Metivier, and W. Zielonka. The power and the limitations of local computations on graphs. Lecture Notes in Computer Science, 657: 333–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blundo:1993:GDS

[86] Carlo Blundo, Alfredo De Santis, Douglas R. Stinson, and Ugo Vaccaro. Graph decompositions and secret sharing schemes. Lecture Notes in Computer Science, 658:1– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580001.
pdf.

Frankel:1993:CIH

[87] Yair Frankel and Yvo Desmedt. Classification of ideal homomorphic threshold schemes over finite Abelian groups (extended abstract). Lecture Notes in Computer Science, 658:25– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580025.
pdf.

Baritaud:1993:FFH

[88] Thierry Baritaud, Henri Gilbert, and Marc Girault. F.F.T. hashing is not collision-free. Lecture Notes in Computer Science, 658:35– 44, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580035.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580035.
                                        pdf.

Schnorr:1993:FHI

[89] Claus P. Schnorr. FFT-hash II, efficient cryptographic hashing. Lecture Notes in Computer Science, 658:45– 54, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580045.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580045.
                                        pdf.

Lai:1993:HFB

[90] Xuejia Lai and James L. Massey. Hash functions based on block ciphers. Lecture Notes in Computer Science, 658: 55–70, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580055.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580055.
                                        pdf.

Berson:1993:DCA

[91] Thomas A. Berson. Differential cryptanalysis mod 2^{32} with applications to MD5. Lecture Notes in Computer Science, 658:71–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580071.
pdf.

Matsui:1993:LCM

[92] M. Matsui. Linear cryptanalysis method for DES cipher. Lecture Notes in Computer Science, 765:386–397, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsui:1993:NMK

[93] Mitsuru Matsui and Atsuhiro Yamagishi. A new method for known plaintext attack of FEAL cipher. Lecture Notes in Computer Science, 658:81– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580081.
pdf.

Nyberg:1993:CHN

[94] Kaisa Nyberg. On the construction of highly nonlinear permutations. Lecture Notes in Computer Science, 658: 92–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580092.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/0658/06580092.
pdf.

Wernsdorf:1993:ORF

[95] Ralph Wernsdorf. The one-round functions of the DES generate the alternating group. Lecture Notes in Computer Science, 658:99–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580099.
pdf.

Golic:1993:CLS

[96] Jovan Dj. Golic. Correlation via linear sequential circuit approximation of combiners with memory. Lecture Notes in Computer Science, 658:113– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580113.
pdf.

Mihaljevic:1993:CBI

[97] Miodrag J. Mihaljevic and Jovan Dj. Golic. Convergence of a Bayesian iterative error-correction procedure on a noisy shift register sequence. Lecture Notes in Computer Science, 658:124– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580124.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/0658/06580124.
pdf.

OConnor:1993:STS

[98] Luke O’Connor and Tim Snider. Suffix trees and string complexity. Lecture Notes in Computer Science, 658:138– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580138.
pdf.

Pfitzmann:1993:APS

[99] Birgit Pfitzmann and Michael Waidner. Attacks on protocols for serveraided RSA computation. Lecture Notes in Computer Science, 658:153– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580153.
pdf.

Harper:1993:PKC

[100] Greg Harper, Alfred Menezes, and Scott A. Vanstone. Public-key cryptosystems with very small key lengths. Lecture Notes in Computer Science, 658:163–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0658/06580163.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0658/06580163.
pdf.

Sauerbrey:1993:RRA

[101] Jörg Sauerbrey and Andreas Dietel. Resource requirements for the application of addition chains in modulo exponentiation. Lecture Notes in Computer Science, 658:174–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580174.
                                        pdf.

Dixon:1993:MPE

[102] Brandon Dixon and Arjen K. Lenstra. Massively parallel elliptic curve factoring. Lecture Notes in Computer Science, 658:183–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0658/06580183.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0658/06580183.pdf.

Anonymous:1993:PR

[103] Anonymous. Panel report. Lecture Notes in Computer Science, 658:194– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desmedt:1993:TPM

[104] Yvo Desmedt, Peter Landrock, Arjen K. Lenstra, Kevin S. McCurley, Andrew M. Odlyzko, Rainer A. Rueppel, and Miles E. Smid. Trapdoor primes and moduli (panel report). Lecture Notes in Computer Science, 658:194– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580194.
pdf.

Brickell:1993:FEP

[105] Ernest F. Brickell, Daniel M. Gordon, Kevin S. McCurley, and David B. Wilson. Fast exponentiation with precomputation (extended abstract). Lecture Notes in Computer Science, 658:200– 207, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580200.
pdf.

Beller:1993:BDH

[106] Michael J. Beller and Yacov Yacobi. Batch Diffie–Hellman key agreement systems and their application to portable communications. Lecture Notes in Computer Science, 658:208– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580208.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580208.
pdf.

Iwamura:1993:HSI

[107] Keiichi Iwamura, Tsutomu Matsumoto, and Hideki Imai. High-speed implementation methods for RSA scheme. Lecture Notes in Computer Science, 658:221–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0658/06580221.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0658/06580221.pdf.

Maurer:1993:SGT

[108] Ueli M. Maurer. A simplified and generalized treatment of Luby-Rackoff pseudorandom permutation generators. Lecture Notes in Computer Science, 658:239–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0658/06580239.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0658/06580239.pdf.

Patarin:1993:HCP

[109] Jacques Patarin. How to construct pseudorandom and super pseudorandom permutations from one single pseudorandom function. Lecture Notes in Computer Science, 658:256– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580256.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580256.
                                        pdf.

Sadeghiyan:1993:CSP

[110] Babak Sadeghiyan and Józef P. Pieprzyk. A construction of super pseudorandom permutations from a single pseudorandom function. Lecture Notes in Computer Science, 658:267– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580267.
pdf.

Beaver:1993:HBS

[111] Donald Beaver. How to break a “secure” oblivious transfer protocol. Lecture Notes in Computer Science, 658: 285–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580285.
pdf.

Barbaroux:1993:URP

[112] Paul Barbaroux. Uniform results in polynomial-time security. Lecture Notes in Computer Science, 658:297– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580297.
pdf.

Beaver:1993:CPP

[113] Donald Beaver and Stuart Haber. Cryptographic protocols provably secure against dynamic adversaries. Lecture Notes in Computer Science, 658: 307–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580307.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580307.
                                        pdf.

Ohta:1993:SBC

[114] Kazuo Ohta, Tatsuaki Okamoto, and Atsushi Fujioka. Secure bit commitment function against divertibility. Lecture Notes in Computer Science, 658: 324–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580324.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580324.
                                        pdf.

Damgaard:1993:NIC

[115] Ivan Damgård. Non-interactive circuit based proofs and non-interactive perfect zero-knowledge with preprocessing. Lecture Notes in Computer Science, 658:341–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0658/06580341.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0658/06580341.pdf.

Damgard:1993:NIC

[116] I. Damgard. Non-interactive circuit based proofs and non-interactive perfect zero-knowledge with preprocessing. Lecture Notes in Computer Science, 658:341–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biehl:1993:TPZ

[117] Ingrid Biehl, Johannes A. Buchmann, Bernd Meyer, Christian Thiel, and Christoph Thiel. Tools for proving zero knowledge. Lecture Notes in Computer Science, 658:356–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580356.
pdf.

Heyst:1993:HME

[118] Eugène van Heyst and Torben Pryds Pedersen. How to make efficient fail-stop signatures. Lecture Notes in Computer Science, 658:366–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580366.
pdf.

vanHeyst:1993:HME

[119] Eugène van Heyst and Torben Pryds Pedersen. How to make efficient fail-stop signatures. Lecture Notes in Computer Science, 658:366–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580366.
pdf.

Evertse:1993:WNR

[120] Jan-Hendrik Evertse and Eugène van Heyst. Which new RSA signatures can be computed from RSA signatures, obtained in a specific interactive protocol? Lecture Notes in Computer Science, 658:378–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580378.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580378.
                                        pdf.

Chaum:1993:TCG

[121] David Chaum and Torben Pryds Pedersen. Transferred cash grows in size. Lecture Notes in Computer Science, 658:390–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0658/06580390.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580390.
                                        pdf.

Niederreiter:1993:LRC

[122] Harald Niederreiter and Claus P. Schnorr. Local randomness in candidate one-way functions. Lecture Notes in Computer Science, 658:408– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580408.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580408.
                                        pdf.

Okamoto:1993:HID

[123] Tatsuaki Okamoto, Kouichi Sakurai, and Hiroki Shizuya. How intractable is the discrete logarithm for a general finite group. Lecture Notes in Computer Science, 658:420– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580420.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580420.
pdf.

Maurer:1993:FO

[124] Ueli M. Maurer. Factoring with an oracle. Lecture Notes in Computer Science, 658:429–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0658/06580429.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0658/06580429.pdf.

Heiman:1993:SAT

[125] Rafi Heiman. Secure audio teleconferencing: A practical solution. Lecture Notes in Computer Science, 658:437– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580437.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580437.
pdf.

Koyama:1993:SCK

[126] Kenji Koyama. Secure conference key distribution schemes for conspiracy attack. Lecture Notes in Computer Science, 658:449–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580449.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580449.
pdf.

Heiman:1993:NDL

[127] Rafi Heiman. A note on discrete logarithms with special structure. Lecture Notes in Computer Science, 658:454– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580454.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580454.
pdf.

Maurer:1993:RNI

[128] Ueli M. Maurer and Yacov Yacobi. A remark on a non-interactive publickey distribution system. Lecture Notes in Computer Science, 658:458– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580458.
pdf.

Chen:1993:SBP

[129] Lidong Chen and Ivan Damgård. Security bounds for parallel versions of identification protocols (extended abstract). Lecture Notes in Computer Science, 658:461–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580461.
pdf.

Sgarro:1993:ITB

[130] Andrea Sgarro. Information-theoretic bounds for authentication frauds. Lecture Notes in Computer Science, 658: 467–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580467.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580467.
pdf.

Golic:1993:GCA

[131] Jovan Dj. Golic and Slobodan V. Petrovic. A generalized correlation attack with a probabilistic constrained edit distance. Lecture Notes in Computer Science, 658:472–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0658/06580472.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0658/06580472.
pdf.

Iwamura:1993:SAM

[132] Keiichi Iwamura, Tsutomu Matsumoto, and Hideki Imai. Systolicarrays for modular exponentiation using Montgomery method (extended abstract). Lecture Notes in Computer Science, 658:477–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580477.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580477.
                                        pdf.

Agnew:1993:DFE

[133] Gordon B. Agnew, R. C. Mullin, and Scott A. Vanstone. On the development of a fast elliptic curve cryptosystem. Lecture Notes in Computer Science, 658:482–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580482.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580482.
                                        pdf.

Naccache:1993:MSF

[134] David Naccache. A Montgomerysuitable Fiat–Shamir-like authentication scheme. Lecture Notes in Computer Science, 658:488–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0658/06580488.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0658/06580488.
                                        pdf.

Anonymous:1993:AIa

[135] Anonymous. Author index. Lecture Notes in Computer Science, 658:493– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/0658/0658ind.pdf.

Brewka:1993:TNR

[136] Gerd Brewka, Jurgen Dix, and Kurt Konolige. A tutorial on nonmonotonic reasoning. Lecture Notes in Computer Science, 659:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abe:1993:SCV

[137] Naoki Abe. On the sample complexity of various learning strategies in the probabilistic PAC learning paradigms. Lecture Notes in Computer Science, 659:89–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arikawa:1993:MAL

[138] Setsuo Arikawa, Takeshi Shinohara, Satoru Miyano, and Ayumi Shinohara. More about learning elementary formal systems. Lecture Notes in Computer Science, 659:107–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arimura:1993:PTA

[139] Hiroki Arimura, Takeshi Shinohara, and Setsuko Otsuki. A polynomial time algorithm for finding finite unions of tree pattern languages. Lecture Notes in Computer Science, 659:118– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barzdins:1993:TEIa

[140] Janis Barzdins and Guntis Barzdins. Towards efficient inductive synthesis: Rapid construction of local regularities. Lecture Notes in Computer Science, 659:132–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besnard:1993:DGD

[141] Ph. Besnard and E. Gregoire. Deductive generalization in a default logic setting. Lecture Notes in Computer Science, 659:141–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brass:1993:DSD

[142] Stefan Brass. Deduction with supernormal defaults. Lecture Notes in Computer Science, 659:153–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daley:1993:MAL

[143] Robert Daley. Multi-agent learning: Theoretical and empirical studies. Lecture Notes in Computer Science, 659: 175–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franova:1993:PSF

[144] Marta Franova and Yves Kodratoff. Predicate synthesis from formal specifications: Using mathematical induction for finding the preconditions of theorems. Lecture Notes in Computer Science, 659:184–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1993:DTH

[145] Rusins Freivalds, Efim B. Kinber, and Rolf Wiehagen. Dual types of hypotheses in inductive inference. Lecture Notes in Computer Science, 659: 209–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakemeyer:1993:AKA

[146] Gerhard Lakemeyer. All I know about Tweety. Lecture Notes in Computer Science, 659:241–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1993:MVN

[147] Steffer Lange and Thomas Zeugmann. Monotonic versus nonmonotonic language learning. Lecture Notes in Computer Science, 659:254–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marek:1993:NFR

[148] V. Wiktor Marek and Miroslaw Truszczynski. Normal form results for default logic. Lecture Notes in Computer Science, 659:270–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehl:1993:RCB

[149] Michael Mehl. Retrieval in casebased reasoning using preferred subtheories. Lecture Notes in Computer Science, 659:284–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takada:1993:ISP

[150] Yuji Takada. Interactive synthesis of process flowcharts. Lecture Notes in Computer Science, 659:298–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viksna:1993:PIA

[151] Juris Viksna. Probabilistic inference of approximations. Lecture Notes in Computer Science, 659:323–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1993:STP

[152] L. M. Pereira, L. Caires, and J. Alferes. SLWV — A theorem prover for logic programming. Lecture Notes in Computer Science, 660:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Satoh:1993:CGD

[153] K. Satoh and N. Iwayama. A correct goal-directed proof procedure for a general logic program with integrity constraints. Lecture Notes in Computer Science, 660:24–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dung:1993:DSH

[154] P. M. Dung. Declarative semantics of hypothetical logic programming with negation as failure. Lecture Notes in Computer Science, 660:45–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramirez:1993:CNC

[155] M. J. Ramirez and M. Falaschi. Conditional narrowing with constructive negation. Lecture Notes in Computer Science, 660:59–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertino:1993:CAD

[156] E. Bertino, M. Martelli, and D. Montesi. CLP (AD) as a deductive database language with updates. Lecture Notes in Computer Science, 660:80–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ait-Kaci:1993:LPF

[157] H. Ait-Kaci and A. Podelski. Logic programming with functions over ordersorted feature terms. Lecture Notes in Computer Science, 660:100–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boley:1993:DSC

[158] H. Boley. A direct semantic characterization of RELFUN. Lecture Notes in Computer Science, 660:120–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dovier:1993:EFS

[159] A. Dovier, E. G. Omodeo, E. Pontelli, and G. Rossi. Embedding finite sets in a logic programming language. Lecture Notes in Computer Science, 660:150–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordano:1993:MFS

[160] L. Giordano and A. Martelli. A modal framework for structured logic programs. Lecture Notes in Computer Science, 660:168–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1993:MSO

[161] A. Brogi and F. Turini. Metalogic for state oriented programming. Lecture Notes in Computer Science, 660:187– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bugliesi:1993:SIL

[162] M. Bugliesi. On the semantics of inheritance in logic programming: Compositionality and full abstraction. Lecture Notes in Computer Science, 660: 205–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Porto:1993:ASC

[163] A. Porto and P. Rosado. The AbstrAct scheme for concurrent programming. Lecture Notes in Computer Science, 660:216–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1993:PCT

[164] D. Miller. The pi-calculus as a theory in linear logic: Preliminary results. Lecture Notes in Computer Science, 660:242–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keronen:1993:NDP

[165] S. Keronen. Natural deduction proof theory for logic programming. Lecture Notes in Computer Science, 660:265– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reddy:1993:TFD

[166] U. S. Reddy. A typed foundation for directional logic programming. Lecture Notes in Computer Science, 660:282– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meier:1993:APE

[167] M. Meier and J. Schimpf. An architecture for Prolog extensions. Lecture Notes in Computer Science, 660:319– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denti:1993:TIC

[168] E. Denti, E. Lamma, P. Mello, A. Natali, and A. Omicini. Techniques for implementing contexts in logic programming. Lecture Notes in Computer Science, 660:339–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwon:1993:INM

[169] K. Kwon, G. Nadathur, and D. S. Wilson. Implementing a notion of modules in the logic programming language lambda Prolog. Lecture Notes in Computer Science, 660:359–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aronsson:1993:IIG

[170] M. Aronsson. Implementational issues in GCLA: A-sufficiency and the Definiens operation. Lecture Notes in Computer Science, 660:394–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akama:1993:MRC

[171] Y. Akama. On Mints’ reduction for ccc-calculus. Lecture Notes in Computer Science, 664:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altenkirch:1993:FSN

[172] Th. Altenkirch. A formalization of the strong normalization proof for system F in LEGO. Lecture Notes in Computer Science, 664:13–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanBakel:1993:PIT

[173] S. van Bakel. Partial intersection type assignment in applicative term rewriting systems. Lecture Notes in Computer Science, 664:29–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbanera:1993:ECC

[174] F. Barbanera and S. Berardi. Extracting constructive content from classical logic via control-like reductions. Lecture Notes in Computer Science, 664: 45–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbanera:1993:CFH

[175] F. Barbanera and M. Fernandez. Combining first and higher order rewrite systems with type assignment systems. Lecture Notes in Computer Science, 664:60–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benton:1993:TCI

[176] N. Benton, G. Bierman, V. de Paiva, and M. Hyland. A term calculus for intuitionistic linear logic. Lecture Notes in Computer Science, 664:75–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1993:PEN

[177] U. Berger. Program extraction from normalization proofs. Lecture Notes in Computer Science, 664:91–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castagna:1993:SLE

[178] G. Castagna, G. Ghelli, and G. Longo. A semantics for lambda &-early: a calculus with overloading and early binding. Lecture Notes in Computer Science, 664:107–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiGianantonio:1993:ANA

[179] P. Di Gianantonio and F. Honsell. An abstract notion of application. Lecture Notes in Computer Science, 664:124– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dowek:1993:UTL

[180] G. Dowek. The undecidability of typability in the lambda-pi-calculus. Lecture Notes in Computer Science, 664: 139–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghelli:1993:RTC

[181] G. Ghelli. Recursive types are not conservative over Flsim. Lecture Notes in Computer Science, 664:146–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deGroote:1993:CTR

[182] Ph. de Groote. The conservation theorem revisited. Lecture Notes in Computer Science, 664:163–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hyland:1993:MRT

[183] J. M. E. Hyland and C.-H. L. Ong. Modified realizability toposes and strong normalization proofs. Lecture Notes in Computer Science, 664:179– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1993:SLO

[184] B. Jacobs. Semantics of lambda-I and of other substructure lambda calculi. Lecture Notes in Computer Science, 664:195–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1993:TDT

[185] B. Jacobs and T. Melham. Translating dependent type theory into higher order logic. Lecture Notes in Computer Science, 664:209–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jung:1993:SFA

[186] A. Jung and A. Stoughton. Studying the fully abstract model of PCF within its continuous function model. Lecture Notes in Computer Science, 664:230– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jung:1993:NCL

[187] A. Jung and J. Tiuryn. A new characterization of lambda definability. Lecture Notes in Computer Science, 664: 245–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leiss:1993:CRD

[188] H. Leiß. Combining recursive and dynamic types. Lecture Notes in Computer Science, 664:258–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leivant:1993:LCC

[189] D. Leivant and J.-Y. Marion. Lambda calculus characterizations of poly-time. Lecture Notes in Computer Science, 664:274–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKinna:1993:PTS

[190] J. McKinna and R. Pollack. Pure type systems formalized. Lecture Notes in Computer Science, 664:289–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:1993:OHO

[191] T. Nipkow. Orthogonal higher-order rewrite systems are confluent. Lecture Notes in Computer Science, 664:306– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otth:1993:MVA

[192] D. F. Otth. Monotonic versus antimonotonic exponentiation. Lecture Notes in Computer Science, 664:318– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulin-Mohring:1993:IDS

[193] Chr. Paulin-Mohring. Inductive definitions in the system Coq; rules and properties. Lecture Notes in Computer Science, 664:328–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierce:1993:ITB

[194] B. C. Pierce. Intersection types and bounded polymorphism. Lecture Notes in Computer Science, 664:346– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plotkin:1993:LPP

[195] G. Plotkin and M. Abadi. A logic for parametric polymorphism. Lecture Notes in Computer Science, 664:361– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sieber:1993:CVN

[196] K. Sieber. Call-by-value and nondeterminism. Lecture Notes in Computer Science, 664:376–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Springintveld:1993:LUB

[197] J. Springintveld. Lower and upper bounds for reductions of types in lambda omega and lambda P. Lecture Notes in Computer Science, 664:391– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahashi:1993:LCC

[198] M. Takahashi. lambda-calculi with conditional rules. Lecture Notes in Computer Science, 664:406–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Urzyczyn:1993:TRU

[199] P. Urzyczyn. Type reconstruction in F_{ω} is undecidable. Lecture Notes in Computer Science, 664:418–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castellani:1993:CDS

[200] I. Castellani. Causal and distributed semantics for concurrent processes. Lecture Notes in Computer Science, 665: 1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:AHS

[201] Anonymous. The alternation hierarchy for sublogarithmic space: an exciting race to STACS’93. Lecture Notes in Computer Science, 665:2–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonBraunmuhl:1993:ATW

[202] B. von Braunmuhl. Alternation for two-way machines with sublogarithmic space. Lecture Notes in Computer Science, 665:5–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liskiewicz:1993:SLL

[203] M. Liskiewicz and R. Reischuk. Separating the lower levels of the sublogarithmic space hierarchy. Lecture Notes in Computer Science, 665:16–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobler:1993:LPP

[204] J. Kobler. Locating P/poly optimally in the extended low hierarchy. Lecture Notes in Computer Science, 665:28–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lutz:1993:MSD

[205] J. H. Lutz and E. Mayordomo. Measure, stochasticity, and the density of hard languages. Lecture Notes in Computer Science, 665:38–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devienne:1993:HPO

[206] P. Devienne, P. Lebegue, and J.-C. Routier. Halting problem of one binary Horn clause is undecidable. Lecture Notes in Computer Science, 665:48–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaochen:1993:DUR

[207] Z. Chaochen, M. R. Hansen, and P. Sestoft. Decidability and undecidability results for duration calculus. Lecture Notes in Computer Science, 665:58–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deGroote:1993:DLT

[208] P. de Groote. Defining lambdatyped lambda-calculi by axiomatizing the typing relation. Lecture Notes in Computer Science, 665:69–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1993:CLB

[209] T. Eiter and G. Gottlob. The complexity of logic-based abduction. Lecture Notes in Computer Science, 665:70–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1993:TCB

[210] T. Kloks and D. Kratsch. Treewidth of chordal bipartite graphs. Lecture Notes in Computer Science, 665:80–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huckenbeck:1993:PNV

[211] U. Huckenbeck. On paths in networks with valves. Lecture Notes in Computer Science, 665:90–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sunder:1993:SIO

[212] S. Sunder and X. He. Scheduling interval ordered tasks in parallel. Lecture Notes in Computer Science, 665: 100–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pietracaprina:1993:SWC

[213] A. Pietracaprina and F. P. Preparata. An O(\sqrt{n})-worst-case-time solution to the granularity problem. Lecture Notes in Computer Science, 665:110– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desel:1993:SPP

[214] J. Desel and W. Reisig. The synthesis problem of Petri nets. Lecture Notes in Computer Science, 665:120–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1993:GRR

[215] E. Best, R. Devillers, and J. Esparza. General refinement and recursion operators for the Petri box calculus. Lecture Notes in Computer Science, 665:130–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonacina:1993:FDA

[216] M. P. Bonacina and J. Hsiang. On fairness in distributed automated deduction. Lecture Notes in Computer Science, 665:141–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1993:DCA

[217] E. W. Mayr and R. Werchner. Divideand-conquer algorithms on the hypercube. Lecture Notes in Computer Science, 665:153–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allender:1993:FOI

[218] E. Allender, J. Balcazar, and N. Immerman. A first-order isomorphism theorem. Lecture Notes in Computer Science, 665:163–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhrman:1993:SRS

[219] H. Buhrman, A. Hoene, and L. Torenvliet. Splittings, robustness and structure of complete sets. Lecture Notes in Computer Science, 665:175–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemachandra:1993:DUD

[220] L. A. Hemachandra and S. K. Jha. Defying upward and downward separation. Lecture Notes in Computer Science, 665:185–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoene:1993:CSS

[221] A. Hoene and A. Nickelsen. Counting, selecting, and sorting by querybounded machines. Lecture Notes in Computer Science, 665:196–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jantzen:1993:CCF

[222] M. Jantzen and H. Petersen. Cancellation in context-free languages: enrichment by reduction. Lecture Notes in Computer Science, 665:206–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cassaigne:1993:COF

[223] J. Cassaigne. Counting overlap-free binary words. Lecture Notes in Computer Science, 665:216–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narbel:1993:LSR

[224] P. Narbel. The limit set of recognizable substitution systems. Lecture Notes in Computer Science, 665:226–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krob:1993:PCL

[225] D. Krob and P. Lalonde. Partially commutative Lyndon words. Lecture Notes in Computer Science, 665:237– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monien:1993:PAD

[226] B. Monien, R. Feldmann, R. Klasing, and R. Luling. Parallel architectures: design and efficient use. Lecture Notes in Computer Science, 665: 247–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Formann:1993:WCP

[227] M. Formann. Weighted closest pairs. Lecture Notes in Computer Science, 665:270–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuierer:1993:RPQ

[228] S. Schuierer. Rectilinear path queries in a simple rectilinear polygon. Lecture Notes in Computer Science, 665: 282–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czumaj:1993:PAM

[229] A. Czumaj. Parallel algorithm for the matrix chain product and the optimal triangulation problems. Lecture Notes in Computer Science, 665:294– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dessmark:1993:MLR

[230] A. Dessmark, A. Lingas, and A. Maheshwari. Multi-list ranking: complexity and applications. Lecture Notes in Computer Science, 665:306–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kucera:1993:EAG

[231] L. Kucera, K. Mehlhorn, B. Preis, and E. Schwarzenecker. Exact algorithms for a geometric packing problem. Lecture Notes in Computer Science, 665:317–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maler:1993:DTP

[232] O. Maler. A decomposition theorem for probabilistic transition systems. Lecture Notes in Computer Science, 665:323–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montalbano:1993:LAC

[233] R. Montalbano. Local automata and completion. Lecture Notes in Computer Science, 665:333–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culik:1993:ECW

[234] K. Culik and S. Dube. Efficient compression of wavelet coefficients for smooth and fractal-like data. Lecture Notes in Computer Science, 665:343– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra:1993:ETW

[235] O. H. Ibarra, T. Jiang, N. Tran, and H. Wang. On the equivalence of two-way pushdown automata and counter machines over bounded languages. Lecture Notes in Computer Science, 665:354–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cosnard:1993:CPL

[236] M. Cosnard, M. Garzon, and P. Koiran. Computability properties of lowdimensional dynamical systems. Lecture Notes in Computer Science, 665: 365–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abrahamson:1993:FPI

[237] K. A. Abrahamson, R. G. Downey, and M. R. Fellows. Fixed-parameter intractability II. Lecture Notes in Computer Science, 665:374–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fich:1993:LPP

[238] F. E. Fich, R. Impagliazzo, B. Kapron, V. King, and M. Kutylowski. Limits on the power of parallel random access machines with weak forms of write conflict resolution. Lecture Notes in Computer Science, 665:386–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fenner:1993:UOC

[239] S. Fenner, S. Ogiwara, M. Homer, and A. L. Selman. On using oracles that compute values. Lecture Notes in Computer Science, 665:398–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gengler:1993:MAS

[240] R. Gengler. Multicounter automata with sublogarithmic reversal bounds. Lecture Notes in Computer Science, 665:408–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uselton:1993:SOS

[241] A. C. Uselton. Structured operational semantics for concurrency and hierarchy. Lecture Notes in Computer Science, 665:418–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hungar:1993:CVF

[242] H. Hungar. The complexity of verifying functional programs. Lecture Notes in Computer Science, 665:428–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lentfert:1993:TFD

[243] P. J. A. Lentfert and S. D. Swierstra. Towards the formal design of selfstabilizing distributed algorithms. Lecture Notes in Computer Science, 665: 440–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Penczek:1993:ATL

[244] W. Penczek. Axiomatizations of temporal logics on trace systems. Lecture Notes in Computer Science, 665:452– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lurwer-Bruggemeier:1993:CCC

[245] K. Lurwer-Bruggemeier and F. Meyer auf der Heide. Capabilities and complexity of computations with integer division. Lecture Notes in Computer Science, 665:463–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niedermeier:1993:ELD

[246] R. Niedermeier and P. Rossmanith. Extended locally definable acceptance types. Lecture Notes in Computer Science, 665:473–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fenner:1993:GDC

[247] S. Fenner, L. Fortnow, and L. Li. Gapdefinability as a closure property. Lecture Notes in Computer Science, 665: 484–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choffrut:1993:LDS

[248] C. Choffrut and L. Guerra. On the logical definability of some rational trace languages. Lecture Notes in Computer Science, 665:494–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilleron:1993:SSS

[249] R. Gilleron, S. Tison, and M. Tommasi. Solving systems of set constraints using tree automata. Lecture Notes in Computer Science, 665:505–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lugiez:1993:CPT

[250] D. Lugiez and J. L. Moysset. Complement problems and tree automata in AC-like theories. Lecture Notes in Computer Science, 665:515–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babai:1993:THP

[251] L. Babai. Transparent (holographic) proofs. Lecture Notes in Computer Science, 665:525–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1993:CSF

[252] Z.-L. Zhang, D. A. Mix Barrington, and J. Tarui. Computing symmetric functions with AND/OR circuits and a single MAJORITY gate. Lecture Notes in Computer Science, 665:535– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maciel:1993:TCI

[253] A. Maciel and D. Therien. Threshold circuits for iterated multiplication: using AC^{O} for free. Lecture Notes in Computer Science, 665:545–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaudry:1993:CMG

[254] M. Beaudry, P. McKenzie, and P. Peladeau. Circuits with monoidal gates. Lecture Notes in Computer Science, 665:555–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Istrail:1993:NPS

[255] S. Istrail and D. Zivkovic. A nonprobabilistic switching lemma for the Sipser function. Lecture Notes in Computer Science, 665:566–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gergov:1993:FFP

[256] J. Gergov and C. Meinel. Frontiers of feasible and probabilistic feasible Boolean manipulation with branching programs. Lecture Notes in Computer Science, 665:576–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maler:1993:SCO

[257] O. Maler and L. Staiger. On syntactic congruences for omega-languages. Lecture Notes in Computer Science, 665: 586–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varricchio:1993:PTA

[258] S. Varricchio. A polynomial time algorithm for the equivalence of two morphisms on omega-regular languages. Lecture Notes in Computer Science, 665:595–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilke:1993:LTT

[259] T. Wilke. Locally threshold testable languages of infinite words. Lecture Notes in Computer Science, 665:607– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekert:1993:DAA

[260] V. Diekert and A. Muscholl. Deterministic asynchronous automata for infinite traces. Lecture Notes in Computer Science, 665:617–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staiger:1993:RAI

[261] L. Staiger. Recursive automata on infinite words. Lecture Notes in Computer Science, 665:629–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sairam:1993:CTA

[262] S. Sairam, J. S. Vitter, and R. Tamassia. A complexity theoretic approach to incremental computation. Lecture Notes in Computer Science, 665:640– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reischuk:1993:PAC

[263] R. Reischuk and C. Schindelhauer. Precise average case complexity. Lecture Notes in Computer Science, 665:650– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miltersen:1993:BPC

[264] P. B. Miltersen. The bit probe complexity measure revisited. Lecture Notes in Computer Science, 665:662–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baliga:1993:LLS

[265] G. Baliga, J. Case, and S. Jain. Language learning with some negative information. Lecture Notes in Computer Science, 665:672–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1993:LLB

[266] S. Lange and T. Zeugmann. Language learning with a bounded number of mind changes. Lecture Notes in Computer Science, 665:682–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blundo:1993:ESM

[267] C. Blundo, A. De Santis, and U. Vaccaro. Efficient sharing of many secrets. Lecture Notes in Computer Science, 665:692–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Drexler:1993:KST

[268] R. Drexler, W. Reif, G. Schellhorn, K. Stenzel, W. Stephan, and A. Wolpers. The KIV system — a tool for formal program development. Lecture Notes in Computer Science, 665:704–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofting:1993:GSI

[269] F. Hofting, E. Wanke, A. Balmosan, and C. Bergmann. 1st grade — a system for implementation, testing and animation of graph algorithms. Lecture Notes in Computer Science, 665: 706–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufl:1993:PVT

[270] T. Kaufl. The program verifier Tatzelwurm. Lecture Notes in Computer Science, 665:708–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naher:1993:LLE

[271] S. Naher. LEDA — a library of efficient data types and algorithms. Lecture Notes in Computer Science, 665: 710–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Backhouse:1993:CPA

[272] R. C. Backhouse, J. P. H. W. Van den Eijnde, and A. J. M. Van Gasteren. Calculating path algorithms. Lecture Notes in Computer Science, 669:3–20, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gardiner:1993:ACP

[273] P. H. B. Gardiner, C. E. Martin, and O. De Moor. An algebraic construction of predicate transformers. Lecture Notes in Computer Science, 669:21–44, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofstee:1993:DCS

[274] H. P. Hofstee. Distributing a class of sequential programs. Lecture Notes in Computer Science, 669:45–66, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogendijk:1993:RPL

[275] P. F. Hoogendijk and R. C. Backhouse. Relational programming laws in the tree, list, bag, set hierarchy. Lecture Notes in Computer Science, 669:67– 106, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1993:DAC

[276] G. Jones and M. Sheeran. Designing arithmetic circuits by refinement in Ruby. Lecture Notes in Computer Science, 669:107–136, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lukkien:1993:OSG

[277] J. J. Lukkien. Operational semantics and generalized weakest preconditions. Lecture Notes in Computer Science, 669:137–156, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moeller:1993:SPG

[278] B. Moeller and M. Russling. Shorter paths to graph algorithms. Lecture Notes in Computer Science, 669:157– 180, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbay:1993:HCL

[279] D. M. Gabbay. How to construct a logic for your application. Lecture Notes in Computer Science, 671:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumgartner:1993:MEC

[280] Peter Baumgartner. A model elimination calculus with built-in theories. Lecture Notes in Computer Science, 671:30–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weidenbach:1993:NSL

[281] Christoph Weidenbach. A new sorted logic. Lecture Notes in Computer Science, 671:43–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xiaorong:1993:EFH

[282] Huang Xiaorong. An explanatory framework for human theorem proving. Lecture Notes in Computer Science, 671:55–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Posegga:1993:TFO

[283] Joachim Posegga and Bertram Ludascher. Towards first-order deduction based on Shannon graphs. Lecture Notes in Computer Science, 671:67–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmann:1993:SFE

[284] Reinhard Bachmann, Thomas Malsch, and Susanne Ziegler. Success and failure of expert systems in different fields of industrial application. Lecture Notes in Computer Science, 671:77–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linster:1993:VKE

[285] Marc Linster. Viewing knowledge engineering as a symbiosis of “modeling to make sense” and “modeling to implement systems”. Lecture Notes in Computer Science, 671:87–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branskat:1993:CBK

[286] Sonja Branskat and Marc Linster. Cases as a basis for knowledge acquisition in the pre-formal phases of knowledge engineering. Lecture Notes in Computer Science, 671:100–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coulon:1993:CGT

[287] Carl-Helmut Coulon, Frank van Harmelen, Werner Karbach, and Angi Vob. Controlling generate and test in any time. Lecture Notes in Computer Science, 671:110–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weigel:1993:ECS

[288] Achim Weigel and Rainer Bleisinger. Efficient computation of solutions for contradictory time interval networks. Lecture Notes in Computer Science, 671:122–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1993:ECL

[289] Franz Baader and Philipp Hanschke. Extensions of concept languages for a mechanical engineering application. Lecture Notes in Computer Science, 671:132–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanschke:1993:CTR

[290] Philipp Hanschke and Knut Hinkelmann. Combining terminological and rule-based reasoning for abstraction processes. Lecture Notes in Computer Science, 671:144–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinkelmann:1993:FLE

[291] Knut Hinkelmann. Forward logic evaluation: Compiling a partially evaluated meta-interpreter into the WAM. Lecture Notes in Computer Science, 671: 156–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prem:1993:CSM

[292] Erich Prem, Markus Mackinger, Georg Dorffner, G. Porenta, and H. Sochor. Concept support as a method for programming neural networks with symbolic knowledge. Lecture Notes in Computer Science, 671:166–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrmann:1993:HIG

[293] Jurgen Herrmann and Renate Beckmann. A heuristic inductive generalization method and its application to VLSI-Design. Lecture Notes in Computer Science, 671:176–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergmann:1993:LPA

[294] Ralph Bergmann. Learning plan abstractions. Lecture Notes in Computer Science, 671:187–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linden:1993:DQF

[295] Alexander Linden. On discontinuous Q-functions in reinforcement learning. Lecture Notes in Computer Science, 671:199–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poeck:1993:ITS

[296] Karsten Poeck and Martin Tins. An intelligent tutoring system for classification problem solving. Lecture Notes in Computer Science, 671:210–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroder:1993:KBP

[297] Martin Schroder. Knowledge-based processing of medical language: A language engineering approach. Lecture Notes in Computer Science, 671:221– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kranzdorf:1993:TPI

[298] Heike Kranzdorf and Ulrike Griefahn. Text planning in ITEX: A hybrid approach. Lecture Notes in Computer Science, 671:235–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1993:YNQ

[299] Marion Schulz and Daniela Schmidt. Yes/no questions with negation: Towards integrating semantics and pragmatics. Lecture Notes in Computer Science, 671:247–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konig:1993:EDA

[300] Esther Konig. An efficient decision algorithm for feature logic. Lecture Notes in Computer Science, 671:255– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Decker:1993:UQQ

[301] Stefan Decker and Christoph Lingenfelder. Universally quantified queries in languages with order-sorted logics. Lecture Notes in Computer Science, 671:267–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meier:1993:SVE

[302] Justus Meier. A semantic view of explanation. Lecture Notes in Computer Science, 671:272–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janetzko:1993:GDS

[303] Dietmar Janetzko, Stefan Wess, and Erica Melis. Goal-driven similarity assessment. Lecture Notes in Computer Science, 671:283–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenting:1993:DNR

[304] Jacques H. J. Lenting and Peter J. Braspenning. Delegated negotiation for resource re-allocation. Lecture Notes in Computer Science, 671:299–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deGreef:1993:TSL

[305] Paul de Greef, Keith Clark, and Frank McCabe. Towards a specification language for cooperation methods. Lecture Notes in Computer Science, 671: 313–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramer:1993:IOS

[306] Martin Kramer. Improving operating system usage. Lecture Notes in Computer Science, 671:320–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolln:1993:RUM

[307] Matthias E. Kolln. The role of user models for conflicts in a constraintbased model of generation. Lecture Notes in Computer Science, 671:334– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchberger:1993:CNL

[308] Ernst Buchberger. Criteria in natural language generation: Minimal criteria and their impacts. Lecture Notes in Computer Science, 671:347–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1993:TRN

[309] Renate A. Schmidt. Terminological representation, natural language and relation algebra. Lecture Notes in Computer Science, 671:357–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lux:1993:LHI

[310] Andreas Lux and Michael Kolb. Linking humans and intelligent systems or: What are user agents good for? Lecture Notes in Computer Science, 671: 372–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kindler:1993:AMA

[311] H. Kindler, D. Densow, and T. M. Fliedner. An advisor for the management of the acute radiation syndrome. Lecture Notes in Computer Science, 671:386–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barak:1993:MDO

[312] Amnon Barak, Shai Guday, and Richard G. Wheeler. The MOSIX distributed operating system: Load balancing for UNIX. Lecture Notes in Computer Science, 672:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chehaibar:1993:ROI

[313] G. Chehaibar. Replacement of open interface subnets and stable state transformation equivalence. Lecture Notes in Computer Science, 674:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherkasova:1993:BSS

[314] L. Cherkasova, R. R. Howell, and L. E. Rosier. Bounded self-stabilizing Petri nets. Lecture Notes in Computer Science, 674:26–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Couvreur:1993:GFP

[315] J. M. Couvreur, S. Haddad, and J. F. Peyre. Generative families of positive invariants in coloured nets subclasses. Lecture Notes in Computer Science, 674:51–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1993:ATC

[316] R. David and H. Alla. Autonomous and timed continuous Petri ness. Lecture Notes in Computer Science, 674:71–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1993:DPN

[317] J. B. Evans. The devnet: a Petri net for discrete event simulation. Lecture Notes in Computer Science, 674:91–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ezpeleta:1993:NTF

[318] J. Ezpeleta, J. M. Couvreur, and M. Silva. A new technique for finding a generating family of siphons, traps and ST-Components. application to colored Petri nets. Lecture Notes in Computer Science, 674:126–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fehling:1993:CHP

[319] R. Fehling. A concept of hierarchical Petri nets with building blocks. Lecture Notes in Computer Science, 674: 148–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldbrugge:1993:PNT

[320] F. Feldbrugge. Petri net tool overview 1992. Lecture Notes in Computer Science, 674:169–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finkel:1993:MCG

[321] A. Finkel. The minimal coverability graph for Petri nets. Lecture Notes in Computer Science, 674:210–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorrieri:1993:DIC

[322] R. Gorrieri and U. Montanari. Distributed implementation of CCS. Lecture Notes in Computer Science, 674: 244–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gouet:1993:SCS

[323] R. Gouet, F. Plo, M. San Miguel, and B. Ycart. Saturation conditions for stochastic Petri nets. Lecture Notes in Computer Science, 674:267–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laftit:1993:MOT

[324] S. Laftit, J. M. Proth, and X. L. Xie. Marking optimization in timed event graphs. Lecture Notes in Computer Science, 674:281–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindqvist:1993:PRT

[325] M. Lindqvist. Parameterized reachability trees for predicate/transition nets. Lecture Notes in Computer Science, 674:301–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1993:IDA

[326] Y. Li and C. Murray Woodside. Iterative decomposition and aggregation of stochastic marked graph Petri nets. Lecture Notes in Computer Science, 674:325–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMichelis:1993:CEC

[327] G. De Michelis. Constraints and extensions in a calculus of EN systems. Lecture Notes in Computer Science, 674:350–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1993:COM

[328] E. Smith. Comparability orders and measurement. Lecture Notes in Computer Science, 674:371–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Souissi:1993:DSS

[329] Y. Souissi. Deterministic systems of sequential processes: a class of structured Petri nets. Lecture Notes in Computer Science, 674:406–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valmari:1993:CSS

[330] A. Valmari. Compositional state space generation. Lecture Notes in Computer Science, 674:427–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mulkers:1993:LDS

[331] Anne Mulkers. Live data structures in logic programs: Derivation by means of abstract interpretation. Lecture Notes in Computer Science, 675:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reiss:1993:RPO

[332] Thomas H. Reiss. Recognizing planar objects using invariant image features. Lecture Notes in Computer Science, 676:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makanin:1993:GSE

[333] G. S. Makanin. On general solution of equations in a free semigroup. Lecture Notes in Computer Science, 677:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choffrut:1993:CFI

[334] C. Choffrut. Conjugacy in free inverse monoids. Lecture Notes in Computer Science, 677:6–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1993:GAU

[335] F. Baader and K. U. Schulz. General A- and AX-Unification via optimized combination procedures. Lecture Notes in Computer Science, 677:23–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charatonik:1993:WET

[336] W. Charatonik and L. Pacholski. Word equations with two variables. Lecture Notes in Computer Science, 677:43–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dehornoy:1993:NPL

[337] P. Dehornoy. The naming problem for left distributivity. Lecture Notes in Computer Science, 677:57–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Enjalbert:1993:CTA

[338] P. Enjalbert and F. Clerin-Debart. A case of termination for associative unification. Lecture Notes in Computer Science, 677:79–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennaceur:1993:TPC

[339] H. Bennaceur and G. Plateau. Theorem proving by combinatorial optimization. Lecture Notes in Computer Science, 677:90–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auer:1993:SSE

[340] P. Auer. Solving string equations with constant restrictions. Lecture Notes in Computer Science, 677:103–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdulrab:1993:LTN

[341] H. Abdulrab. LOP: Toward a new implementation of Makanin’s algorithm.

Lecture Notes in Computer Science, 677:133–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1993:WUT

[342] K. U. Schulz. Word unification and transformation of generalized equations. Lecture Notes in Computer Science, 677:150–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auer:1993:UCD

[343] P. Auer. Unification in the combination of disjoint theories. Lecture Notes in Computer Science, 677:177–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neraud:1993:SRT

[344] J. Neraud. On the subsets of rank two in a free monoid: A fast decision algorithm. Lecture Notes in Computer Science, 677:187–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kounalis:1993:SCP

[345] E. Kounalis, D. Lugiez, and L. Pottier. A solution of the complement problem in associative-commutative theories. Lecture Notes in Computer Science, 677:197–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valiant:1993:CMP

[346] Leslie G. Valiant. A combining mechanism for parallel computers. Lecture Notes in Computer Science, 678:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vishkin:1993:CPS

[347] Uzi Vishkin. A case for the PRAM as a standard programmer’s model. Lecture Notes in Computer Science, 678:11–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AufDerHeide:1993:HSS

[348] Friedhelm Meyer Auf Der Heide. Hashing strategies for simulating shared memory on distributed memory machines. Lecture Notes in Computer Science, 678:20–29, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosenberg:1993:BPA

[349] Arnold L. Rosenberg. Better parallel architectures via emulations. Lecture Notes in Computer Science, 678:30–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cosnard:1993:RBM

[350] Michel Cosnard and Pascal Koiran. Relations between models of parallel abstract machines. Lecture Notes in Computer Science, 678:37–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preparata:1993:FPC

[351] Franco P. Preparata. Frontiers of parallel computing. Lecture Notes in Computer Science, 678:47–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Formella:1993:CED

[352] Arno Formella, Werner Massonne, and Wolfgang J. Paul. Cost effectiveness of data flow machines and vector processors. Lecture Notes in Computer Science, 678:48–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leiserson:1993:NCM

[353] Charles E. Leiserson. The networks of the connection machine CM-5. Lecture Notes in Computer Science, 678:66–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnsson:1993:MPC

[354] S. Lennart Johnsson. Massively parallel computing: Data distribution and communication. Lecture Notes in Computer Science, 678:68–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monien:1993:REP

[355] Burkhard Monien, Reinhard Luling, and Falk Langhammer. A realizable efficient parallel architecture. Lecture Notes in Computer Science, 678:93–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rudolph:1993:HSC

[356] Larry Rudolph. Hardware support for collective communication operations. Lecture Notes in Computer Science, 678:110–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertin:1993:PAM

[357] Patrice Bertin, Didier Roncin, and Jean Vuillemin. Programmable active memories: A performance assessment. Lecture Notes in Computer Science, 678:119–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leighton:1993:RRD

[358] Tom Leighton and Bruce Maggs. The role of randomness in the design of interconnection networks. Lecture Notes in Computer Science, 678:131– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snyder:1993:CRF

[359] Lawrence Snyder. Chaos router: Finally, a practical adaptive router? Lecture Notes in Computer Science, 678: 146–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Felperin:1993:ESW

[360] Sergio Felperin, Prabhakar Raghavan, and Eli Upfal. An experimental study of wormhole routing in parallel computers. Lecture Notes in Computer Science, 678:156–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luccio:1993:TNC

[361] Fabrizio Luccio, Linda Pagli, and Geppino Pucci. Three non conventional paradigms of parallel computation. Lecture Notes in Computer Science, 678:166–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snir:1993:SPC

[362] Marc Snir. Scalable parallel computers and scalable parallel codes: From theory to practice. Lecture Notes in Computer Science, 678:176–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranade:1993:FAL

[363] Abhiram Ranade. A framework for analyzing locality and portability issues in parallel computing. Lecture Notes in Computer Science, 678:185–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1993:OIG

[364] Ernst W. Mayr and Ralph Werchner. Optimal implementation of general divide-and-conquer on the hypercube and related networks. Lecture Notes in Computer Science, 678:195–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Annexstein:1993:PIG

[365] Fred S. Annexstein. Parallel implementations of graph embeddings. Lecture Notes in Computer Science, 678:207– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heydari:1993:SPR

[366] Mohammad H. Heydari and I. Hal Sudborough. On sorting by prefix reversals and the diameter of pancake networks. Lecture Notes in Computer Science, 678:218–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:I

[367] Anonymous. Introduction. Lecture Notes in Computer Science, 679:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:T

[368] Anonymous. Terminology. Lecture Notes in Computer Science, 679:6–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:SCR

[369] Anonymous. Semantic clash resolution as decision procedure. Lecture Notes in Computer Science, 679:17–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:COR

[370] Anonymous. Completeness of ordering refinements. Lecture Notes in Computer Science, 679:60–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:STB

[371] Anonymous. Semantic tree based resolution variants. Lecture Notes in Computer Science, 679:93–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:DCK

[372] Anonymous. Deciding the class K by an ordering refinement. Lecture Notes in Computer Science, 679:130–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:RBM

[373] Anonymous. A resolution based method for building finite models. Lecture Notes in Computer Science, 679: 151–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:A

[374] Anonymous. Applications. Lecture Notes in Computer Science, 679:180– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:R

[375] Anonymous. References. Lecture Notes in Computer Science, 679:198– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1993:PDS

[376] Berthold Hoffmann and Bernd Krieg-Bruckner. Program development by specification and transformation. the PROSPECTRA methodology, language family, and system. Lecture Notes in Computer Science, 680:3–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wansing:1993:LIS

[377] Heinrich Wansing. The logic of information structures. Lecture Notes in Computer Science, 681:1–158, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1993:PA

[378] D. Dubois and H. Prade. Possibilistic abduction. Lecture Notes in Computer Science, 682:3–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cayrol:1993:MPA

[379] C. Cayrol, V. Royer, and C. Saurel. Management of preferences in assumptionbased reasoning. Lecture Notes in Computer Science, 682:13–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coupey:1993:DEK

[380] P. Coupey. Default exclusion in a KL-ONE-Like terminological component. Lecture Notes in Computer Science, 682:23–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moinard:1993:UVA

[381] Y. Moinard. Unifying various approaches to default logic. Lecture Notes in Computer Science, 682:33–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cussens:1993:UME

[382] J. Cussens and A. Hunter. Using maximum entropy in a defeasible logic with probabilistic semantics. Lecture Notes in Computer Science, 682:43–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madani:1993:LIN

[383] N. Madani. Legality in inheritance networks. Lecture Notes in Computer Science, 682:53–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monserrat:1993:NIS

[384] M. Monserrat, F. Rossello, and J. Torrens. A note on information systems associated to termal algebras. Lecture Notes in Computer Science, 682:63–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferney:1993:BCR

[385] D. Ferney and A. Quilliot. A backward chaining resolution process involving non-monotonic operators. Lecture Notes in Computer Science, 682: 73–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trillas:1993:FCG

[386] E. Trillas. On fuzzy conditionals generalising the material conditional. Lecture Notes in Computer Science, 682: 85–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tamburrini:1993:IRL

[387] G. Tamburrini and S. Termini. Integrating resolution-like procedures with Łukasiewicz implication. Lecture Notes in Computer Science, 682:101– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krause:1993:DLA

[388] P. Krause, S. Ambler, and J. Fox. The development of a “logic of argumentation”. Lecture Notes in Computer Science, 682:109–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnould:1993:AO

[389] T. Arnould and A. Ralescu. From “and” to “or”. Lecture Notes in Computer Science, 682:119–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eberbach:1993:RST

[390] E. Eberbach and A. Trudel. Representing spatial and temporal uncertainty. Lecture Notes in Computer Science, 682:129–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodriguez:1993:ATR

[391] R. V. Rodriguez and F. D. Anger. An analysis of the temporal relations of intervals in relativistic space-time. Lecture Notes in Computer Science, 682: 139–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casasnovas:1993:AIF

[392] J. Casasnovas and J. J. Miro-Julia. Accumulation and inference over finitegenerated algebras for mapping approximations. Lecture Notes in Computer Science, 682:149–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonissone:1993:SMC

[393] P. P. Bonissone and S. Ayub. Similarity measures for case-based reasoning systems. Lecture Notes in Computer Science, 682:161–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutherland:1993:SML

[394] A. Sutherland, R. Henery, R. Molina, C. C. Taylor, and R. King. Statistical methods in learning. Lecture Notes in Computer Science, 682:173–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kacprzyk:1993:LEE

[395] J. Kacprzyk and C. Iwanski. Learning from erroneous examples using fuzzy logic and “textbook” knowledge. Lecture Notes in Computer Science, 682: 183–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arigoni:1993:ILR

[396] A. O. Arigoni, C. Furlanello, and V. Maniezzo. Incremental learning of roughly represented concepts. Lecture Notes in Computer Science, 682:192– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagib:1993:SOQ

[397] G. Nagib, W. Gharieb, and Z. Binder. Self-organizing qualitative multimodel control. Lecture Notes in Computer Science, 682:199–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bordeaux:1993:MHL

[398] F. Bordeaux, F. Forest, and B. Grau. MoHA, an hybrid learning model: A model based on the perception of the environment by an individual. Lecture Notes in Computer Science, 682:209– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiol:1993:NPI

[399] G. Fiol, J. Miro-Nicolau, and J. Miro-Julia. A new perspective in the inductive acquisition of knowledge from examples. Lecture Notes in Computer Science, 682:219–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruiz:1993:KRT

[400] I. Luque Ruiz, J. L. Cruz Soto, and M. A. Gomez-Nieto. Knowledge representation through object in the development of expert system chemical synthesis and reaction. Lecture Notes in Computer Science, 682:229–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yager:1993:HRF

[401] R. R. Yager. Hierarchical representation of fuzzy If-Then rules. Lecture Notes in Computer Science, 682:239– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spilski-Petieau:1993:ARE

[402] A. M. Spilski-Petieau and D. Willaeys. Approximate reasoning in expert systems: Inference and combination tools. Lecture Notes in Computer Science, 682:248–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohout:1993:MIB

[403] L. J. Kohout and W. Bandler. Modes of interval-based plausible reasoning viewed via the checklist paradigm. Lecture Notes in Computer Science, 682: 256–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardona:1993:RBS

[404] L. Cardona, J. Kohlas, and P. A. Monney. Rule-based systems with unreliable conditions. Lecture Notes in Computer Science, 682:265–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quevedo:1993:FSE

[405] J. Quevedo, C. Gilbert Oliveras, and J. Aguilar-Martin. Fuzzy semantics in expert process control. Lecture Notes in Computer Science, 682:275–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rakoto-Ravalontsalama:1993:QOP

[406] N. Rakoto-Ravalontsalama, A. Missier, and J. S. Kikkert. Qualitative operators and process engineer semantics of uncertainty. Lecture Notes in Computer Science, 682:284–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sawadogo:1993:FUM

[407] S. Sawadogo, A. K. Achaibou, and F. Mora-Camino. Facing uncertainty in the management of large irrigation systems: Qualitative approach. Lecture Notes in Computer Science, 682: 294–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oleary:1993:SAE

[408] D. E. O’leary. Semantic ambiguity in expert systems: The case of deterministic systems. Lecture Notes in Computer Science, 682:304–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Casasnovas:1993:DRA

[409] J. Casasnovas. A deduction rule for the approximated knowledge of a mapping. Lecture Notes in Computer Science, 682:314–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feruglio:1993:KBR

[410] G. Valiente Feruglio. On knowledge base redundancy under uncertain reasoning. Lecture Notes in Computer Science, 682:321–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguilar-Crespo:1993:FLA

[411] J. A. Aguilar-Crespo, E. De Pablo, and X. Alaman. A fuzzy logic approach for sensor validation in real time expert systems. Lecture Notes in Computer Science, 682:330–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glorennec:1993:ANF

[412] P. Y. Glorennec, C. Barret, and M. Brunet. Application of neurofuzzy networks to the identification and control of nonlinear dynamical systems. Lecture Notes in Computer Science, 682:341–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinez:1993:CBA

[413] C. Hervas Martinez, E. J. Romero Soto, N. Garcia Pedrajas, and R. Medina Carnicer. Comparison between artificial neural networks and classical statistical methods in pattern recognition. Lecture Notes in Computer Science, 682:351–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amghar:1993:LMO

[414] S. Amghar, H. Paugam-Moisy, and J. P. Royet. Learning methods for odor recognition modeling. Lecture Notes in Computer Science, 682:361–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poncelet:1993:CSU

[415] P. Poncelet and L. Lakhal. Consistent structural updates for object database design. Lecture Notes in Computer Science, 685:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ewald:1993:PAS

[416] C. A. Ewald and M. E. Orlowska. A procedural approach to schema evolution. Lecture Notes in Computer Science, 685:22–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellahsene:1993:AMM

[417] Z. Bellahsene. An active meta-model for knowledge evolution in an objectoriented database. Lecture Notes in Computer Science, 685:39–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thieme:1993:SIO

[418] C. Thieme and A. Siebes. Schema integration in object-oriented databases. Lecture Notes in Computer Science, 685:54–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johannesson:1993:STA

[419] P. Johannesson. Schema transformations as an aid in view integration. Lecture Notes in Computer Science, 685:71–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perrin:1993:TMP

[420] O. Perrin and N. Boudjlida. Towards a model for persistent data integration. Lecture Notes in Computer Science, 685:93–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gulla:1993:UEI

[421] J. A. Gulla and G. Willumsen. Using explanations to improve the validation of executable models. Lecture Notes in Computer Science, 685:118–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lalioti:1993:VV

[422] V. Lalioti and P. Loucopoulos. Visualisation for validation. Lecture Notes in Computer Science, 685:143–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindland:1993:VCM

[423] O. I. Lindland and J. Krogstie. Validating conceptual models by transformational prototyping. Lecture Notes in Computer Science, 685:165–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blazy:1993:PES

[424] S. Blazy and P. Facon. Partial evaluation and symbolic computation for the understanding of Fortran programs. Lecture Notes in Computer Science, 685:184–198, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piprani:1993:MMA

[425] C. B. Piprani and R. B. Morris. A multi-model approach for deriving requirements specifications for a megaproject. Lecture Notes in Computer Science, 685:199–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breu:1993:BMA

[426] M. Breu, G. Leonardi, B. Percie du Sert, L. Barengo, M. Pfeiffer, and J.- C. Utter. The BOS-Method architecture: An improved structured approach for the development of distributed information systems. Lecture Notes in Computer Science, 685:221–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chung:1993:DSR

[427] L. Chung. Dealing with security requirements during the development of information systems. Lecture Notes in Computer Science, 685:234–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1993:EFR

[428] E. Dubois, P. Du Bois, and M. Petit. Elicitating and formalising requirements for C.I.M information systems. Lecture Notes in Computer Science, 685:252–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pohl:1993:TDR

[429] K. Pohl. The three dimensions of requirements engineering. Lecture Notes in Computer Science, 685:275– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brunet:1993:OOA

[430] J. Brunet, C. Cauvet, D. Meddahi, and F. Semmak. Object-oriented analysis in practice. Lecture Notes in Computer Science, 685:293–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Opdahl:1993:CRW

[431] A. L. Opdahl and G. Sindre. Concepts for real-world modelling. Lecture Notes in Computer Science, 685:309– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seltveit:1993:ABR

[432] A. H. Seltveit. An abstraction-based rule approach to large-scale information systems development. Lecture Notes in Computer Science, 685:328– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Motschnig-Pitrik:1993:SPV

[433] R. Motschnig-Pitrik. The semantics of parts versus aggregates in data/ knowledge modelling. Lecture Notes in Computer Science, 685:352–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quer:1993:OIO

[434] C. Quer and A. Olive. Object interaction in object-oriented deductive conceptual models. Lecture Notes in Computer Science, 685:374–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edelweiss:1993:OOT

[435] N. Edelweiss, J. P. M. de Oliveira, and B. Pernici. An object-oriented temporal model. Lecture Notes in Computer Science, 685:397–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kraiem:1993:OOD

[436] N. Kraiem, F. Gargouri, and F. Boufares. From object-oriented design towards object-oriented programming. Lecture Notes in Computer Science, 685:416–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCormack:1993:AMC

[437] J. J. McCormack, T. A. Halpin, and P. R. Ritson. Automated mapping of conceptual schemas to relational schemas. Lecture Notes in Computer Science, 685:432–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andonoff:1993:NOO

[438] E. Andonoff. Normalization of objectoriented conceptual schemes. Lecture Notes in Computer Science, 685:449– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vlachantonis:1993:TRI

[439] N. Vlachantonis, R. Herzig, M. Gogolla, G. Denker, S. Conrad, and H.-D. Ehrich. Towards reliable information systems: The KorSo approach. Lecture Notes in Computer Science, 685:463–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spanoudakis:1993:SAS

[440] G. Spanoudakis and P. Constantopoulos. Similarity for analogical software reuse: A conceptual modelling approach. Lecture Notes in Computer Science, 685:483–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAntonellis:1993:TAR

[441] V. De Antonellis and L. Vandoni. Temporal aspects in reuse of requirement specifications. Lecture Notes in Computer Science, 685:504–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brockers:1993:CAV

[442] A. Brockers and V. Gruhn. Computeraided verification of software process model properties. Lecture Notes in Computer Science, 685:521–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Croshere:1993:AFS

[443] R. Croshere, R. Van de Riet, and A. Blom. An animation facility to simulate an information and communication system. Lecture Notes in Computer Science, 685:547–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palanque:1993:DUD

[444] P. A. Palanque, R. Bastide, L. Dourte, and C. Sibertin-Blanc. Design of userdriven interfaces using Petri nets and objects. Lecture Notes in Computer Science, 685:569–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prevelakis:1993:PSD

[445] V. Prevelakis and D. Tsichritzis. Perspectives on software development environments. Lecture Notes in Computer Science, 685:586–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1993:OOD

[446] B. Wang, P. Hitchcock, and T. Holden. An object-oriented database approach for supporting hypertext. Lecture Notes in Computer Science, 685:601– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puigjaner:1993:EPP

[447] R. Puigjaner, A. Benzekri, S. Ayache, et al. Estimation process of performance constraints during the design of real-time and embedded systems. Lecture Notes in Computer Science, 685:629–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breu:1993:WDM

[448] M. Breu, S. Brinkkemper, M. Franckson, H. Habrias, T. Halpin, and P. Loucopoulos. Will IS development methods be completely incorporated in CASE tools in the future? Lecture Notes in Computer Science, 685:649–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leibovic:1993:BNC

[449] K. N. Leibovic. Biophysics of neural computation (invited paper). Lecture Notes in Computer Science, 686:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corbacho:1993:ILR

[450] F. J. Corbacho and M. A. Arbib. Integrated learning in Rana Computatrix. Lecture Notes in Computer Science, 686:12–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murciano:1993:MCV

[451] A. Murciano, J. Zamora, and M. Reviriego. A model for centering visual stimuli through adaptive value learning. Lecture Notes in Computer Science, 686:20–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andrade:1993:MDN

[452] M. A. Andrade and F. Moran. A model for the development of neurons selective to visual stimulus size. Lecture Notes in Computer Science, 686:24–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno-Diaz:1993:IRM

[453] R. Moreno-Diaz and O. Bolivar Toledo. An invariant representation mechanism after presynaptic inhibition. Lecture Notes in Computer Science, 686:30–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez-Andres:1993:PBC

[454] J. V. Sanchez-Andres and B. Soria. The pancreatic B-cell as a voltagecontrolled oscillator. Lecture Notes in Computer Science, 686:37–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoekstra:1993:ASD

[455] J. Hoekstra. Approximation of the solution of the dendritic cable equation by a small series of coupled differential equations. Lecture Notes in Computer Science, 686:43–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pelaez:1993:NNM

[456] J. Ropero Pelaez. A neural network model inspired in global appreciations about the thalamic reticular nucleus and cerebral cortex connectivity. Lecture Notes in Computer Science, 686:49–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mira:1993:TMR

[457] J. Mira, A. E. Delgado, J. R. Alvarez, A. P. de Madrid, and M. Santos. Towards more realistic self contained models of neurons: High-order, recurrence and local learning. Lecture Notes in Computer Science, 686:55–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scott:1993:MNR

[458] R. J. Scott. McCulloch’s neurons revisited. Lecture Notes in Computer Science, 686:63–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petkov:1993:BMA

[459] N. Petkov, P. Kruizinga, and T. Lourens. Biologically motivated approach to face recognition. Lecture Notes in Computer Science, 686:68–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vico:1993:LRP

[460] F. J. Vico, F. Sandoval, and J. Almaraz. Learning by reinforcement: A psychobiological model. Lecture Notes in Computer Science, 686:78–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aleksander:1993:NSM

[461] I. Aleksander and H. Morton. A neural state machine for iconic language representation. Lecture Notes in Computer Science, 686:84–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez-Moliner:1993:VBU

[462] J. Lopez-Moliner and J. M. Sopena. Variable binding using serial order in recurrent neural networks. Lecture Notes in Computer Science, 686:90–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castillo:1993:RIR

[463] F. Castillo, J. Cabestany, and J. M. Moreno. Region of influence (ROI) networks. model and implementation. Lecture Notes in Computer Science, 686: 96–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huning:1993:NSA

[464] H. Huning. A node splitting algorithm that reduces the number of connections in a Hamming distance classifying network. Lecture Notes in Computer Science, 686:102–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aligue:1993:HON

[465] F. J. Lopez Aligue, M. I. Acevedo Sotoca, and M. A. Jaramillo Moran. A high order neural model. Lecture Notes in Computer Science, 686:108– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bofill:1993:HON

[466] P. Bofill and C. Torras. Higher-order networks for the optimization of block designs. Lecture Notes in Computer Science, 686:114–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jutten:1993:NBC

[467] C. Jutten and P. Comon. Neural Bayesian classifier. Lecture Notes in Computer Science, 686:119–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gentric:1993:CMN

[468] P. Gentric and H. C. A. M. Withagen. Constructive methods for a new classifier based on a radial-basis-function neural network accelerated by a tree. Lecture Notes in Computer Science, 686:125–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lemarie:1993:PRR

[469] B. Lemarie. Practical realization of a radial basis function network for handwritten digit recognition. Lecture Notes in Computer Science, 686:131– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hubert:1993:DFP

[470] C. Hubert. Design of fully and partially connected random neural networks for pattern completion. Lecture Notes in Computer Science, 686:137–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alquezar:1993:RRR

[471] R. Alquezar and A. Sanfeliu. Representation and recognition of regular grammars by means of second-order recurrent neural networks. Lecture Notes in Computer Science, 686:143–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1993:CMS

[472] J. Santos and R. P. Otero. Connectionist models for syllabic recognition in the time domain. Lecture Notes in Computer Science, 686:149–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1993:SIA

[473] D. Liu and A. N. Michel. Sparsely interconnected artificial neural networks for associative memories. Lecture Notes in Computer Science, 686:155–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arenas:1993:DAN

[474] A. Arenas and C. J. Perez Vicente. Dynamic analysis of networks of neural oscillators. Lecture Notes in Computer Science, 686:161–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkitt:1993:OAN

[475] A. N. Burkitt. Optimised attractor neural networks with external inputs. Lecture Notes in Computer Science, 686:167–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weigl:1993:NOB

[476] K. Weigl and M. Berthod. Nonorthogonal bases and metric tensors: An application to artificial neural networks. Lecture Notes in Computer Science, 686:173–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marin:1993:GSD

[477] F. J. Marin and F. Sandoval. Genetic synthesis of discrete-time recurrent neural network. Lecture Notes in Computer Science, 686:179–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merelo:1993:OCL

[478] J. J. Merelo, M. Paton, A. Canas, A. Prieto, and F. Moran. Optimization of a competitive learning neural network by genetic algorithms. Lecture Notes in Computer Science, 686: 185–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ligomenides:1993:AMN

[479] P. A. Ligomenides. Adaptive models in neural networks. Lecture Notes in Computer Science, 686:193–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mannes:1993:SOG

[480] C. Mannes. Self-organizing grammar induction using a neural network model. Lecture Notes in Computer Science, 686:198–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tambouratzis:1993:RFE

[481] G. Tambouratzis and T. J. Stonham. The role of forgetting in efficient learning strategies for self-organising discriminator-based systems. Lecture Notes in Computer Science, 686:204– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castano:1993:SSR

[482] M. A. Castano, F. Casacuberta, and E. Vidal. Simulation of stochastic regular grammars through simple recurrent networks. Lecture Notes in Computer Science, 686:210–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grana:1993:LSC

[483] M. Grana, A. D’Anjou, F. X. Albizuri, F. J. Torrealdea, and M. C. Hernandez. Local stochastic competition and vector quantization. Lecture Notes in Computer Science, 686:216–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramirez:1993:MEC

[484] J. M. Ramirez. MHC — an evolutive connectionist model for hybrid training. Lecture Notes in Computer Science, 686:223–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aizenberg:1993:FCL

[485] N. N. Aizenberg and I. N. Aizenberg. Fast-convergence learning algorithms for multi-level and binary neurons and solution of some image processing problems. Lecture Notes in Computer Science, 686:230–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1993:IOR

[486] K. Ito, M. Hamamoto, J. Kamruzzaman, and Y. Kumagai. Invariant object recognition using Fahlman and Lebiere’s learning algorithm. Lecture Notes in Computer Science, 686:237– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamamoto:1993:RSC

[487] M. Hamamoto, K. Ito, J. Kamruzzaman, and Y. Kumagai. Realization of surjective correspondence in artificial neural network trained by Fahlman and Lebiere’s learning algorithm. Lecture Notes in Computer Science, 686:243–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slade:1993:BDR

[488] P. Slade and T. D. Gedeon. Bimodal distribution removal. Lecture Notes in Computer Science, 686:249–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guazzelli:1993:SAA

[489] A. Guazzelli, D. Barone, and E. C. de B. Carvalho Filho. A simplified Artmap architecture for real-time learning. Lecture Notes in Computer Science, 686:255–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langlois:1993:BLR

[490] T. Langlois and S. Canu. B-learning: A reinforcement learning algorithm, comparison with dynamic programming. Lecture Notes in Computer Science, 686:261–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cloete:1993:ICT

[491] I. Cloete and J. Ludik. Increased complexity training. Lecture Notes in Computer Science, 686:267–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno:1993:OLI

[492] J. M. Moreno, F. Castillo, and J. Cabestany. Optimized learning for improving the evolution of piecewise linear separation incremental algorithms. Lecture Notes in Computer Science, 686:272–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pelillo:1993:MPL

[493] M. Pelillo and A. M. Fanelli. A method of pruning layered feed-forward neural networks. Lecture Notes in Computer Science, 686:278–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crespo:1993:TDR

[494] J. L. Crespo and E. Mora. Tests of different regularization terms in small networks. Lecture Notes in Computer Science, 686:284–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yin:1993:DFS

[495] H. Yin and N. M. Allinson. On the distribution of feature space in self-organising mapping and convergence accelerating by a Kalman algorithm. Lecture Notes in Computer Science, 686:291–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin-Smith:1993:LAO

[496] P. Martin-Smith, F. J. Pelayo, A. Diaz, J. Ortega, and A. Prieto. A learning algorithm to obtain self-organizing maps using fixed neighbourhood Kohonen networks. Lecture Notes in Computer Science, 686:297–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cottrell:1993:ACT

[497] M. Cottrell, P. Letremy, and E. Roy. Analysing a contingency table with Kohonen maps: A factorial correspondence analysis. Lecture Notes in Computer Science, 686:305–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Der:1993:DSO

[498] R. Der and T. Villmann. Dynamics of self-organized feature mapping. Lecture Notes in Computer Science, 686: 312–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wann:1993:CSS

[499] C.-D. Wann and S. C. A. Thomopoulos. Comparative study of self-organizing neural networks. Lecture Notes in Computer Science, 686:316–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

White:1993:GGA

[500] D. White and P. Ligomenides. GAN-Net: A genetic algorithm for optimizing topology and weights in neural network design. Lecture Notes in Computer Science, 686:322–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demartines:1993:VQP

[501] P. Demartines and J. Herault. Vector quantization and projection neural network. Lecture Notes in Computer Science, 686:328–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perez:1993:CDL

[502] J.-C. Perez and E. Vidal. Constructive design of LVQ and DSM classifiers. Lecture Notes in Computer Science, 686:334–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verleysen:1993:LVC

[503] M. Verleysen, P. Thissen, and J.-D. Legat. Linear vector classification: An improvement on LVQ algorithms to create classes of patterns. Lecture Notes in Computer Science, 686:340– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1993:NGA

[504] Z. Wang. Non-greedy adaptive vector quantizers. Lecture Notes in Computer Science, 686:346–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treleaven:1993:HPE

[505] P. C. Treleaven and P. V. Rocha. Hybrid programming environments (invited paper). Lecture Notes in Computer Science, 686:351–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dreiseitl:1993:AGC

[506] S. Dreiseitl and D. Wang. Automatic generation of C++ code for neural network simulation. Lecture Notes in Computer Science, 686:358–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuentes:1993:UOO

[507] L. Fuentes, J. F. Aldana, and J. M. Troya. URANO: An object-oriented artificial neural network simulation tool. Lecture Notes in Computer Science, 686:364–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaudot:1993:RST

[508] W. Beaudot, P. Palagi, and J. Herault. Realistic simulation tool for early visual processing including space, time and colour data. Lecture Notes in Computer Science, 686:370–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burdorf:1993:LSS

[509] C. Burdorf. Language supported storage and reuse of persistent neural network objects. Lecture Notes in Computer Science, 686:376–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taylor:1993:FOE

[510] J. C. Taylor, M. L. Recce, and A. S. Mangat. Flexible operating environment for matrix based neurocomputers. Lecture Notes in Computer Science, 686:382–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filippi:1993:PIK

[511] E. Filippi and J. C. Lawson. A parallel implementation of Kohonen’s selforganizing maps on the smart neurocomputer. Lecture Notes in Computer Science, 686:388–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilke:1993:SNN

[512] P. Wilke. Simulation of neural networks in a distributed computing environment using NeuroGraph. Lecture Notes in Computer Science, 686:394– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alba:1993:FAA

[513] E. Alba, J. F. Aldana, and J. M. Troya. Full automatic ANN design: A genetic approach. Lecture Notes in Computer Science, 686:399–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelCorso:1993:HIA

[514] D. Del Corso. Hardware implementations of artificial neural networks (invited paper). Lecture Notes in Computer Science, 686:405–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiaberge:1993:NNC

[515] M. Chiaberge, D. Del Corso, F. Gregoretti, and L. M. Reyneri. A neural network chip using CPWM modulation. Lecture Notes in Computer Science, 686:420–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carrabina:1993:HIN

[516] J. Carrabina, F. Lisa, V. Gaitan, L. Garrido, and E. Valderrama. Hardware implementation of a neural network for high energy physics application. Lecture Notes in Computer Science, 686:426–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ortega:1993:MAP

[517] J. Ortega, F. J. Pelayo, A. Prieto, B. Pino, and C. G. Puntonet. MapA: An array processor architecture for neural networks. Lecture Notes in Computer Science, 686:432–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Regueiro:1993:LCM

[518] C. V. Regueiro, S. Barro, and A. Yanez. Limitation of connectionism in MLP. Lecture Notes in Computer Science, 686:441–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nigri:1993:HLS

[519] M. E. Nigri and P. C. Treleaven. High level synthesis of neural network chips. Lecture Notes in Computer Science, 686:448–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sumpter:1993:NNS

[520] B. G. Sumpter, R. E. Guenther, C. Halloy, C. Getino, and D. W. Noid. Neural network simulations on massively parallel computers: Applications in chemical physics. Lecture Notes in Computer Science, 686:454–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naylor:1993:MBA

[521] D. Naylor and S. Jones. A model based approach to the performance analysis of multi-layer networks realised in linear systolic arrays. Lecture Notes in Computer Science, 686:459–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christodoulou:1993:TNL

[522] C. Christodoulou, G. Bugmann, T. G. Clarkson, and J. G. Taylor. The temporal noisy-leaky integrator neuron with additional inhibitory inputs. Lecture Notes in Computer Science, 686:465– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarkson:1993:ASL

[523] T. G. Clarkson and C. K. Ng. Architectures for self-learning neural network modules. Lecture Notes in Computer Science, 686:471–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vellasco:1993:GNA

[524] M. M. B. R. Vellasco and P. C. Treleaven. The generic neuron architectural framework for the automatic generation of ASICs. Lecture Notes in Computer Science, 686:476–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacheco:1993:RAS

[525] M. Pacheco and P. C. Treleaven. A RISC architecture to support neural net simulation. Lecture Notes in Computer Science, 686:482–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruping:1993:HDS

[526] S. Ruping, U. Ruckert, and K. Goser. Hardware design for self-organizing feature maps with binary input vectors. Lecture Notes in Computer Science, 686:488–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lagunas:1993:KSP

[527] M. A. Lagunas, A. Perez-Neira, M. Najar, and A. Pages. The Kolmogorov signal processor (invited paper). Lecture Notes in Computer Science, 686: 494–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joya:1993:PIP

[528] G. Joya and F. Sandoval. Projectivity invariant pattern recognition with high-order neural networks. Lecture Notes in Computer Science, 686:513– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smieja:1993:RIA

[529] F. J. Smieja. Rejection of incorrect answers from a neural net classifier. Lecture Notes in Computer Science, 686:519–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pantaleon-Prieto:1993:NTS

[530] C. J. Pantaleon-Prieto and A. R. Figueiras-Vidal. Nonlinear time series modeling by competitive segmentation of state space. Lecture Notes in Computer Science, 686:525–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adam:1993:IPN

[531] O. Adam, J.-L. Zarader, and M. Milgram. Identification and prediction of non-linear models with recurrent neural network. Lecture Notes in Computer Science, 686:531–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodriguez:1993:UUN

[532] M. J. Rodriguez, F. del Pozo, and M. T. Arredondo. Use of unsupervised neural networks for classification of blood pressure time series. Lecture Notes in Computer Science, 686:536–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Rodriguez:1993:AAN

[533] J. J. Fernandez-Rodriguez, A. Canas, E. Roca, F. J. Pelayo, J. Fernandez-Mena, and A. Prieto. Application of artificial neural networks to chest image classification. Lecture Notes in Computer Science, 686:542–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuya:1993:CSO

[534] J. Tuya, E. Arias, L. Sanchez, and J. A. Corrales. Combination of selforganizing maps and multilayer perceptrons for speaker independent isolated word recognition. Lecture Notes in Computer Science, 686:550–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yahiaoui:1993:IAN

[535] G. Yahiaoui and B. Borocco. An industrial application of neural networks to natural textures classification. Lecture Notes in Computer Science, 686:556–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garrido:1993:ULN

[536] L. Garrido, V. Gaitan, M. Serra-Ricart, and X. Calbet. Use of a layered neural nets as a display method for N-dimensional distributions. Lecture Notes in Computer Science, 686:563– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lecourtier:1993:MMV

[537] Y. Lecourtier, B. Dorizzi, P. Sebire, and A. Ennaji. MLP modular versus YPREL classifiers. Lecture Notes in Computer Science, 686:569–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patinel:1993:HMH

[538] J. Patinel, G. Leone, and M. Milgram. How many hidden neurons are needed to recognize a symmetrical pattern? Lecture Notes in Computer Science, 686:575–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cavalieri:1993:HNN

[539] S. Cavalieri, A. Di Stefano, and O. Mirabella. Hopfield neural network for routing. Lecture Notes in Computer Science, 686:583–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia-Lopera:1993:NNR

[540] A. Garcia-Lopera, A. Diaz Estrella, F. Garcia Oller, and F. Sandoval. Neural network routing controller for communication parallel multistage interconnection networks. Lecture Notes in Computer Science, 686:593–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minot:1993:ARU

[541] J. Minot. Adaptive routing using cellular automata. Lecture Notes in Computer Science, 686:599–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cid-Sueiro:1993:OBE

[542] J. Cid-Sueiro, L. Weruaga, and A. R. Figueiras-Vidal. Optimal blind equalization of Gaussian channels. Lecture Notes in Computer Science, 686:605– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cammarata:1993:NPU

[543] G. Cammarata, S. Cavalieri, A. Fichera, and L. Marletta. Noise prediction in urban traffic by a neural approach. Lecture Notes in Computer Science, 686:611–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sako:1993:CAC

[544] H. Sako and H. J. Avi-Itzhak. A connectionist approach to the correspondence problem in computer vision. Lecture Notes in Computer Science, 686:620–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Natowicz:1993:SOF

[545] R. Natowicz and R. Sokol. Selforganizing feature maps for image segmentation. Lecture Notes in Computer Science, 686:626–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freisleben:1993:RFI

[546] B. Freisleben, J.-H. Greve, and J. Lober. Recognition of fractal images using a neural network. Lecture Notes in Computer Science, 686:632– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lisa:1993:FFN

[547] F. Lisa, J. Carrabina, C. Perez-Vicente, N. Avellana, and E. Valderrama. Feed forward network for vehicle license character recognition. Lecture Notes in Computer Science, 686:638–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyauchi:1993:IOF

[548] M. Miyauchi, M. Seki, A. Watanabe, and A. Miyauchi. Interpretation of optical flow through complex neural network. Lecture Notes in Computer Science, 686:645–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cabello:1993:CIS

[549] D. Cabello, M. G. Penedo, S. Barro, J. M. Pardo, and J. Heras. CT image segmentation by self-organizing learning. Lecture Notes in Computer Science, 686:651–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mosquera:1993:TIS

[550] A. Mosquera, D. Cabello, M. J. Carreira, and M. G. Penedo. Texture image segmentation using a modified Hopfield network. Lecture Notes in Computer Science, 686:657–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freisleben:1993:ICS

[551] B. Freisleben and M. Mengel. Image compression with self-organizing networks. Lecture Notes in Computer Science, 686:664–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bahrami:1993:NND

[552] M. Bahrami. Neural networks as direct adaptive controllers. Lecture Notes in Computer Science, 686:670–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barret:1993:NAC

[553] C. Barret, M. Houkari, P. Meyne, J. M. Martinez, A. Garassino, and P. Tormo. A neural adaptive controller for a turbofan exhaust nozzle. Lecture Notes in Computer Science, 686:676–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bulsari:1993:FFN

[554] A. Bulsari, B. Saxen, and H. Saxen. Feed-forward neural networks for bioreactor control. Lecture Notes in Computer Science, 686:682–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borland:1993:LNP

[555] L. Borland and H. Haken. Learning networks for process identification and associative action. Lecture Notes in Computer Science, 686:688–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pimentel:1993:LPE

[556] J. R. Pimentel, D. Gachet, L. Moreno, and M. A. Salichs. On-line performance enhancement of a behavioral neural network controller. Lecture Notes in Computer Science, 686:694–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palmer:1993:AIC

[557] R. P. Palmer and P. A. Rounce. An architecture for implementing control and signal processing neural networks. Lecture Notes in Computer Science, 686:702–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mrsic-Flogel:1993:PAP

[558] J. Mrsic-Flogel. Planlite: Adaptive planning using weightless systems. Lecture Notes in Computer Science, 686: 708–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Margarita:1993:SPV

[559] S. Margarita and A. Beltratti. Stock prices and volume in an artificial adaptive stock market. Lecture Notes in Computer Science, 686:714–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deMiguel:1993:AFA

[560] L. J. de Miguel, E. Revilla, J. M. Rodriguez, and J. M. Cano. Application of the fuzzy Artmap neural network architecture to bank failure predictions. Lecture Notes in Computer Science, 686:720–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ros:1993:CNN

[561] F. Ros, A. Brons, F. Sevila, G. Rabatel, and C. Touzet. Combination of neural network and statistical methods for sensory evaluation of biological products: On-line beauty selection of flowers. Lecture Notes in Computer Science, 686:726–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crestani:1993:AIR

[562] F. Crestani. An adaptive information retrieval system based on neural networks. Lecture Notes in Computer Science, 686:732–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clochon:1993:SPE

[563] P. Clochon, D. Clarencon, R. Caterini, and V. Roman. Software pattern EEG recognition after a wavelet transform by a neural network. Lecture Notes in Computer Science, 686:738–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bookstein:1993:FSD

[564] F. L. Bookstein and W. D. K. Green. A feature space for derivatives of deformations. Lecture Notes in Computer Science, 687:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nastar:1993:NRM

[565] C. Nastar and N. Ayache. Non-rigid motion analysis in medical images: a physically based approach. Lecture Notes in Computer Science, 687:17–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cootes:1993:UAS

[566] T. F. Cootes, A. Hill, C. J. Taylor, and J. Haslam. The use of active shape models for locating structures in medical images. Lecture Notes in Computer Science, 687:33–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Worring:1993:PFB

[567] M. Worring, A. W. M. Smeulders, L. H. Staib, and J. S. Duncan. Parameterized feasible boundaries in gradient vector fields. Lecture Notes in Computer Science, 687:48–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vemuri:1993:MRS

[568] B. C. Vemuri, A. Radisavljevic, and C. M. Leonard. Multi-resolution stochastic 3D shape models for image segmentation. Lecture Notes in Computer Science, 687:62–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Romeny:1993:HOD

[569] B. M. ter Haar Romeny, L. M. J. Florack, A. H. Salden, and M. A. Viergever. Higher order differential structure of images. Lecture Notes in Computer Science, 687:77–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerig:1993:SDD

[570] G. Gerig, Th. Koller, G. Szekely, Ch. Brechbuhler, and O. Kubler. Symbolic description of 3-D structures applied to cerebral vessel tree obtained from MR angiography volume data. Lecture Notes in Computer Science, 687:94–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morse:1993:MMA

[571] B. S. Morse, S. M. Pizer, and A. Liu. Multiscale medial analysis of medical images. Lecture Notes in Computer Science, 687:112–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tagare:1993:ASR

[572] H. D. Tagare, F. Vos, C. C. Jaffe, and J. S. Duncan. Arrangement: A spatial relation comparing part embeddings and its use in medical image comparisons. Lecture Notes in Computer Science, 687:132–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whitaker:1993:CFS

[573] R. T. Whitaker. Characterizing first and second order patches using geometry-limited diffusion. Lecture Notes in Computer Science, 687:149– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schubert:1993:SKR

[574] R. Schubert, K. H. Hohne, A. Pommert, M. Riemer, Th. Schiemann, and U. Tiede. Spatial knowledge representation for visualization of human anatomy and function. Lecture Notes in Computer Science, 687:168–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1993:SAM

[575] D. L. G. Hill, D. J. Hawkes, N. Harrison, and C. F. Ruff. A strategy for automated multimodality image registration incorporating anatomical knowledge and imager characteristics. Lecture Notes in Computer Science, 687:182–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:1993:MBR

[576] G. P. Robinson, A. C. F. Colchester, and L. D. Griffin. Model-based recognition of anatomical objects from medical images. Lecture Notes in Computer Science, 687:197–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haring:1993:MAI

[577] S. Haring, M. A. Viergever, and J. N. Kok. A multiscale approach to image segmentation using Kohonen networks. Lecture Notes in Computer Science, 687:212–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Worth:1993:SMR

[578] A. J. Worth and D. N. Kennedy. Segmentation of magnetic resonance brain images using analog constraint satisfaction neural networks. Lecture Notes in Computer Science, 687:225–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulson:1993:FNL

[579] K. Paulson, W. Lionheart, and M. Pidcock. Fast, non-linear inversion for electrical impedance tomography. Lecture Notes in Computer Science, 687:244– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arridge:1993:IMO

[580] S. R. Arridge and M. Schweiger. Inverse methods for optical tomography. Lecture Notes in Computer Science, 687:259–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cao:1993:FGA

[581] Y. Cao and D. N. Levin. Featureguided acquisition and reconstruction of MR images. Lecture Notes in Computer Science, 687:278–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1993:RTD

[582] W. E. Smith, J. K. Riek, and A. Murat Tekalp. Reconstruction of a threedimensional volume from a motioncorrupted two-dimensional data set in magnetic resonance imaging. Lecture Notes in Computer Science, 687:293– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnson:1993:FIS

[583] V. E. Johnson. A framework for incorporating structural prior information into the estimation of medical images. Lecture Notes in Computer Science, 687:307–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gindi:1993:BRE

[584] G. Gindi, A. Rangarajan, M. Lee, P. J. Hong, and I. G. Zubal. Bayesian reconstruction for emission tomography via deterministic annealing. Lecture Notes in Computer Science, 687:322– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pan:1993:ACP

[585] X. Pan, C.-T. Chen, J. N. Aarsvold, and W. H. Wong. Analytical considerations of photon attenuation and system response function in SPECT reconstruction. Lecture Notes in Computer Science, 687:339–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1993:MIR

[586] Z. Wu. MAP image reconstruction using wavelet decomposition. Lecture Notes in Computer Science, 687:354– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fessler:1993:TRU

[587] J. A. Fessler. Tomographic reconstruction using information-weighted spline smoothing. Lecture Notes in Computer Science, 687:372–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1993:DFB

[588] C. Wu, M. N. Wernick, and C.-T. Chen. A 3-D filtered-backprojection reconstruction algorithm for combined parallel- and cone-beam SPECT data. Lecture Notes in Computer Science, 687:387–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benali:1993:FFA

[589] H. Benali, I. Buvat, F. Frouin, J. P. Bazin, and R. Di Paola. Foundations of factor analysis of medical image sequences: A unified approach and some practical implications. Lecture Notes in Computer Science, 687:401–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samal:1993:BIP

[590] M. Samal, M. Karny, and D. Zahalka. Bayesian identification of a physiological model in dynamic scintigraphic data. Lecture Notes in Computer Science, 687:422–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shields:1993:IRI

[591] K. Shields, D. C. Barber, and S. B. Sherriff. Image registration for the investigation of atherosclerotic plaque movement. Lecture Notes in Computer Science, 687:438–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoo:1993:USP

[592] T. S. Yoo and J. M. Coggins. Using statistical pattern recognition techniques to control variable conductance diffusion. Lecture Notes in Computer Science, 687:459–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karssemeijer:1993:ANE

[593] N. Karssemeijer. Adaptive noise equalization and image analysis in mammography. Lecture Notes in Computer Science, 687:472–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vandermeulen:1993:CVC

[594] D. Vandermeulen, R. Verbeeck, L. Berben, P. Suetens, and G. Marchal. Continuous voxel classification by stochastic relaxation: Theory and application to MR imaging and MR angiography. Lecture Notes in Computer Science, 687:487–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1993:MGP

[595] R. F. Wagner, D. G. Brown, J.-P. Guedon, K. J. Myers, and K. A. Wear. Multivariate Gaussian pattern classification: Effects of finite sample size and the addition of correlated or noisy features on summary measures of goodness. Lecture Notes in Computer Science, 687:507–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swensson:1993:MDL

[596] R. G. Swensson. Measuring detection and localization performance. Lecture Notes in Computer Science, 687:525– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Insana:1993:MEE

[597] M. F. Insana and T. J. Hall. Methods for estimating the efficiency of human and computational observers in ultrasonography. Lecture Notes in Computer Science, 687:542–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1993:GFB

[598] M. P. Anderson, M. H. Loew, and D. G. Brown. Gabor function based medical image compression. Lecture Notes in Computer Science, 687:553–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minker:1993:EDL

[599] J. Minker and C. Ruiz. On extended disjunctive logic programs. Lecture Notes in Computer Science, 689:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu:1993:MFS

[600] H. Chu and D. A. Plaisted. Model finding stategies in semantically guided instance-based theorem proving. Lecture Notes in Computer Science, 689: 19–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busch:1993:ETV

[601] D. R. Busch. An expressive threevalued logic with two negations. Lecture Notes in Computer Science, 689: 29–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Posegga:1993:CPS

[602] J. Posegga. Compiling proof search in semantic tableaux. Lecture Notes in Computer Science, 689:39–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahnle:1993:SCF

[603] R. Hahnle. Short CNF in finetelyvalued logics. Lecture Notes in Computer Science, 689:49–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordano:1993:DVD

[604] L. Giordano. Defining variants of default logic: a modal approach. Lecture Notes in Computer Science, 689:59–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shue:1993:AHS

[605] L.-Y. Shue and R. Zamani. An admissible heuristic search algorithm. Lecture Notes in Computer Science, 689:69–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1993:BES

[606] S.-J. Lee and C.-H. Wu. Building an expert system language interpreter with the rule network technique. Lecture Notes in Computer Science, 689:76–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valiente:1993:IDC

[607] G. Valiente. Input-driven control of rule-based expert systems. Lecture Notes in Computer Science, 689:86–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez:1993:CBP

[608] B. Lopez and E. Plaza. Case-based planning for medical diagnosis. Lecture Notes in Computer Science, 689:96–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klut:1993:MME

[609] J. P. Klut and J. H. P. Eloff. MethoDex: A methodology for expert systems development. Lecture Notes in Computer Science, 689:106–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bry:1993:TID

[610] F. Bry. Towards intelligent databases. Lecture Notes in Computer Science, 689:116–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padgham:1993:CCN

[611] L. Padgham and B. Nebel. Combining classification and nonmonotonic inheritance reasoning: A first step. Lecture Notes in Computer Science, 689: 132–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rasiowa:1993:MPS

[612] H. Rasiowa and V. W. Marek. Mechanical proof systems for logic II, consensus programs and their processing (extended abstract). Lecture Notes in Computer Science, 689:142–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1993:LOK

[613] J. Chen. The logic of only knowing as a unified framework for nonmonotonic reasoning. Lecture Notes in Computer Science, 689:152–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lambrix:1993:TLI

[614] P. Lambrix and R. Ronnquist. Terminological logic involving time and evolution: A preliminary report. Lecture Notes in Computer Science, 689:162– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orman:1993:KME

[615] L. V. Orman. Knowledge management by example. Lecture Notes in Computer Science, 689:172–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dewan:1993:SRL

[616] H. M. Dewan and S. J. Stolfo. System reorganization and load balancing of parallel database rule processing. Lecture Notes in Computer Science, 689:186–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaasterland:1993:USI

[617] T. Gaasterland and J. Lobo. Using semantic information for processing negation and disjunction in logic programs. Lecture Notes in Computer Science, 689:198–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosc:1993:ISO

[618] P. Bosc, L. Lietard, and O. Pivert. On the interpretation of set-oriented fuzzy quantified queries and their evaluation in a database management system. Lecture Notes in Computer Science, 689:209–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lowry:1993:MKB

[619] M. R. Lowry. Methodologies for knowledge-based software engineering. Lecture Notes in Computer Science, 689:219–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leone:1993:ULP

[620] N. Leone, L. Palopoli, and M. Romeo. Updating logic programs. Lecture Notes in Computer Science, 689:235– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robertson:1993:EPR

[621] D. Robertson, J. Agusti, J. Hesketh, and J. Levy. Expressing program requirements using refinement lattices. Lecture Notes in Computer Science, 689:245–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chadha:1993:FLC

[622] R. Chadha and D. A. Plaisted. Finding logical consequences using unskolemization. Lecture Notes in Computer Science, 689:255–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajasekar:1993:CES

[623] A. Rajasekar. Controlled explanation systems. Lecture Notes in Computer Science, 689:265–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murray:1993:SFL

[624] N. V. Murray and E. Rosenthal. Signed formulas: A liftable meta-logic for multiple-valued logics. Lecture Notes in Computer Science, 689:275–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tano:1993:NDC

[625] S. Tano, W. Okamoto, and T. Iwatani. New design concepts for the FLINS-Fuzzy lingual system: Text-based and fuzzy-centered architectures. Lecture Notes in Computer Science, 689:285– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skowron:1993:BRD

[626] A. Skowron. Boolean reasoning for decision rules generation. Lecture Notes in Computer Science, 689:295– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chau:1993:ULE

[627] C. W. R. Chau, P. Lingras, and S. K. M. Wong. Upper and lower entropies of belief functions using compatible probability functions. Lecture Notes in Computer Science, 689:306– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liau:1993:RAH

[628] C. J. Liau and B. I-P. Lin. Reasoning about higher order uncertainty in possibilistic logic. Lecture Notes in Computer Science, 689:316–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauszer:1993:AMK

[629] C. M. Rauszer. Approximation methods for knowledge representation systems. Lecture Notes in Computer Science, 689:326–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ljung:1993:MIS

[630] L. Ljung. Modelling of industrial systems. Lecture Notes in Computer Science, 689:338–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puget:1993:SSC

[631] J.-F. Puget. On the satisfiability of symmetrical constrained satisfaction problems. Lecture Notes in Computer Science, 689:350–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1993:LRC

[632] A. L. Brown, S. Mantha, and T. Wakayama. A logical reconstruction of constraint relaxation hierarchies in logic programming. Lecture Notes in Computer Science, 689:362–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berlandier:1993:PEB

[633] P. Berlandier. A performance evaluation of backtrack-bounded search methods for N-ary constraint networks. Lecture Notes in Computer Science, 689:375–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1993:FDC

[634] M. A. Meyer and J. P. Muller. Finite domain consistency techniques: Their combination and application in computer-aided process planning. Lecture Notes in Computer Science, 689: 385–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imam:1993:SDT

[635] I. F. Imam and R. S. Michalski. Should decision trees be learned from examples or from decision rules? Lecture Notes in Computer Science, 689:395– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lounis:1993:IML

[636] H. Lounis. Integrating machinelearning techniques in knowledge-based systems verification. Lecture Notes in Computer Science, 689:405–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bagai:1993:ATG

[637] R. Bagai, V. Shanbhogue, J. M. Zytkow, and S. C. Chou. Automatic theorem generation in plane geometry. Lecture Notes in Computer Science, 689:415–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordana:1993:LSR

[638] A. Giordana, L. Saitta, and C. Baroglio. Learning simple recursive theories. Lecture Notes in Computer Science, 689:425–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeRaedt:1993:MFI

[639] L. De Raedt and N. Lavrac. The many faces of inductive logic programming. Lecture Notes in Computer Science, 689:435–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bateman:1993:CMD

[640] M. Bateman, S. Martin, and A. Slade. Consensus: A method for the development of distributed intelligent systems. Lecture Notes in Computer Science, 689:450–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gan:1993:SFM

[641] H. Gan. Script and frame: Mixed natural language understanding system with default theory. Lecture Notes in Computer Science, 689:466–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franova:1993:CMM

[642] M. Franova, Y. Kodratoff, and M. Gross. Constructive matching methodology: Formally creative or intelligent inductive theorem proving? Lecture Notes in Computer Science, 689:476–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosz:1993:RKU

[643] G. Grosz and C. Rolland. Representing the knowledge used during the requirement engineering activity with generic structures. Lecture Notes in Computer Science, 689:486–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caselli:1993:DPE

[644] S. Caselli, A. Natali, and F. Zanichelli. Development of a programming environment for intelligent robotics. Lecture Notes in Computer Science, 689: 496–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaerf:1993:CIC

[645] A. Schaerf. On the complexity of the instance checking problem in concept languages with existential quantification. Lecture Notes in Computer Science, 689:508–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambroszkiewicz:1993:MK

[646] S. Ambroszkiewicz. Mutual knowledge. Lecture Notes in Computer Science, 689:518–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thirunarayan:1993:EEI

[647] K. Thirunarayan. Expressive extensions to inheritance networks. Lecture Notes in Computer Science, 689:528– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bittencourt:1993:CSC

[648] G. Bittencourt. A connectionistsymbolic cognitive model. Lecture Notes in Computer Science, 689:538– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiManzo:1993:MCS

[649] M. Di Manzo and E. Giunchiglia. Multi-context systems as a tool to model temporal evolution. Lecture Notes in Computer Science, 689:548– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandewall:1993:SAT

[650] E. Sandewall. Systematic assessment of temporal reasoning methods for use in autonomous agents. Lecture Notes in Computer Science, 689:558–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klauck:1993:GGG

[651] Ch. Klauck and J. Schwagereit. GGD: Graph grammar developer for features in CAD/CAM. Lecture Notes in Computer Science, 689:571–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1993:KBA

[652] K. Wang. A knowledge-based approach to group analysis in automated manufacturing systems. Lecture Notes in Computer Science, 689:581–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu:1993:CSA

[653] B.-T. B. Chu. Center: A system architecture for matching design and manufacturing. Lecture Notes in Computer Science, 689:591–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sobolewski:1993:KBS

[654] M. Sobolewski. Knowledge-based system integration in a concurrent engineering environment. Lecture Notes in Computer Science, 689:601–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charlton:1993:RSP

[655] P. Charlton. A reflective strategic problem solving model. Lecture Notes in Computer Science, 689:612–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wuthrich:1993:LRU

[656] B. Wuthrich. On the learning of rule uncertainties and their integration into probabilistic knowledge bases. Lecture Notes in Computer Science, 689:622– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zembowicz:1993:RFD

[657] R. Zembowicz and J. M. Zytkow. Recognition of functional dependencies in data. Lecture Notes in Computer Science, 689:632–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slowinski:1993:RSL

[658] R. Slowinski. Rough set learning of preferential attitude in multi-criteria decision making. Lecture Notes in Computer Science, 689:642–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freeston:1993:BCI

[659] Michael Freeston. Begriffsverzeichnis: A concept index. Lecture Notes in Computer Science, 696:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaniolo:1993:UAD

[660] Carlo Zaniolo. On the unification of active databases and deductive databases. Lecture Notes in Computer Science, 696:23–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goble:1993:SCM

[661] Carole A. Goble, Andrzej Glowinski, and Keith G. Jeffery. Semantic constraints in a medical information system. Lecture Notes in Computer Science, 696:40–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castellanos:1993:MSE

[662] Malu Castellanos. A methodology for semantically enriching interoperable databases. Lecture Notes in Computer Science, 696:58–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ringwood:1993:DDT

[663] Daniel Ringwood and Graem Cohen. Distributed databases tied with StrIng. Lecture Notes in Computer Science, 696:76–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barclay:1993:VO

[664] Peter J. Barclay and Jessie B. Kennedy. Viewing objects. Lecture Notes in Computer Science, 696:93–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carnduff:1993:FMT

[665] T. W. Carnduff and W. A. Gray. Function materialization through object versioning in object-oriented databases. Lecture Notes in Computer Science, 696:111–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hohenstein:1993:CDI

[666] Uwe Hohenstein and Erik Odberg. A C++ database interface based on the entity-relationship approach. Lecture Notes in Computer Science, 696:129– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tagg:1993:OOD

[667] Roger Tagg and Benny Liew. Objectoriented database methodology — state of the art. Lecture Notes in Computer Science, 696:147–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sunderraman:1993:DDC

[668] Rajshekhar Sunderraman. Deductive databases with conditional facts. Lecture Notes in Computer Science, 696: 162–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandes:1993:DOO

[669] Alvaro A. A. Fernandes, Maria L. Barja, Norman W. Paton, and M. Howard Williams. A deductive object-oriented database for data intensive application development. Lecture Notes in Computer Science, 696: 176–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singleton:1993:SRF

[670] Paul Singleton and O. Pearl Brereton. Storage and retrieval of first-order terms using a relational database. Lecture Notes in Computer Science, 696: 199–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mcbrien:1993:PIH

[671] Peter Mcbrien. Principles of implementing historical databases in RDBMS. Lecture Notes in Computer Science, 696:220–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reddi:1993:ICE

[672] Swarup Reddi. Integrity constraint enforcement in the functional database language PFL. Lecture Notes in Computer Science, 696:238–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eaglestone:1993:IVM

[673] Barry Eaglestone, Geoff Davies, Mick Ridley, and Nigel Hulley. Implementation of a version model for artists using extended relational technology. Lecture Notes in Computer Science, 696:258– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ait-Kaci:1993:EDO

[674] Hassan Ait-Kaci and Andreas Podelski. Entailment and disentailment of ordersorted feature constraints. Lecture Notes in Computer Science, 698:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1993:CED

[675] Grigoris Antoniou, Elmar Langetepe, and Volker Sperschneider. Computing extensions of default logic — preliminary report. Lecture Notes in Computer Science, 698:19–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barklund:1993:PAB

[676] Jonas Barklund and Johan Bevemyr. Prolog with arrays and bounded quantifications. Lecture Notes in Computer Science, 698:28–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barth:1993:LIE

[677] Peter Barth. Linear 0-1 inequalities and extended clauses. Lecture Notes in Computer Science, 698:40–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruning:1993:SSP

[678] Stefan Bruning. Search space pruning by checking dynamic term growth. Lecture Notes in Computer Science, 698: 52–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiu:1993:PSS

[679] Fai Chau Hiu. A proof search system for a modal substructural logic based on labelled deductive systems. Lecture Notes in Computer Science, 698:64–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chebotarev:1993:CCA

[680] Anatoli N. Chebotarev and Marina K. Morokhovets. Consistency checking of automata functional specifications. Lecture Notes in Computer Science, 698:76–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corsini:1993:YAA

[681] Marc-Michel Corsini, Alain Griffault, and Antoine Rauzy. Yet another application for toupie: Verification of mutual exclusion algorithms. Lecture Notes in Computer Science, 698:86–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davison:1993:PDT

[682] Andrew Davison. Parsing with DCG-Terms. Lecture Notes in Computer Science, 698:98–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1993:FOR

[683] Uwe Egly. A first order resolution calculus with symmetries. Lecture Notes in Computer Science, 698:110– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fermuller:1993:OPR

[684] Christian G. Fermuller and Gernot Salzer. Ordered paramodulation and resolution as decision procedure. Lecture Notes in Computer Science, 698: 122–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

File:1993:SAP

[685] Gilberto File and Sabina Rossi. Static analysis of Prolog with cut. Lecture Notes in Computer Science, 698:134– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gardneri:1993:NTT

[686] Philippa Gardneri. A new type theory for representing logics. Lecture Notes in Computer Science, 698:146–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahnle:1993:VSL

[687] Reiner Hahnle and Werner Kernig. Verification of switch-level designs with many-valued logic. Lecture Notes in Computer Science, 698:158–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hibti:1993:DHT

[688] M. Hibti, H. Lombardi, and B. Legeard. Deciding in HFS-Theory via linear integer programming. Lecture Notes in Computer Science, 698:170– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1993:CTL

[689] Patricia M. Hill. The completion of typed logic programs and SLDNF-Resolution. Lecture Notes in Computer Science, 698:182–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manning:1993:IVH

[690] Alistair Manning, Andrew Ireland, and Alan Bundy. Increasing the versatility of heuristic based theorem provers. Lecture Notes in Computer Science, 698: 194–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massey:1993:SPL

[691] Bart C. Massey and Evan Tick. Sequentialization of parallel logic programs with mode analysis. Lecture Notes in Computer Science, 698:205– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1993:REM

[692] Klaus Mayr. Refinements and extensions of model elimination. Lecture Notes in Computer Science, 698:217– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappel:1993:ESB

[693] Angelica Maria Kappel. Executable specifications based on dynamic algebras. Lecture Notes in Computer Science, 698:229–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNivelle:1993:GRP

[694] H. De Nivelle. Generic resolution in propositional modal systems. Lecture Notes in Computer Science, 698:241– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohlbach:1993:OTM

[695] Hans Jurgen Ohlbach. Optimized translation of multi modal logic into predicate logic. Lecture Notes in Computer Science, 698:253–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panitz:1993:DRC

[696] Sven Eric Panitz. Default reasoning with a constraint resolution principle. Lecture Notes in Computer Science, 698:265–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramesh:1993:NCD

[697] Anavai Ramesh and Neil V. Murray. Non-clausal deductive techniques for computing prime implicants and prime implicates. Lecture Notes in Computer Science, 698:277–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt-Schaub:1993:UUO

[698] Manfred Schmidt-Schaub. Unification under one-sided distributivity with a multiplicative unit. Lecture Notes in Computer Science, 698:289–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Socher-Ambrosius:1993:UOS

[699] Rolf Socher-Ambrosius. Unification in order-sorted logic with term declarations. Lecture Notes in Computer Science, 698:301–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solnon:1993:EIH

[700] Christine Solnon and Michel Rueher. Extracting inheritance hierarchies from Prolog programs: A system based on the inference of type relations. Lecture Notes in Computer Science, 698: 309–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutcliffe:1993:CMA

[701] Geoff Sutcliffe. A comparison of mechanisms for avoiding repetition of subdeductions in chain format linear deduction systems. Lecture Notes in Computer Science, 698:321–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1993:NPE

[702] Gerd Wagner. Neutralization and preemption in extended logic programs. Lecture Notes in Computer Science, 698:333–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1993:MSA

[703] Matthias Baaz, Christian G. Fermuller, Arie Ovrutcki, and Richard Zach. Multlog: A system for axiomatizing many-valued logics (system description). Lecture Notes in Computer Science, 698:345–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galmiche:1993:SSP

[704] D. Galmiche and O. Hermann. SKIL: A system for programming with proofs (system description). Lecture Notes in Computer Science, 698:348–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1993:RAR

[705] John Harrison and Laurent Thery. Reasoning about the reals: The marriage of HOL and Maple. Lecture Notes in Computer Science, 698:351–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yexuan:1993:SDL

[706] Gui Yexuan and Mitsuhiro Okada. System description of LAMBDALG — A higher order algebraic specification language. Lecture Notes in Computer Science, 698:354–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nuallain:1993:MM

[707] Breanndan O. Nuallain. Mixing Metafor. Lecture Notes in Computer Science, 698:357–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soloviev:1993:CAS

[708] Sergei V. Soloviev. A complete axiom system for isomorphism of types in closed categories. Lecture Notes in Computer Science, 698:360–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wegner:1993:RMC

[709] Peter Wegner. Reasoning, modeling, and component-based technology. Lecture Notes in Computer Science, 698: 372–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sowa:1993:RDL

[710] John F. Sowa. Relating diagrams to logic. Lecture Notes in Computer Science, 699:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaines:1993:RDL

[711] Brian R. Gaines. Representation, discourse, logic and truth: situating knowledge technology. Lecture Notes in Computer Science, 699:36–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barwise:1993:HR

[712] Jon Barwise. Heterogeneous reasoning. Lecture Notes in Computer Science, 699:64–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biebow:1993:CBC

[713] B. Biebow and G. Chaty. A comparison between conceptual graphs and KL-one. Lecture Notes in Computer Science, 699:75–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mineau:1993:TDO

[714] G. W. Mineau. The term definition operators of ontolingua and of the conceptual graph formalism: a comparison. Lecture Notes in Computer Science, 699:90–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bowen:1993:ICG

[715] B. Bowen and P. Kocura. Implementing conceptual graphs in a RDBMS. Lecture Notes in Computer Science, 699:106–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Creasy:1993:CGA

[716] P. Creasy and G. Ellis. A conceptual graphs approach to conceptual schema integration. Lecture Notes in Computer Science, 699:126–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boksenbaum:1993:CGR

[717] C. Boksenbaum, B. Carbonneill, O. Haemmerle, and T. Libourel. Conceptual graphs for relational databases. Lecture Notes in Computer Science, 699:142–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basili:1993:ICR

[718] R. Basili and M. T. Pazienza. Inducing a CG representation for basic-level categorization of verbs. Lecture Notes in Computer Science, 699:162–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulin:1993:RLI

[719] B. Moulin. The representation of linguistic information in an approach used for modelling temporal knowledge in discourses. Lecture Notes in Computer Science, 699:182–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nazarenko:1993:RNL

[720] A. Nazarenko. Representing natural language causality in conceptual graphs: the higher order conceptual relation problem. Lecture Notes in Computer Science, 699:205–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lukose:1993:ECS

[721] D. Lukose. Executable conceptual structures. Lecture Notes in Computer Science, 699:223–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pagnucco:1993:IRC

[722] M. Pagnucco and N. Foo. Inverting resolution with conceptual graphs. Lecture Notes in Computer Science, 699: 238–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levinson:1993:TDI

[723] R. Levinson. Towards domainindependent machine intelligence. Lecture Notes in Computer Science, 699: 254–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellis:1993:ERH

[724] G. Ellis. Efficient retrieval from hierarchies of objects using lattice operations. Lecture Notes in Computer Science, 699:274–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mugnier:1993:CAR

[725] M. L. Mugnier and M. Chein. Characterization and algorithmic recognition of canonical conceptual graphs. Lecture Notes in Computer Science, 699: 294–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willems:1993:CSO

[726] M. Willems. A conceptual semantics ontology for conceptual graphs. Lecture Notes in Computer Science, 699: 312–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cyre:1993:ATK

[727] W. Cyre. Acquiring temporal knowledge from schedules. Lecture Notes in Computer Science, 699:328–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wermelinger:1993:CSM

[728] M. Wermelinger and A. Bejan. Conceptual structures for modeling in CIM. Lecture Notes in Computer Science, 699:345–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aimeur:1993:ETB

[729] E. Aimeur and J. G. Ganascia. Elicitation of taxonomies based on the use of conceptual graphs operators. Lecture Notes in Computer Science, 699: 361–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heaton:1993:PPL

[730] J. E. Heaton and P. Kocura. Presenting a Peirce logic based inference engine and theorem prover for conceptual graphs. Lecture Notes in Computer Science, 699:381–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emond:1993:OCS

[731] B. Emond. Operations on conceptual structures and Peirce’s system of existential graphs. Lecture Notes in Computer Science, 699:401–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDenBerg:1993:MLC

[732] H. Van Den Berg. Modal logics for conceptual graphs. Lecture Notes in Computer Science, 699:411–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wuwongse:1993:FCG

[733] V. Wuwongse and M. Manzano. Fuzzy conceptual graphs. Lecture Notes in Computer Science, 699:430–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1993:PRC

[734] M. Blum. Program result checking: A new approach to making programs more reliable. Lecture Notes in Computer Science, 700:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersson:1993:DIS

[735] A. Andersson and C. Mattsson. Dynamic interpolation search in O(log log n) time. Lecture Notes in Computer Science, 700:15–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frederickson:1993:SAI

[736] G. N. Frederickson. Searching among intervals and compact routing tables. Lecture Notes in Computer Science, 700:28–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lund:1993:AMS

[737] C. Lund and M. Yannakakis. The approximation of maximum subgraph problems. Lecture Notes in Computer Science, 700:40–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kann:1993:PBM

[738] V. Kann. Polynomially bounded minimization problems which are hard to approximate. Lecture Notes in Computer Science, 700:52–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1993:PDA

[739] N. Garg, V. V. Vazirani, and M. Yannakakis. Primal-dual approximation algorithms for integral flow and multicut in trees, with applications to matching and set cover. Lecture Notes in Computer Science, 700:64–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marathe:1993:CAP

[740] M. V. Marathe, H. B. Hunt, III, and S. S. Ravi. The complexity of approximating PSPACE-Complete problems for hierarchical specifications. Lecture Notes in Computer Science, 700:76–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czumaj:1993:PPT

[741] A. Czumaj and A. Gibbons. Problems on pairs of trees and the four colour problem of planar graphs. Lecture Notes in Computer Science, 700:88–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalyanasundaram:1993:CCT

[742] B. Kalyanasundaram and K. R. Pruhs. Constructing competitive tours from local information. Lecture Notes in Computer Science, 700:102–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1993:TPP

[743] H. Bodlaender, T. Kloks, and D. Kratsch. Treewidth and pathwidth of permutation graphs. Lecture Notes in Computer Science, 700:114– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaromczyk:1993:TEF

[744] J. W. Jaromczyk and G. Swiatek. A theory of even functionals and their algorithmic applications. Lecture Notes in Computer Science, 700:126– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flajolet:1993:EAD

[745] P. Flajolet and M. Golin. Exact asymptotics of divide-and-conquer recurrences. Lecture Notes in Computer Science, 700:137–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kozen:1993:OBC

[746] D. Kozen and S. Zaks. Optimal bounds for the change-making problem. Lecture Notes in Computer Science, 700: 150–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reif:1993:CBS

[747] J. H. Reif and S. R. Tate. The complexity of N-body simulation. Lecture Notes in Computer Science, 700:162– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dillencourt:1993:SMR

[748] M. B. Dillencourt and W. D. Smith. A simple method for resolving degeneracies in Delaunay triangulations. Lecture Notes in Computer Science, 700: 177–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemachandra:1993:FTC

[749] L. A. Hemachandra. Fault-tolerance and complexity. Lecture Notes in Computer Science, 700:189–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamamoto:1993:RST

[750] H. Yamamoto. Reversal-space tradeoffs for simultaneous resource-bounded nondeterministic Turing machines. Lecture Notes in Computer Science, 700:203–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orponen:1993:CPD

[751] P. Orponen. On the computational power of discrete Hopfield nets. Lecture Notes in Computer Science, 700: 215–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karpinski:1993:RVD

[752] M. Karpinski and R. Verbeek. On randomized versus deterministic computation. Lecture Notes in Computer Science, 700:227–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ablayev:1993:LBO

[753] F. Ablayev. Lower bounds for oneway probabilistic communication complexity. Lecture Notes in Computer Science, 700:241–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagerup:1993:MDP

[754] T. Hagerup, K. Mehlhorn, and I. Munro. Maintaining discrete probability distributions optimally. Lecture Notes in Computer Science, 700:253– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franklin:1993:SEL

[755] M. Franklin and M. Yung. Secure and efficient off-line digital money. Lecture Notes in Computer Science, 700:265– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juedes:1993:CDR

[756] D. W. Juedes, J. I. Lathrop, and J. H. Lutz. Computational depth and reducibility. Lecture Notes in Computer Science, 700:277–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baliga:1993:LAC

[757] G. Baliga and J. Case. Learnability: Admissible, co-finite, and hypersimple languages. Lecture Notes in Computer Science, 700:289–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1993:IUP

[758] T. Jiang, A. Salomaa, K. Salomaa, and S. Yu. Inclusion is undecidable for pattern languages. Lecture Notes in Computer Science, 700:301–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra:1993:NDR

[759] O. H. Ibarra, T. Jiang, N. Tran, and H. Wang. New decidability results concerning two-way counter machines and applications. Lecture Notes in Computer Science, 700:313–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michaux:1993:CTS

[760] C. Michaux and R. Villemaire. Cobham’s theorem seen through Buchi’s theorem. Lecture Notes in Computer Science, 700:325–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ebinger:1993:LDI

[761] W. Ebinger and A. Muscholl. Logical definability on infinite traces. Lecture Notes in Computer Science, 700:335– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilke:1993:ACR

[762] T. Wilke. Algebras for classifying regular tree languages and an application to frontier testability. Lecture Notes in Computer Science, 700:347–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1993:FAC

[763] A. Gupta. Finite automata as characterizations of minor closed tree families. Lecture Notes in Computer Science, 700:359–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dolev:1993:DAB

[764] D. Dolev and D. Malki. On distributed algorithms in a broadcast domain. Lecture Notes in Computer Science, 700:371–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chlebus:1993:SNS

[765] B. S. Chlebus, K. Diks, and A. Pelc. Sparse networks supporting efficient reliable broadcasting. Lecture Notes in Computer Science, 700:388–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AufDerHeide:1993:SAT

[766] F. Meyer Auf Der Heide, B. Oesterdiekhoff, and R. Wanka. Strongly adaptive token distribution. Lecture Notes in Computer Science, 700:398– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schonhage:1993:FPC

[767] A. Schonhage. Fast parallel computation of characteristic polynomials by Leverrier’s power sum method adapted to fields of finite characteristic. Lecture Notes in Computer Science, 700:410–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirousis:1993:FPC

[768] L. M. Kirousis. Fast parallel constraint satisfaction. Lecture Notes in Computer Science, 700:418–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1993:PRL

[769] I. Simon. The product of rational languages. Lecture Notes in Computer Science, 700:430–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ochmanski:1993:RCS

[770] E. Ochmanski and P.-A. Wacrenier. On regular compatibility of semicommutations. Lecture Notes in Computer Science, 700:445–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dumas:1993:AAB

[771] P. Dumas. Algebraic aspects of B-regular series. Lecture Notes in Computer Science, 700:457–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1993:PFS

[772] D. M. Cohen and M. L. Fredman. Products of finite state machines with full coverage. Lecture Notes in Computer Science, 700:469–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Senizergues:1993:EVS

[773] G. Senizergues. An effective version of Stallings’ theorem in the case of context-free groups. Lecture Notes in Computer Science, 700:478–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lepisto:1993:PPI

[774] A. Lepisto. On the power of periodic iteration of morphisms. Lecture Notes in Computer Science, 700:496–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mignosi:1993:IDL

[775] F. Mignosi and P. Seebold. If a DOL language is k-power free than it is circular. Lecture Notes in Computer Science, 700:507–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jategaonkar:1993:DTC

[776] L. Jategaonkar and A. R. Meyer. Deciding true concurrency equivalences on finite safe nets. Lecture Notes in Computer Science, 700:519–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1993:TTC

[777] W. Vogler. Timed testing of concurrent systems. Lecture Notes in Computer Science, 700:532–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havelund:1993:FC

[778] K. Havelund and K. G. Larsen. The fork calculus. Lecture Notes in Computer Science, 700:544–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inverardi:1993:ETS

[779] P. Inverardi, C. Priami, and D. Yankelevich. Extended transition systems for parametric bisimulation. Lecture Notes in Computer Science, 700:558– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1993:TLC

[780] C. Brown and D. Gurr. Temporal logic and categories of Petri nets. Lecture Notes in Computer Science, 700:570– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lodaya:1993:DPO

[781] K. Lodaya, R. Ramanujam, and P. S. Thiagarajan. Decidability of a partial order based temporal logic. Lecture Notes in Computer Science, 700: 582–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hungar:1993:LMC

[782] H. Hungar and B. Steffen. Local model checking for context-free processes. Lecture Notes in Computer Science, 700:593–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abiteboul:1993:CS

[783] S. Abiteboul and V. Vianu. Computing on structures. Lecture Notes in Computer Science, 700:606–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Contejean:1993:PSD

[784] E. Contejean. A partial solution for D-unification based on a reduction to AC1-Unification. Lecture Notes in Computer Science, 700:621–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codish:1993:EAC

[785] M. Codish, M. Falaschi, K. Marriott, and W. Winsborough. Efficient analysis of concurrent constraint logic programs. Lecture Notes in Computer Science, 700:633–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiCosmo:1993:CRE

[786] R. Di Cosmo and D. Kesner. A confluent reduction for the extensional typed lambda-calculus with pairs, sums, recursion and terminal object. Lecture Notes in Computer Science, 700:645– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbanera:1993:MTC

[787] F. Barbanera and M. Fernandez. Modularity of termination and confluence in combinations of rewrite systems with lambda omega. Lecture Notes in Computer Science, 700:657–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bracho:1993:DAC

[788] F. Bracho and M. Droste. From domains to automata with concurrency. Lecture Notes in Computer Science, 700:669–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanneganti:1993:WUH

[789] R. Kanneganti and R. Cartwright. What is a universal higher-order programming language? Lecture Notes in Computer Science, 700:682–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atzeni:1993:LO

[790] Paolo Atzeni. LOGIDATA+: Overview. Lecture Notes in Computer Science, 701:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ceri:1993:BOL

[791] Stefano Ceri and Letizia Tanca. Bridging objects with logical rules: Towards object-oriented deductive databases. Lecture Notes in Computer Science, 701:7–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atzeni:1993:LM

[792] Paolo Atzeni, Filippo Cacace, Stefano Ceri, and Letizia Tanca. The LOGIDATA+ model. Lecture Notes in Computer Science, 701:20–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atzeni:1993:LLS

[793] Paolo Atzeni, Luca Cabibbo, Giansalvatore Mecca, and Letizia Tanca. The LOGIDATA+ language and semantics. Lecture Notes in Computer Science, 701:30–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cabibbo:1993:TAL

[794] Luca Cabibbo and Giansalvatore Mecca. Travel agency: A LOGI-DATA+ application. Lecture Notes in Computer Science, 701:42–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palopoli:1993:MEU

[795] Luigi Palopoli and Riccardo Torlone. Management of extended update operations. Lecture Notes in Computer Science, 701:60–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beneventano:1993:TRL

[796] Domenico Beneventano, Sonia Bergamaschi, Claudio Sartori, Alessandro Artale, Francesca Cesarini, and Giovanni Soda. Taxonomic reasoning in LOGIDATA+. Lecture Notes in Computer Science, 701:79–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artale:1993:ITR

[797] Alessandro Artale, Francesca Cesarini, and Giovanni Soda. Introducing taxonomic reasoning in LOGIDATA+. Lecture Notes in Computer Science, 701: 85–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beneventano:1993:TRC

[798] Domenico Beneventano, Sonia Bergamaschi, and Claudio Sartori. Taxonomic reasoning with cycles in LOGI-DATA+. Lecture Notes in Computer Science, 701:105–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Formica:1993:MSI

[799] Anna Formica and Michele Missikoff. Modeling semantic integrity constraints in object-oriented database schemas. Lecture Notes in Computer Science, 701:129–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greco:1993:ENL

[800] Sergio Greco, Massimo Romeo, and Domenico Sacca. Evaluation of negative logic programs. Lecture Notes in Computer Science, 701:148–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leone:1993:EIN

[801] Nicola Leone, Massimo Romeo, Pasquale Rullo, and Domenico Sacca. Effective implementation of negation in database logic query languages. Lecture Notes in Computer Science, 701: 159–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fabrizio:1993:MLP

[802] Annalina Fabrizio, Maurizio Capaccioli, and Sandra Valeri. Modules in logic programming: A framework for knowledge management. Lecture Notes in Computer Science, 701:176–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nanni:1993:LLO

[803] Umberto Nanni, Silvio Salza, and Mario Terranova. LOA: The LOGI-DATA+ object algebra. Lecture Notes in Computer Science, 701:195– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nanni:1993:LPS

[804] Umberto Nanni, Silvio Salza, and Mario Terranova. The LOGIDATA+ prototype system. Lecture Notes in Computer Science, 701:211–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lugli:1993:MAO

[805] Marco Lugli, Luca Nini, and Stefano Ceri. MOODsup *: An architecture for object-oriented access to a relational database. Lecture Notes in Computer Science, 701:233–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artale:1993:PLP

[806] Alessandro Artale, Jean-Pierre Ballerini, Sonia Bergamaschi, Filippo Cacace, Stefano Ceri, Francesca Cesarini, Anna Formica, Herman Lam, Sergio Greco, and Giorgio Marrella. Prototypes in the LOGIDATA+ project. Lecture Notes in Computer Science, 701:252–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aanderaa:1993:UTM

[807] Stal Aanderaa. A universal Turing machine. Lecture Notes in Computer Science, 702:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aanderaa:1993:RIL

[808] Stal Aanderaa and Herman Ruge Jervell. Recursive inseparability in linear logic. Lecture Notes in Computer Science, 702:5–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artemov:1993:BLP

[809] Sergei Artemov and Tyko Straben. The basic logic of proofs. Lecture Notes in Computer Science, 702:14–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1993:ASC

[810] Matthias Baaz and Richard Zach. Algorithmic structuring of cut-free proofs. Lecture Notes in Computer Science, 702:29–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Behrendt:1993:OPE

[811] Thomas Behrendt, Kevin Compton, and Erich Gradel. Optimization problems: Expressibility, approximation properties and expected asymptotic growth of optimal solutions. Lecture Notes in Computer Science, 702:43–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benton:1993:LLC

[812] Nick Benton, Gavin Bierman, Valeria De Paiva, and Martin Hyland. Linear lambda-calculus and categorical models revisited. Lecture Notes in Computer Science, 702:61–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berarducci:1993:SIL

[813] Alessandro Berarducci and Corrado Bohm. A self-interpreter of lambda calculus having a normal form. Lecture Notes in Computer Science, 702:85–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosse:1993:EFG

[814] Uwe Bosse. An “ehrenfeucht-fraisse game” for fixpoint logic and stratified fixpoint logic. Lecture Notes in Computer Science, 702:100–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Creignou:1993:CPL

[815] Nadia Creignou. The class of problems that are linearly equivalent to satisfiability, or a uniform method for proving NP-completeness. Lecture Notes in Computer Science, 702:115–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fermuller:1993:MBR

[816] Christian G. Fermuller and Alexander Leitsch. Model building by resolution. Lecture Notes in Computer Science, 702:134–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernando:1993:CTS

[817] Tim Fernando. Comparative transition system semantics. Lecture Notes in Computer Science, 702:149–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gavilanes-Franco:1993:RHO

[818] Antonio Gavilanes-Franco, Francisca Lucio-Carrasco, and Mario Rodriguez-Artalejo. Reasoning with higher order partial functions. Lecture Notes in Computer Science, 702:167–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glavan:1993:CEA

[819] Paola Glavan and Dean Rosenzweig. Communicating evolving algebras. Lecture Notes in Computer Science, 702: 182–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez-Moreno:1993:CNO

[820] Juan Carlos Gonzalez-Moreno, Maria Teresa Hortala-Gonzalez, and Mario Rodriguez-Artalejo. On the completeness of narrowing as the operational semantics of functional logic programming. Lecture Notes in Computer Science, 702:216–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gradel:1993:IDC

[821] Erich Gradel and Martin Otto. Inductive definability with counting on finite structures. Lecture Notes in Computer Science, 702:231–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grandjean:1993:LTA

[822] Etienne Grandjean. Linear time algorithms and NP-Complete problems. Lecture Notes in Computer Science, 702:248–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gurevich:1993:SCP

[823] Yuri Gurevich and James K. Huggins. The semantics of the C programming language. Lecture Notes in Computer Science, 702:274–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honsell:1993:TCF

[824] Furio Honsell, Ian A. Mason, Scott Smith, and Carolyn Talcott. A theory of classes for a functional language with effects. Lecture Notes in Computer Science, 702:309–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lautemann:1993:LDN

[825] Clemens Lautemann. Logical definability of NP-Optimization problems with monadic auxiliary predicates. Lecture Notes in Computer Science, 702:327– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marzetta:1993:UTT

[826] Markus Marzetta. Universes in the theories of types and names. Lecture Notes in Computer Science, 702:340– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1993:NSR

[827] John C. Mitchell and Andre Scedrov. Notes on sconing and relators. Lecture Notes in Computer Science, 702: 352–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schiermeyer:1993:SSL

[828] Ingo Schiermeyer. Solving 3-Satisfiability in less than 1, 579^{n} steps. Lecture Notes in Computer Science, 702:379–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1993:KSE

[829] Jan M. Smith. Kleene’s slash and existence of values of open terms in type theory. Lecture Notes in Computer Science, 702:395–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stark:1993:NCL

[830] Robert F. Stark. Negation-complete logic programs. Lecture Notes in Computer Science, 702:403–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stewart:1993:LCB

[831] Iain A. Stewart. Logical characterizations of bounded query classes II: Polynomial-time oracle machines. Lecture Notes in Computer Science, 702: 410–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tyszkiewicz:1993:APM

[832] Jerzy Tyszkiewicz. On asymptotic probabilities of monadic second order properties. Lecture Notes in Computer Science, 702:425–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBerg:1993:RSD

[833] Mark de Berg. Ray shooting, depth orders and hidden surface removal. Lecture Notes in Computer Science, 703: 1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulisch:1993:DEG

[834] Frances Newbery Paulisch. The design of an extendible graph editor. Lecture Notes in Computer Science, 704:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biehl:1993:OCA

[835] Gunter Biehl. Overview of complex array-based PLDs. Lecture Notes in Computer Science, 705:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Isoaho:1993:TUF

[836] Jouni Isoaho, Arto Nummela, and Hannu Tenhunen. Technologies and utilisation of field programmable gate arrays. Lecture Notes in Computer Science, 705:11–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sangiovanni-Vincentelli:1993:SCF

[837] Alberto Sangiovanni-Vincentelli. Some considerations on field-programmable gate arrays and their impact on system design. Lecture Notes in Computer Science, 705:26–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fawcett:1993:SBF

[838] Bradly K. Fawcett. SRAM-Based FP-GAs ease system verification. Lecture Notes in Computer Science, 705:35–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hauck:1993:MFS

[839] Scott Hauck, Gaetano Boriello, Steven Burns, and Carl Ebeling. MONTAGE:

An FPGA for synchronous and asynchronous circuits. Lecture Notes in Computer Science, 705:44–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hill:1993:ONA

[840] Dwight Hill, Barry Britton, William Oswald, Woo Nam-Sung, Satwant Singh, Chen Che-Tsung, and Bob Krambeck. ORCA: A new architecture for high-performance FPGAs. Lecture Notes in Computer Science, 705:52–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1993:PML

[841] Masahiro Fujita and Yuji Kukimoto. Patching method for lookup-table type FPGAs. Lecture Notes in Computer Science, 705:61–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1993:AOH

[842] Dave Allen. Automatic one-hot reencoding for FPGAs. Lecture Notes in Computer Science, 705:71–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li-Fei:1993:MPR

[843] Wu Li-Fei and Marek A. Perkowski. Minimization of permuted Reed– Muller trees for cellular logic programmable gate arrays. Lecture Notes in Computer Science, 705:78–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blight:1993:SOK

[844] David C. Blight and Robert D. McLeod. Self-organizing Kohonen maps for FPGA placement. Lecture Notes in Computer Science, 705:88–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poechmueller:1993:HLS

[845] Peter Poechmueller, Hans-Jurgen Herpel, Manfred Glesner, and Fang Longsen. High level synthesis in an FPGA-Based computer aided prototyping environment. Lecture Notes in Computer Science, 705:96–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohta:1993:NAF

[846] Naohisa Ohta, Kazuhisa Yamada, Akihiro Tsutsui, and Hiroshi Nakada. New application of FPGAs to programmable digital communication circuits. Lecture Notes in Computer Science, 705:106–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kempa:1993:FBL

[847] Georg Kempa and Peter Jung. FPGA based logic synthesis of squarers using VHDL. Lecture Notes in Computer Science, 705:112–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Surmann:1993:OFC

[848] Hartmut Surmann, Ansgar Ungering, and Karl Goser. Optimized fuzzy controller architecture for field programmable gate arrays. Lecture Notes in Computer Science, 705:124– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindh:1993:RTK

[849] Lennart Lindh, Klaus Muller-Glaser, Hans Rauch, and Frank Stanischewski. A real-time kernel — rapid prototyping with VHDL and FPGAs. Lecture Notes in Computer Science, 705:134– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grunbacher:1993:JBM

[850] Herbert Grunbacher and Alexander Jaud. JAPROC — an 8-bit micro controller design and its test environment. Lecture Notes in Computer Science, 705:146–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heeb:1993:CWD

[851] Beat Heeb and Cuno Pfister. Chameleon: A workstation of a different colour. Lecture Notes in Computer Science, 705:152–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shaw:1993:HPF

[852] Paul Shaw and George Milne. A highly parallel FPGA-Based machine and its formal verification. Lecture Notes in Computer Science, 705:162–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kunzmann:1993:FBS

[853] Arno Kunzmann. FPGA based self-test with deterministic test patterns. Lecture Notes in Computer Science, 705: 174–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoang:1993:FIS

[854] Dzung T. Hoang and Daniel P. Lopresti. FPGA implementation of systolic sequence alignment. Lecture Notes in Computer Science, 705:183– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brunvand:1993:UFP

[855] Eric Brunvand. Using FPGAs to prototype a self-timed computer. Lecture Notes in Computer Science, 705:192– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linde:1993:UFI

[856] Arne Linde, Tomas Nordstrom, and Mikael Taveniku. Using FPGAs to implement a reconfigurable highly parallel computer. Lecture Notes in Computer Science, 705:199–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ast:1993:NHP

[857] Andreas Ast, Reiner Hartenstein, Rainer Kress, Helmut Reinig, and Karin Schmidt. Novel high performance machine paradigms and fastturnaround ASIC design methods. Lecture Notes in Computer Science, 705: 211–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basili:1993:KEP

[858] Victor R. Basili. Keynote: The experimental paradigm in software engineering. Lecture Notes in Computer Science, 706:3–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agresti:1993:PAA

[859] William W. Agresti. Profile of an artifact assessment capability. Lecture Notes in Computer Science, 706:13–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belady:1993:EMS

[860] Les A. Belady. Experiments and measurements for systems integration. Lecture Notes in Computer Science, 706: 17–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1993:SES

[861] Norbert Fuchs. Software engineering still on the way to an engineering discipline. Lecture Notes in Computer Science, 706:19–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludewig:1993:PMS

[862] Jochen Ludewig. Problems in modeling the software development process as an adventure game. Lecture Notes in Computer Science, 706:23–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scacchi:1993:QTT

[863] Walt Scacchi. Qualitative techniques and tools for measuring, analyzing, and simulating software processes. Lecture Notes in Computer Science, 706:27–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tichy:1993:ECS

[864] Walter F. Tichy. On experimental computer science. Lecture Notes in Computer Science, 706:30–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agresti:1993:DS

[865] William Agresti. Discussion summary. Lecture Notes in Computer Science, 706:33–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cusumano:1993:KOC

[866] Michael A. Cusumano. Keynote: Objectives and context of software measurement, analysis and control. Lecture Notes in Computer Science, 706:41–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fenton:1993:PP

[867] Norman Fenton. Position paper. Lecture Notes in Computer Science, 706: 60–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koch:1993:SEO

[868] Gunter R. Koch. Software engineering as an organisational challenge. Lecture Notes in Computer Science, 706:62–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madhavji:1993:QMB

[869] Nazim H. Madhavji, John Botsford, Tilman F. W. Bruckhaus, and Khaled El Emam. Quantitative measurements based on process and context models. Lecture Notes in Computer Science, 706:67–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moller:1993:SIM

[870] Karl-Heinrich Moller. Selecting, implementing, and measuring methods to improve the software development process. Lecture Notes in Computer Science, 706:73–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Porter:1993:RMS

[871] Adam A. Porter. Rethinking measurement to support incremental process improvement. Lecture Notes in Computer Science, 706:77–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fenton:1993:DS

[872] Norman Fenton. Discussion summary. Lecture Notes in Computer Science, 706:82–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Selby:1993:KSM

[873] Richard W. Selby. Keynote: Software measurement and experimentation frameworks, mechanisms, and infrastructure. Lecture Notes in Computer Science, 706:89–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1993:TWD

[874] Warren Harrison. Towards welldefined, shareable product data. Lecture Notes in Computer Science, 706: 107–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffery:1993:VUT

[875] Ross Jeffery. A view on the use of three research philosophies to address empirically determined weaknesses of the software engineering process. Lecture Notes in Computer Science, 706: 111–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemerer:1993:BGB

[876] Chris F. Kemerer. Bridging the gap between research and practice in software engineering management: Reflections on the staffing factors paradox. Lecture Notes in Computer Science, 706:116–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitchenham:1993:MES

[877] Barbara Kitchenham. A methodology for evaluating software engineering methods and tools. Lecture Notes in Computer Science, 706:121–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller:1993:ESE

[878] Hausi A. Muller. Experimental software engineering should concentrate on software evolution. Lecture Notes in Computer Science, 706:125–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sumner:1993:YAL

[879] Eric Sumner, Jr. Yet another laboratory for software engineering. Lecture Notes in Computer Science, 706:129– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zelkowitz:1993:AMP

[880] Marvin V. Zelkowitz. An axiomatic model for program complexity. Lecture Notes in Computer Science, 706: 133–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zuse:1993:SEM

[881] Horst Zuse. Support of experimentation by measurement theory. Lecture Notes in Computer Science, 706:137– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitchenham:1993:DS

[882] Barbara Kitchenham. Discussion summary. Lecture Notes in Computer Science, 706:141–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thebaut:1993:KTS

[883] Stephen M. Thebaut and Vincent Y. Shen. Keynote: Task-specific utility assessment models and their role in the development of software engineering handbooks. Lecture Notes in Computer Science, 706:151–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Briand:1993:QEM

[884] Lionel C. Briand. Quantitative empirical modeling for managing software development: Constraints, needs and solutions. Lecture Notes in Computer Science, 706:158–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cantone:1993:SBC

[885] Giovanni Cantone. Software business, concurrent engineering and experience factory relationships. Lecture Notes in Computer Science, 706:164–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffman:1993:EFS

[886] Daniel Hoffman. Establishing the fundamentals of software engineering. Lecture Notes in Computer Science, 706: 171–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Littlewood:1993:MBM

[887] Bev Littlewood. Measurement-based modelling issues — the problem of assuring ultra-high dependability. Lecture Notes in Computer Science, 706: 173–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonMayrhauser:1993:RSS

[888] Anneliese von Mayrhauser. The role of simulation in software engineering experimentation. Lecture Notes in Computer Science, 706:177–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oivo:1993:MVS

[889] Markku Oivo. Multiple viewpoints of software models. Lecture Notes in Computer Science, 706:180–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonMayrhauser:1993:DS

[890] Anneliese von Mayrhauser. Discussion summary. Lecture Notes in Computer Science, 706:183–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wentzel:1993:KSE

[891] Kevin D. Wentzel. Keynote: Software engineering models, using and reusing. Lecture Notes in Computer Science, 706:191–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Endres:1993:MRT

[892] Albert Endres. Model reuse and technology transfer. Lecture Notes in Computer Science, 706:202–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldman:1993:PRR

[893] Stuart I. Feldman. Packaging for reuse and reuse of models. Lecture Notes in Computer Science, 706:206–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewerentz:1993:RCS

[894] Claus Lewerentz. A reuse culture for software construction. Lecture Notes in Computer Science, 706:210–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McGarry:1993:ESE

[895] Frank McGarry. Experimental software engineering; packaging for reuse. Lecture Notes in Computer Science, 706: 213–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneidewind:1993:EDV

[896] Norman F. Schneidewind. Experimental designs for validating metrics and applying them across multiple projects. Lecture Notes in Computer Science, 706:216–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McGarry:1993:DS

[897] Frank McGarry. Discussion summary. Lecture Notes in Computer Science, 706:221–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehman:1993:KPP

[898] Manny M. Lehman. Keynote: Position paper. Lecture Notes in Computer Science, 706:229–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marciniak:1993:TT

[899] John J. Marciniak. Technology transfer. Lecture Notes in Computer Science, 706:234–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rombach:1993:SST

[900] H. Dieter Rombach. Systematic software technology transfer. Lecture Notes in Computer Science, 706:239– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zweben:1993:EUM

[901] Stu Zweben. Effective use of measurement and experimentation in computing curricula. Lecture Notes in Computer Science, 706:247–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rombach:1993:DS

[902] H. Dieter Rombach. Discussion summary. Lecture Notes in Computer Science, 706:252–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamming:1993:ICM

[903] Michael G. Lamming. Intimate computing and the memory prosthesis: A challenge for computer systems research? Lecture Notes in Computer Science, 707:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0707/07070001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0707/07070001.pdf.

Rosson:1993:APS

[904] Mary Beth Rosson and John M. Carroll. Active programming strategies in reuse. Lecture Notes in Computer Science, 707:4–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0707/07070004.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0707/07070004.pdf.

Birrer:1993:FFE

[905] Andreas Birrer and Thomas Eggenschwiler. Frameworks in the financial engineering domain: An experience report. Lecture Notes in Computer Science, 707:21–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0707/07070021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0707/07070021.
pdf.

Holzle:1993:IID

[906] Urs Hölzle. Integrating independentlydeveloped components in objectoriented languages. Lecture Notes in Computer Science, 707:36–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0707/07070036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0707/07070036.
pdf.

Black:1993:EP

[907] Andrew P. Black and Mark P. Immel. Encapsulating plurality. Lecture Notes in Computer Science, 707:57– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0707/07070057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0707/07070057.
pdf.

Konstantas:1993:OOI

[908] Dimitri Konstantas. Object oriented interoperability. Lecture Notes in Computer Science, 707:80–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070080.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070080.
                                        pdf.

Achauer:1993:IDT

[909] Bruno Achauer. Implementation of distributed trellis. Lecture Notes in Computer Science, 707:103–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070103.
                                        pdf.

Liskov:1993:NDS

[910] Barbara Liskov and Jeannette M. Wing. A new definition of the subtype relation. Lecture Notes in Computer Science, 707:118–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070118.
                                        pdf.

Caseau:1993:ASO

[911] Yves Caseau and Laurent Perron. Attaching second-order types to methods in an object-oriented language. Lecture Notes in Computer Science, 707: 142–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0707/07070142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0707/07070142.
pdf.

Balsters:1993:TSB

[912] Herman Balsters, Rolf A. De By, and Roberto Zicari. Typed sets as a basis for object-oriented database schemas. Lecture Notes in Computer Science, 707:161–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0707/07070161.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0707/07070161.
pdf.

Ashford:1993:OMO

[913] Colin Ashford. The OSI managedobject model. Lecture Notes in Computer Science, 707:185–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/0707/07070185.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0707/07070185.pdf.

Steyaert:1993:NMM

[914] Patrick Steyaert, Wim Codenie, Theo D’hondt, Koen De Hondt, Carine Lucas, and Marc Van Limberghen. Nested mixin-methods in agora. Lecture Notes in Computer Science, 707:

197–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070197.
                                        pdf.

Meseguer:1993:SIA

[915] José Meseguer. Solving the inheritance anomaly in concurrent objectoriented programming. Lecture Notes in Computer Science, 707:220–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070220.
                                        pdf.

Agesen:1993:TIS

[916] Ole Agesen, Jens Palsberg, and Michael I. Schwartzbach. Type inference of SELF: Analysis of objects with dynamic and multiple inheritance. Lecture Notes in Computer Science, 707: 247–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070247.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070247.
                                        pdf.

Chambers:1993:PC

[917] Craig Chambers. Predicate classes. Lecture Notes in Computer Science, 707:268–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0707/07070268.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0707/07070268.
pdf.

Rose:1993:TTO

[918] Ellen Rose and Arie Segev. TOOA: A temporal object-oriented algebra. Lecture Notes in Computer Science, 707: 297–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0707/07070297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0707/07070297.
pdf.

Satoh:1993:TCD

[919] Ichiro Satoh and Mario Tokoro. A timed calculus for distributed objects with clocks. Lecture Notes in Computer Science, 707:326–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0707/07070326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0707/07070326.
pdf.

Frolund:1993:LFM

[920] Svend Frølund and Gul Agha. A language framework for multiobject coordination. Lecture Notes in Computer Science, 707:346–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/0707/07070346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070346.
                                        pdf.

Assenmacher:1993:PSD

[921] Holger Assenmacher, Thomas Breitbach, Peter Buhler, Volker Hübsch, and Reinhard Schwarz. PANDA — supporting distributed programming in C++. Lecture Notes in Computer Science, 707:361–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070361.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070361.
                                        pdf.

Jezequel:1993:TPT

[922] Jean-Marc Jézéquel. Transparent parallelisation through reuse: Between a compiler and a library approach. Lecture Notes in Computer Science, 707: 384–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070384.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070384.
                                        pdf.

Gamma:1993:DPA

[923] Erich Gamma, Richard Helm, Ralph Johnson, and John Vlissides. Design patterns: Abstraction and reuse of object-oriented design. Lecture Notes in Computer Science, 707:406– 431, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0707/07070406.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0707/07070406.
pdf.

Gangopadhyay:1993:OTS

[924] Dipayan Gangopadhyay and Subrata Mitra. ObjChart: Tangible specification of reactive object behavior. Lecture Notes in Computer Science, 707: 432–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0707/07070432.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0707/07070432.
pdf.

Dubois:1993:RAA

[925] Eric Dubois, Philippe Du Bois, and Michaël Petit. O-O requirements analysis: An agent perspective. Lecture Notes in Computer Science, 707:458– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0707/07070458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0707/07070458.
pdf.

Chiba:1993:DED

[926] Shigeru Chiba and Takashi Masuda. Designing an extensible distributed language with a metalevel architecture. Lecture Notes in Computer Science, 707:482–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070482.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070482.
                                        pdf.

Johnson:1993:MFM

[927] Richard Johnson and Murugappan Palaniappan. MetaFlex: A flexible metaclass generator. Lecture Notes in Computer Science, 707:502–527, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070502.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070502.
                                        pdf.

Banahan:1993:PAM

[928] Mike Banahan, L. Peter Deutsch, Boris Magnusson, and Jens Palsberg. Panel: Aims, means, and futures of object-oriented languages. Lecture Notes in Computer Science, 707:528– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0707/07070528.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0707/07070528.
                                        pdf.

Anonymous:1993:AIb

[929] Anonymous. Author index. Lecture Notes in Computer Science, 707:531– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
papers/0707/0707ind.pdf.

Boissonnat:1993:SPB

[930] J. D. Boissonnat, A. Cerezo, and J. Leblond. Shortest paths of bounded curvature in the plane. Lecture Notes in Computer Science, 708:3–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraichard:1993:KPS

[931] Th. Fraichard and C. Laugier. Kinodynamic planning in a structured and time-varying workspace. Lecture Notes in Computer Science, 708:19–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simeon:1993:MPN

[932] T. Simeon. Motion planning for a non-holonomic mobile robot on 3dimensional terrains. Lecture Notes in Computer Science, 708:38–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liegeois:1993:OMP

[933] A. Liegeois and C. Moignard. Optimal motion planning of a mobile robot on a triangulated terrain model. Lecture Notes in Computer Science, 708:51–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lazanas:1993:LBR

[934] A. Lazanas and J. C. Latombe. Landmark-based robot motion planning. Lecture Notes in Computer Science, 708:69–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahuactzin:1993:UGA

[935] J. M. Ahuactzin, El-Ghazali Talbi, P. Bessiere, and E. Mazer. Using genetic algorithms for robot motion planning. Lecture Notes in Computer Science, 708:84–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zapata:1993:FMR

[936] R. Zapata and P. Lepinay. Fast mobile robots in unstructured environments. Lecture Notes in Computer Science, 708:94–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Espiau:1993:NAV

[937] B. Espiau, F. Chaumette, and P. Rives. A new approach to visual servoing in robotics. Lecture Notes in Computer Science, 708:106–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohr:1993:GSS

[938] R. Mohr and L. Morin. Geometric solutions to some 3D vision problems. Lecture Notes in Computer Science, 708:139–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chassery:1993:GRS

[939] J. M. Chassery and A. Montanvert. Geometrical representation of shapes and objects for visual perception. Lecture Notes in Computer Science, 708:163– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zoppis:1993:PGSa

[940] B. Zoppis, G. Calvary, and J. L. Crowley. Perceptual grouping for scene interpretation in an active vision system. Lecture Notes in Computer Science, 708:183–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moutarlier:1993:IFS

[941] P. Moutarlier and R. Chatila. Incremental free-space modelling from uncertain data by an autonomous mobile robot. Lecture Notes in Computer Science, 708:200–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lavallee:1993:MDS

[942] S. Lavallee, R. Szeliski, and L. Brunie. Matching 3-D smooth surfaces with their 2-D projections using 3-D distance maps. Lecture Notes in Computer Science, 708:217–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nastar:1993:NPB

[943] C. Nastar and N. Ayache. A new physically based model for efficient tracking and analysis of deformations. Lecture Notes in Computer Science, 708: 239–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leitner:1993:SSS

[944] F. Leitner and P. Cinquin. From splines and snakes to SNAKE SPLINES. Lecture Notes in Computer Science, 708: 264–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atallah:1993:CAP

[945] M. J. Atallah and D. Z. Chen. Computing the all-pairs longest chains in the plane. Lecture Notes in Computer Science, 709:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borodin:1993:TBU

[946] A. Borodin. Towards a better understanding of pure packet routing. Lecture Notes in Computer Science, 709: 14–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cole:1993:TFM

[947] R. Cole. Tolerating faults in meshes and other networks (abstract). Lecture Notes in Computer Science, 709:26–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karp:1993:GBS

[948] R. M. Karp. A generalization of binary search. Lecture Notes in Computer Science, 709:27–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1993:GAC

[949] A. C. Yao. Groups and algebraic complexity (abstract). Lecture Notes in Computer Science, 709:35–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1993:CCS

[950] P. K. Agarwal and M. Van Kreveld. Connected component and simple polygon intersection searching. Lecture Notes in Computer Science, 709:36–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amato:1993:OAF

[951] N. M. Amato. An optimal algorithm for finding the separation of simple polygons. Lecture Notes in Computer Science, 709:48–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersson:1993:BST

[952] A. Andersson. Balanced search trees made simple. Lecture Notes in Computer Science, 709:60–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aoki:1993:PSH

[953] Y. Aoki, H. Imai, K. Imai, and D. Rappaport. Probing a set of hyperplanes by lines and related problems. Lecture Notes in Computer Science, 709:72–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arge:1993:GLB

[954] L. Arge, M. Knudsen, and K. Larsen. A general lower bound on the I/Ocomplexity of comparison-based algorithms. Lecture Notes in Computer Science, 709:83–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arkin:1993:PPD

[955] E. M. Arkin, M. T. Goodrich, J. S. B. Mitchell, D. Mount, C. D. Piatko, and S. S. Skiena. Point probe decision trees for geometric concept classes. Lecture Notes in Computer Science, 709:95–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armon:1993:DSA

[956] D. Armon and J. Reif. A dynamic separator algorithm. Lecture Notes in Computer Science, 709:107–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Azar:1993:OLB

[957] Y. Azar, B. Kalyanasundaram, S. Plotkin, K. R. Pruhs, and O. Waarts. Online load balancing of temporary tasks. Lecture Notes in Computer Science, 709:119–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balakrishnan:1993:CDS

[958] H. Balakrishnan, A. Rajaraman, and C. P. Rangan. Connected domination and Steiner set on asteroidal triple-free graphs. Lecture Notes in Computer Science, 709:131–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balasubramanian:1993:CFC

[959] R. Balasubramanian, V. Raman, and G. Srinivasaraghavan. The complexity of finding certain trees in tournaments. Lecture Notes in Computer Science, 709:142–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiBattista:1993:SOR

[960] G. Di Battista, G. Liotta, and F. Vargiu. Spirality of orthogonal representations and optimal drawings of series-parallel graphs and 3-Planar graphs. Lecture Notes in Computer Science, 709:151–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beame:1993:SPE

[961] P. Beame, F. E. Fich, and R. K. Sinha. Separating the power of EREW and CREW PRAMs with small communication width. Lecture Notes in Computer Science, 709:163–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berkman:1993:TLU

[962] O. Berkman, Y. Matias, and P. Ragde. Triply-logarithmic upper and lower bounds for minimum, range minima, and related problems with integer inputs. Lecture Notes in Computer Science, 709:175–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bern:1993:PCQ

[963] M. Bern, D. Eppstein, and S.-H. Teng. Parallel construction of quadtrees and quality triangulations. Lecture Notes in Computer Science, 709:188–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1993:PMP

[964] P. Bose, J. F. Buss, and A. Lubiw. Pattern matching for permutations. Lecture Notes in Computer Science, 709:200–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1993:FPM

[965] P. Bose, M. Van Kreveld, and G. Toussaint. Filling polyhedral molds. Lecture Notes in Computer Science, 709: 210–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1993:DQE

[966] M.-S. Chang, S.-L. Peng, and J.-L. Liaw. Deferred-query — an efficient approach for problems on interval and circular-arc graphs. Lecture Notes in Computer Science, 709:222–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1993:CGE

[967] J. Chen, S. P. Kanchi, and A. Kanevsky. On the complexity of graph embeddings. Lecture Notes in Computer Science, 709:234–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarkson:1993:APC

[968] K. L. Clarkson. Algorithms for polytope covering and approximation. Lecture Notes in Computer Science, 709: 246–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codenotti:1993:GSA

[969] B. Codenotti, G. Manzini, L. Margara, and G. Resta. Global strategies for augmenting the efficiency of TSP heuristics. Lecture Notes in Computer Science, 709:253–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Datta:1993:SDA

[970] A. Datta, H.-P. Lenhof, C. Schwarz, and M. Smid. Static and dynamic algorithms for k-point clustering problems. Lecture Notes in Computer Science, 709:265–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devillers:1993:SAB

[971] O. Devillers and A. Fabri. Scalable algorithms for bichromatic line segment intersection problems on coarse grained multicomputers. Lecture Notes in Computer Science, 709:277–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1993:PRP

[972] P. F. Dietz and R. Raman. Persistence, randomization and parallelization: On some combinatorial games and their applications. Lecture Notes in Computer Science, 709:289–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ding:1993:KDH

[973] Y. Ding and M. A. Weiss. The k-D heap: An efficient multi-dimensional priority queue. Lecture Notes in Computer Science, 709:302–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobrindt:1993:CEA

[974] K. Dobrindt, K. Mehlhorn, and M. Yvinec. A complete and efficient algorithm for the intersection of a general and a convex polyhedron. Lecture Notes in Computer Science, 709:314– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Efrat:1993:CSE

[975] A. Efrat, M. Sharir, and A. Ziv. Computing the smallest k-enclosing circle and related problems. Lecture Notes in Computer Science, 709:325–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giancarlo:1993:IDS

[976] R. Giancarlo. An index data structure for matrices, with applications to fast two-dimensional pattern matching. Lecture Notes in Computer Science, 709:337–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1993:PSA

[977] T. Graf and K. Hinrichs. A planesweep algorithm for the all-nearest-neighbors problem for a set of convex planar objects. Lecture Notes in Computer Science, 709:349–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1993:FRG

[978] P. Gupta, R. Janardan, and M. Smid. Further results on generalized intersection searching problems: Counting, reporting, and dynamization. Lecture Notes in Computer Science, 709:361– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heffernan:1993:GAA

[979] P. J. Heffernan. Generalized approximate algorithms for point set congruence. Lecture Notes in Computer Science, 709:373–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1993:ASS

[980] T. Jiang and M. Li. Approximating shortest superstrings with constraints. Lecture Notes in Computer Science, 709:385–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kannan:1993:TRP

[981] S. Kannan and T. Warnow. Tree reconstruction from partial orders. Lecture Notes in Computer Science, 709: 397–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1993:IPD

[982] M.-Y. Kao, S.-H. Teng, and K. Toyama. Improved parallel depth-first search in undirected planar graphs. Lecture Notes in Computer Science, 709:409– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karger:1993:ALP

[983] D. Karger, R. Motwani, and G. D. S. Ramkumar. On approximating the longest path in a graph. Lecture Notes in Computer Science, 709:421– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khuller:1993:DMC

[984] S. Khuller, B. Raghavachari, and N. Young. Designing multi-commodity flow trees. Lecture Notes in Computer Science, 709:433–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:1993:FDA

[985] P. N. Klein and S. Subramanian. A fully dynamic approximation scheme for all-pairs paths in planar graphs. Lecture Notes in Computer Science, 709:442–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanKreveld:1993:FPF

[986] M. Van Kreveld. On fat partitioning, fat covering, and the union size of polygons. Lecture Notes in Computer Science, 709:452–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krizanc:1993:TRT

[987] D. Krizanc. A time-randomness tradeoff for selection in parallel. Lecture Notes in Computer Science, 709:464– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1993:DRC

[988] H.-I. Lu, P. N. Klein, and R. H. B. Netzer. Detecting race conditions in parallel programs that use one semaphore. Lecture Notes in Computer Science, 709:471–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maggs:1993:AFP

[989] B. M. Maggs and M. Rauch. An algorithm for finding predecessors in integer sets. Lecture Notes in Computer Science, 709:483–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maier:1993:ESM

[990] R. S. Maier and R. Schott. The exhaustion of shared memory: Stochastic results. Lecture Notes in Computer Science, 709:494–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mirzaian:1993:MWE

[991] A. Mirzaian. Minimum weight Euclidean matching and weighted relative neighborhood graphs. Lecture Notes in Computer Science, 709:506–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitra:1993:EAS

[992] P. Mitra and B. Bhattacharya. Efficient approximate shortest-path queries among isothetic rectangular obstacles. Lecture Notes in Computer Science, 709:518–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palazzi:1993:CRR

[993] L. Palazzi and J. Snoeyink. Counting and reporting red/blue segment intersections. Lecture Notes in Computer Science, 709:530–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pellegrini:1993:RHS

[994] M. Pellegrini. Repetitive hiddensurface-removal for polyhedral scenes. Lecture Notes in Computer Science, 709:541–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DePrisco:1993:RVL

[995] R. De Prisco and A. Monti. On reconfigurability of VLSI linear arrays. Lecture Notes in Computer Science, 709:553–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skiena:1993:RSS

[996] S. S. Skiena and G. Sundaram. Reconstructing strings from substrings. Lecture Notes in Computer Science, 709: 565–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Souvaine:1993:CCS

[997] D. L. Souvaine and C.-K. Yap. Combinatorial complexity of signed discs. Lecture Notes in Computer Science, 709:577–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stallmann:1993:FAO

[998] M. F. M. Stallmann and T. A. Hughes. Fast algorithms for one-dimensional compaction with jog insertion. Lecture Notes in Computer Science, 709:589– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swanson:1993:OAR

[999] K. Swanson. An optimal algorithm for roundness determination on convex polygons. Lecture Notes in Computer Science, 709:601–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Telle:1993:PAP

[1000] J. A. Telle and A. Proskurowski. Practical algorithms on partial k-trees with an application to domination-like problems. Lecture Notes in Computer Science, 709:610–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Westbrook:1993:GAL

[1001] J. Westbrook and D. C. K. Yan. Greedy algorithms for the on-line Steiner tree and generalized Steiner problems. Lecture Notes in Computer Science, 709: 622–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekert:1993:RMF

[1002] V. Diekert. Rewriting, mobius functions and semi-commutations. Lecture Notes in Computer Science, 710:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietzfelbinger:1993:SBD

[1003] M. Dietzfelbinger. Simulations between different models of parallel computers. Lecture Notes in Computer Science, 710:16–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1993:DDP

[1004] M. Ito. Dense and disjunctive properties of languages. Lecture Notes in Computer Science, 710:31–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jurgensen:1993:HC

[1005] H. Jurgensen and S. Konstantinidis.

The hierarchy of codes. Lecture Notes in Computer Science, 710:50–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kreowski:1993:FFH

[1006] H.-J. Kreowski. Five facets of hyperedge replacement beyond contextfreeness. Lecture Notes in Computer Science, 710:69–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milner:1993:ASS

[1007] R. Milner. An action structure for synchronous pi-calculus. Lecture Notes in Computer Science, 710:87–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pudlak:1993:CCS

[1008] R. Pudlak. AC^{0} circuit complexity. A survey of lower bounds. Lecture Notes in Computer Science, 710:106– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomaa:1993:PLP

[1009] A. Salomaa. Pattern languages: Problems of decidability and generation. Lecture Notes in Computer Science, 710:121–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdulrab:1993:GSM

[1010] H. Abdulrab and M. Maksimenko. General solution of mirror equation. Lecture Notes in Computer Science, 710:133–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andre:1993:DEL

[1011] Y. Andre and F. Bossut. Decidability of equivalence for linear letter to letter top-down tree transducers. Lecture Notes in Computer Science, 710:142–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergstra:1993:TBF

[1012] J. A. Bergstra and Gh. Stefanescu.

Translations between flowchart schemes and process graphs. Lecture Notes in Computer Science, 710:152–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cazanescu:1993:LEL

[1013] V. E. Cazanescu. Local equational logic. Lecture Notes in Computer Science, 710:162–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chrzastowski-Wachtel:1993:LWC

[1014] P. Chrzastowski-Wachtel and M. Raczunas. Liveness of weighted circuits and the Diophantine problem of Frobenius. Lecture Notes in Computer Science, 710:171–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1993:CFG

[1015] B. Courcelle. Context-free graph grammars: Separating vertex replacement from hyperedge replacement. Lecture Notes in Computer Science, 710:181– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domosi:1993:FLC

[1016] P. Domosi, S. Horvath, M. Ito, L. Kaszonyi, and M. Katsura. Formal languages consisting of primitive words. Lecture Notes in Computer Science, 710:194–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand:1993:USP

[1017] B. Durand. Undecidability of the surjectivity problem for 2D cellular automata: A simplified proof. Lecture Notes in Computer Science, 710:204– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ebert:1993:EIS

[1018] J. Ebert. Efficient interpretation of state charts. Lecture Notes in Computer Science, 710:212–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fassbender:1993:IUU

[1019] H. Fassbender. Implementation of a universal unification algorithm for macro tree transducers. Lecture Notes in Computer Science, 710:222– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1993:FMC

[1020] P. Fischer. Finding maximum convex polygons. Lecture Notes in Computer Science, 710:234–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1993:AAA

[1021] P. Fischer, K.-U. Hoffgen, H. Lefmann, and T. Luczak. Approximations with axis-aligned rectangles. Lecture Notes in Computer Science, 710:244– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girkar:1993:VSA

[1022] M. Girkar and R. Moll. Vector sequence analysis and full weak safety for concurrent systems. Lecture Notes in Computer Science, 710:256–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groger:1993:DTH

[1023] H. D. Groger. Does transitivity help?

on the complexity of poset properties. Lecture Notes in Computer Science, 710:266–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagerup:1993:GTS

[1024] T. Hagerup and M. Maas. Generalized topological sorting in linear time. Lecture Notes in Computer Science, 710:279–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemachandra:1993:ECS

[1025] L. A. Hemachandra and R. Silvestri.

Easily checked self-reducibility. Lecture Notes in Computer Science, 710:289– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzer:1993:CLL

[1026] M. Holzer and K.-J. Lange. On the complexities of linear LL(1) and LR(1) grammars. Lecture Notes in Computer Science, 710:299–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hung:1993:RBF

[1027] D. V. Hung and T. V. Dung. On the relation between firing sequences and processes of Petri nets. Lecture Notes in Computer Science, 710:309–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jansen:1993:MCD

[1028] K. Jansen, P. Scheffler, and G. Woeginger. Maximum covering with D cliques. Lecture Notes in Computer Science, 710:319–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemp:1993:MLO

[1029] R. Kemp. Monotonically labelled ordered trees and multidimensional binary trees. Lecture Notes in Computer Science, 710:329–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuske:1993:MPL

[1030] S. Kuske. Maximum path length pumping lemma for edge-replacement languages. Lecture Notes in Computer Science, 710:342–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maier:1993:RAS

[1031] R. S. Maier and R. Schott. Regular approximations to shuffle products of context-free languages, and convergence of their generating functions. Lecture Notes in Computer Science, 710:352–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manes:1993:ETB

[1032] E. G. Manes. The equational theory of a Boolean monad. Lecture Notes in Computer Science, 710:363–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Margenstern:1993:NET

[1033] M. Margenstern. Non erasing Turing machines: A frontier between a decidable halting problem and universality. Lecture Notes in Computer Science, 710:375–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mateescu:1993:SSM

[1034] A. Mateescu. On scattered syntactic monoids. Lecture Notes in Computer Science, 710:386–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Potthoff:1993:RTL

[1035] A. Potthoff and W. Thomas. Regular tree languages without unary symbols are star-free. Lecture Notes in Computer Science, 710:396–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roka:1993:OWC

[1036] Z. Roka. One-way cellular automata on Cayley graphs. Lecture Notes in Computer Science, 710:406–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruzicka:1993:TPU

[1037] P. Ruzicka and I. Privara. On tree pattern unification problems. Lecture Notes in Computer Science, 710:418– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomaa:1993:SEE

[1038] K. Salomaa, D. Wood, and S. Yu. Structural equivalence and ETOL grammars. Lecture Notes in Computer Science, 710:430–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slutzki:1993:HDT

[1039] G. Slutzki and S. Vagvolgyi. A hierarchy of deterministic top-down tree transformations. Lecture Notes in Computer Science, 710:440–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1993:STT

[1040] H. Wu. Synthesis of O(lgn) testable trees. Lecture Notes in Computer Science, 710:452–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yin:1993:LRP

[1041] J. Yin and Z. Hong. On the learnability of a restricted predicate formulae. Lecture Notes in Computer Science, 710:462–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apt:1993:KRA

[1042] K. R. Apt and S. Etalle. Krzysztof R. Apt: On the unification free Prolog programs. Lecture Notes in Computer Science, 711:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arnold:1993:AAE

[1043] A. Arnold and A. Dicky. Andre Arnold: Equivalences and preorders of transition systems. Lecture Notes in Computer Science, 711:20–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKinna:1993:RBD

[1044] J. McKinna and R. Burstall. Rod Burstall: Deliverables: a categorical approach to program development in type theory. Lecture Notes in Computer Science, 711:32–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:VDC

[1045] Anonymous. Volker Diekert: Complex and complex-like traces. Lecture Notes in Computer Science, 711:68–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:MHS

[1046] Anonymous. Matthew Hennessy: Symbolic bisimulations (abstract). Lecture Notes in Computer Science, 711:83–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honsell:1993:FHS

[1047] F. Honsell and M. Lenisa. Furio Honsell: Some results on the full abstraction problem for restricted lambda calculi. Lecture Notes in Computer Science, 711:84–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:RMA

[1048] Anonymous. Robin Milner: Action calculi, or syntactic action structures. Lecture Notes in Computer Science, 711:105–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pitts:1993:AMP

[1049] A. M. Pitts and I. D. B. Stark. Andrew M. Pitts: Observable properties of higher order functions that dynamically create local names, or: What’s new? Lecture Notes in Computer Science, 711:122–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:VPS

[1050] Anonymous. Vaughan Pratt: The second calculus of binary relations. Lecture Notes in Computer Science, 711: 142–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrenfeucht:1993:GRI

[1051] A. Ehrenfeucht and G. Rozenberg. Grzegorz Rozenberg: An introduction to dynamic labeled 2-structures. Lecture Notes in Computer Science, 711: 156–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mateescu:1993:AKS

[1052] A. Mateescu and A. Salomaa. Arto K. Salomaa: Post correspondence problem: Primitivity and interrelations with complexity classes. Lecture Notes in Computer Science, 711:174– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:PWT

[1053] Anonymous. Philip Wadler: A taste of linear logic. Lecture Notes in Computer Science, 711:185–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1993:TIP

[1054] L. Alonso and R. Schott. On the tree inclusion problem. Lecture Notes in Computer Science, 711:211–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amadio:1993:APM

[1055] R. Amadio. On the adequacy of per models. Lecture Notes in Computer Science, 711:222–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1993:HRS

[1056] V. Arvind, J. Kobler, and M. Mundhenk. Hausdorff reductions to sparse sets and to sets of high information content. Lecture Notes in Computer Science, 711:232–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astesiano:1993:SHT

[1057] E. Astesiano, G. Reggio, and E. Zucca.

Stores as homomorphisms and their transformations. Lecture Notes in Computer Science, 711:242–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBakker:1993:CSL

[1058] J. W. de Bakker, F. Van Breugel, and A. de Bruin. Comparative semantics for linear arrays of communicating processes, a study of the UNIX fork and pipe commands. Lecture Notes in Computer Science, 711:252–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beauquier:1993:RTA

[1059] D. Beauquier and A. Podelski. Rabin tree automata and finite monoids. Lecture Notes in Computer Science, 711: 262–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benke:1993:ETR

[1060] M. Benke. Efficient type reconstruction in the presence of inheritance. Lecture Notes in Computer Science, 711:272– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berstel:1993:CSM

[1061] J. Berstel and P. Seebold. A characterization of Sturmian morphisms. Lecture Notes in Computer Science, 711: 281–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1993:CSI

[1062] H. L. Bodlaender and K. Jansen. On the complexity of scheduling incompatible jobs with unit-times. Lecture Notes in Computer Science, 711:291– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonsangue:1993:IBP

[1063] M. Bonsangue and J. N. Kok. Isomorphisms between predicate and state transformers. Lecture Notes in Computer Science, 711:301–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liming:1993:ANP

[1064] Cai Liming and Chen Jianer. On the amount of nondeterminism and the power of verifying. Lecture Notes in Computer Science, 711:311–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castellani:1993:ODP

[1065] I. Castellani. Observing distribution in processes. Lecture Notes in Computer Science, 711:321–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerin:1993:SRT

[1066] C. Cerin and A. Petit. Speedup of recognizable trace languages. Lecture Notes in Computer Science, 711:332– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerioli:1993:MBY

[1067] M. Cerioli and J. Meseguer. May I borrow your logic? Lecture Notes in Computer Science, 711:342–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhuri:1993:AED

[1068] S. Chaudhuri, T. Hagerup, and R. Raman. Approximate and exact deterministic parallel selection. Lecture Notes in Computer Science, 711:352–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jian:1993:DSS

[1069] Chen Jian and J. Staples. Defining soft sortedness by abstract interpretation. Lecture Notes in Computer Science, 711:362–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liang:1993:MRT

[1070] Chen Liang. A model for real-time process algebras. Lecture Notes in Computer Science, 711:372–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costa:1993:DEM

[1071] J. F. Costa, A. Sernadas, and C. Sernadas. Data encapsulation and modularity: Three views of inheritance. Lecture Notes in Computer Science, 711: 382–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culik:1993:ICU

[1072] K. Culik II and J. Kari. Image compression using weighted finite automata. Lecture Notes in Computer Science, 711:392–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dezani-Ciancaglini:1993:FMP

[1073] M. Dezani-Ciancaglini, U. De’liguoro, and A. Piperno. Filter models for a parallel and non deterministic lambdacalculus. Lecture Notes in Computer Science, 711:403–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

diGianantonio:1993:RNC

[1074] P. di Gianantonio. Real number computability and domain theory. Lecture Notes in Computer Science, 711:413– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diskin:1993:LSA

[1075] Z. B. Diskin and I. D. Beylin. Lambda substitution algebras. Lecture Notes in Computer Science, 711:423–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand:1993:GPC

[1076] B. Durand. Global properties of 2D cellular automata: Some complexity results. Lecture Notes in Computer Science, 711:433–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engberg:1993:CRL

[1077] U. Engberg and G. Winskel. Completeness results for linear logic on Petri nets. Lecture Notes in Computer Science, 711:442–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fantechi:1993:ELB

[1078] A. Fantechi, S. Gnesi, and V. Perticaroli. An expressive logic for basic process algebra. Lecture Notes in Computer Science, 711:453–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frandsen:1993:CFR

[1079] G. S. Frandsen, P. B. Miltersen, and S. Skyum. The complexity of finding replicas using equality tests. Lecture Notes in Computer Science, 711:463– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGlabbeek:1993:CAB

[1080] R. J. Van Glabbeek. A complete axiomatization for branching bisimulation congruence of finite-state behaviours. Lecture Notes in Computer Science, 711:473–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deHaas:1993:OOA

[1081] E. de Haas and P. Van Emde Boas. Object oriented application flow graphs and their semantics. Lecture Notes in Computer Science, 711:485–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1993:SHC

[1082] J. Hromkovic, J. Kari, and L. Kari. Some hierarchies for the communication complexity measures of cooperating grammar systems. Lecture Notes in Computer Science, 711:495–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra:1993:EPG

[1083] L. Ibarra and D. Richards. Efficient parallel graph algorithms based on open ear decomposition. Lecture Notes in Computer Science, 711:506– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra:1993:CCP

[1084] O. H. Ibarra and N. Tran. On the communication complexity of parallel computation. Lecture Notes in Computer Science, 711:517–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jancar:1993:TFA

[1085] P. Jancar, F. Mraz, and M. Platek. A taxonomy of forgetting automata. Lecture Notes in Computer Science, 711: 527–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1993:HPP

[1086] Jifeng He. Hybrid parallel programming and implementation of synchronised communication. Lecture Notes in Computer Science, 711:537–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiehn:1993:PSC

[1087] A. Kiehn. Proof systems for cause based equivalences. Lecture Notes in Computer Science, 711:547–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1993:UUC

[1088] B. Martin. A uniform universal CREW PRAM. Lecture Notes in Computer Science, 711:557–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murphy:1993:OLC

[1089] D. Murphy. Observing located concurrency. Lecture Notes in Computer Science, 711:566–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narbel:1993:BSS

[1090] P. Narbel. The boundary of substitution systems. Lecture Notes in Computer Science, 711:577–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neraud:1993:NAD

[1091] J. Neraud. New algorithms for detecting morphic images of a word. Lecture Notes in Computer Science, 711: 588–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

PAczkowski:1993:INI

[1092] P. P/Aczkowski. Ignoring nonessential interleavings in assertional reasoning on concurrent programs. Lecture Notes in Computer Science, 711:598– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parigot:1993:CTR

[1093] M. Parigot and P. Roziere. Constant time reductions in lambda-calculus. Lecture Notes in Computer Science, 711:608–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parisi-Presicce:1993:HUA

[1094] F. Parisi-Presicce and S. Veglioni. Heterogeneous unified algebra. Lecture Notes in Computer Science, 711:618– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pigozzi:1993:RTL

[1095] D. Pigozzi and A. Salibra. A representation theorem for lambda abstraction algebras. Lecture Notes in Computer Science, 711:629–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskevicius:1993:SCL

[1096] R. Pliuskevicius. On saturated calculi for a linear temporal logic. Lecture Notes in Computer Science, 711:640– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puhlmann:1993:SPD

[1097] H. Puhlmann. The snack powerdomain for database semantics. Lecture Notes in Computer Science, 711:650– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reus:1993:VPM

[1098] B. Reus and T. Streicher. Verifying properties of module construction in type theory. Lecture Notes in Computer Science, 711:660–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruzicka:1993:TST

[1099] P. Ruzicka and J. Waczulik. On timespace trade-offs in dynamic graph pebbling. Lecture Notes in Computer Science, 711:671–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sassone:1993:DBM

[1100] V. Sassone, M. Nielsen, and G. Winskel. Deterministic behavioural models for concurrency. Lecture Notes in Computer Science, 711:682–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scholefield:1993:RTR

[1101] D. Scholefield, H. Zedan, and Jifeng He. Real-time refinement: Semantics and application. Lecture Notes in Computer Science, 711:693–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steffen:1993:DTE

[1102] B. Steffen and C. Weise. Deciding testing equivalence for real-time processes with dense time. Lecture Notes in Computer Science, 711:703–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stephan:1993:CHO

[1103] W. Stephan and A. Wolpers. A calculus for higher order procedures with global variables. Lecture Notes in Computer Science, 711:714–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stevens:1993:VSI

[1104] D. Stevens. Variable substitution with iconic combinators. Lecture Notes in Computer Science, 711:724–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treinen:1993:FCF

[1105] R. Treinen. Feature constraints with first-class features. Lecture Notes in Computer Science, 711:734–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1993:BMC

[1106] D. Wagner and F. Wagner. Between min cut and graph bisection. Lecture Notes in Computer Science, 711:744– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wanke:1993:PCF

[1107] E. Wanke. Paths and cycles in finite periodic graphs. Lecture Notes in Computer Science, 711:751–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jilei:1993:LDL

[1108] Yin Jilei and Hong Zhu. Learning decision lists from noisy examples. Lecture Notes in Computer Science, 711:761–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zabel:1993:ATF

[1109] N. Zabel. Analytic tableaux for finite and infinite Post logics. Lecture Notes in Computer Science, 711:767– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:Sa

[1110] Anonymous. Session 1. Lecture Notes in Computer Science, 712:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffay:1993:ABE

[1111] K. Jeffay, D. L. Stone, T. Talley, and F. D. Smith. Adaptive, best-effort delivery of digital audio and video across packet-switched networks. Lecture Notes in Computer Science, 712:3–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smit:1993:ARR

[1112] Gerard J. M. Smit and Paul J. M. Havinga. The architecture of rattlesnake: A real-time multimedia network. Lecture Notes in Computer Science, 712:15–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrtwich:1993:BSI

[1113] Ralf Guido Herrtwich and Luca Delgrossi. Beyond ST-II: Fulfilling the requirements of multimedia communication. Lecture Notes in Computer Science, 712:25–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Needham:1993:ART

[1114] Roger Needham and Akira Nakamura. An approach to real-time scheduling — but is it really a problem for multimedia? Lecture Notes in Computer Science, 712:32–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:Sb

[1115] Anonymous. Session 2. Lecture Notes in Computer Science, 712:40–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1993:DAG

[1116] Philip S. Yu, Chen Mon-Song, and Dilip D. Kandlur. Design and analysis of a grouped sweeping scheme for multimedia storage management. Lecture Notes in Computer Science, 712:44–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vin:1993:ACA

[1117] Harrick M. Vin and P. Venkat Rangan. Admission control algorithms for multimedia on-demand servers. Lecture Notes in Computer Science, 712:56–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lougher:1993:DIC

[1118] Phillip Lougher and Doug Shepherd.

The design and implementation of a continuous media storage server. Lecture Notes in Computer Science, 712: 69–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nichols:1993:PSD

[1119] Kathleen M. Nichols. Performance studies of digital video in a client/ server environment. Lecture Notes in Computer Science, 712:81–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:Sc

[1120] Anonymous. Session 3. Lecture Notes in Computer Science, 712:92–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carmo:1993:BSC

[1121] Luiz F. Rust Da Costa Carmo, Pierre De Saqui-Sannes, and Jean-Pierre Courtiat. Basic synchronization concepts in muldimedia systems. Lecture Notes in Computer Science, 712:94–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rothermel:1993:SJV

[1122] Kurt Rothermel and Gabriel Dermler. Synchronization in joint-viewing environments. Lecture Notes in Computer Science, 712:106–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bulterman:1993:SMS

[1123] Dick C. A. Bulterman. Synchronization of multi-sourced multimedia data for heterogeneous target systems. Lecture Notes in Computer Science, 712: 119–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Little:1993:ISC

[1124] T. D. C. Little and F. Kao. An intermedia skew control system for multimedia data presentation. Lecture Notes in Computer Science, 712:130–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:Sd

[1125] Anonymous. Session 4. Lecture Notes in Computer Science, 712:142– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moran:1993:SSE

[1126] Mark Moran and Riccardo Gusella.

System support for efficient dynamicallyconfigurable multi-party interactive multimedia applications. Lecture Notes in Computer Science, 712:143–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

New:1993:RND

[1127] Darren New, Jonathan Rosenberg, Gil Cruz, and Tom Judd. Requirements for network delivery of stored interactive multimedia. Lecture Notes in Computer Science, 712:157–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baker:1993:MPM

[1128] Rusti Baker, Alan Downing, Kate Finn, Earl Rennison, Doohyun David Kim, and Hwan Lim Young. Multimedia processing model for a distributed multimedia I/O system. Lecture Notes in Computer Science, 712: 164–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arango:1993:ETM

[1129] Mauricio Arango, Michael Kramer, Steven L. Rohall, Lillian Ruston, and Abel Weinrib. Enhancing the touring machine API to support integrated digital transport. Lecture Notes in Computer Science, 712:176–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:Se

[1130] Anonymous. Session 5. Lecture Notes in Computer Science, 712:183– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hui:1993:PMR

[1131] Zhang Hui and Tom Fisher. Preliminary measurement of the RMTP/ RTIP. Lecture Notes in Computer Science, 712:185–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pasquale:1993:MMC

[1132] Joseph C. Pasquale, George C. Plyzos, Eric W. Anderson, and Vachaspathi P. Kompella. The multimedia multicast channel. Lecture Notes in Computer Science, 712:197–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liming:1993:ARM

[1133] Wei Liming, Liaw Fongching, Deborah Estrin, Allyn Romanow, and Tom Lyon. Analysis of a resequencer model for multicast over ATM networks. Lecture Notes in Computer Science, 712:209–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:Sf

[1134] Anonymous. Session 6. Lecture Notes in Computer Science, 712:221– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blair:1993:IPC

[1135] G. Blair, G. Coulson, P. Auzimour, L. Hazard, F. Horn, and J. B. Stefani. An integrated platform and computational model for open distributed multimedia applications. Lecture Notes in Computer Science, 712:223–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchanan:1993:SMD

[1136] M. Cecelia Buchanan and Polle T. Zellweger. Scheduling multimedia documents using temporal constraints. Lecture Notes in Computer Science, 712: 237–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1993:SSD

[1137] Thomas G. Aguierre Smith and Glorianna Davenport. The stratification system: A design environment for random access video. Lecture Notes in Computer Science, 712:250–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koegel:1993:DMI

[1138] John F. Koegel. On the design of multimedia interchange formats. Lecture Notes in Computer Science, 712:262– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:Sg

[1139] Anonymous. Session 7. Lecture Notes in Computer Science, 712:272– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wall:1993:BBM

[1140] Gerard A. Wall, James G. Hanko, and J. Duane Northcutt. Bus bandwidth management in a high resolution video workstation. Lecture Notes in Computer Science, 712:274–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Druschel:1993:ASD

[1141] Peter Druschel, Mark B. Abbott, Michael Pagels, and Larry L. Peterson. Analysis of I/O subsystem design for multimedia workstation. Lecture Notes in Computer Science, 712: 289–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dannenberg:1993:TTL

[1142] Roger B. Dannenberg, Tom Neuendorffer, Joseph M. Newcomer, and Dean Rubine. Tactus: Toolkit-level support for synchronized interactive multimedia. Lecture Notes in Computer Science, 712:302–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:SPSa

[1143] Anonymous. Short paper session I. Lecture Notes in Computer Science, 712: 314–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yee:1993:AMR

[1144] James Yee and Pravin Varaiya. An analytical model for real-time multimedia disk scheduling. Lecture Notes in Computer Science, 712:315–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1993:RTS

[1145] Tom Fisher. Real-time scheduling support in Ultrix-4.2 for multimedia communication. Lecture Notes in Computer Science, 712:321–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramanathan:1993:CMS

[1146] Srinivas Ramanathan and P. Venkat Rangan. Continuous media synchronization in distributed multimedia systems. Lecture Notes in Computer Science, 712:328–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:SPSb

[1147] Anonymous. Short paper session II.

Lecture Notes in Computer Science, 712:336–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherry:1993:HSN

[1148] Guy Cherry, Jim Nussbaum, and Mayer Schwartz. High speed networks and the digital TV studio. Lecture Notes in Computer Science, 712:337– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schooler:1993:ISM

[1149] Eve M. Schooler. The impact of scaling on a multimedia connection architecture. Lecture Notes in Computer Science, 712:341–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:SPSc

[1150] Anonymous. Short paper session III.

Lecture Notes in Computer Science, 712:347–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jamin:1993:ACA

[1151] Sugih Jamin, Scott Shenker, Lixia Zhang, and David D. Clark. An admission control algorithm for predictive real-time service (extended abstract). Lecture Notes in Computer Science, 712:349–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ingold:1993:MMA

[1152] Daniel P. Ingold. MEGAPHONE: A multimedia application based on object-oriented communication. Lecture Notes in Computer Science, 712: 357–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chou:1993:SSD

[1153] Stephen T.-C. Chou and Hideyuki Tokuda. System support for dynamic QOS control of continuous media communication. Lecture Notes in Computer Science, 712:363–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:SPSd

[1154] Anonymous. Short paper session IV. Lecture Notes in Computer Science, 712:369–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patel:1993:NNM

[1155] Sameer Patel, Ghaleb Abdulla, Marc Abrams, and Edward A. Fox. NMFS: Network multimedia file system protocol. Lecture Notes in Computer Science, 712:370–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rowe:1993:CMP

[1156] Lawrence A. Rowe and Brian C. Smith.

A continuous media player. Lecture Notes in Computer Science, 712:376– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swanberg:1993:AMI

[1157] Deborah Swanberg, Fe Shu Chiao, and Ramesh Jain. Architecture of a multimedia information system for contentbased retrieval. Lecture Notes in Computer Science, 712:387–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:SPSe

[1158] Anonymous. Short paper session V.

Lecture Notes in Computer Science, 712:393–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gibbs:1993:ACC

[1159] Simon Gibbs. Application construction and component design in an objectoriented multimedia framework. Lecture Notes in Computer Science, 712: 394–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamakawa:1993:AVE

[1160] Rei Hamakawa, Hidekazu Sakagami, and Jun Rekimoto. Audio and video extensions to graphical user interface toolkits. Lecture Notes in Computer Science, 712:399–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappner:1993:IHH

[1161] Thomas Kappner, Dietmar Hehmann, and Ralf Steinmetz. An introduction to HeiMAT: The Heidelberg Multimedia Application Toolkit. Lecture Notes in Computer Science, 712:405–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borger:1993:MSP

[1162] Egon Borger and Dean Rosenzweig. The mathematics of set predicates in Prolog. Lecture Notes in Computer Science, 713:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cowen:1993:SCB

[1163] Robert Cowen. Some connections between set theory and computer science. Lecture Notes in Computer Science, 713:14–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feferman:1993:GDI

[1164] Solomon Feferman. Godel’s dialectica interpretation and its two-way stretch. Lecture Notes in Computer Science, 713:23–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hajek:1993:EEA

[1165] Petr Hajek. Epistemic entrenchment and arithmetical hierarchy (abstract). Lecture Notes in Computer Science, 713:41–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halpern:1993:CRD

[1166] Joseph Y. Halpern. A critical reexamination of default logic, autoepistemic logic, and only knowing. Lecture Notes in Computer Science, 713:43–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marek:1993:CIN

[1167] V. Wiktor Marek. Complexity issues in nonmonotonic logic and logic programming (abstract). Lecture Notes in Computer Science, 713:61–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mints:1993:SRM

[1168] Grigori Mints. Strategies for resolution method in non-classical logics (abstract). Lecture Notes in Computer Science, 713:62–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacholski:1993:UIP

[1169] Leszek Pacholski. Undecidability of implication problems in logic programming, data base theory and classical logic. Lecture Notes in Computer Science, 713:63–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sambin:1993:BTB

[1170] Giovanni Sambin and Silvio Valentini.

Building up a tool-box for Martin-Löf’s type theory (abstract). Lecture Notes in Computer Science, 713:69–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artemov:1993:LGP

[1171] Sergei Artemov and Tyko Straben. The logic of the godel proof predicate. Lecture Notes in Computer Science, 713:71–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1993:SSD

[1172] Leo Bachmair, Harald Ganzinger, and Uwe Waldmann. Superposition with simplification as a decision procedure for the monadic class with equality. Lecture Notes in Computer Science, 713:83–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bamber:1993:CAR

[1173] Nicholas P. Bamber. Computation with access to the reals, but using only classical machines. Lecture Notes in Computer Science, 713:97–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckert:1993:EML

[1174] Bernhard Beckert, Reiner Hahnle, and Peter H. Schmitt. The even more liberalized delta-rule in free variable semantic tableaux. Lecture Notes in Computer Science, 713:108–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benevides:1993:DAE

[1175] Mario R. F. Benevides. Differentiating assumptions from extra-logical axioms in natural deduction. Lecture Notes in Computer Science, 713:120–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bottoni:1993:IFF

[1176] Alberto Bottoni and Giorgio Levi. The inverse of Fitting’s functional. Lecture Notes in Computer Science, 713:132– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruning:1993:LDC

[1177] Stefan Bruning. On loop detection in connection calculi. Lecture Notes in Computer Science, 713:144–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DaCosta:1993:AHS

[1178] Newton C. A. Da Costa and Francisco A. Doria. On Arnold’s Hilbert symposium problems. Lecture Notes in Computer Science, 713:152–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danos:1993:SEU

[1179] Vincent Danos, Jean-Baptiste Joinet, and Harold Schellinx. The structure of exponentials: Uncovering the dynamics of linear logic proofs. Lecture Notes in Computer Science, 713:159–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egly:1993:DCF

[1180] Uwe Egly. On different concepts of function introduction. Lecture Notes in Computer Science, 713:172–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faglia:1993:DEI

[1181] Giovanni Faglia. Double exponential inseparability of Robinson subsystem Q_{+} from the unsatisfiable sentences in the language of addition. Lecture Notes in Computer Science, 713:184– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faglia:1993:MEU

[1182] Giovanni Faglia and Paul Young. On the meaning of essentially unprovable theorems in the Presburger theory of addition (abstract). Lecture Notes in Computer Science, 713:187–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilmore:1993:SCP

[1183] Paul C. Gilmore. A syntactic consistency proof for NaDSet. Lecture Notes in Computer Science, 713:190– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault:1993:RBA

[1184] Jean Goubault. A rule-based algorithm for rigid E-unification. Lecture Notes in Computer Science, 713:202–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janhunen:1993:SWN

[1185] Tomi Janhunen and Ilkka Niemela. A scheme for weakened negative introspection in autoepistemic reasoning. Lecture Notes in Computer Science, 713:211–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johannsen:1993:WSB

[1186] Jan Johannsen. On the weakness of sharply bounded polynomial induction. Lecture Notes in Computer Science, 713:223–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolany:1993:LH

[1187] Adam Kolany. On the logic of hypergraphs. Lecture Notes in Computer Science, 713:231–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kummer:1993:RTP

[1188] Martin Kummer and Frank Stephan.

Recursion theoretic properties of frequency computation and bounded queries (extended abstract). Lecture Notes in Computer Science, 713:243– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nies:1993:ITA

[1189] Andre Nies. Interpreting true arithmetic in degree structures. Lecture Notes in Computer Science, 713:255– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parigot:1993:CPP

[1190] Michel Parigot. Classical proofs as programs. Lecture Notes in Computer Science, 713:263–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petermann:1993:CPC

[1191] Uwe Petermann. Completeness of the pool calculus with an open built-in theory. Lecture Notes in Computer Science, 713:277–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskevicius:1993:SPL

[1192] Regimantas Pliuskevicius. On the saturation principle for a linear temporal logic. Lecture Notes in Computer Science, 713:289–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sazonov:1993:CTL

[1193] Vladimir Sazonov and Andrei Voronkov. A construction of typed lambda models related to feasible computability. Lecture Notes in Computer Science, 713:301–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwarz:1993:NRS

[1194] Grigori Schwarz and Miroslaw Truszczynski. Nonmonotonic reasoning is sometimes simpler. Lecture Notes in Computer Science, 713:313–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willard:1993:SVA

[1195] Dan E. Willard. Self-verifying axiom systems. Lecture Notes in Computer Science, 713:325–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zlatuska:1993:CCC

[1196] Jiri Zlatuska. Committed-choice concurrent logic programming in linear logic. Lecture Notes in Computer Science, 713:337–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kastens:1993:ESL

[1197] U. Kastens. Executable specifications for language implementation. Lecture Notes in Computer Science, 714:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boye:1993:ADD

[1198] J. Boye. Avoiding dynamic delays in functional logic programs. Lecture Notes in Computer Science, 714:12–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanus:1993:DMF

[1199] M. Hanus and B. Josephs. A debugging model for functional logic programs. Lecture Notes in Computer Science, 714:28–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1993:CAM

[1200] P. Lim and J. Schimpf. A conservative approach to meta-programming in constraint logic programming. Lecture Notes in Computer Science, 714:44–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jourdan:1993:VHD

[1201] J. Jourdan and T. Sola. The versatility of handling disjunctions as constraints. Lecture Notes in Computer Science, 714:60–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corsini:1993:EBA

[1202] M.-M. Corsini, K. Musumbu, A. Rauzy, and B. L. Charlier. Efficient bottomup abstract interpretation of Prolog by means of constraint solving over symbolic finite domains. Lecture Notes in Computer Science, 714:75–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kagedal:1993:ICT

[1203] A. Kagedal. Improvements in compiletime analysis for ground Prolog. Lecture Notes in Computer Science, 714: 92–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nederhof:1993:NTP

[1204] M.-J. Nederhof. A new top-down parsing algorithm for left-recursive DCGs. Lecture Notes in Computer Science, 714:108–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jourdan:1993:SIG

[1205] M. Jourdan, C. L. Bellec, D. Parigot, and G. Roussel. Specification and implementation of grammar couplings using attribute grammars. Lecture Notes in Computer Science, 714:123– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poetzsch-Heffter:1993:PLS

[1206] A. Poetzsch-Heffter. Programming language specification and prototyping using the MAX system. Lecture Notes in Computer Science, 714:137–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mantsivoda:1993:FI

[1207] A. Mantsivoda. Flang and its implementation. Lecture Notes in Computer Science, 714:151–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno-Navarro:1993:ELN

[1208] J. J. Moreno-Navarro, H. Kuchen, J. Marino-Carballo, S. Winkler, and W. Hans. Efficient lazy narrowing using demandedness analysis. Lecture Notes in Computer Science, 714:167– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loogen:1993:DDC

[1209] R. Loogen, F. L. Fraguas, and M. Rodrigues-Artalejo. A demand driven computation strategy for lazy narrowing. Lecture Notes in Computer Science, 714:184–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darlington:1993:FPL

[1210] J. Darlington, Y. Guo, and M. Kohler.

Functional programming languages with logical variables: A linear logic view. Lecture Notes in Computer Science, 714:201–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Natali:1993:OSC

[1211] A. Natali and A. Omicini. Objects with state in contextual logic programming. Lecture Notes in Computer Science, 714:220–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ali:1993:NMP

[1212] K. A. M. Ali and R. Karlsson. A novel method for parallel implementation of findall. Lecture Notes in Computer Science, 714:235–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moolenaar:1993:PIA

[1213] R. Moolenaar and B. Demoen. A parallel implementation for AKL. Lecture Notes in Computer Science, 714:246– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaser:1993:IRS

[1214] O. Kaser. Inlining to reduce stack space. Lecture Notes in Computer Science, 714:262–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dovier:1993:WBI

[1215] A. Dovier and E. Pontelli. A WAM-Based implementation of a logic language with sets. Lecture Notes in Computer Science, 714:275–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benjumea:1993:PPM

[1216] V. Benjumea and J. M. Troya. An OR parallel Prolog model for distributed memory systems. Lecture Notes in Computer Science, 714:291–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barklund:1993:EBQ

[1217] J. Barklund and J. Bevemyr. Executing bounded quantifications on shared memory multiprocessors. Lecture Notes in Computer Science, 714: 302–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clark:1993:LAG

[1218] D. Clark and C. Hankin. A lattice of abstract graphs. Lecture Notes in Computer Science, 714:318–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosendahl:1993:HOC

[1219] M. Rosendahl. Higher-order chaotic iteration sequences. Lecture Notes in Computer Science, 714:332–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burn:1993:PCC

[1220] G. Burn and D. L. Metayer. Proving the correctness of compiler optimisations based on strictness analysis. Lecture Notes in Computer Science, 714:346–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dikovsky:1993:ACP

[1221] A. Y. Dikovsky. Abstract complexity of Prolog based on WAM. Lecture Notes in Computer Science, 714:365– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Letichevsky:1993:DRS

[1222] A. Letichevsky. Development of rewriting strategies. Lecture Notes in Computer Science, 714:378–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alpuente:1993:NAO

[1223] M. Alpuente, M. Falaschi, M. J. Ramis, and G. Vidal. Narrowing approximations as an optimization for equational logic programs. Lecture Notes in Computer Science, 714:391–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canalda:1993:PBE

[1224] P. Canalda, L. Cognard, A. Despland, and M. Mazaud. PAGODE: A back end generator. Lecture Notes in Computer Science, 714:410–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bugliesi:1993:SLI

[1225] M. Bugliesi and G. Nardiello. SelfLog: Language and implementation. Lecture Notes in Computer Science, 714:412– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jourdan:1993:EDS

[1226] M. Jourdan, F. Lagnier, F. Maraninchi, and P. Raymond. Embedding declarative subprograms into imperative constructs. Lecture Notes in Computer Science, 714:414–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hogen:1993:SMR

[1227] G. Hogen and R. Loogen. Stack management of runtime structures in distributed implementations. Lecture Notes in Computer Science, 714:416– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kebler:1993:ERA

[1228] C. W. Kebler and T. Rauber. Efficient register allocation for large basic blocks. Lecture Notes in Computer Science, 714:418–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:1993:GSC

[1229] M. Alt, G. Sander, and R. Wilhelm.

Generation of synchronization code for parallel compilers. Lecture Notes in Computer Science, 714:420–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudol:1993:ITL

[1230] Gerard Boudol. Invited talk: The lambda-calculus with multiplicities (extended abstract). Lecture Notes in Computer Science, 715:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1993:MSP

[1231] Joost Engelfriet. A multiset semantics for the pi-calculus with replication. Lecture Notes in Computer Science, 715:7–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dam:1993:MCM

[1232] Mads Dam. Model checking mobile processes. Lecture Notes in Computer Science, 715:22–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruns:1993:PTP

[1233] Glenn Bruns. A practical technique for process abstraction. Lecture Notes in Computer Science, 715:37–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fokkinga:1993:MCC

[1234] Maarten Fokkinga, Mannes Poel, and Job Zwiers. Modular completeness for communication closed layers. Lecture Notes in Computer Science, 715:50–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGlabbeek:1993:LTB

[1235] Rob J. Van Glabbeek. The linear time — branching time spectrum II (the semantics of sequential systems with silent moves). Lecture Notes in Computer Science, 715:66–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sassone:1993:CMC

[1236] Vladimiro Sassone, Mogens Nielsen, and Glynn Winskel. A classification of models for concurrency. Lecture Notes in Computer Science, 715:82–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aceto:1993:ITW

[1237] Luca Aceto and David Murphy. On the ill-timed but well-caused. Lecture Notes in Computer Science, 715:97–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amadio:1993:RCB

[1238] Roberto M. Amadio. On the reduction of chocs bisimulation to picalculus bisimulation. Lecture Notes in Computer Science, 715:112–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sangiorgi:1993:TBP

[1239] Davide Sangiorgi. A theory of bisimulation for the pi-calculus. Lecture Notes in Computer Science, 715:127– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1993:BED

[1240] Soren Christensen, Yoram Hirshfeld, and Faron Moller. Bisimulation equivalence is decidable for basic parallel processes. Lecture Notes in Computer Science, 715:143–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1993:ITP

[1241] Cliff B. Jones. Invited talk: A picalculus semantics for an object-based design notation. Lecture Notes in Computer Science, 715:158–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prasad:1993:PB

[1242] K. V. S. Prasad. Programming with broadcasts. Lecture Notes in Computer Science, 715:173–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holmer:1993:IBC

[1243] Uno Holmer. Interpreting broadcast communication in SCCS. Lecture Notes in Computer Science, 715:188– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hennessy:1993:PSM

[1244] Matthew Hennessy and Lin Huimin. Proof systems for message-passing process algebras. Lecture Notes in Computer Science, 715:202–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1993:RDV

[1245] Michael J. Butler. Refinement and decomposition of value-passing action systems. Lecture Notes in Computer Science, 715:217–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolper:1993:ITP

[1246] Pierre Wolper and Patrice Godefroid. Invited talk: Partial-order methods for temporal verification. Lecture Notes in Computer Science, 715:233–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1993:MCI

[1247] Ole Hogh Jensen, Christian Jeppesen, Jarl Tuxen Lang, and Kim G. Larsen. Model construction for implicit specifications in modal logic. Lecture Notes in Computer Science, 715:247–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernholtz:1993:BTT

[1248] Orna Bernholtz and Orna Grumberg. Branching time temporal logic and amorphous tree automata. Lecture Notes in Computer Science, 715:262– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gunawardena:1993:GES

[1249] Jeremy Gunawardena. A generalized event structure for the Muller unfolding of a safe net. Lecture Notes in Computer Science, 715:278–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goubault:1993:DHD

[1250] Eric Goubault. Domains of higherdimensional automata. Lecture Notes in Computer Science, 715:293–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1993:ITN

[1251] Jos C. M. Baeten and Jan A. Bergstra.

Invited talk: Non interleaving process algebra. Lecture Notes in Computer Science, 715:308–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Segala:1993:QFT

[1252] Roberto Segala. Quiescence, fairness, testing, and the notion of implementation. Lecture Notes in Computer Science, 715:324–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Udink:1993:TFA

[1253] Rob T. Udink and Joost N. Kok. Two fully abstract models for UNITY. Lecture Notes in Computer Science, 715: 339–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shengzong:1993:TPP

[1254] Zhou Shengzong, Rob Gerth, and Ruurd Kuiper. Transformations preserving properties and properties preserved by transformations in fair transition systems. Lecture Notes in Computer Science, 715:353–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cubric:1993:MMS

[1255] Marija Cubric and Prakash Panangaden. Minimal memory schedules for dataflow networks. Lecture Notes in Computer Science, 715:368–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yates:1993:NRT

[1256] Robert Kim Yates. Networks of realtime processes. Lecture Notes in Computer Science, 715:384–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lengauer:1993:ITL

[1257] Christian Lengauer. Invited talk: Loop parallelization in the polytope model. Lecture Notes in Computer Science, 715:398–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bremond-Gregoire:1993:AAC

[1258] Patrice Bremond-Gregoire, Insup Lee, and Richard Gerber. ACSR: An algebra of communicating shared resources with dense time and priorities. Lecture Notes in Computer Science, 715:417–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fokkink:1993:ETR

[1259] Willem Jan Fokkink. An elimination theorem for regular behaviours with integration. Lecture Notes in Computer Science, 715:432–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vergauwen:1993:LLM

[1260] Bart Vergauwen and Johan Lewi. A linear local model checking algorithm for CTL. Lecture Notes in Computer Science, 715:447–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogers:1993:LES

[1261] P. W. Hoogers, H. C. M. Kleijn, and P. S. Thiagarajan. Local event structures and Petri nets. Lecture Notes in Computer Science, 715:462–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baeten:1993:CTS

[1262] Jos C. M. Baeten and Chris Verhoef. A congruence theorem for structured operational semantics with predicates. Lecture Notes in Computer Science, 715:477–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielson:1993:CPA

[1263] Flemming Nielson and Hanne Riis Nielson. From CML to process algebras. Lecture Notes in Computer Science, 715:493–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honda:1993:TDI

[1264] Kohei Honda. Types for dyadic interaction. Lecture Notes in Computer Science, 715:509–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasconcelos:1993:PTS

[1265] Vasco T. Vasconcelos and Kohei Honda. Principal typing schemes in a polyadic pi–calculus. Lecture Notes in Computer Science, 715:524–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloom:1993:ITS

[1266] Bard Bloom. Invited talk: Structured operational semantics for process algebras and equational axiom systems (abstract). Lecture Notes in Computer Science, 715:539–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maab:1993:CMP

[1267] Wolfgang Maab. A cognitive model for the process of multimodal, incremental route descriptions. Lecture Notes in Computer Science, 716:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tversky:1993:CMC

[1268] Barbara Tversky. Cognitive maps, cognitive collages, and spatial mental models. Lecture Notes in Computer Science, 716:14–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vieu:1993:LFR

[1269] Laure Vieu. A logical framework for reasoning about space. Lecture Notes in Computer Science, 716:25–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hernandez:1993:MQS

[1270] Daniel Hernandez. Maintaining qualitative spatial knowledge. Lecture Notes in Computer Science, 716:36–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ligozat:1993:QTS

[1271] Gerard F. Ligozat. Qualitative triangulation for spatial reasoning. Lecture Notes in Computer Science, 716:54–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zimmermann:1993:EQS

[1272] Kai Zimmermann. Enhancing qualitative spatial reasoning — combining orientation and distance. Lecture Notes in Computer Science, 716:69–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pratt:1993:MS

[1273] Ian Pratt. Map semantics. Lecture Notes in Computer Science, 716:77–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramirez:1993:DCL

[1274] J. Raul Ramirez. Development of a cartographic language. Lecture Notes in Computer Science, 716:92–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deFloriani:1993:SQD

[1275] Leila de Floriani, Paola Marzano, and Enrico Puppo. Spatial queries and data models. Lecture Notes in Computer Science, 716:113–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deHoop:1993:TQM

[1276] Sylvia de Hoop, Peter Van Oosterom, and Martien Molenaar. Topological querying of multiple map layers. Lecture Notes in Computer Science, 716: 139–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheylan:1993:TCD

[1277] Jean-Paul Cheylan and Sylvie Lardon. Towards a conceptual data model for the analysis of spatio-temporal processes: The example of the search for optimal grazing strategies. Lecture Notes in Computer Science, 716:158– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirtle:1993:CSS

[1278] Stephen C. Hirtle, Thea Ghiselli-Crippa, and Michael B. Spring. The cognitive structure of space: An analysis of temporal sequences. Lecture Notes in Computer Science, 716:177– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whigham:1993:HST

[1279] Peter A. Whigham. Hierarchies of space and time. Lecture Notes in Computer Science, 716:190–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edwards:1993:VMC

[1280] Geoffrey Edwards. The Voronoi model and cultural space: Applications to the social sciences and humanities. Lecture Notes in Computer Science, 716:

202–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Volta:1993:IGA

[1281] Gary S. Volta and Max J. Egenhofer.

Interaction with GIS attribute data based on categorical coverage. Lecture Notes in Computer Science, 716:215– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papadias:1993:SRS

[1282] Dimitris Papadias and Timos Sellis.

The semantics of relations in 2D space using representative points: Spatial indexes. Lecture Notes in Computer Science, 716:234–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deFloriani:1993:CVM

[1283] Leila de Floriani and Paola Magillo.

Computing visibility maps on a digital terrain model. Lecture Notes in Computer Science, 716:248–269, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mark:1993:TTF

[1284] David M. Mark. Toward a theoretical framework for geographic entity types. Lecture Notes in Computer Science, 716:270–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guttenberg:1993:LSS

[1285] Albert Z. Guttenberg. Land, space and spatial planning in three time regions. Lecture Notes in Computer Science, 716:284–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pederson:1993:GMS

[1286] Eric Pederson. Geographic and manipulable space in two tamil linguistic systems. Lecture Notes in Computer Science, 716:294–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montello:1993:SMP

[1287] Daniel R. Montello. Scale and multiple psychologies of space. Lecture Notes in Computer Science, 716:312–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Getis:1993:GMP

[1288] Arthur Getis. GIS and modeling prerequisites. Lecture Notes in Computer Science, 716:322–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rigaux:1993:MEK

[1289] Philippe Rigaux, Michel Scholl, and Agnes Voisard. A map editing kernel implementation: Application to multiple scale display. Lecture Notes in Computer Science, 716:341–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuhn:1993:MCT

[1290] Werner Kuhn. Metaphors create theories for users. Lecture Notes in Computer Science, 716:366–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chalmers:1993:ULM

[1291] Matthew Chalmers. Using a landscape metaphor to represent a corpus of documents. Lecture Notes in Computer Science, 716:377–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erickson:1993:IIS

[1292] Thomas Erickson. From interface to interplace: The spatial environment as a medium for interaction. Lecture Notes in Computer Science, 716:391– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haunold:1993:KLA

[1293] Peter Haunold and Werner Kuhn. A keystroke level analysis of manual map digitizing. Lecture Notes in Computer Science, 716:406–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yue-Hong:1993:CIE

[1294] Chou Yue-Hong. Critical issues in the evaluation of spatial autocorrelation. Lecture Notes in Computer Science, 716:421–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhan:1993:DPD

[1295] Cixiang Zhan, Sudhakar Menon, and Gao Peng. A directional path distance model for raster distance mapping. Lecture Notes in Computer Science, 716:434–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jungert:1993:SSR

[1296] Erland Jungert. Symbolic spatial reasoning in object shapes for qualitative reasoning. Lecture Notes in Computer Science, 716:444–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saunders:1993:RAS

[1297] Claude W. Saunders. Reasoning about spatial structure in landscapes with Geographic Information Systems. Lecture Notes in Computer Science, 716: 463–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:XP

[1298] Anonymous. XII. posters. Lecture Notes in Computer Science, 716:478– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buxton:1993:DCP

[1299] J. N. Buxton. On the decline of classical programming. Lecture Notes in Computer Science, 717:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harel:1993:CO

[1300] D. Harel. Computers are not omnipotent. Lecture Notes in Computer Science, 717:10–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghezzi:1993:RTS

[1301] C. Ghezzi, M. Felder, and C. Bellettini. Real-time systems: A survey of approaches to formal specification and verification. Lecture Notes in Computer Science, 717:11–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denert:1993:SEB

[1302] E. Denert. Software engineering in business and academia: How wide is the gap? Lecture Notes in Computer Science, 717:37–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perry:1993:SFE

[1303] D. E. Perry and C. S. Stieg. Software faults in evolving a large, real-time system. Lecture Notes in Computer Science, 717:48–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basili:1993:EFR

[1304] V. R. Basili. The experience factory and its relationship to other improvement paradigm. Lecture Notes in Computer Science, 717:68–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finkelstein:1993:IHM

[1305] A. Finkelstein, D. Gabbay, A. Hunter, J. Kramer, and B. Nuseibeh. Inconsistency handling in multi-perspective specifications. Lecture Notes in Computer Science, 717:84–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarke:1993:REI

[1306] M. Jarke, K. Pohl, S. Jacobs, J. Bubenko, P. Assenova, P. Holm, B. Wangler, C. Rolland, V. Plihon, and J.-R. Schmitt. Requirements engineering: An integrated view of representation, process and domain. Lecture Notes in Computer Science, 717:100– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bustard:1993:MCF

[1307] D. W. Bustard and A. C. Winstanley. Making changes to formal specifications: Requirements and an example. Lecture Notes in Computer Science, 717:115–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagelstein:1993:FRM

[1308] J. Hagelstein, D. Roelants, and P. Wodon. Formal requirements made practical. Lecture Notes in Computer Science, 717:127–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emmerich:1993:DSE

[1309] W. Emmerich, W. Schafer, and J. Welsh. Databases for software engineering environments the goal has not yet been attained. Lecture Notes in Computer Science, 717:145–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1993:RTD

[1310] L. Liu, D. J. Robson, and R. Ellis. A regression testing database model. Lecture Notes in Computer Science, 717:163–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1993:EFE

[1311] A. W. Brown, E. J. Morris, P. F.

Zarrella, F. W. Long, and W. M. Caldwell. Experiences with a federated environment testbed. Lecture Notes in Computer Science, 717:175–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1993:OOM

[1312] I. Thomas. Observations on object management systems and process support in environments. Lecture Notes in Computer Science, 717:197–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eggers:1993:STD

[1313] H. Eggers. Software technology for a distributed telecommunication system. Lecture Notes in Computer Science, 717:213–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cederling:1993:ISD

[1314] U. Cederling. Industrial software development. Lecture Notes in Computer Science, 717:226–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuntzmann-Combelles:1993:QAS

[1315] A. Kuntzmann-Combelles. Quantitative approach to software management: The ami method. Lecture Notes in Computer Science, 717:238–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reichardt:1993:PSE

[1316] J. Reichardt. Preventative software engineering. Lecture Notes in Computer Science, 717:251–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aue:1993:DIS

[1317] A. Aue and M. Breu. Distributed information systems: An advanced methodology. Lecture Notes in Computer Science, 717:263–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheung:1993:TFA

[1318] S. C. Cheung and J. Kramer. Tractable flow analysis for anomaly detection in distributed programs. Lecture Notes in Computer Science, 717:283–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuya:1993:PTD

[1319] J. Tuya, L. Sanchez, R. Zurita, and J. A. Corrales. A pragmatic task design approach base on a ward/mellor real-time structured specification. Lecture Notes in Computer Science, 717: 301–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonDerBeeck:1993:ISA

[1320] M. von Der Beeck. Integration of structured analysis and timed statecharts for real-time and concurrency specification. Lecture Notes in Computer Science, 717:313–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Depaoli:1993:LCC

[1321] F. Depaoli and F. Tisato. Language constructs for cooperative systems design. Lecture Notes in Computer Science, 717:329–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gintell:1993:SCI

[1322] J. Gintell, J. Arnold, M. Houde, J. Kruszelnicki, R. McKenney, and G. Memmi. Scrutiny: A collaborative inspection and review system. Lecture Notes in Computer Science, 717:344– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godart:1993:CTM

[1323] C. Godart. COO: A transaction model to support COOperating software developers COOrdination. Lecture Notes in Computer Science, 717: 361–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adams:1993:ESR

[1324] R. Adams. An experiment in software retrieval. Lecture Notes in Computer Science, 717:380–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeng:1993:UFM

[1325] J.-J. Jeng and B. H. C. Cheng. Using formal methods to construct a software component library. Lecture Notes in Computer Science, 717:397–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gall:1993:COR

[1326] H. Gall and R. Klosch. Capsule oriented reverse engineering for software reuse. Lecture Notes in Computer Science, 717:418–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1993:ART

[1327] C. Liu and R. Conradi. Automatic replanning of task networks for process model evolution in EPOS. Lecture Notes in Computer Science, 717: 434–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnamurthy:1993:PPV

[1328] B. Krishnamurthy and N. S. Barghouti. Provence: A process visualisation and enactment environment. Lecture Notes in Computer Science, 717: 451–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breiteneder:1993:PPA

[1329] C. J. Breiteneder and T. A. Mueck. Process programming with active and passive components. Lecture Notes in Computer Science, 717:466–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coen-Porisini:1993:FFA

[1330] A. Coen-Porisini, R. A. Kemmerer, and D. Mandrioli. A formal framework for ASTRAL intra-level proof obligations. Lecture Notes in Computer Science, 717:483–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourdoncle:1993:ABD

[1331] F. Bourdoncle. Assertion-based debugging of imperative programs by abstract interpretation. Lecture Notes in Computer Science, 717:501–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desmedt:1993:TC

[1332] Y. Desmedt. Threshold cryptosystems. Lecture Notes in Computer Science, 718:3–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tombak:1993:ACP

[1333] L. Tombak and R. Safavi-Naini. Authentication codes with perfect protection. Lecture Notes in Computer Science, 718:15–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desmedt:1993:PPS

[1334] Y. Desmedt and J. Seberry. Practical proven secure authentication with arbitration. Lecture Notes in Computer Science, 718:27–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Safavi-Naini:1993:ACU

[1335] R. Safavi-Naini and L. Tombak. Authentication codes under impersonation attack. Lecture Notes in Computer Science, 718:35–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1993:CAG

[1336] W.-A. Jackson and K. M. Martin.

Cumulative arrays and geometric secret sharing schemes. Lecture Notes in Computer Science, 718:48–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ogata:1993:NSS

[1337] W. Ogata, K. Kurosawa, and S. Tsujii. Nonperfect secret sharing schemes. Lecture Notes in Computer Science, 718:56–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertilsson:1993:CPS

[1338] M. Bertilsson and I. Ingemarsson. A construction of practical secret sharing schemes using linear block codes. Lecture Notes in Computer Science, 718: 67–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1993:HOW

[1339] Y. Zheng, J. Pieprzyk, and J. Seberry. HAVAL — a one-way hashing algorithm with variable length of output. Lecture Notes in Computer Science, 718:83–104, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preneel:1993:PMD

[1340] B. Preneel, R. Govaerts, and J. Vandewalle. On the power of memory in the design of collision resistant hash functions. Lecture Notes in Computer Science, 718:105–121, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hardjono:1993:PDM

[1341] T. Hardjono and Y. Zheng. A practical digital multisignature scheme based on discrete logarithms. Lecture Notes in Computer Science, 718:122–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harn:1993:GOU

[1342] L. Harn and S. Yang. Group-oriented undeniable signature schemes without the assistance of a mutually trusted party. Lecture Notes in Computer Science, 718:133–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seberry:1993:HNB

[1343] J. Seberry and X.-M. Zhang. Highly nonlinear 0-1 balanced Boolean functions satisfying strict avalanche criterion. Lecture Notes in Computer Science, 718:145–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charnes:1993:LNV

[1344] C. Charnes and J. Pieprzyk. Linear nonequivalence versus nonlinearity. Lecture Notes in Computer Science, 718:156–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Detombe:1993:CLC

[1345] J. Detombe and S. E. Tavares.

Constructing large cryptographically strong S-boxes. Lecture Notes in Computer Science, 718:165–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korzhik:1993:NEI

[1346] V. Korzhik and V. Yakovlev. Nonasymptotic estimates of information protection efficiency for the wire-tape channel concept. Lecture Notes in Computer Science, 718:185–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knudsen:1993:CLa

[1347] L. R. Knudsen. Cryptanalysis of LOKI91. Lecture Notes in Computer Science, 718:196–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawson:1993:CSG

[1348] E. Dawson. Cryptanalysis of summation generator. Lecture Notes in Computer Science, 718:209–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laih:1993:SAS

[1349] C.-S. Laih and S.-M. Yen. Secure addition sequence and its applications on the server-aided secret computation protocols. Lecture Notes in Computer Science, 718:219–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakurai:1993:SCS

[1350] K. Sakurai and T. Itoh. Subliminal channels for signature transfer and their application to signature distribution schemes. Lecture Notes in Computer Science, 718:231–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujioka:1993:PSV

[1351] A. Fujioka, T. Okamoto, and K. Ohta. A practical secret voting scheme for large scale elections. Lecture Notes in Computer Science, 718:244–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Satoh:1993:PMP

[1352] T. Satoh, K. Kurosawa, and S. Tsujii. Privacy for multi-party protocols. Lecture Notes in Computer Science, 718:252–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pailles:1993:NPE

[1353] J. C. Pailles. New protocols for electronic money. Lecture Notes in Computer Science, 718:263–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nieh:1993:MAC

[1354] B. B. Nieh and S. E. Tavares. Modelling and analyzing cryptographic protocols using Petri nets. Lecture Notes in Computer Science, 718:275–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsumoto:1993:VIA

[1355] T. Matsumoto, H. Imai, C.-S. Laih, and S.-M. Yen. On verifiable implicit asking protocols for RSA computation. Lecture Notes in Computer Science, 718:296–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1993:MMY

[1356] C. H. Lim and P. J. Lee. Modified Maurer-Yacobi’s scheme and its applications. Lecture Notes in Computer Science, 718:308–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klapper:1993:VGS

[1357] A. Klapper. The vulnerability of geometric sequences based on fields of odd characteristic. Lecture Notes in Computer Science, 718:327–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1993:FCC

[1358] X. Lai, R. A. Rueppel, and J. Woollven. A fast cryptographic checksum algorithm based on stream ciphers. Lecture Notes in Computer Science, 718:339–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mihaljevic:1993:AIS

[1359] M. J. Mihaljevic. An approach to the initial state reconstruction of a clock-controlled shift register based on a novel distance measure. Lecture Notes in Computer Science, 718:349– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1993:CAB

[1360] J.-H. Yang and Z.-D. Dai. Construction of m-ary de Bruijn sequences. Lecture Notes in Computer Science, 718: 357–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snare:1993:ITS

[1361] J. Snare. Information technology security standards — an Australian perspective. Lecture Notes in Computer Science, 718:367–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerecedo:1993:NIG

[1362] M. Cerecedo, T. Matsumoto, and H. Imai. Non-interactive generation of shared pseudorandom sequences. Lecture Notes in Computer Science, 718: 385–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Portz:1993:GDB

[1363] M. Portz. A generalized description of DES-based and Benes-based permutation generators. Lecture Notes in Computer Science, 718:397–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Condie:1993:PGD

[1364] L. Condie. Prime generation with the Demytko-Miller-Trbovich algorithm. Lecture Notes in Computer Science, 718:413–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiltgen:1993:CFO

[1365] A. P. L. Hiltgen. Construction of feebly-one-way families of permutations. Lecture Notes in Computer Science, 718:422–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakurai:1993:BCA

[1366] K. Sakurai and T. Itoh. On bit correlations among preimages of “many to one” one-way functions. Lecture Notes in Computer Science, 718:435– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yen:1993:FCE

[1367] S.-M. Yen and C.-S. Laih. The fast cascade exponentiation algorithm and its applications on cryptography. Lecture Notes in Computer Science, 718: 447–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1993:DCK

[1368] C.-C. Chang, T.-C. Wu, and C.-P. Chen. The design of a conference key distribution system. Lecture Notes in Computer Science, 718:459–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zuk:1993:RDC

[1369] E. Zuk. Remarks on “the design of a conference key distribution system”. Lecture Notes in Computer Science, 718:467–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harn:1993:PKC

[1370] L. Harn and S. Yang. Public-key cryptosystem based on the discrete logarithm problem. Lecture Notes in Computer Science, 718:469–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyaji:1993:ECF

[1371] A. Miyaji. Elliptic curves over fsub p suitable for cryptosystems. Lecture Notes in Computer Science, 718:479– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waldvogel:1993:PDD

[1372] C. P. Waldvogel and J. L. Massey.

The probability distribution of the Diffie–Hellman key. Lecture Notes in Computer Science, 718:492–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sauerbrey:1993:MEU

[1373] J. Sauerbrey. A modular exponentiation unit based on systolic arrays. Lecture Notes in Computer Science, 718: 505–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OKeefe:1993:CKD

[1374] C. M. O’Keefe. A comparison of key distribution patterns constructed from circle geometries. Lecture Notes in Computer Science, 718:517–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rezny:1993:BCM

[1375] M. Rezny and E. Trimarchi. A block cipher method using combinations of different methods under the control of the user key. Lecture Notes in Computer Science, 718:531–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preneel:1993:ATH

[1376] B. Preneel, R. Govaerts, and J. Vandewalle. An attack on two hash functions by Zheng-Matsumoto-Imai. Lecture Notes in Computer Science, 718:535– 538, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lidl:1993:PTL

[1377] R. Lidl and W. B. Muller. Primality testing with Lucas functions. Lecture Notes in Computer Science, 718:539– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haralick:1993:PCC

[1378] R. M. Haralick. Performance characterization in computer vision. Lecture Notes in Computer Science, 719:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimelfarb:1993:LLC

[1379] G. L. Gimel’farb. Low-level computational mono and stereo vision: A Bayesian approach. Lecture Notes in Computer Science, 719:10–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Musin:1993:TSI

[1380] O. R. Musin. Topographic structure of image. Lecture Notes in Computer Science, 719:24–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kropatsch:1993:DIP

[1381] W. G. Kropatsch, Ch. Reither, D. Willersinn, and G. Wlaschitz. The dual irregular pyramid. Lecture Notes in Computer Science, 719:31–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vassilakopoulos:1993:ARQ

[1382] M. Vassilakopoulos and Y. Manolopoulos. Analytical results on the quadtree storage-requirements. Lecture Notes in Computer Science, 719:41–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kertesz:1993:PQA

[1383] M. Kertesz, A. Kummert, and F. Csillag. From pyramids to quadtrees: Approximation of heterogeneous surfaces by fixing complexity. Lecture Notes in Computer Science, 719:49–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kummert:1993:CES

[1384] A. Kummert and S. Kabos. Calculation and estimation of sample statistics of binary images using quadtree data representations. Lecture Notes in Computer Science, 719:58–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evans:1993:TSR

[1385] A. N. Evans and M. S. Nixon. Temporal speckle reduction for feature extraction in ultrasound images. Lecture Notes in Computer Science, 719:65–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sziranyi:1993:NES

[1386] T. Sziranyi. Noise effects in statistical subpixel pattern recognition. Lecture Notes in Computer Science, 719:74–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feichtinger:1993:FIR

[1387] H. G. Feichtinger and T. Strohmer.

Fast iterative reconstruction of bandlimited images from non-uniform sampling values. Lecture Notes in Computer Science, 719:82–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bajla:1993:AFM

[1388] I. Bajla, M. Marusiak, and M. Sramek. Anisotropic filtering of MRI data based upon image gradient histogram. Lecture Notes in Computer Science, 719: 90–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasperovich:1993:FDC

[1389] L. V. Kasperovich and V. F. Babkin. Fast discrete cosine transform approximation for JPEG image compression. Lecture Notes in Computer Science, 719:98–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skarbek:1993:EDB

[1390] W. Skarbek and A. Pietrowcew. Error diffusion in block truncation coding. Lecture Notes in Computer Science, 719:105–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bojkovic:1993:BSN

[1391] Z. Bojkovic, D. Milanovic, and A. Samcovic. On a bound on signal-to-noise ratio in subband coding of Gaussian image process. Lecture Notes in Computer Science, 719:113–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bundschuh:1993:LPR

[1392] B. Bundschuh. A linear predictor as a regularization function in adaptive image restoration and reconstruction. Lecture Notes in Computer Science, 719:118–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suesse:1993:ICS

[1393] H. Suesse and K. Voss. Inversion of convolution by small kernels. Lecture Notes in Computer Science, 719:123– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valev:1993:MBI

[1394] V. Valev. A model-based image quantization technique for supervised image recognition. Lecture Notes in Computer Science, 719:128–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ottenberg:1993:BCD

[1395] K. Ottenberg. Brightness-contrast diffusion and the grouping of missing angles. Lecture Notes in Computer Science, 719:133–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziou:1993:RIE

[1396] D. Ziou and J.-P. Fabre. Rotation invariance in edge detection. Lecture Notes in Computer Science, 719:141– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alshatti:1993:UEV

[1397] W. Alshatti and P. Lambert. Using eigenvectors of a vector field for deriving a second directional derivative operator for color images. Lecture Notes in Computer Science, 719:149– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Selmaoui:1993:CLD

[1398] N. Selmaoui, C. Leschi, and H. Emptoz. Crest lines detection in grey level images: Studies of different approaches and proposition of a new one. Lecture Notes in Computer Science, 719:157–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBeauville:1993:CDU

[1399] J.-P. Asselin de Beauville and M. C. Mraghni. Contour-detection using the shape of the nearest neighbors set (NN-Set). Lecture Notes in Computer Science, 719:165–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shen:1993:CSP

[1400] J. Shen, S. Castan, and J. Zhao. A comparative study of performance for noisy roof edge detection. Lecture Notes in Computer Science, 719:170–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1993:HLP

[1401] L. T. S. Lam, W. C. Y. Lam, and D. N. K. Leung. A Hough-like prediction/correction approach for ellipse detection. Lecture Notes in Computer Science, 719:175–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1993:FPP

[1402] W. C. Y. Lam, K. S. Y. Yuen, and D. N. K. Leung. Fourier parameterization provide uniform bounded Hough space. Lecture Notes in Computer Science, 719:183–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalviainen:1993:MER

[1403] H. Kalviainen. Motion estimation and the randomized Hough transform (RHT): New methods with gradient information. Lecture Notes in Computer Science, 719:191–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nixon:1993:CEL

[1404] M. S. Nixon. Circle extraction via least squares and the Kalman filter. Lecture Notes in Computer Science, 719: 199–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

diBaja:1993:MSD

[1405] G. Sanniti di Baja and E. Thiel. A multiresolution shape description algorithm. Lecture Notes in Computer Science, 719:208–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choras:1993:ICM

[1406] R. S. Choras. Image coding by morphological skeleton transformation. Lecture Notes in Computer Science, 719: 216–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1993:NLS

[1407] G. Brown, P. Forte, R. Malyan, and P. Barnwell. A non-linear shape abstraction technique. Lecture Notes in Computer Science, 719:223–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korneenko:1993:MST

[1408] N. Korneenko. Minimum-space timeoptimal convex hull algorithms. Lecture Notes in Computer Science, 719: 231–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guo:1993:DCP

[1409] J.-K. Guo, R.-L. Hsu, C.-H. Chen, and Y.-N. Sun. Detecting corners of polygonal and polyhedral objects. Lecture Notes in Computer Science, 719:237– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melen:1993:FAD

[1410] T. Melen and T. Ozanian. A fast algorithm for dominant point detection on chain-coded contours. Lecture Notes in Computer Science, 719:245–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jourlin:1993:SM

[1411] M. Jourlin, I. Fillere, J.-M. Becker, and M.-J. Laboure. Shapes and metrics. Lecture Notes in Computer Science, 719:254–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Starovoitov:1993:STS

[1412] V. V. Starovoitov. Simplified technique of structure extraction from textural images. Lecture Notes in Computer Science, 719:259–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1993:APA

[1413] S. W. C. Lam and H. H. S. Ip. Adaptive pyramid approach to texture segmentation. Lecture Notes in Computer Science, 719:267–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimelfarb:1993:MRF

[1414] G. L. Gimel’farb and A. V. Zalesny.

Markov random fields with short-and long-range interaction for modelling gray-scale textured images. Lecture Notes in Computer Science, 719:275– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fioravanti:1993:TRO

[1415] S. Fioravanti and D. D. Giusto. Texture recognition by the q-th order fractal analysis. Lecture Notes in Computer Science, 719:283–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1993:PML

[1416] E. Tanaka, H. Awano, and S. Masuda.

A proximity measure of line drawings for comparison of chemical compounds. Lecture Notes in Computer Science, 719:291–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bunke:1993:JPS

[1417] H. Bunke and G. Kaufmann. Jigsaw puzzle solving using approximate string matching and best-first search. Lecture Notes in Computer Science, 719: 299–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgis:1993:MAG

[1418] N. Georgis, M. Petrou, and J. Kittler. A morphological approach to the generalised 2-stage stock-cutting problem. Lecture Notes in Computer Science, 719:309–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambroszkiewicz:1993:PCP

[1419] S. Ambroszkiewicz. Primitive and compound patterns. Lecture Notes in Computer Science, 719:317–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Washio:1993:SMB

[1420] N. Washio, E. Tanaka, and S. Masuda. A similarity measure between 3-D objects and its parallel computation. Lecture Notes in Computer Science, 719:322–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liedtke:1993:ALR

[1421] C.-E. Liedtke, Th. Schnier, and A. Blomer. Automated learning of rules using genetic operators. Lecture Notes in Computer Science, 719:327– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desachy:1993:SII

[1422] J. Desachy and E. H. Zahzah. Symbolic and iconic information combination for satellite imagery interpretation. Lecture Notes in Computer Science, 719:335–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howard:1993:MCC

[1423] C. G. Howard and P. Bock. Multiclass classification and symbolic cognitive processing with ALISA. Lecture Notes in Computer Science, 719:343– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rios:1993:CIF

[1424] H. Rios. Computing image flow using a coarse-to-fine strategy for spatiotemporal filters. Lecture Notes in Computer Science, 719:355–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasprzak:1993:VME

[1425] W. Kasprzak and H. Niemann. Visual motion estimation from image contour tracking. Lecture Notes in Computer Science, 719:363–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Irani:1993:RRE

[1426] M. Irani, B. Rousso, and S. Peleg. Robust recovery of ego-motion. Lecture Notes in Computer Science, 719:371– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bulas-Cruz:1993:TST

[1427] J. Bulas-Cruz, A. T. Ali, and E. L. Dagless. A temporal smoothing technique for real-time motion detection. Lecture Notes in Computer Science, 719:379– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikolaidis:1993:CEM

[1428] N. Nikolaidis, I. Pitas, and M. G. Strintzis. Combined evaluation of motion and disparity vector fields for stereoscopic sequence coding. Lecture Notes in Computer Science, 719:387– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heikkonen:1993:RTM

[1429] J. Heikkonen. Recovering translational motion parameters from image sequences using randomized Hough transform. Lecture Notes in Computer Science, 719:395–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1993:EOF

[1430] R. Agarwal and J. Sklansky. Estimating optical flow for large interframe displacements. Lecture Notes in Computer Science, 719:403–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pajdla:1993:SDR

[1431] T. Pajdla and V. Hlavac. Surface discontinuities in range images. Lecture Notes in Computer Science, 719:412– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klette:1993:ASS

[1432] R. Klette and V. Rodehorst. Algorithms for shape from shading, lighting direction and motion. Lecture Notes in Computer Science, 719:420–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaklic:1993:SDS

[1433] A. Jaklic and F. Solina. Separating diffuse and specular component of image irradiance by translating a camera. Lecture Notes in Computer Science, 719:428–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1993:NSC

[1434] X. Hu and N. Ahuja. Necessary and sufficient conditions for a unique solution of plane motion and structure. Lecture Notes in Computer Science, 719:436–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schluns:1993:PSN

[1435] K. Schluns. Photometric stereo for non-Lambertian surfaces using color information. Lecture Notes in Computer Science, 719:444–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brunnstrom:1993:AFJ

[1436] K. Brunnstrom and J.-O. Eklundh.

Active fixation for junction classification. Lecture Notes in Computer Science, 719:452–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1993:PMR

[1437] Z. Zhang. Point matching for registration of free-form surfaces. Lecture Notes in Computer Science, 719:460– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bost:1993:TPA

[1438] J.-M. Bost, R. Milanese, and T. Pun. Temporal precedence in asynchronous visual indexing. Lecture Notes in Computer Science, 719:468–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monks:1993:ISM

[1439] T. P. Monks and J. N. Carter. Improved stripe matching for colour encoded structured light. Lecture Notes in Computer Science, 719:476– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chebaro:1993:FST

[1440] B. Chebaro, A. Crouzil, L. Massip-Pailhes, and S. Castan. Fusion of the stereoscopic and temporal matching results by an algorithm of coherence control and conflicts management. Lecture Notes in Computer Science, 719:486–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lutton:1993:MBS

[1441] E. Lutton. 3D model based stereo reconstruction using coupled Markov random fields. Lecture Notes in Computer Science, 719:494–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinhage:1993:OSV

[1442] V. Steinhage. Occlusions and special views within the reconstruction of polyhedral scenes. Lecture Notes in Computer Science, 719:502–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gudmundsson:1993:FSV

[1443] B. Gudmundsson and C. Tegenfeldt.

Fast shadowing of volume data. Lecture Notes in Computer Science, 719: 510–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guo:1993:TDM

[1444] X. Guo. Three-dimensional moment invariants under rigid transformation. Lecture Notes in Computer Science, 719:518–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemus:1993:PSG

[1445] M. D. Hemus and J. N. Carter. Parameterisation of simple geometrical lambertian surfaces using photometric stereo. Lecture Notes in Computer Science, 719:523–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawato:1993:HTE

[1446] S. Kawato. Hough transform to extract 3D information from images of different view points. Lecture Notes in Computer Science, 719:528–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolesnik:1993:FAS

[1447] M. I. Kolesnik. Fast algorithm for the stereo pair matching with parallel computation. Lecture Notes in Computer Science, 719:533–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koschan:1993:DSC

[1448] A. Koschan. Dense stereo correspondence using polychromatic block matching. Lecture Notes in Computer Science, 719:538–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maver:1993:PNV

[1449] J. Maver, A. Leonardis, and F. Solina. Planning the next view using the max-min principle. Lecture Notes in Computer Science, 719:543–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shapiro:1993:CCR

[1450] V. Shapiro. Cross-correlation with reconstruction: A new approach to pattern matching. Lecture Notes in Computer Science, 719:548–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nowinski:1993:INB

[1451] W. L. Nowinski. The iterated normalized backprojection method of image reconstruction. Lecture Notes in Computer Science, 719:556–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trofimov:1993:CBR

[1452] O. Trofimov. Cone beam reconstruction and Fourier transform of distributions. Lecture Notes in Computer Science, 719:564–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flusser:1993:CRA

[1453] J. Flusser and T. Suk. Character recognition by affine moment invariants. Lecture Notes in Computer Science, 719:572–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fageth:1993:UFP

[1454] R. Fageth, W. G. Allen, and U. Jager.

Ultra fast pattern classification by fuzzy logic. Lecture Notes in Computer Science, 719:578–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klauer:1993:OOP

[1455] B. Klauer and K. Waldschmidt. An object-oriented pen-based recognizer for handprinted characters. Lecture Notes in Computer Science, 719:586– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shchepin:1993:MCP

[1456] E. V. Shchepin and G. M. Nepomnyashchii. On the method of critical points in character recognition. Lecture Notes in Computer Science, 719: 594–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ablameyko:1993:PTS

[1457] S. Ablameyko and O. Okun. Printed text segmentation using distance transform. Lecture Notes in Computer Science, 719:599–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takamura:1993:RHL

[1458] S. Takamura and M. Takagi. Recognition of handwritten layout drawings. Lecture Notes in Computer Science, 719:604–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krause:1993:ARS

[1459] F.-L. Krause, H. Jansen, and N. Luth. Automatic recognition of scanned technical drawings. Lecture Notes in Computer Science, 719:609–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belikova:1993:DIP

[1460] T. P. Belikova. Digital image processing in radiology. an experience of development of technology for computeraided diagnosis. Lecture Notes in Computer Science, 719:614–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vargas:1993:MQP

[1461] J. R. Vargas, R. D. Appel, D. F. Hochstrasser, and C. Pellegrini. Modeling and quantification of protein maps by Gaussian fitting. Lecture Notes in Computer Science, 719:627–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kovalev:1993:CAL

[1462] V. A. Kovalev and N. V. Mytzik. Computer analysis of the large intestine contours for the recognition of diseases. Lecture Notes in Computer Science, 719:634–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radeva:1993:RBA

[1463] P. I. Radeva. A rule-based approach to hand X-ray image segmentation. Lecture Notes in Computer Science, 719:641–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1993:DCV

[1464] Chun-Han Chen, Yen-Ching Yu, Jinn-Kwei Guo, Chin-Hsing Chen, Yung-Nien Sun, and C. I. Yu. 3-D cerebral vessel reconstruction from angiograms. Lecture Notes in Computer Science, 719:649–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jan:1993:CDA

[1465] J. Jan, B. Gabrhel, J. Haluzikova, P. Kilian, R. Kubak, Jr., and I. Provaznik. Comparison of different approaches to suppress speckles in ultrasonic tomograms. Lecture Notes in Computer Science, 719:657–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikolajczyk:1993:TBC

[1466] K. Mikolajczyk, J. Owczarczyk, and W. Recko. A test-bed for computerassisted fusion of multi-modality medical images. Lecture Notes in Computer Science, 719:664–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paulus:1993:OOV

[1467] D. W. R. Paulus and M. Wolf. Objectoriented volume segmentation. Lecture Notes in Computer Science, 719:669– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jankovic:1993:SMT

[1468] V. Jankovic, E. Ruzicky, and L. Niepel. Smooth morphological transformation of CT and MR medical data. Lecture Notes in Computer Science, 719:674–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smutny:1993:EPF

[1469] V. Smutny, T. Dostalova, J. Duskova, and V. Hlavac. Evaluation of plaque formation — surface reflectance measurement. Lecture Notes in Computer Science, 719:679–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szekely:1993:AMA

[1470] G. Szekely, G. Gerig, Th. Koller, Ch. Brechbuhler, and O. Kubler. Analysis of MR angiography volume data leading to the structural description of the cerebral vessel tree. Lecture Notes in Computer Science, 719:687–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiffen:1993:ACN

[1471] L. N. Wiffen and J. T. Kent. The automatic classification of normal and abnormal chromosomes using image analysis. Lecture Notes in Computer Science, 719:693–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanajik:1993:RTI

[1472] M. Hanajik and R. G. Van Vliet. Real time image processing for fast seam tracking. Lecture Notes in Computer Science, 719:698–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boccignone:1993:EVT

[1473] G. Boccignone, L. Esposito, and A. Marcelli. An experimental vision tool for real time quality control. Lecture Notes in Computer Science, 719: 706–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wen:1993:HTA

[1474] J. Wen. High-tech approaches of computer vision in industry. Lecture Notes in Computer Science, 719:711– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ivanov:1993:IPS

[1475] E. Ivanov, D. Holm, W. Osten, and W. Juptner. An image processing system for material flow control in coal mining industry. Lecture Notes in Computer Science, 719:716–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bres:1993:TQA

[1476] S. Bres, H. Emptoz, and G. Merle.

Transparency quantification. application to composite materials quality controls by image processing. Lecture Notes in Computer Science, 719:721– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulecane:1993:CCM

[1477] H. Boulecane, N. Vincent, M. Ruffier, and H. Emptoz. Control of composite material structure by fractal methods. Lecture Notes in Computer Science, 719:726–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palenichka:1993:MBA

[1478] R. M. Palenichka and R. T. Mysak. Model-based adaptive preprocessing of images in automated visual inspection. Lecture Notes in Computer Science, 719:732–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perner:1993:AKB

[1479] P. Perner. Application of knowledgebased image inspection system for diagnosis of misprints in offsetprinting. Lecture Notes in Computer Science, 719: 738–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toth:1993:MIM

[1480] Ch. K. Toth and T. Schenk. Multiple image matching in an automatic aerotriangulation system. Lecture Notes in Computer Science, 719:750–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simberova:1993:DPS

[1481] S. Simberova. Digital processing of Skylab X-ray images of the solar corona. Lecture Notes in Computer Science, 719:759–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1993:FHF

[1482] G. Yang and T. S. Huang. Finding human faces in a picture. Lecture Notes in Computer Science, 719:766–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hegedus:1993:CCA

[1483] Gy. Cs. Hegedus. CADI: Computer assisted educational package for digital image processing. Lecture Notes in Computer Science, 719:770–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobrusin:1993:CIA

[1484] Y. Dobrusin, Y. Edan, I. Grinshpun, U. M. Peiper, I. Wolf, and A. Hetzroni. Computer image analysis to locate targets for an agricultural robot. Lecture Notes in Computer Science, 719:775–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berke:1993:ADI

[1485] J. Berke, K. Gyorffy, G. Fischl, L. Karpati, and J. Bakonyi. The application of digital image processing in the evaluation of agricultural experiments. Lecture Notes in Computer Science, 719:780–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hattori:1993:GAL

[1486] K. Hattori, M. Yamaguchi, N. Iwama, and M. Hayakawa. GCV-Aided linear image regularization for the reconstruction of wave distribution function of magnetospheric VLF/ELF waves. Lecture Notes in Computer Science, 719:

788–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alekseev:1993:CDA

[1487] A. S. Alekseev, V. P. Pyatkin, and G. I. Salov. Crater detection in aero-space imagery using simple nonparametric statistical tests. Lecture Notes in Computer Science, 719:793–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honisch:1993:DRS

[1488] U. Honisch. Digital retina — simulating dynamic behavior of visual perception. Lecture Notes in Computer Science, 719:800–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

diGesu:1993:NHR

[1489] V. di Gesu, G. Gerardi, S. Impedovo, B. Lenzitti, G. Parodi, and D. Tegolo. A new heterogeneous and reconfigurable architecture for image analysis. Lecture Notes in Computer Science, 719:805–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roska:1993:ASC

[1490] T. Roska. The analogic single-chip CNN visual supercomputer — a review. Lecture Notes in Computer Science, 719:813–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kotropoulos:1993:VLV

[1491] C. Kotropoulos and I. Pitas. A variant of learning vector quantizer based on split-merge statistical tests. Lecture Notes in Computer Science, 719:822– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mascarilla:1993:NNC

[1492] L. Mascarilla, E. H. Zahzah, and J. Desachy. Neural networks classifiers based on geocoded data and MultiSpectral images for satellite image interpretation. Lecture Notes in Computer Science, 719:830–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michaelis:1993:ANN

[1493] B. Michaelis and G. Krell. Artificial neural networks for image improvement. Lecture Notes in Computer Science, 719:838–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radvanyi:1993:UCN

[1494] A. G. Radvanyi. Using cellular neural network to “see” random-dot stereograms. Lecture Notes in Computer Science, 719:846–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busse:1993:IHC

[1495] R. Busse, A. Muller, and E. J. Neuhold.

Information handling — A challenge for databases and expert systems. Lecture Notes in Computer Science, 720: 1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Al-Jadir:1993:CVO

[1496] L. Al-Jadir, G. Falquet, and M. Leonard. Context versions in an objectoriented model. Lecture Notes in Computer Science, 720:24–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gross-Hardt:1993:TCL

[1497] M. Gross-Hardt and G. Vossen. Towards class-less object models for engineering design applications. Lecture Notes in Computer Science, 720:36–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pokorny:1993:SRC

[1498] J. Pokorny. Semantic relativism in conceptual modelling. Lecture Notes in Computer Science, 720:48–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmann:1993:ASC

[1499] T. Hartmann, R. Jungclaus, and G. Saake. Animation support for a conceptual modelling language. Lecture Notes in Computer Science, 720:56–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duong:1993:UMD

[1500] T. Duong, J. Hiller, and U. Srinivasan. A unifying model of data, metadata and context. Lecture Notes in Computer Science, 720:68–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbara:1993:IBS

[1501] D. Barbara and Ch. Clifton. Information brokers: Sharing knowledge in a heterogeneous distributed system. Lecture Notes in Computer Science, 720:80–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1993:CMT

[1502] J. Chen, O. A. Bukhres, and J. Sharif-Askary. A customized multidatabase transaction management strategy. Lecture Notes in Computer Science, 720: 92–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danes:1993:IBD

[1503] A. Danes, F. Exertier, and S. Haj Houssain. Interoperability between a distributed system and a database system. Lecture Notes in Computer Science, 720:104–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mullen:1993:RCU

[1504] J. G. Mullen, J. Jing, and J. Sharif-Askary. Reservation commitment and its use in multidatabase systems. Lecture Notes in Computer Science, 720: 116–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meng:1993:PQP

[1505] W. Meng, Ch. Liu, W. Sun, and C. Yu.

Predict query processing cost in a distributed database system. Lecture Notes in Computer Science, 720:122– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu:1993:CCQ

[1506] W. W. Chu. CoBase: A cooperative query answering facility for database systems. Lecture Notes in Computer Science, 720:134–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdelguerfi:1993:DDR

[1507] M. Abdelguerfi, K. Grant, E. Murphy, and W. Patterson. Duplicate deletion in a ring connected, sharednothing, parallel database system. Lecture Notes in Computer Science, 720: 146–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurutach:1993:TFT

[1508] W. Kurutach and J. Franklin. On temporal-fuzziness in temporal fuzzy databases. Lecture Notes in Computer Science, 720:154–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Springsteel:1993:OBS

[1509] F. N. Springsteel. Object-based schema integration for heterogeneous databases: A logical approach. Lecture Notes in Computer Science, 720: 166–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veijalainen:1993:HMT

[1510] J. Veijalainen. Heterogeneous multilevel transaction management with multiple substransactions. Lecture Notes in Computer Science, 720:181– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ling:1993:ICO

[1511] T. W. Ling and P. K. Teo. Inheritance conflicts in object-oriented systems. Lecture Notes in Computer Science, 720:189–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhao:1993:MDD

[1512] J. L. Zhao. Managing derived data in intelligent database systems: An implementation study. Lecture Notes in Computer Science, 720:201–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1993:ICM

[1513] N. Zhong and S. Ohsuga. An integrated calculation model for discovering functional relations from databases. Lecture Notes in Computer Science, 720: 213–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishakbeyoglu:1993:MII

[1514] N. S. Ishakbeyoglu and Z. M. Ozsoyoglu. On the maintenance of implication integrity constraints. Lecture Notes in Computer Science, 720:221–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naqvi:1993:RAD

[1515] W. Naqvi and M. T. Ibrahim. RE-FLEX active database model: Application of Petri-nets. Lecture Notes in Computer Science, 720:233–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1993:RAA

[1516] J. Wu and L. Harbird. Road accident analysis using a functional database language. Lecture Notes in Computer Science, 720:241–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Revell:1993:DPE

[1517] N. Revell and M. W. Youssef. Database performance evaluation: a methodological approach. Lecture Notes in Computer Science, 720:253–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kersten:1993:DID

[1518] M. L. Kersten and F. Kwakkel. Design and implementation of a DBMS performance assessment tool. Lecture Notes in Computer Science, 720:265– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Du:1993:MDQ

[1519] K. Du and G. Ozsoyoglu. Modifying database queries and error constraints. Lecture Notes in Computer Science, 720:277–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabitti:1993:PES

[1520] F. Rabitti, R. S. Sferrazza, M. G. Tori, and P. Zezula. Performance evaluation system for object stores. Lecture Notes in Computer Science, 720:289– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hameurlain:1993:OMD

[1521] A. Hameurlain and F. Morvan. An optimization method of data communication and control for parallel execution of SQL queries. Lecture Notes in Computer Science, 720:301–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Son:1993:DDS

[1522] S. H. Son, D. W. George, and Y.-K.

Kim. Developing a database system for time-critical applications. Lecture Notes in Computer Science, 720:313– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keim:1993:OOQ

[1523] D. A. Keim, H.-P. Kriegel, and A. Miethsam. Object-oriented querrying of existing relational databases. Lecture Notes in Computer Science, 720:325– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kornatzky:1993:PSD

[1524] Y. Kornatzky and S. E. Shimony. A probabilistic spatial data model. Lecture Notes in Computer Science, 720: 337–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriegel:1993:QPG

[1525] H.-P. Kriegel, S. Heep, A. Fahldiek, and N. Mysliwitz. Query processing of geometric objects with free form boundaries in spatial databases. Lecture Notes in Computer Science, 720: 349–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anogianakis:1993:BDB

[1526] G. Anogianakis, A. Krotopoulou, P. Spirakis, D. Terpou, and A. Tsakalidis. Brain data base (BDB). Lecture Notes in Computer Science, 720:361– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gardarin:1993:ICR

[1527] G. Gardarin. Integrating classes and relations to model and query geographical databases. Lecture Notes in Computer Science, 720:365–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemerly:1993:TCG

[1528] A. S. Hemerly, A. L. Furtado, and M. A. Casanova. Towards cooperativeness in geographic databases. Lecture Notes in Computer Science, 720:373– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1993:OOC

[1529] B. David, L. Raynal, and G. Schorter. GeO_{2}: Object-oriented contribution for a geographical DBMS? Lecture Notes in Computer Science, 720:377– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Song:1993:GES

[1530] I.-Y. Song, H. M. Godsey, J. Newton, and B. Bargmeyer. GemCode: An expert system generating mnemonic codes for data elements and data items. Lecture Notes in Computer Science, 720:384–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Webster:1993:APK

[1531] C. A. G. Webster, M. Weller, M. M.

Sfantsikopoulos, and V. D. Tsoukalas. ALEXSYS — A prototype knowledge based expert system for the quality assurance of high pressure die castings. Lecture Notes in Computer Science, 720:396–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finch:1993:VFE

[1532] I. Finch. Viewpoints — facilitating expert systems for multiple users. Lecture Notes in Computer Science, 720: 401–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guan:1993:ISL

[1533] J. W. Guan and D. A. Bell. Improving Shafer-Logan’s algorithm for handling hierarchical evidence. Lecture Notes in Computer Science, 720:413–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Popper:1993:LLH

[1534] M. Popper. From low-level to high-level operations in expert systems. Lecture Notes in Computer Science, 720:424– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1993:CEK

[1535] E. Wilson. Corpora as expert knowledge domains: the Oxford Advanced Learner’s Dictionary. Lecture Notes in Computer Science, 720:428–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehner:1993:MKB

[1536] F. Lehner, H. F. Hofmann, R. Setzer, and R. Maier. Maintenance of knowledge bases. Lecture Notes in Computer Science, 720:436–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zdrahal:1993:UCS

[1537] Z. Zdrahal. Using candidate space structure to propose the next measurement in model based diagnosis. Lecture Notes in Computer Science, 720: 448–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debenham:1993:DFC

[1538] J. Debenham. Decomposition of four component items. Lecture Notes in Computer Science, 720:457–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brayshaw:1993:IID

[1539] M. Brayshaw. Intelligent inference for debugging concurrent systems. Lecture Notes in Computer Science, 720:461– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botti:1993:STK

[1540] V. Botti, F. Barber, A. Crespo, D. Gallardo, I. Ripoll, E. Onaindia, and L. Hernandez. Sharing temporal knowledge by multiple agents. Lecture Notes in Computer Science, 720:470– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hung:1993:QEL

[1541] H.-K. Hung, P. Martin, J. Glasgow, Ch. Walmsley, and M. Jenkins. Querying and exploring large knowledge bases. Lecture Notes in Computer Science, 720:474–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watt:1993:MTO

[1542] S. Watt. Managing text objectively.

Lecture Notes in Computer Science, 720:478–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schweighofer:1993:LES

[1543] E. Schweighofer and W. Winiwarter.

Legal expert system KONTERM — automatic representation of document structure and contents. Lecture Notes in Computer Science, 720:486– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munoz:1993:MME

[1544] J. F. Munoz and F. Galindo. Matrim, man expert system on marital law. Lecture Notes in Computer Science, 720: 498–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poulin:1993:CC

[1545] D. Poulin, P. St-Vincent, and P. Bratley. Contradiction and confirmation. Lecture Notes in Computer Science, 720:502–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiscornia:1993:MRL

[1546] D. Tiscornia. Meta-reasoning in law: A computational model. Lecture Notes in Computer Science, 720:514–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1993:AKT

[1547] M. Baaz, F. Galindo, G. Quirchmayr, and M. Vazqez. The application of Kripke-type structures to regional development programs. Lecture Notes in Computer Science, 720:523–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markowitz:1993:DMT

[1548] V. M. Markowitz and I.-M. A. Chen. Data management tools for genomic applications: A progress report. Lecture Notes in Computer Science, 720: 529–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rotterdam:1993:RCI

[1549] E. Rotterdam, S. Van Denneheuvel, P. Hennis, and P. Van Emde Boas. Resolution of constraint inconsistency with the aim to provide support in anaesthesia. Lecture Notes in Computer Science, 720:541–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toure:1993:OOI

[1550] F. Toure and M. Schneider. An objectoriented implementation for a semantic system (CANDID). Lecture Notes in Computer Science, 720:553–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Afsarmanesh:1993:DSM

[1551] H. Afsarmanesh, F. Tuijnman, M. Wiedijk, and L. O. Hertzberger. Distributed schema management in a cooperation network of autonomous agents. Lecture Notes in Computer Science, 720:565–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dilger:1993:DAS

[1552] W. Dilger and S. Kassel. A distributed AI system for job shop control. Lecture Notes in Computer Science, 720:577–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gospodarowicz:1993:ESP

[1553] A. Gospodarowicz, E. Kania, S. Krawczyk, M. Rymarczyk, and A. M. Tjoa. Expert system for production planning of perishable goods. Lecture Notes in Computer Science, 720:583–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mekras:1993:ESM

[1554] N. D. Mekras, A. G. Malama, G. P.

Parnassas, and I. P. Tatsiopoulos. An expert system as a manager in the application of production planning and control software in CIM environments. Lecture Notes in Computer Science, 720:593–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djeraba:1993:CDR

[1555] C. Djeraba, A. A. Hssain, and B. Descotes-Genon. Composition and dependency relationships in production information system design. Lecture Notes in Computer Science, 720: 605–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takizawa:1993:VT

[1556] M. Takizawa, S. Hamada, and S. M. Deen. Vehicle transactions. Lecture Notes in Computer Science, 720:611– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gevers:1993:AIR

[1557] T. Gevers and A. W. M. Smeulders. An approach to image retrieval for image databases. Lecture Notes in Computer Science, 720:615–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Candela:1993:FPC

[1558] S. Candela, C. Garcia, J. Munoz, and F. Alayon. Facilitatory process for contrast detection. Lecture Notes in Computer Science, 720:627–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emmerich:1993:OOD

[1559] W. Emmerich, P. Kroha, and W. Schafer. Object-oriented database management systems for construction of CASE environments. Lecture Notes in Computer Science, 720:631– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1993:SDR

[1560] T.-L. Hwang. Summary data representations in application developments. Lecture Notes in Computer Science, 720:643–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rolland:1993:RPC

[1561] C. Rolland and N. Prakash. Reusable process chunks. Lecture Notes in Computer Science, 720:655–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez:1993:ADD

[1562] O. P. Lopez, I. Ramos, and J. H.

Canos. From analysis to design in a deductive and object-oriented environment. Lecture Notes in Computer Science, 720:667–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vlachantonis:1993:CSO

[1563] N. Vlachantonis. A case study for an open CASE system: The TROLL light development environment. Lecture Notes in Computer Science, 720: 673–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welzer:1993:MDM

[1564] T. Welzer and J. Eder. Meta data model for database design. Lecture Notes in Computer Science, 720:677– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1993:EPO

[1565] X. Wu and J. Neuhaus. Extending PCTE with object-oriented capabilities. Lecture Notes in Computer Science, 720:681–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1993:NHD

[1566] H. Maurer, N. Scherbakov, and P. Srinivasan. A new hypermedia data model. Lecture Notes in Computer Science, 720:685–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bench-Capon:1993:LSH

[1567] T. J. M. Bench-Capon, P. E. S. Dunne, and G. Staniford. Linearisation schemata for hypertext. Lecture Notes in Computer Science, 720:697– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amann:1993:HOI

[1568] B. Amann, V. Christophides, and M. Scholl. HyperPATH/ osub 2: Integrating hypermedia systems with object-oriented database systems. Lecture Notes in Computer Science, 720: 709–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nanard:1993:IKB

[1569] J. Nanard, M. Nanard, A.-M. Massotte, A. Djemaa, A. Joubert, H. Betaille, and J. Chauche. Integrating knowledge-based hypertext and database for task-oriented access to documents. Lecture Notes in Computer Science, 720:721–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karagiannis:1993:RUI

[1570] D. Karagiannis, E. Ortwein, and J. Gag. Reengineering of user interfaces for the migration of database applications. Lecture Notes in Computer Science, 720:733–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klein:1993:UIK

[1571] M. R. Klein and R. Traunmuller.

User interface of knowledge Based-DSS development environment. Lecture Notes in Computer Science, 720: 746–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qutaishat:1993:HCS

[1572] M. A. Qutaishat, W. A. Gray, and N. J.

Fiddian. A highly-customisable schema meta-visualisation system for objectoriented (O-O) database schemas — overview. Lecture Notes in Computer Science, 720:756–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuroki:1993:WUA

[1573] S. Kuroki, K. Kikkawa, K. Kaneko, and A. Makinouchi. Walkthrough using animation database system MOVE. Lecture Notes in Computer Science, 720: 760–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buchberger:1993:MSD

[1574] B. Buchberger. Mathematica: A system for doing mathematics by computer? Lecture Notes in Computer Science, 722:1, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauer:1993:PCA

[1575] B. Bauer and R. Hennicker. Proving the correctness of algebraic implementations by the ISAR system. Lecture Notes in Computer Science, 722:2–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1993:SCC

[1576] M. Gogolla, S. Conrad, and R. Herzig. Sketching concepts and computational model of TROLL light. Lecture Notes in Computer Science, 722:17–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1993:ATT

[1577] B. Yi and J. Xu. Analogical type theory. Lecture Notes in Computer Science, 722:33–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jebelean:1993:IME

[1578] T. Jebelean. Improving the multiprecision Euclidean algorithm. Lecture Notes in Computer Science, 722:45–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maeder:1993:SAK

[1579] R. Maeder. Storage allocation for the Karatsuba integer multiplication algorithm. Lecture Notes in Computer Science, 722:59–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz:1993:PSD

[1580] A. Diaz, M. Hitz, E. Kaltofen, A. Lobo, and T. Valente. Process scheduling in DSC and the large sparse linear systems challenge. Lecture Notes in Computer Science, 722:66–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monagan:1993:GPD

[1581] Michael B. Monagan. Gauss: A parameterized domain of computation system with support for signature functions. Lecture Notes in Computer Science, 722:81–94, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1993:CCA

[1582] A. Weber. On coherence in computer algebra. Lecture Notes in Computer Science, 722:95–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

diBlasio:1993:SIL

[1583] P. di Blasio and M. Temperini. Subtyping inheritance in languages for symbolic computation systems. Lecture Notes in Computer Science, 722:107– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calmet:1993:UAB

[1584] J. Calmet and I. A. Tjandra. A unified-algebra-based specification language for symbolic computing. Lecture Notes in Computer Science, 722:122– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hearn:1993:OSA

[1585] A. C. Hearn and E. Schrufer. An ordersorted approach to algebraic computation. Lecture Notes in Computer Science, 722:134–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fritzson:1993:VHI

[1586] P. Fritzson, V. Engelson, and L. Viklund. Variant handling, inheritance and composition in the ObjectMath computer algebra environment. Lecture Notes in Computer Science, 722:145–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grivas:1993:MUO

[1587] G. Grivas and R. E. Maeder. Matching and unification for the objectoriented symbolic computation system AlgBench. Lecture Notes in Computer Science, 722:164–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santas:1993:TSC

[1588] P. S. Santas. A type system for computer algebra. Lecture Notes in Computer Science, 722:177–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Omodeo:1993:DPS

[1589] E. G. Omodeo and A. Policriti. Decision procedures for set/hyperset contexts. Lecture Notes in Computer Science, 722:192–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farmer:1993:RC

[1590] W. M. Farmer, J. D. Guttman, and F. J. Thayer. Reasoning with contexts. Lecture Notes in Computer Science, 722:216–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caferra:1993:GGF

[1591] R. Caferra and M. Herment. GLEFAT-INF: A graphic framework for combining theorem provers and editing proofs for different logics. Lecture Notes in Computer Science, 722:229–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caprotti:1993:ERC

[1592] O. Caprotti. Extending RISC-CLP (real) to handle symbolic functions. Lecture Notes in Computer Science, 722:241–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feng:1993:DTR

[1593] S. Feng, T. Sakabe, and Y. Inagaki.

Dynamic term rewriting calculus and its application to inductive equational reasoning. Lecture Notes in Computer Science, 722:256–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonacina:1993:DDC

[1594] M. P. Bonacina and J. Hsiang. Distributed deduction by clause-diffusion: The Aquarius Prover. Lecture Notes in Computer Science, 722:272–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:1993:DSP

[1595] H. Hong, A. Neubacher, and W. Schreiner. The design of the SACLIB/PACLIB kernels. Lecture Notes in Computer Science, 722:288–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zippel:1993:WCA

[1596] R. Zippel. The Weyl computer algebra substrate. Lecture Notes in Computer Science, 722:303–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Limongelli:1993:URM

[1597] C. Limongelli and M. Temperini. On the uniform representation of mathematical data structures. Lecture Notes in Computer Science, 722:319– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norman:1993:CDS

[1598] A. C. Norman. Compact delivery support for REDUCE. Lecture Notes in Computer Science, 722:331–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fournier:1993:IPL

[1599] R. Fournier, N. Kajler, and B. Mourrain. IZIC: A portable language-driven tool for mathematical surfaces visualization. Lecture Notes in Computer Science, 722:341–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meshveliani:1993:ACC

[1600] S. D. Meshveliani. The algebraic constructor CAC: Computing in construction-defined domains. Lecture Notes in Computer Science, 722:354– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Missura:1993:EAT

[1601] S. A. Missura. Extending AlgBench with a type system. Lecture Notes in Computer Science, 722:359–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vantaggiato:1993:MFF

[1602] A. Vantaggiato. Modeling finite fields with Mathematica — applications to the computation of exponential sums and to the solution of equations over finite fields. Lecture Notes in Computer Science, 722:364–368, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonamico:1993:ESC

[1603] S. Bonamico, G. Cioni, and A. Colagrossi. An enhanced sequent calculus for reasoning in a given domain. Lecture Notes in Computer Science, 722:369–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nepomniaschy:1993:POM

[1604] V. A. Nepomniaschy and A. A.

Sulimov. Problem-oriented means of program specification and verification in project SPECTRUM. Lecture Notes in Computer Science, 722:374– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walsh:1993:GPP

[1605] T. Walsh. General purpose proof plans. Lecture Notes in Computer Science, 722:379–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaines:1993:MEE

[1606] Brian R. Gaines. Modelling and extending expertise. Lecture Notes in Computer Science, 723:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boy:1993:KAD

[1607] Guy Boy. Knowledge acquisition in dynamic systems: How can logicism and situatedness go together? Lecture Notes in Computer Science, 723:23–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akkermans:1993:SCP

[1608] Hans Akkermans, Bob Wielinga, and Guus Schreiber. Steps in constructing problem solving methods. Lecture Notes in Computer Science, 723:45–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breuker:1993:MAL

[1609] Joost Breuker. Modelling artificial legal reasoning. Lecture Notes in Computer Science, 723:66–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zackland:1993:SBC

[1610] Manuel Zackland and Dominique Fontaine. Systematic building of conceptual classification systems with C-KAT. Lecture Notes in Computer Science, 723:79–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gappa:1993:MRL

[1611] Karsten Gappa and Ute Poeck. Making role limiting shells more flexible. Lecture Notes in Computer Science, 723: 103–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1993:MLT

[1612] Jerome Thomas, Philippe Laublet, and Jean-Gabriel Ganascia. A machine learning tool designed for a modelbased knowledge acquisition approach. Lecture Notes in Computer Science, 723:123–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alllemang:1993:GTK

[1613] Dean Alllemang and Gertjan Van Heijst. Generic tasks in KEW. Lecture Notes in Computer Science, 723:139– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linster:1993:RSM

[1614] Marc Linster. A review of Sisyphus 91 and 92: Models of problem-solving knowledge. Lecture Notes in Computer Science, 723:159–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Causse:1993:HCK

[1615] Karine Causse. Heuristic control knowledge. Lecture Notes in Computer Science, 723:183–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neubert:1993:MCM

[1616] Susanne Neubert. Model construction in MIKE (model-based and incremental knowledge engineering). Lecture Notes in Computer Science, 723:200– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paris:1993:EIS

[1617] Cecile Paris and Yolanda Gil. EX-PECT: Intelligent support for knowledge base refinement. Lecture Notes in Computer Science, 723:220–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vicat:1993:CCA

[1618] Catherine Vicat, Alain Busac, and Jean-Gabriel Ganascia. CERISE: A cyclic approach for knowledge acquisition. Lecture Notes in Computer Science, 723:237–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shaw:1993:PCP

[1619] Mildred L. G. Shaw and Brian R. Gaines. Personal construct psychology foundations for knowledge acquisition and representation. Lecture Notes in Computer Science, 723:256–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Compton:1993:KAA

[1620] Paul Compton, Kang Byeong, Philip Preston, and Mary Mulholland. Knowledge acquisition without analysis. Lecture Notes in Computer Science, 723: 277–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mengshoel:1993:AMU

[1621] Ole Jakob Mengshoel and Ingeborg Solvberg. Acquisition and modelling of uncertain, incomplete and time-varying knowledge. Lecture Notes in Computer Science, 723:300–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domingue:1993:EVW

[1622] John Domingue, Enrico Motta, and Stuart Watt. The emerging VITAL workbench. Lecture Notes in Computer Science, 723:320–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tijerino:1993:MIE

[1623] Yuri Adrian Tijerino and Riichiro Mizoguchi. MULTIS II: Enabling endusers to design problem-solving engines via two-level task ontologies. Lecture Notes in Computer Science, 723:340– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muzard:1993:PDC

[1624] Joel Muzard, Aziza Hanafi, Andre Dupont, and Benoit Tiffou. The participatory design of a computer assisted knowledge engineering methodology and tool: The ALADIN+Project. Lecture Notes in Computer Science, 723:360–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Addis:1993:KAV

[1625] Tom R. Addis, David C. Goodong, and Jan J. Townsend. Knowledge acquisition with visual functional programming. Lecture Notes in Computer Science, 723:379–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dieng:1993:AGK

[1626] Rose Dieng, Olivier Corby, and Stephane Lapalut. Acquisition of gradual knowledge. Lecture Notes in Computer Science, 723:407–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biebow:1993:AVT

[1627] Brigitte Biebow and Sylvie Szulman. Acquisition and validation: From text to semantic network. Lecture Notes in Computer Science, 723:427–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanHentenryck:1993:IGA

[1628] P. Van Hentenryck, O. Degimbe, B. Le Charlier, and L. Michel. The impact of granularity in abstract interpretation of Prolog. Lecture Notes in Computer Science, 724:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeCharlier:1993:OTG

[1629] B. Le Charlier, O. Degimbe, L. Michel, and P. Van Hentenryck. Optimization techniques for general purpose fixpoint algorithms: practical efficiency for the abstract interpretation of Prolog. Lecture Notes in Computer Science, 724: 15–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jorgensen:1993:CFI

[1630] N. Jorgensen. Chaotic fixpoint iteration guided by dynamic dependency. Lecture Notes in Computer Science, 724:27–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferguson:1993:FAI

[1631] A. Ferguson and J. Hughes. Fast abstract interpretation using sequential algorithms. Lecture Notes in Computer Science, 724:45–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1993:AIV

[1632] J. C. Fernandez. Abstract interpretation and verification of reactive systems. Lecture Notes in Computer Science, 724:60–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cridlig:1993:SAL

[1633] R. Cridlig and E. Goubault. Semantics and analysis of Linda-based languages. Lecture Notes in Computer Science, 724:72–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talpin:1993:CFC

[1634] J-P Talpin and P. Jouvelot. Compiling FX on the CM-2. Lecture Notes in Computer Science, 724:87–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maffeis:1993:CDA

[1635] O. Maffeis and P. Le Guernic. Combining dependability with architectural adaptability by means of the SIGNAL language. Lecture Notes in Computer Science, 724:99–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanRoy:1993:CDU

[1636] P. Van Roy. Challenges in developing useful and practical static analysis for logic programs. Lecture Notes in Computer Science, 724:111–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gluck:1993:ORM

[1637] R. Gluck and A. Klimov. Occam’s razor in metacomputation: the notion of a perfect process tree. Lecture Notes in Computer Science, 724:112–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chin:1993:TFM

[1638] W-N. Chin and S-C. Khoo. Tupling functions with multiple recursion parameters. Lecture Notes in Computer Science, 724:124–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thiemann:1993:ART

[1639] P. Thiemann. Avoiding repeated tests in pattern matching. Lecture Notes in Computer Science, 724:141–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruynooghe:1993:FSL

[1640] M. Bruynooghe and M. Codish. Freeness, sharing, linearity and correctness — all at once. Lecture Notes in Computer Science, 724:153–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boye:1993:SDI

[1641] J. Boye, J. Paakki, and J. Maluszynski. Synthesis of directionality information for functional logic programs. Lecture Notes in Computer Science, 724:165–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bert:1993:AR

[1642] D. Bert, R. Echahed, and B. M.

Ostvold. Abstract rewriting. Lecture Notes in Computer Science, 724:178– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hudak:1993:RPO

[1643] P. Hudak. Reflections on program optimization. Lecture Notes in Computer Science, 724:193–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielson:1993:FCS

[1644] F. Nielson and H. R. Nielson. Finiteness conditions for strictness analysis. Lecture Notes in Computer Science, 724:194–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benton:1993:SPL

[1645] P. N. Benton. Strictness properties of lazy algebraic datatypes. Lecture Notes in Computer Science, 724:206– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amtoft:1993:MT

[1646] T. Amtoft. Minimal thunkification. Lecture Notes in Computer Science, 724:218–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kato:1993:EAR

[1647] S. Kato, H. Seki, and H. Itoh. An efficient abductive reasoning system based on program analysis. Lecture Notes in Computer Science, 724:230–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Errington:1993:CGP

[1648] L. Errington, C. Hankin, and T. Jensen. A congruence for gamma programs. Lecture Notes in Computer Science, 724:242–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wright:1993:UAN

[1649] D. A. Wright and C. A. Baker-Finch. Usage analysis with natural reduction types. Lecture Notes in Computer Science, 724:254–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monsuez:1993:PTW

[1650] B. Monsuez. Polymorphic types and widening operators. Lecture Notes in Computer Science, 724:267–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sahlin:1993:DSA

[1651] D. Sahlin and T. Sjoland. Demonstration: static analysis of AKL. Lecture Notes in Computer Science, 724:282– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Israeli:1993:EWF

[1652] A. Israeli and L. Rappoport. Efficient wait-free implementation of a concurrent priority queue. Lecture Notes in Computer Science, 725:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoepman:1993:BS

[1653] J.-H. Hoepman and J. Tromp. Binary snapshots. Lecture Notes in Computer Science, 725:18–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Israeli:1993:LTS

[1654] A. Israeli, A. Shaham, and A. Shirazi. Linear-time snapshot protocols for unbalanced systems. Lecture Notes in Computer Science, 725:26–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1993:TNS

[1655] J. H. Anderson and M. Moir. Towards a necessary and sufficient condition for wait-free synchronization. Lecture Notes in Computer Science, 725:39–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirousis:1993:EAC

[1656] L. M. Kirousis and A. G. Veneris. Efficient algorithms for checking the atomicity of a run of read and write operations. Lecture Notes in Computer Science, 725:54–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Afek:1993:BFM

[1657] Y. Afek, M. Merritt, and G. Taubenfeld. Benign failure models for shared memory. Lecture Notes in Computer Science, 725:69–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burns:1993:GAB

[1658] J. E. Burns, R. I. Cruz, and M. C. Loui. Generalized agreement between concurrent fail-stop processes. Lecture Notes in Computer Science, 725:84–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanellakis:1993:CMA

[1659] P. C. Kanellakis, D. Michailidis, and A. A. Shvartsman. Controlling memory access concurrency in efficient faulttolerant parallel algorithms. Lecture Notes in Computer Science, 725:99–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabinovich:1993:AEM

[1660] M. Rabinovich and E. D. Lazowska. Asynchronous epoch management in replicated databases. Lecture Notes in Computer Science, 725:115–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dolev:1993:CRC

[1661] S. Dolev and J. L. Welch. Crash resilient communication in dynamic networks. Lecture Notes in Computer Science, 725:129–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choy:1993:DJS

[1662] M. Choy and A. K. Singh. Distributed job scheduling using snapshots. Lecture Notes in Computer Science, 725:145–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dolev:1993:OTS

[1663] S. Dolev. Optimal time self stabilization in dynamic systems. Lecture Notes in Computer Science, 725:160– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anagnostou:1993:TTP

[1664] E. Anagnostou and V. Hadzilacos. Tolerating transient and permanent failures. Lecture Notes in Computer Science, 725:174–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1993:QAB

[1665] P. Berman and A. A. Bharali. Quick atomic broadcast. Lecture Notes in Computer Science, 725:189–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attiya:1993:TBD

[1666] H. Attiya and T. Djerassi-Shintel.

Time bounds for decision problems in the presence of timing uncertainty and failures. Lecture Notes in Computer Science, 725:204–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flammini:1993:BR

[1667] M. Flammini, G. Gambosi, and S. Salomone. Boolean routing. Lecture Notes in Computer Science, 725:219– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hardavellas:1993:NSC

[1668] N. Hardavellas, D. Karakos, and M. Mavronicolas. Notes on sorting and counting networks. Lecture Notes in Computer Science, 725:234–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Higham:1993:SEA

[1669] L. Higham and T. Przytycka. A simple, efficient algorithm for maximum finding on rings. Lecture Notes in Computer Science, 725:249–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Syrotiuk:1993:WTD

[1670] V. R. Syrotiuk, C. J. Colbourn, and J. Pachl. Wang tilings and distributed orientation on anonymous torus networks. Lecture Notes in Computer Science, 725:264–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1993:FPS

[1671] C. Wu, G. V. Bochmann, and M. Yao. Fairness of N-party synchronization and its implementation in a distributed environment. Lecture Notes in Computer Science, 725:279–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boniol:1993:PDR

[1672] F. Boniol and M. Adelantado. Programming distributed reactive systems: a strong and weak synchronous coupling. Lecture Notes in Computer Science, 725:294–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leong:1993:UMS

[1673] H. V. Leong, D. Agrawal, and J. R.

Agre. Using message semantics to reduce rollback in the time warp mechanism. Lecture Notes in Computer Science, 725:309–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1993:ILC

[1674] S. Albers. The influence of lookahead in competitive paging algorithms. Lecture Notes in Computer Science, 726:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atallah:1993:OAS

[1675] M. J. Atallah, D. Z. Chen, and D. T.

Lee. An optimal algorithm for shortest paths on weighted interval and circulararc graphs, with applications. Lecture Notes in Computer Science, 726:13–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-Asher:1993:ESS

[1676] Y. Ben-Asher, D. Gordon, and A. Schuster. Efficient self simulation algorithms for reconfigurable arrays. Lecture Notes in Computer Science, 726: 25–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertolazzi:1993:OUP

[1677] P. Bertolazzi, G. di Battista, C. Mannino, and R. Tamassia. Optimal upward planarity testing of single-source digraphs. Lecture Notes in Computer Science, 726:37–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhatt:1993:BRV

[1678] S. N. Bhatt, G. Bilardi, G. Pucci, A. Ranade, A. L. Rosenberg, and E. J. Schwabe. On bufferless routing of variable-length messages in leveled networks. Lecture Notes in Computer Science, 726:49–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breslauer:1993:SCC

[1679] D. Breslauer. Saving comparisons in the Crochemore-Perrin string matching algorithm. Lecture Notes in Computer Science, 726:61–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruggemann-Klein:1993:UER

[1680] A. Brüggemann-Klein. Unambiguity of extended regular expressions in SGML document grammars. Lecture Notes in Computer Science, 726:73–84, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bshouty:1993:DSC

[1681] N. H. Bshouty. On the direct sum conjecture in the straight line model. Lecture Notes in Computer Science, 726: 85–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1993:CCG

[1682] R. F. Cohen and R. Tamassia. Combine and conquer: a general technique for dynamic algorithms. Lecture Notes in Computer Science, 726:97–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Datta:1993:OCP

[1683] A. Datta, A. Maheshwari, and J.- R. Sack. Optimal CREW-PRAM algorithms for direct dominance problems. Lecture Notes in Computer Science, 726:109–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deBerg:1993:TAF

[1684] M. de Berg and M. Van Kreveld.

Trekking in the Alps without freezing or getting tired. Lecture Notes in Computer Science, 726:121–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devillers:1993:DBP

[1685] O. Devillers and M. Golin. Dog bites postman: Point location in the moving Voronoi diagram and related problems. Lecture Notes in Computer Science, 726:133–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz:1993:PAS

[1686] J. Diaz, M. J. Serna, and J. Toran. Parallel approximation schemes for problems on planar graphs. Lecture Notes in Computer Science, 726:145–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fellows:1993:DPM

[1687] M. R. Fellows, M. T. Hallett, and H. T. Wareham. DNA physical mapping: Three ways difficult. Lecture Notes in Computer Science, 726:157–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flajolet:1993:CRG

[1688] P. Flajolet, P. Zimmermann, and B. Van Cutsem. A calculus of random generation. Lecture Notes in Computer Science, 726:169–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerstel:1993:BCD

[1689] O. Gerstel and S. Zaks. The bit complexity of distributed sorting. Lecture Notes in Computer Science, 726:181– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagauer:1993:TCP

[1690] J. Hagauer and G. Rote. Threeclustering of points in the plane. Lecture Notes in Computer Science, 726: 192–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1993:GVD

[1691] J. Hromkovic, R. Klasing, E. A. Stohr, and H. Wagener. Gossiping in vertexdisjoint paths mode in d-dimensional grids and planar graphs. Lecture Notes in Computer Science, 726:200– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Italiano:1993:FDP

[1692] G. F. Italiano, J. A. La Poutre, and M. H. Rauch. Fully dynamic planarity testing in planar embedded graphs. Lecture Notes in Computer Science, 726:212–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ivkovic:1993:FDA

[1693] Z. Ivkovic and E. L. Lloyd. Fully dynamic algorithms for bin packing: Being (mostly) myopic helps. Lecture Notes in Computer Science, 726:224– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jordan:1993:IVC

[1694] T. Jordan. Increasing the vertexconnectivity in directed graphs. Lecture Notes in Computer Science, 726: 236–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klinz:1993:RPB

[1695] B. Klinz, R. Rudolf, and G. J. Woeginger. On the recognition of permuted bottleneck Monge matrices. Lecture Notes in Computer Science, 726:248– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1993:CTM

[1696] T. Kloks, H. Bodlaender, H. Muller, and D. Kratsch. Computing treewidth and minimum fill-in: All you need are the minimal separators. Lecture Notes in Computer Science, 726:260– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kunde:1993:BGG

[1697] M. Kunde. Block gossiping on grids and tori: Deterministic sorting and routing match the bisection bound. Lecture Notes in Computer Science, 726: 272–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenstra:1993:CST

[1698] J. K. Lenstra, M. Veldhorst, and B. Veltman. The complexity of scheduling trees with communication delays. Lecture Notes in Computer Science, 726:284–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1993:OTC

[1699] E. W. Mayr and R. Werchner. Optimal tree contraction on the hypercube and related networks. Lecture Notes in Computer Science, 726:295–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paterson:1993:EA

[1700] M. Paterson. Evolution of an algorithm. Lecture Notes in Computer Science, 726:306–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajasekaran:1993:MCC

[1701] S. Rajasekaran. Mesh connected computers with fixed and reconfigurable buses: Packet routing, sorting, and selection. Lecture Notes in Computer Science, 726:309–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramachandran:1993:EPA

[1702] V. Ramachandran and H. Yang. An efficient parallel algorithm for the layered planar monotone circuit value problem. Lecture Notes in Computer Science, 726:321–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1993:RRM

[1703] J. F. Sibeyn, M. Kaufmann, and R. Raman. Randomized routing on meshes with buses. Lecture Notes in Computer Science, 726:333–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1993:DTC

[1704] K. Simon, D. Crippa, and F. Collenberg. On the distribution of the transitive closure in a random acyclic digraph. Lecture Notes in Computer Science, 726:345–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schrijver:1993:CDP

[1705] A. Schrijver. Complexity of disjoint paths problems in planar graphs. Lecture Notes in Computer Science, 726: 357–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srivastav:1993:IMF

[1706] A. Srivastav and P. Stangier. Integer multicommodity flows with reduced demands. Lecture Notes in Computer Science, 726:360–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subramanian:1993:FDD

[1707] S. Subramanian. A fully dynamic data structure for reachability in planar digraphs. Lecture Notes in Computer Science, 726:372–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1993:LTA

[1708] D. Wagner and K. Weihe. A lineartime algorithm for edge-disjoint paths in planar graphs. Lecture Notes in Computer Science, 726:384–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waterman:1993:SCS

[1709] M. S. Waterman. Sequence comparison and statistical significance in molecular biology. Lecture Notes in Computer Science, 726:396–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welzl:1993:SRB

[1710] E. Welzl and B. Wolfers. Surface reconstruction between simple polygons via angle criteria. Lecture Notes in Computer Science, 726:397–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1993:LAE

[1711] X. Zhou, S. Nakano, and T. Nishizeki. A linear algorithm for edge-coloring partial k-trees. Lecture Notes in Computer Science, 726:409–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasser:1993:OCD

[1712] Les Gasser, Ingemar Hulthage, Brian Leverich, Jon Lieb, and Ann Majchrzak. Organizations as complex, dynamic design problems. Lecture Notes in Computer Science, 727:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1993:SVD

[1713] Michael Fisher and Michael Wooldridge. Specifying and verifying distributed intelligent systems. Lecture Notes in Computer Science, 727:13–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horstmann:1993:LAD

[1714] Thilo C. Horstmann. A logical approach for distributed truth maintenance. Lecture Notes in Computer Science, 727:29–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodrigues:1993:BTT

[1715] Irene Pimenta Rodrigues and Jose G.

Lopes. Building text temporal structure. Lecture Notes in Computer Science, 727:45–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damas:1993:RCA

[1716] Luis Damas, Nelma Moreira, and Sabine Broda. Resolution of constraints in algebras of rational trees. Lecture Notes in Computer Science, 727:61–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1993:IHT

[1717] Gabriel David and Antonio Porto. Inheritance in a hierarchy of theories. Lecture Notes in Computer Science, 727:77–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reichwein:1993:SIK

[1718] Georg Reichwein and Jose Fiadeiro. Semantics of interworking knowledge based systems. Lecture Notes in Computer Science, 727:93–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schild:1993:CTL

[1719] Klaus Schild. Combining terminological logics with tense logic. Lecture Notes in Computer Science, 727:105– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mamede:1993:TCA

[1720] Margarida Mamede and Luis Monteiro. Towards complete answers in concept languages. Lecture Notes in Computer Science, 727:121–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1993:VSA

[1721] Grigoris Antoniou. Version space algorithms on hierarchies with exceptions. Lecture Notes in Computer Science, 727:136–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matos:1993:RLC

[1722] Armando B. Matos. Regular languages and a class of logic programs. Lecture Notes in Computer Science, 727:150– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matos:1993:SRC

[1723] Armando B. Matos. Some results on the complexity of SLD-Derivations. Lecture Notes in Computer Science, 727:158–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silva:1993:PPE

[1724] Fernando M. A. Silva. An or-parallel Prolog execution model for a distributed shared memory machine. Lecture Notes in Computer Science, 727: 170–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1993:DDC

[1725] Luis Moniz Pereira, Carlos Viegas Damasio, and Jose J. Alferes. Diagnosis and debugging as contradiction removal in logic programs. Lecture Notes in Computer Science, 727:183– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalinski:1993:WFA

[1726] Jurgen Kalinski. Well-founded approximations of autoepistemic logic. Lecture Notes in Computer Science, 727: 198–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1993:PMD

[1727] Grigoris Antoniou and Elmar Langetepe. A process model for default logic and its realization in logic programming. Lecture Notes in Computer Science, 727:213–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cravo:1993:UAD

[1728] Maria R. Cravo and Joao P. Martins. A unified approach to default reasoning and belief revision. Lecture Notes in Computer Science, 727:226–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curien:1993:SOM

[1729] Regis Curien. Second order E-matching as a tool for automated theorem proving. Lecture Notes in Computer Science, 727:242–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haselbock:1993:ASI

[1730] Alois Haselbock and Markus Stumptner. Attribute-specific interchangeability in constraint satisfaction problems. Lecture Notes in Computer Science, 727:258–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berlandier:1993:UIM

[1731] Pierre Berlandier. The use and interpretation of meta level constraints. Lecture Notes in Computer Science, 727: 271–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menezes:1993:PFI

[1732] Francisco Menezes and Pedro Barahona. Preliminary formalization of an incremental hierarchical constraint solver. Lecture Notes in Computer Science, 727:281–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filgueiras:1993:FMS

[1733] Miguel Filgueiras and Ana P. Tomas.

Fast methods for solving linear Diophantine equations. Lecture Notes in Computer Science, 727:297–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1993:NCM

[1734] Maria Fox and Derek Long. A note on Chapman’s modal truth criterion. Lecture Notes in Computer Science, 727:307–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferreira:1993:LOW

[1735] Jose L. S. Ferreira and Ernesto J. F.

Costa. Learning operators while planning. Lecture Notes in Computer Science, 727:311–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luck:1993:EEM

[1736] Michael M. Luck. Evaluating evidence for motivated discovery. Lecture Notes in Computer Science, 727:324– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auriol:1993:HLI

[1737] Eric Auriol. How to learn in an incomplete knowledge environment. Lecture Notes in Computer Science, 727: 340–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giesl:1993:SRC

[1738] Jurgen Giesl and Ingrid Neumann. The semantics of rational contractions. Lecture Notes in Computer Science, 727: 356–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demolombe:1993:UFD

[1739] Robert Demolombe. A uniform framework for deductive database derivation strategies. Lecture Notes in Computer Science, 727:357–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aparicio:1993:ASS

[1740] Joaquim Nunes Aparicio. Assumption set semantics (the procedures). Lecture Notes in Computer Science, 727: 358–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bussmann:1993:BA

[1741] Stefan Bussmann and Jurgen Muller. Bargaining agents. Lecture Notes in Computer Science, 727:359–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trombettoni:1993:CMA

[1742] Gilles Trombettoni. A constraint maintenance algorithm based on value propagation. Lecture Notes in Computer Science, 727:360–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deCampos:1993:TPT

[1743] Geraldo Lima de Campos and Dimas Trevisan Chbane. A text-to-phoneme translator for the Portuguese language. Lecture Notes in Computer Science, 727:361–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diday:1993:LPM

[1744] E. Diday, P. Brito, and E. Mfoumoune.

Learning probabilistic models by conceptual pyramidal clustering. Lecture Notes in Computer Science, 727:362– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

diManzo:1993:PFT

[1745] M. di Manzo, E. Giunchiglia, A. Armando, and P. Pecchiari. Proving formulas through reduction to decidable classes. Lecture Notes in Computer Science, 728:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armando:1993:BEP

[1746] A. Armando, A. Cimatti, and L. Vigano. Building and executing proof strategies in a formal metatheory. Lecture Notes in Computer Science, 728:11–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamma:1993:CVS

[1747] E. Lamma and P. Mello. Computing 3-valued stable models by using the ATMS. Lecture Notes in Computer Science, 728:23–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1993:APC

[1748] G. Antoniou. Abstract properties for the choice provability relation in nonmonotonic logics. Lecture Notes in Computer Science, 728:35–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pirri:1993:CPI

[1749] F. Pirri and C. Pizzuti. Characterizing prime implicants as projective spaces. Lecture Notes in Computer Science, 728:41–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conati:1993:ESM

[1750] C. Conati and J. F. Lehman. EFH-Soar: Modeling education in highly interactive microworlds. Lecture Notes in Computer Science, 728:47–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelfranchi:1993:FID

[1751] C. Castelfranchi, A. Cesta, R. Conte, and M. Miceli. Foundations for interaction: the dependence theory. Lecture Notes in Computer Science, 728:59–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McGraw:1993:LSA

[1752] G. McGraw and D. Hofstadter. Letter spirit: an architecture for creativity in a microdomain. Lecture Notes in Computer Science, 728:65–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ardizzone:1993:NSE

[1753] E. Ardizzone, A. Chella, and R. Pirrone. New systems for extracting 3-D shape information from images. Lecture Notes in Computer Science, 728:71–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gori:1993:PSS

[1754] M. Gori and G. Soda. Projecting subsymbolic onto symbolic representations in artificial neural networks. Lecture Notes in Computer Science, 728:84–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braggiotti:1993:ISS

[1755] A. Braggiotti, G. Chemello, C. Sossai, and G. Trainito. Integrating the symbolic and the sub-symbolic level in sonar-based navigation. Lecture Notes in Computer Science, 728:90–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Terna:1993:RIR

[1756] P. Terna. Randomness, imitation or reason explain agents’ behaviour into an artificial stock market? Lecture Notes in Computer Science, 728:96–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monfroglio:1993:NNC

[1757] A. Monfroglio. Neural networks for constraint satisfaction. Lecture Notes in Computer Science, 728:102– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaerf:1993:RIC

[1758] A. Schaerf. Reasoning with individuals in concept languages. Lecture Notes in Computer Science, 728:108–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettini:1993:FTT

[1759] C. Bettini. A family of temporal terminological logics. Lecture Notes in Computer Science, 728:120–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaerf:1993:LPA

[1760] M. Schaerf. Logic programming and autoepistemic logics: new relations and complexity results. Lecture Notes in Computer Science, 728:132–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaspari:1993:ILL

[1761] M. Gaspari, E. Motta, and A. Stutt. Inferring in Lego-land: an architecture for the integration of heterogeneous inference modules. Lecture Notes in Computer Science, 728:142–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adorni:1993:MLM

[1762] G. Adorni and A. Poggi. MAP — a language for the modelling of multi-agent systems. Lecture Notes in Computer Science, 728:154–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vossos:1993:DCO

[1763] G. Vossos, J. Zeleznikow, and D. Hunter. Developing co-operating legal knowledge based systems. Lecture Notes in Computer Science, 728:160– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esposito:1993:NSO

[1764] F. Esposito, D. Malerba, and G. Semeraro. Negation as a specializing operator. Lecture Notes in Computer Science, 728:166–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nienhuys-Cheng:1993:CRO

[1765] S. H. Nienhuys-Cheng, P. R. J. van der Laag, and L. W. N. van der Torre. Constructing refinement operators by decomposing logical implication. Lecture Notes in Computer Science, 728:178–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergadano:1993:LRB

[1766] F. Bergadano and D. Gunetti. Learning relations: basing top-down methods on inverse resolution. Lecture Notes in Computer Science, 728:190–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guvenir:1993:CCM

[1767] H. Altay Guvenir and I. Sirin. Complexity of the CFP, a method for classification based on feature partitioning. Lecture Notes in Computer Science, 728:202–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larranaga:1993:GAE

[1768] P. Larranaga, M. Grana, A. D’Anjou, and F. J. Torrealdea. Genetic algorithms elitist probabilistic of degree 1, a generalization of simulated annealing. Lecture Notes in Computer Science, 728:208–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordana:1993:LRU

[1769] A. Giordana, L. Saitta, M. E. Campidoglio, and G. Lo Bello. Learning relations using genetic algorithms. Lecture Notes in Computer Science, 728: 218–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pelillo:1993:ELR

[1770] M. Pelillo, F. Abbattista, and A. Maffione. Evolutionary learning for relaxation labeling processes. Lecture Notes in Computer Science, 728:230– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maier:1993:ICA

[1771] E. A. Maier and E. Not. Increasing cohesion in automatically generated natural language texts. Lecture Notes in Computer Science, 728:242–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ardissono:1993:PCA

[1772] L. Ardissono, L. Lesmo, A. Lombardo, and D. Sestero. Production of cooperative answers on the basis of partial knowledge in information-seeking dialogues. Lecture Notes in Computer Science, 728:254–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciravegna:1993:CMR

[1773] F. Ciravegna and E. Giorda. Coping with modifiers in a restricted domain. Lecture Notes in Computer Science, 728:266–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1993:EST

[1774] G. Ferrari, M. Carenini, and P. Moreschini. Explanation strategies in a tutoring system. Lecture Notes in Computer Science, 728:272–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cervoni:1993:MCQ

[1775] R. Cervoni, A. Cesta, and A. Oddi.

Maintaining consistency in quantitative temporal constraint networks for planning and scheduling. Lecture Notes in Computer Science, 728:278– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badaloni:1993:MAR

[1776] S. Badaloni, E. Pagello, L. Stocchiero, and A. Zanardi. Making an autonomous robot plan temporally constrained maintenance operations. Lecture Notes in Computer Science, 728:

290–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stumptner:1993:GCF

[1777] M. Stumptner and A. Haselbock. A generative constraint formalism for configuration problems. Lecture Notes in Computer Science, 728:302–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Portinale:1993:SOT

[1778] L. Portinale. Selecting observation time in the monitoring and interpretation of time-varying data. Lecture Notes in Computer Science, 728:314–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varzi:1993:SRH

[1779] A. C. Varzi. Spatial reasoning in a holey world. Lecture Notes in Computer Science, 728:326–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayani:1993:PS

[1780] R. Ayani. Parallel simulation. Lecture Notes in Computer Science, 729:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balsamo:1993:PAQ

[1781] S. Balsamo. Properties and analysis of queueing network models with finite capacities. Lecture Notes in Computer Science, 729:21–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blanc:1993:PAO

[1782] H. Blanc. Performance analysis and optimization with the power-series algorithm. Lecture Notes in Computer Science, 729:53–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conti:1993:MAN

[1783] M. Conti, E. Gregori, and L. Lenzini.

Metropolitan area networks (MANs): Protocols, modeling and performance evaluation. Lecture Notes in Computer Science, 729:81–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gotz:1993:MDS

[1784] N. Gotz, U. Herzog, and M. Rettelbach. Multiprocessor and distributed system design: The integration of functional specification and performance analysis using stochastic process algebras. Lecture Notes in Computer Science, 729:121–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1993:RTD

[1785] P. G. Harrison. Response time distributions in queueing network models. Lecture Notes in Computer Science, 729:147–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heidelberger:1993:FSR

[1786] P. Heidelberger. Fast simulation of rare events in queueing and reliability models. Lecture Notes in Computer Science, 729:165–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hellerstein:1993:IMD

[1787] J. L. Hellerstein. An introduction to modeling dynamic behavior with time series analysis. Lecture Notes in Computer Science, 729:203–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaeli:1993:ITD

[1788] D. R. Kaeli. Issues in trace-driven simulation. Lecture Notes in Computer Science, 729:224–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kouvatsos:1993:MEA

[1789] D. Kouvatsos. Maximum entropy analysis of queueing network models. Lecture Notes in Computer Science, 729: 245–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindemann:1993:PMU

[1790] C. Lindemann. Performance modeling using DSPNexpress. Lecture Notes in Computer Science, 729:291–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lubachevsky:1993:RMP

[1791] B. D. Lubachevsky. Relaxation for massively parallel discrete event simulation. Lecture Notes in Computer Science, 729:307–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucantoni:1993:BGQ

[1792] D. M. Lucantoni. The BMAP/G/1 QUEUE: A tutorial. Lecture Notes in Computer Science, 729:330–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melamed:1993:OTP

[1793] B. Melamed. An overview of tes processes and modeling methodology. Lecture Notes in Computer Science, 729: 359–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woodside:1993:PEC

[1794] C. Murray Woodside. Performance engineering of client-server systems. Lecture Notes in Computer Science, 729: 394–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Onvural:1993:QNF

[1795] R. O. Onvural. Queueing networks with finite capacities. Lecture Notes in Computer Science, 729:411–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panzieri:1993:RTS

[1796] F. Panzieri and R. Davoli. Real time systems: A tutorial. Lecture Notes in Computer Science, 729:435–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reed:1993:PIT

[1797] D. A. Reed. Performance instrumentation techniques for parallel systems. Lecture Notes in Computer Science, 729:463–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schweitzer:1993:SBA

[1798] P. J. Schweitzer, G. Serazzi, and M. Broglia. A survey of bottleneck analysis in closed networks of queues. Lecture Notes in Computer Science, 729:491–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1993:SPE

[1799] C. U. Smith. Software performance engineering. Lecture Notes in Computer Science, 729:509–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thurner:1993:PMU

[1800] E. M. Thurner. Performance measurement using system monitors. Lecture Notes in Computer Science, 729:537– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Towsley:1993:PQS

[1801] D. Towsley. Providing quality of service packet switched networks. Lecture Notes in Computer Science, 729:560– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trivedi:1993:DPA

[1802] K. S. Trivedi, G. Ciardo, M. Malhotra, and R. A. Sahner. Dependability and performability analysis. Lecture Notes in Computer Science, 729: 587–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rangan:1993:AAD

[1803] P. Venkat Rangan. Architectures and algorithms for digital multimedia ondemand servers. Lecture Notes in Computer Science, 729:613–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yechiali:1993:ACP

[1804] U. Yechiali. Analysis and control of polling systems. Lecture Notes in Computer Science, 729:630–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1993:MAT

[1805] P. S. Yu. Modeling and analysis of transaction processing systems. Lecture Notes in Computer Science, 729: 651–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maier:1993:SSA

[1806] David Maier, Jonathan Walpole, and Richard Staehli. Storage system architectures for continuous media data. Lecture Notes in Computer Science, 730:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang-Jen:1993:PSO

[1807] Chen Huang-Jen and T. D. C. Little.

Physical storage organizations for timedependent multimedia data. Lecture Notes in Computer Science, 730:19–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohtsu:1993:UIM

[1808] Takashi Ohtsu and Michael A. Harrison. User interface management system embedded in a multimedia document editor framework. Lecture Notes in Computer Science, 730:35–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghandeharizadeh:1993:MPR

[1809] Shahram Ghandeharizadeh and Cyrus Shahabi. Management of physical replicas in parallel multimedia information systems. Lecture Notes in Computer Science, 730:51–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agrawal:1993:ESS

[1810] Rakesh Agrawal, Christos Faloutsos, and Arun Swami. Efficient similarity search in sequence databases. Lecture Notes in Computer Science, 730:69–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahn:1993:FH

[1811] Ilsoo Ahn. Filtered hashing. Lecture Notes in Computer Science, 730:85– 100, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devine:1993:DID

[1812] Robert Devine. Design and implementation of DDH: A distributed dynamic hashing algorithm. Lecture Notes in Computer Science, 730:101–114, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sung:1993:PAS

[1813] Sam Y. Sung. Performance analysis of superimposing-coded signature files. Lecture Notes in Computer Science, 730:115–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merrett:1993:TMR

[1814] T. H. Merrett and Heping Shang. Trie methods for representing text. Lecture Notes in Computer Science, 730:130– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitagawa:1993:EFD

[1815] Hiroyuki Kitagawa, Yoshiaki Fukushima, Yoshiharu Ishikawa, and Nobuo Ohbo. Estimation of false drops in set-valued object retrieval with signature files. Lecture Notes in Computer Science, 730:146–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kambayashi:1993:DSA

[1816] Yahiko Kambayashi, Hiroki Takakura, and Shintaro Meki. Data structure and algorithms for new hardware technology. Lecture Notes in Computer Science, 730:164–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biskup:1993:EUB

[1817] Joachim Biskup and Holger Stiefeling. Evaluation of upper bounds and least nodes as database operations. Lecture Notes in Computer Science, 730:197–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1993:CBP

[1818] Arthur M. Keller and Catherine Hamon. A C++ binding for penguin: a system for data sharing among heterogeneous object models. Lecture Notes in Computer Science, 730:215– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grosky:1993:DSE

[1819] William I. Grosky, Sun Junping, and Farshad Fotouhi. Dynamic selectivity estimation for multidimensional queries. Lecture Notes in Computer Science, 730:231–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agrawal:1993:RPT

[1820] Rakesh Agrawal and Jerry Kiernan. Reducing page thrashing in recursive query processing. Lecture Notes in Computer Science, 730:247–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Litwin:1993:HJR

[1821] Witold Litwin, Tomasz Imielinski, David Lomet, and Marie-Ann Neimat. Highways and jukeboxes: A revolution for data structures? Lecture Notes in Computer Science, 730:264–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1993:SRS

[1822] Yongdon Wang, Jane Chiao, and Vasu Nori. The Sybase Replication Server Project. Lecture Notes in Computer Science, 730:265–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimada:1993:AIC

[1823] Shigeru Shimada, Toshihisa Aoshima, and Tetsuzo Uehara. Analogical inferred compensation method for heterogeneous database schema mismatch and its application to hypermedia cooperative work system. Lecture Notes in Computer Science, 730:271– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohan:1993:SDR

[1824] C. Mohan. A survey of DBMS research issues in supporting very large tables. Lecture Notes in Computer Science, 730:279–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerlhof:1993:PBC

[1825] Carsten Gerlhof, Alfons Kemper, Christoph Kilger, and Guido Moerkotte. Partition-based clustering in object bases: From theory to practice. Lecture Notes in Computer Science, 730:301–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dathan:1993:AIR

[1826] Brahma Dathan. An algorithm for the implementation of replicated tables. Lecture Notes in Computer Science, 730:317–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orji:1993:ITM

[1827] Cyril U. Orji, Mark A. Weiss, and Jon A. Solworth. Improved traditional mirror. Lecture Notes in Computer Science, 730:329–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheuermann:1993:ALB

[1828] Peter Scheuermann, Gerhard Weikum, and Peter Zabback. Adaptive load balancing in disk arrays. Lecture Notes in Computer Science, 730:345–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grossman:1993:SDC

[1829] Robert Grossman, Yannis Ioannidis, Edward May, Sakti Pramanik, and Ruth Pordes. Scientific databases: Challenges to the database community. Lecture Notes in Computer Science, 730:361–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takada:1993:OOO

[1830] Hideyuki Takada and Yahiko Kambayashi. An object-oriented office space description model and an office view management mechanism for distributed office environment. Lecture Notes in Computer Science, 730:362– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cellary:1993:SLM

[1831] Wojciech Cellary and Waldemar Wieczerzycki. Stamp locking method for multiversion composite objects. Lecture Notes in Computer Science, 730:378–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minohara:1993:QSH

[1832] Tatsuo Minohara, Ryuichi Watanabe, and Mario Tokoro. Queries on structures in hypertext. Lecture Notes in Computer Science, 730:394–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bever:1993:DSO

[1833] M. Bever, K. Geihs, L. Heuser, M. Muhlhauser, and A. Schill. Distributed systems, OSF DCE, and beyond. Lecture Notes in Computer Science, 731:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beitz:1993:CTD

[1834] A. Beitz, P. King, and K. Raymond.

Comparing two distributed environments: DCE and ANSAware. Lecture Notes in Computer Science, 731:21–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabenseifner:1993:CDR

[1835] R. Rabenseifner and A. Schuch. Comparison of DCE RPC, DFN-RPC, ONC and PVM. Lecture Notes in Computer Science, 731:39–46, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dasarathy:1993:SDP

[1836] B. Dasarathy, K. Khalil, and D. E.

Ruddock. Some DCE performance analysis results. Lecture Notes in Computer Science, 731:47–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinka:1993:PSD

[1837] J. Martinka, R. Friedrich, P. Friedenbach, and T. Sienknecht. A performance study of the DCE 1.0.1 cell directory service: Implications for application and tool programmers. Lecture Notes in Computer Science, 731:63–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laifer:1993:FTU

[1838] R. Laifer and A. Knocke. fidl — A tool for using DCE from Fortran. Lecture Notes in Computer Science, 731:78–88, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McDonald:1993:CLF

[1839] T. M. McDonald. Converting legacy Fortran applications to distributed applications. Lecture Notes in Computer Science, 731:89–103, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirsch:1993:UST

[1840] B. S. Hirsch. Using standard tools to build an open, client/server prototype. Lecture Notes in Computer Science, 731:104–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narkiewicz:1993:POD

[1841] J. D. Narkiewicz, M. Girkar, M. Srivastava, A. S. Gaylord, and M. Rahman. Pilgrim’s OSF DCE-based services architecture. Lecture Notes in Computer Science, 731:120–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wei:1993:CMP

[1842] Y.-H. Wei, S. S. B. Shi, and D. D. H.

Lin. Converting monolithic programs for DCE client/server computing given incomplete cutset information. Lecture Notes in Computer Science, 731:135– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tikku:1993:MTO

[1843] S. Tikku, S. Vinter, and S. Bertrand.

Managing the transition to OSF DCE security. Lecture Notes in Computer Science, 731:147–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Obrenic:1993:DCU

[1844] B. Obrenic, K. S. Dibella, and A. S.

Gaylord. DCE cells under megascope: Pilgrim insight into the resource status. Lecture Notes in Computer Science, 731:162–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adcock:1993:SCM

[1845] P. Adcock, N. Davies, and G. S.

Blair. Supporting continuous media in open distributed systems architectures. Lecture Notes in Computer Science, 731:179–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wei:1993:IRM

[1846] Y.-H. Wei and C. Wu. Integrating RPC and message passing for distributed programming. Lecture Notes in Computer Science, 731:192–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mittasch:1993:OSS

[1847] C. Mittasch and S.-I. Diethmann. Optimized selection of servers for reduced response time in RPC. Lecture Notes in Computer Science, 731:207–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heite:1993:EDR

[1848] R. Heite and H. Eberle. Extending DCE RPC by dynamic objects and dynamic typing. Lecture Notes in Computer Science, 731:214–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teng:1993:SOI

[1849] Q. Teng, Y. Xie, and B. Martin. A simple ORB implementation on top of DCE for distributed object-oriented programming. Lecture Notes in Computer Science, 731:229–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mock:1993:DDC

[1850] M. U. Mock. DCE++: Distributing C++-objects using OSF DCE. Lecture Notes in Computer Science, 731: 242–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dilley:1993:OOD

[1851] J. Dilley. Object-oriented distributed computing with C++ and OSF DCE. Lecture Notes in Computer Science, 731:256–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gellersen:1993:GDS

[1852] H.-W. Gellersen. Graphical design support for DCE applications. Lecture Notes in Computer Science, 731:267– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Svobodova:1993:THN

[1853] L. Svobodova. Taming heterogeneity in networked environments. Lecture Notes in Computer Science, 731:282– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwider:1993:OI

[1854] J. Schwider, N. Streibl, and K. Zurl.

Optoelectronic interconnections. Lecture Notes in Computer Science, 732: 1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1993:MME

[1855] F. Hofmann, M. Dal Cin, A. Grygier, H. Hessenauer, U. Hildebrand, C.-U. Linster, T. Thiel, and S. Turowski. MEMSY — A modular expandable multiprocessor system. Lecture Notes in Computer Science, 732:15–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DalCin:1993:FTD

[1856] M. Dal Cin, A. Grygier, H. Hessenauer, U. Hildebrand, J. Honig, W. Hohl, E. Michel, and A. Pataricza. Fault tolerance in distributed shared memory multiprocessors. Lecture Notes in Computer Science, 732:31–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jessen:1993:OMC

[1857] E. Jessen, W. Ertel, and Ch. Suttner. Optimal multiprogramming control for parallel computations. Lecture Notes in Computer Science, 732:49–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1993:DHM

[1858] R. Hofmann. The distributed hardware monitor ZM4 and its interface to MEMSY. Lecture Notes in Computer Science, 732:66–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartleb:1993:GMP

[1859] F. Hartleb. Graph models for performance evaluation of parallel programs. Lecture Notes in Computer Science, 732:80–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludwig:1993:LMM

[1860] Th. Ludwig. Load management on multiprocessor systems. Lecture Notes in Computer Science, 732:87–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hellwagner:1993:RSM

[1861] H. Hellwagner. Randomized shared memory — concept and efficiency of a scalable shared memory scheme. Lecture Notes in Computer Science, 732: 102–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bockle:1993:MEF

[1862] G. Bockle, Ch. Stormann, and I. Wildgruber. Methods for exploitation of fine-grained parallelism. Lecture Notes in Computer Science, 732:118– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomm:1993:CBP

[1863] D. Gomm and E. Kindler. Causality based proof of a distributed shared memory system. Lecture Notes in Computer Science, 732:133–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kleinoder:1993:OMM

[1864] J. Kleinoder. Object- and memorymanagement architecture — A concept for open, object-oriented operating systems. Lecture Notes in Computer Science, 732:150–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faustle:1993:ODL

[1865] M. Faustle. An orthogonal distribution language for uniform object-oriented languages. Lecture Notes in Computer Science, 732:166–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hauck:1993:TIU

[1866] F. Hauck. Towards the implementation of a uniform object model. Lecture Notes in Computer Science, 732: 180–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dederichs:1993:FFD

[1867] F. Dederichs, C. Dendorfer, and R. Weber. Focus: A formal design method for distributed systems. Lecture Notes in Computer Science, 732:190–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1993:PSN

[1868] V. Fischer and H. Niemann. Parallelism in a semantic network for image understanding. Lecture Notes in Computer Science, 732:203–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spruth:1993:APS

[1869] H. Spruth and F. Johannes. Architectures for parallel slicing enumeration in VLSI layout. Lecture Notes in Computer Science, 732:219–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krauss:1993:AFP

[1870] P. Krauss and K. Antreich. Application of fault parallelism to the automatic test pattern generation for sequential circuits. Lecture Notes in Computer Science, 732:234–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pawlowski:1993:PSL

[1871] M. Pawlowski and R. Bayer. Parallel sorting of large data volumes on distributed memory multiprocessors. Lecture Notes in Computer Science, 732: 246–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fruchtl:1993:QMP

[1872] H. Fruchtl and P. Otto. Quantum mechanical programs for distributed systems: Strategies and results. Lecture Notes in Computer Science, 732:265– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griebel:1993:PSP

[1873] M. Griebel, W. Huber, T. Stortkuhl, and C. Zenger. On the parallel solution of 3D PDEs on a network of workstations and on vector computers. Lecture Notes in Computer Science, 732:276–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peric:1993:NSC

[1874] M. Peric, M. Schafer, and E. Schreck.

Numerical simulation of complex fluid flows on MIMD computers. Lecture Notes in Computer Science, 732:292– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eichelberg:1993:IIG

[1875] Dominik Eichelberg and Philipp Ackermann. Integrating interactive 3Dgraphics into an object-oriented application framework. Lecture Notes in Computer Science, 733:3–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haishan:1993:EKB

[1876] Huang Haishan and Ernest Edmonds.

An experimental knowledge-based user interface management system. Lecture Notes in Computer Science, 733:13–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Penz:1993:OFO

[1877] Franz Penz and Luis Carrico. Objects feeling objects in a MultiView object space. Lecture Notes in Computer Science, 733:25–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaananen:1993:SMB

[1878] Kaisa Vaananen. ShareME: A metaphor-based authoring tool for multimedia environments. Lecture Notes in Computer Science, 733:39–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maurer:1993:SBH

[1879] H. Maurer, F. Kappe, N. Scherbakov, and P. Srinivasan. Structured browsing of hypermedia databases. Lecture Notes in Computer Science, 733:51–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gotze:1993:OOD

[1880] Rainer Gotze, Helmut Eirund, and Roland Claaben. Object-oriented dialog control for multimedia user interfaces. Lecture Notes in Computer Science, 733:63–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strauss:1993:CHF

[1881] Friedrich Strauss. Contextsensitive help-facilities in GUIs through situations. Lecture Notes in Computer Science, 733:79–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yetim:1993:UAH

[1882] Fahri Yetim. User-adapted hypertext explanations. Lecture Notes in Computer Science, 733:91–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fach:1993:CDA

[1883] Peter W. Fach, Maria Bannert, and Klaus Kunkel. From conflict to dialogue: On attack and defence in on-line assistance. Lecture Notes in Computer Science, 733:103–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derder:1993:TDM

[1884] A. Derder and C. Garbay. Toward deep modelling and control of human activity at the microscope. Lecture Notes in Computer Science, 733:117–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauterberg:1993:LPA

[1885] Matthias Rauterberg and Christian Cachin. Locating the primary attention focus of the user. Lecture Notes in Computer Science, 733:129–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Girgensohn:1993:MIE

[1886] Andreas Girgensohn. Modifier: Improving an end-user modifiable system through user studies. Lecture Notes in Computer Science, 733:141–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attiogbe:1993:HIS

[1887] Christian Attiogbe and Jean-Louis Durieux. Handling interaction in software specification. Lecture Notes in Computer Science, 733:155–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lauesen:1993:DDT

[1888] Soren Lauesen and Morten Borup Harning. Dialogue design through modified dataflow and data modelling. Lecture Notes in Computer Science, 733:172–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolli:1993:DFR

[1889] Raghu Kolli and Jim Hennessey. Deriving the functional requirements for a concept sketching device. Lecture Notes in Computer Science, 733:184– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adelson:1993:TBN

[1890] Beth Adelson. Theory-based negotiation frameworks for supporting group work. Lecture Notes in Computer Science, 733:199–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berlage:1993:UMU

[1891] Thomas Berlage and Andreas Genau. From undo to multi-user applications. Lecture Notes in Computer Science, 733:213–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edlich:1993:SCS

[1892] Stefan Edlich. Software cooperation with the share-kit: Influences of semantic levels on the working efficiency. Lecture Notes in Computer Science, 733:225–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizunashi:1993:MRC

[1893] Jun Mizunashi, Kumiko Nakagawa, Kazuhito Kojima, and Yutaka Matsushita. Map and room: A concept of spatial user interface for accessing network services. Lecture Notes in Computer Science, 733:235–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quinn:1993:TAE

[1894] Clark N. Quinn and Richard Preston. TestIt: An automated evaluation tool for human-computer interaction projects. Lecture Notes in Computer Science, 733:249–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prumper:1993:SEB

[1895] Jochen Prumper. Software-evaluation based upon ISO 9241 Part 10. Lecture Notes in Computer Science, 733: 255–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordelle:1993:DQL

[1896] D. Cordelle and D. Bronisz. Diagnosing quality of life in working environments. Lecture Notes in Computer Science, 733:266–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lauesen:1993:DIB

[1897] Soren Lauesen, Morten Borup Harning, and Hans Frederik Boving. Dialogue independence based on a structured UIMS interface. Lecture Notes in Computer Science, 733:279–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reiterer:1993:HFB

[1898] Harald Reiterer. A human factors based user interface design. Lecture Notes in Computer Science, 733:291– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koivunnen:1993:AIU

[1899] Marja-Riitta Koivunnen, Ora Lassila, Juha Ahvo, Minna Rankinen, Sirpa Riihiaho, and Bodil Riihinen. ActorStudio: An interactice user interface editor. Lecture Notes in Computer Science, 733:303–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franzke:1993:NTW

[1900] Marita Franzke and John Rieman. Natural training wheels: Learning and transfer between two versions of a computer application. Lecture Notes in Computer Science, 733:317–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Musil:1993:BBH

[1901] Sabine Musil. Beyond bars and hour glasses: Designing performance and progress indicators. Lecture Notes in Computer Science, 733:329–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pohl:1993:PUA

[1902] Margit Pohl and Peter Purgathofer.

The problems of untrained authors creating hypertext documents. Lecture Notes in Computer Science, 733:338– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sonnentag:1993:WSA

[1903] Sabine Sonnentag. What is so special about exceptional software professionals? Lecture Notes in Computer Science, 733:353–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paetau:1993:SAR

[1904] Michael Paetau. System-adaption and reality’s own dynamism. Lecture Notes in Computer Science, 733:364– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heinecke:1993:SER

[1905] Andreas M. Heinecke. Software ergonomics for real-time systems. Lecture Notes in Computer Science, 733: 379–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Canfield:1993:IDC

[1906] Kip Canfield and Kerry Petrucci. Interface design for clinical information systems: An ecological interface design approach. Lecture Notes in Computer Science, 733:391–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Love:1993:EID

[1907] Stephen Love, John C. Foster, and Mervin A. Jack. On the effects of individual differences with reference to spoken language dialogue systems. Lecture Notes in Computer Science, 733: 405–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Al-Khonaizi:1993:UNA

[1908] M. Al-Khonaizi, M. Al-A’ali, and A. Al-Zobaidie. Understanding natural Arabic text. Lecture Notes in Computer Science, 733:407–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schleich:1993:SBE

[1909] R. Schleich. Structured browsers enhancing the interface concept of graphical editors. Lecture Notes in Computer Science, 733:409–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grunfelder:1993:RHT

[1910] Ruth Grunfelder. “RelieF” — Herstellung von tastbaren Bildvorlagen fur blinde oder sehbehinderte Personen. Lecture Notes in Computer Science, 733:411–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhler:1993:AUI

[1911] Christian Buhler, Helmut Heck, and Rainer Wallbruch. Adaptable user interfaces for people with special needs. Lecture Notes in Computer Science, 733:413–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seiler:1993:VAB

[1912] Franz P. Seiler, J. Haider, P. Mayer, A. Zagler, and W. Zagler. VIP — arbeitsplatz fur blinde und sehbehinderte studenten. Lecture Notes in Computer Science, 733:415–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

North:1993:IEV

[1913] Max M. North and Sarah M. North.

An information exploration and visualization approach for direct manipulation of databases. Lecture Notes in Computer Science, 733:417–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pemberton:1993:OBD

[1914] Lyn Pemberton. Offices, balconies, doors and corridors: An experimental interface metaphor for integrating collaborative design styles. Lecture Notes in Computer Science, 733:419– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monger:1993:IDI

[1915] David Monger and Graham Leedham. An interactive document image description for OCR of handwritten forms. Lecture Notes in Computer Science, 733:423–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bumiller:1993:EMA

[1916] Johannes Bumiller and Stefanie Rather. Electronic meeting assistance. Lecture Notes in Computer Science, 733:425–??, 1993. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Soroko:1993:HCM

[1917] Edward M. Soroko and Sergey V. Kirpich. Harmonization criterion for the man-machine interaction. Lecture Notes in Computer Science, 733:427– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1993:DVR

[1918] Peter Thomas and Keith Goss. 3dimensional visual representations for graphical user interfaces: The art of user involvement. Lecture Notes in Computer Science, 733:429–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodart:1993:GCI

[1919] François Bodart, Monique Noirhomme-Fraiture, and Jean Vanderdonckt. Guidelines for choosing interaction objects. Lecture Notes in Computer Science, 733:431–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofkirchner:1993:CMD

[1920] Wolfgang Hofkirchner, Peter Purgathofer, Ralf Pichler, Anton Bruno Trapp, Karl Berger, and Walter Nierer. Eine computergestutzte multimediale Dokumentation uber den Golfkrieg. Lecture Notes in Computer Science, 733:433– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grimstad:1993:GWC

[1921] Terje Grimstad and Riitta Hellman.

The global window as a communication mediator in public administration. Lecture Notes in Computer Science, 733:435–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishimoto:1993:SSH

[1922] Kazushi Nishimoto, Kenji Mochizuki, Tsutomu Miyasato, and Fumio Kishino. A supporting system for human creativity: Computer aided divergent thinking process by provision of associative pieces of information. Lecture Notes in Computer Science, 733:437– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pauwels:1993:ICE

[1923] H. W. J. Borst Pauwels, J. Sousa Pinlo, B. Sousa Santos, and J. A. Martins. An integrated courseware editor based on OLE technology. Lecture Notes in Computer Science, 733:439–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petrovic:1993:EFU

[1924] Otto Petrovic. Empirical findings on using information technology in the creativity stage of problem solving. Lecture Notes in Computer Science, 733:441–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riedling:1993:CLA

[1925] Eveline Riedling. Computers in local administrative offices (Province of Salzburg) — aspects of stress, strain and training. Lecture Notes in Computer Science, 733:443–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Purgathofer:1993:CGS

[1926] Peter Purgathofer and Peter Ruppel. Configurable guides for a seasonal database. Lecture Notes in Computer Science, 733:445–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gill:1993:ODP

[1927] Simon H. Gill and David K. Wright. Offering different perspectives in a learning environment. Lecture Notes in Computer Science, 733:447–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robic:1993:HPC

[1928] B. Robic and J. Silc. High-performance computing on a honeycomb architecture. Lecture Notes in Computer Science, 734:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schepers:1993:RLI

[1929] J. Schepers. Refined local instruction scheduling considering pipeline interlocks. Lecture Notes in Computer Science, 734:14–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoover:1993:MMD

[1930] W. G. Hoover, C. G. Hoover, A. J. de Groot, and T. G. Pierce. Microscopic and macroscopic dynamics. Lecture Notes in Computer Science, 734:26–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neto:1993:FRR

[1931] A. G. Neto. Further results of the relaxed timing model for distributed simulation. Lecture Notes in Computer Science, 734:45–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elgindy:1993:PCP

[1932] H. Elgindy. Pipelining computations on processor arrays with reconfigurable bus systems. Lecture Notes in Computer Science, 734:56–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucka:1993:EAC

[1933] M. Lucka. An effective algorithm for computation of two-dimensional Fourier transform for N × M matrices. Lecture Notes in Computer Science, 734:64–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Limongelli:1993:RNA

[1934] C. Limongelli and H. W. Loidl. Rational number arithmetic by parallel P-adic algorithms. Lecture Notes in Computer Science, 734:72–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stifter:1993:SNS

[1935] S. Stifter. Shortest non-synchronized motions — parallel versions for shared memory CREW models. Lecture Notes in Computer Science, 734:87–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spalt:1993:PAI

[1936] A. Spalt. A pipeline algorithm for interactive volume visualization. Lecture Notes in Computer Science, 734:105– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehrotra:1993:DPP

[1937] P. Mehrotra. Data parallel programming: The promises and limitations of High Performance Fortran. Lecture Notes in Computer Science, 734:114, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Snyder:1993:FPP

[1938] L. Snyder. Foundations of practical parallel programming languages. Lecture Notes in Computer Science, 734: 115–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasselbring:1993:PPA

[1939] W. Hasselbring. Prototyping parallel algorithms with ProSet-Linda. Lecture Notes in Computer Science, 734:135– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalogeropulos:1993:IAP

[1940] S. Kalogeropulos. Identifying the available parallelism using static analysis. Lecture Notes in Computer Science, 734:151–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kebler:1993:APP

[1941] C. W. Kebler and W. J. Paul. Automatic parallelization by patternmatching. Lecture Notes in Computer Science, 734:166–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lampe:1993:P

[1942] J. Lampe. Parallelization. Lecture Notes in Computer Science, 734:182– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geist:1993:PBN

[1943] G. A. Geist. PVM 3 beyond network computing. Lecture Notes in Computer Science, 734:194–203, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiner:1993:DPK

[1944] W. Schreiner and H. Hong. The design of the PACLIB kernel for parallel algebraic computation. Lecture Notes in Computer Science, 734:204–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fritzson:1993:GPC

[1945] P. Fritzson and N. Andersson. Generating parallel code from equations in the ObjectMath programming environments. Lecture Notes in Computer Science, 734:219–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ries:1993:PPM

[1946] B. Ries, R. Anderson, W. Auld, D. Breazeal, K. Callaghan, E. Richards, and W. Smith. The paragon performance monitoring environment. Lecture Notes in Computer Science, 734: 233–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ershov:1993:TDN

[1947] Yu L. Ershov. Theory of domains and nearby. Lecture Notes in Computer Science, 735:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grundy:1993:PPS

[1948] J. Grundy. Predicative programming — A survey. Lecture Notes in Computer Science, 735:8–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grundy:1993:TVL

[1949] J. Grundy. A three-valued logic for refinement. Lecture Notes in Computer Science, 735:26–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sands:1993:CSC

[1950] D. Sands. A compositional semantics of combining forms for gamma programs. Lecture Notes in Computer Science, 735:43–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Futschek:1993:APL

[1951] G. Futschek. Algebraic properties of loop invariants. Lecture Notes in Computer Science, 735:57–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reif:1993:APF

[1952] W. Reif. An approach to parameterized first-order specifications: Semantics, correctness, parameter passing. Lecture Notes in Computer Science, 735:67–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zamulin:1993:AMI

[1953] A. V. Zamulin. Algebraic modelling of imperative languages with pointers. Lecture Notes in Computer Science, 735:81–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cousot:1993:GCB

[1954] P. Cousot and R. Cousot. Galois connection based abstract interpretations for strictness analysis. Lecture Notes in Computer Science, 735:98–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourdoncle:1993:ECI

[1955] F. Bourdoncle. Efficient chaotic iteration strategies with widenings. Lecture Notes in Computer Science, 735: 128–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Masdupuy:1993:SAI

[1956] F. Masdupuy. Semantic analysis of interval congruences. Lecture Notes in Computer Science, 735:142–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monsuez:1993:PTC

[1957] B. Monsuez. Polymorphic typing for call-by-name semantics. Lecture Notes in Computer Science, 735:156– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1993:LPT

[1958] L. Lu and P. Greenfield. Logic program testing based on abstract interpretation. Lecture Notes in Computer Science, 735:170–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sabelfeld:1993:ASS

[1959] V. Sabelfeld. Analysis of some semantic properties for programs of the applicative language AL. Lecture Notes in Computer Science, 735:181–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jorrand:1993:CUP

[1960] Ph. Jorrand. Communication as unification in process algebras: Operational semantics. Lecture Notes in Computer Science, 735:190–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pepper:1993:FDM

[1961] P. Pepper, J. Exner, and M. Sudholt. Functional development of massively parallel programs. Lecture Notes in Computer Science, 735:217–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Virbitskaite:1993:OSP

[1962] I. B. Virbitskaite. Observing some properties of event structures. Lecture Notes in Computer Science, 735:239– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taitslin:1993:OLL

[1963] M. Taitslin and D. Arkhangelsky. The other linear logic. Lecture Notes in Computer Science, 735:251–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhou:1993:DCO

[1964] Ch Zhou. Duration calculi: An overview. Lecture Notes in Computer Science, 735:256–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bahsoun:1993:UFS

[1965] J. P. Bahsoun, C. Servieres, and C. Seguin. A unique formalism for specifying and designing objects in a parallel environment. Lecture Notes in Computer Science, 735:267–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haeberer:1993:SCR

[1966] A. M. Haeberer, G. A. Baum, and G. Schmidt. On the smooth calculation of relational recursive expression out of first-order non-constructive specifications involving quantifiers. Lecture Notes in Computer Science, 735:281– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pliuskevicius:1993:SRI

[1967] R. Pliuskevicius. Saturation replaces induction for a miniscoped linear temporal logic. Lecture Notes in Computer Science, 735:299–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nickl:1993:FAR

[1968] F. Nickl and M. Wirsing. A formal approach to requirements engineering. Lecture Notes in Computer Science, 735:312–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gannod:1993:TPA

[1969] G. C. Gannod and B. H. C. Cheng. A two-phase approach to reverse engineering using formal methods. Lecture Notes in Computer Science, 735: 335–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levin:1993:APS

[1970] V. Levin. Algebraically provable specification of optimized compilations. Lecture Notes in Computer Science, 735: 349–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lutfiyya:1993:FDE

[1971] H. Lutfiyya, B. McMillin, and A. Su. Formal derivation of an error-detecting distributed data scheduler using changeling. Lecture Notes in Computer Science, 735:363–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nepomniaschy:1993:RCS

[1972] V. Nepomniaschy and N. Shilov. REAL92: A combined specification language for real-time concurrent systems and properties. Lecture Notes in Computer Science, 735:377–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moller:1993:ACG

[1973] B. Moller. Algebraic calculation of graph and sorting algorithms. Lecture Notes in Computer Science, 735:394– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beltiukov:1993:ASP

[1974] A. P. Beltiukov. Automatical synthesis of programs with recursions. Lecture Notes in Computer Science, 735: 414–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boiten:1993:PI

[1975] E. A. Boiten. Parsing in ISBES. Lecture Notes in Computer Science, 735: 423–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grossman:1993:I

[1976] R. L. Grossman, A. Nerode, A. P.

Ravn, and H. Rischel. Introduction. Lecture Notes in Computer Science, 736:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manna:1993:VHS

[1977] Z. Manna and A. Pnueli. Verifying hybrid systems. Lecture Notes in Computer Science, 736:4–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaochen:1993:EDC

[1978] Z. Chaochen, A. P. Ravn, and M. R. Hansen. An extended duration calculus for hybrid real-time systems. Lecture Notes in Computer Science, 736:36–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1993:TRT

[1979] T. A. Henzinger, Z. Manna, and A. Pnueli. Towards refining temporal specifications into hybrid systems. Lecture Notes in Computer Science, 736:60–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamport:1993:HS

[1980] L. Lamport. Hybrid systems in TLA^{+}. Lecture Notes in Computer Science, 736:77–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurki-Suonio:1993:HMF

[1981] R. Kurki-Suonio. Hybrid models with fairness and distributed clocks. Lecture Notes in Computer Science, 736: 103–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooman:1993:CAD

[1982] J. Hooman. A compositional approach to the design of hybrid systems. Lecture Notes in Computer Science, 736:121–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicollin:1993:ADA

[1983] X. Nicollin, A. Olivero, J. Sifakis, and S. Yovine. An approach to the description and analysis of hybrid systems. Lecture Notes in Computer Science, 736:149–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesten:1993:IGC

[1984] Y. Kesten, A. Pnueli, J. Sifakis, and S. Yovine. Integration graphs: A class of decidable hybrid systems. Lecture Notes in Computer Science, 736:179– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1993:HAA

[1985] R. Alur, C. Courcoubetis, T. A. Henzinger, and P-H. Ho. Hybrid automata: An algorithmic approach to the specification and verification of hybrid systems. Lecture Notes in Computer Science, 736:209–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benveniste:1993:HSS

[1986] A. Benveniste, M. Le Borgne, and P. Le Guernic. Hybrid systems: The SIGNAL approach. Lecture Notes in Computer Science, 736:230–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Back:1993:DSF

[1987] A. Back, J. Guckenheimer, and M. Myers. A dynamical simulation facility for hybrid systems. Lecture Notes in Computer Science, 736:255–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lemmon:1993:EII

[1988] M. Lemmon, J. A. Stiver, and P. J. Antsaklis. Event identification and intelligent hybrid control. Lecture Notes in Computer Science, 736:268– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nerode:1993:MAH

[1989] A. Nerode and W. Kohn. Multiple agent hybrid control architecture. Lecture Notes in Computer Science, 736: 297–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nerode:1993:MHS

[1990] A. Nerode and W. Kohn. Models for hybrid systems: Automata, topologies, controllability, observability. Lecture Notes in Computer Science, 736:317– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grossman:1993:SRA

[1991] R. L. Grossman and R. G. Larson. Some remarks about flows in hybrid systems. Lecture Notes in Computer Science, 736:357–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antsaklis:1993:HSM

[1992] P. J. Antsaklis, J. A. Stiver, and M. Lemmon. Hybrid system modeling and autonomous control systems. Lecture Notes in Computer Science, 736:366–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blanke:1993:FAF

[1993] M. Blanke, S. B. Nielsen, and R. B. Jorgensen. Fault accommodation in feedback control systems. Lecture Notes in Computer Science, 736:393–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1993:FSI

[1994] T. Anderson, R. de Lemos, J. S.

Fitzgerald, and A. Saeed. On formal support for industrial-scale requirements analysis. Lecture Notes in Computer Science, 736:426–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engel:1993:FAC

[1995] M. Engel, M. Kubica, J. Madey, D. L.

Parnas, A. P. Ravn, and A. J. Van Schouwen. A formal approach to computer systems requirements documentation. Lecture Notes in Computer Science, 736:452–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calmet:1993:AIS

[1996] J. Calmet and J. A. Campbell. Artificial intelligence and symbolic mathematical computations. Lecture Notes in Computer Science, 737:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Struss:1993:QMP

[1997] P. Struss. Qualitative modeling of physical systems in AI research. Lecture Notes in Computer Science, 737:20–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sellen:1993:TSC

[1998] J. Sellen. On the topological structure of configuration spaces. Lecture Notes in Computer Science, 737:50–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ladkin:1993:SAI

[1999] P. Ladkin and A. Reinefeld. A symbolic approach to interval constraint problems. Lecture Notes in Computer Science, 737:65–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwarzler:1993:AAK

[2000] G. Schwarzler. An algebraic approach to knowledge-based modeling. Lecture Notes in Computer Science, 737:85–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clarkson:1993:IUI

[2001] M. E. Clarkson. Intelligent user interfaces for symbolic and numeric computation. Lecture Notes in Computer Science, 737:96–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butler:1993:PTI

[2002] G. Butler. The progress towards an intelligent assistant — A discussion paper. Lecture Notes in Computer Science, 737:107–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfalzgraf:1993:MMR

[2003] J. Pfalzgraf. On mathematical modeling in robotics. Lecture Notes in Computer Science, 737:116–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monfroy:1993:GBS

[2004] E. Monfroy. Gröbner bases: Strategies and applications. Lecture Notes in Computer Science, 737:133–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:1993:HSS

[2005] H. Hong. Heuristic search strategies for cylindrical algebraic decomposition. Lecture Notes in Computer Science, 737:152–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calmet:1993:UDA

[2006] J. Calmet, K. Homann, and I. A. Tjandra. Unified domains and abstract computational structures. Lecture Notes in Computer Science, 737:166–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ardeleanu:1993:CIT

[2007] E. Ardeleanu. Completion and invariant theory in symbolic computation and artificial intelligence. Lecture Notes in Computer Science, 737:178– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1993:TCP

[2008] A. Weber. A type-coercion problem in computer algebra. Lecture Notes in Computer Science, 737:188–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koepf:1993:ADP

[2009] W. Koepf. Algorithmic development of power series. Lecture Notes in Computer Science, 737:195–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerschberg:1993:CAQ

[2010] L. Kerschberg and A. Waisanen. A cooperative approach to query processing: Integrating historical, structural, and behavioral knowledge sources. Lecture Notes in Computer Science, 737: 214–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cioni:1993:DTS

[2011] G. Cioni, A. Colagrossi, and A. Miola. A desk-top sequent calculus machine. Lecture Notes in Computer Science, 737:224–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaza:1993:GSC

[2012] J. A. Plaza. Gentzen-style characterizations of negation as failure. Lecture Notes in Computer Science, 737:237– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahnle:1993:NTD

[2013] R. Hahnle. A new translation from deduction into integer programming. Lecture Notes in Computer Science, 737: 262–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golumbic:1993:RAT

[2014] M. C. Golumbic. Reasoning about time. Lecture Notes in Computer Science, 737:276–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roque:1993:LQP

[2015] W. L. Roque. Learning qualitative physics reasoning from regime analysis. Lecture Notes in Computer Science, 737:277–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garigliano:1993:QMM

[2016] R. Garigliano and D. J. Nettleton.

Qualitative mathematical modeling of genetic algorithm. Lecture Notes in Computer Science, 737:294–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davies:1993:TME

[2017] Donald W. Davies. The transition from mechanisms to electronic computers, 1940 to 1950. Lecture Notes in Computer Science, 739:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knudsen:1993:CLb

[2018] Lars Ramkilde Knudsen. Cryptanalysis of LOKI. Lecture Notes in Computer Science, 739:22–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1993:IRD

[2019] Lawrence Brown, Matthew Kwan, Josef Pieprzyk, and Jennifer Seberry. Improving resistance to differential cryptanalysis and the redesign of LOKI. Lecture Notes in Computer Science, 739:36–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyano:1993:MEN

[2020] Hiroshi Miyano. A method to estimate the number of ciphertext pairs for differential cryptanalysis. Lecture Notes in Computer Science, 739:51–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1993:CLB

[2021] Kwangjo Kim. Construction of DES-Like S-boxes based on Boolean functions satisfying the SAC. Lecture Notes in Computer Science, 739:59–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1993:DBS

[2022] Jun-Hui Yang, Zong-Duo Dai, and Ken-Cheng Zeng. The data base of selected permutations. Lecture Notes in Computer Science, 739:73–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daemen:1993:FDO

[2023] Joan Daemen, Rene Govaerts, and Joos Vandewalle. A framework for the design of one-way hash functions including cryptanalysis of Damgard’s oneway function based on a cellular automaton. Lecture Notes in Computer Science, 739:82–96, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadeghiyan:1993:HCF

[2024] Babak Sadeghiyan, Zheng Yuliang, and Josef Pieprzyk. How to construct a family of strong one-way permutations. Lecture Notes in Computer Science, 739:97–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ogata:1993:CFF

[2025] Wakaha Ogata and Kaoru Kurosawa.

On claw free families. Lecture Notes in Computer Science, 739:111–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuliang:1993:SIF

[2026] Zheng Yuliang, Thomas Hardjono, and Josef Pieprzyk. Sibling intractable function families and their applications. Lecture Notes in Computer Science, 739:124–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohta:1993:DMS

[2027] Kazuo Ohta and Tatsuaki Okamoto.

A digital multisignature scheme based on the Fiat–Shamir scheme. Lecture Notes in Computer Science, 739:139– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1993:GSS

[2028] Hung-Yu Lin and Lein Harn. A generalized secret sharing scheme with cheater detection. Lecture Notes in Computer Science, 739:149–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laih:1993:GTC

[2029] Chi-Sung Laih and Lein Harn. Generalized threshold cryptosystems. Lecture Notes in Computer Science, 739: 159–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Safavi-Naini:1993:FTA

[2030] Reihaneh Safavi-Naini. Feistel type authentication codes. Lecture Notes in Computer Science, 739:170–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rhee:1993:RAC

[2031] Man Y. Rhee. Research activities on cryptology in korea. Lecture Notes in Computer Science, 739:179–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadeghiyan:1993:NSC

[2032] Babak Sadeghiyan and Josef Pieprzyk. On necessary and sufficient conditions for the construction of super pseudorandom permutations. Lecture Notes in Computer Science, 739:194–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Even:1993:CCS

[2033] Shimon Even and Yishay Mansour. A construction of a cipher from a single pseudorandom permutation. Lecture Notes in Computer Science, 739: 210–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pieprzyk:1993:OPR

[2034] Josef Pieprzyk and Babak Sadeghiyan.

Optimal perfect randomizers. Lecture Notes in Computer Science, 739:225– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwan:1993:GPT

[2035] Matthew Kwan and Josef Pieprzyk. A general purpose technique for locating key scheduling weaknesses in DESlike cryptosystems. Lecture Notes in Computer Science, 739:237–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morita:1993:RSC

[2036] Hikaru Morita, Kazuo Ohta, and Shoji Miyaguchi. Results of switchingclosure-test on FEAL. Lecture Notes in Computer Science, 739:247–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizusawa:1993:ICT

[2037] Jun ichi Mizusawa. IC-Cards and telecommunication services. Lecture Notes in Computer Science, 739:253– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimbo:1993:CSC

[2038] Atsushi Shimbo and Shin ichi Kawamura. Cryptanalysis of several conference key distribution schemes. Lecture Notes in Computer Science, 739:265– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klapper:1993:RIP

[2039] Andrew Klapper and Mark Goresky. Revealing information with partial period correlations. Lecture Notes in Computer Science, 739:277–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meijers:1993:EMV

[2040] Joost Meijers and Johan van Tilburg. Extended majority voting and privatekey algebraic-code encryptions. Lecture Notes in Computer Science, 739:288– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldburg:1993:SAS

[2041] B. Goldburg, E. Dawson, and S. Sridharan. A secure analog speech scrambler using the discrete cosine transform. Lecture Notes in Computer Science, 739:299–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harn:1993:OTP

[2042] Lein Harn and Lin Hung-Yu. An oblivious transfer protocol and its application for the exchange of secrets. Lecture Notes in Computer Science, 739: 312–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saito:1993:MPZ

[2043] Takeshi Saito, Kaoru Kurosawa, and Kouichi Sakurai. 4 move perfect ZKIP of knowledge with no assumption. Lecture Notes in Computer Science, 739:321–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Itoh:1993:CCR

[2044] Toshiya Itoh and Kouichi Sakurai. On the complexity of constant round ZKIP of possession of knowledge. Lecture Notes in Computer Science, 739:331– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fortnow:1993:PTL

[2045] Lance Fortnow and Mario Szegedy. On the power of two-local random reductions. Lecture Notes in Computer Science, 739:346–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feigenbaum:1993:NOP

[2046] Joan Feigenbaum and Rafail Ostrovsky. A note on one-prover, instancehiding zero-knowledge proof systems. Lecture Notes in Computer Science, 739:352–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desmedt:1993:EZK

[2047] Yvo Desmedt and Mike Burmester. An efficient zero-knowledge scheme for the discrete logarithm based on smooth numbers. Lecture Notes in Computer Science, 739:360–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okamoto:1993:EZK

[2048] Tatsuaki Okamoto. An extension of zero-knowledge proofs and its applications. Lecture Notes in Computer Science, 739:368–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Itoh:1993:LID

[2049] Toshiya Itoh, Kouichi Sakurai, and Hiroki Shizuya. Any language in IP has a divertible ZKIP. Lecture Notes in Computer Science, 739:382–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaosheng:1993:MPP

[2050] Shu Chaosheng, Tsutomu Matsumoto, and Hideki Imai. A multi-purpose proof system. Lecture Notes in Computer Science, 739:397–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toussaint:1993:FVP

[2051] Marie-Jeanne Toussaint. Formal verification of probabilistic properties in cryptographic protocols. Lecture Notes in Computer Science, 739:412– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rivest:1993:CML

[2052] Ronald L. Rivest. Cryptography and machine learning. Lecture Notes in Computer Science, 739:427–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandt:1993:SPN

[2053] Jorgen Brandt, Ivan Damgard, and Peter Landrock. Speeding up prime number generation. Lecture Notes in Computer Science, 739:440–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laih:1993:TES

[2054] Chi-Sung Laih, Sung-Ming Yen, and Lein Harn. Two efficient server-aided secret computation protocols based on the addition sequence. Lecture Notes in Computer Science, 739:450–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyaji:1993:OEC

[2055] Atsuko Miyaji. On ordinary elliptic curve cryptosystems. Lecture Notes in Computer Science, 739:460–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joux:1993:CAK

[2056] Antoine Joux and Jacques Stern.

Cryptanalysis of another knapsack cryptosystem. Lecture Notes in Computer Science, 739:470–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daemen:1993:CSH

[2057] Joan Daemen, Antoon Bosselaers, Rene Govaerts, and Joos Vandewalle. Collisions for Schnorr’s hash function FFT-Hash presented at Crypto’91. Lecture Notes in Computer Science, 739:477–480, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rivest:1993:NPD

[2058] Ronald L. Rivest. On NIST’s proposed Digital Signature Standard. Lecture Notes in Computer Science, 739:481– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaneko:1993:KPA

[2059] Toshinobu Kaneko. A known-plaintext attack of FEAL-4 based on the system of linear equations on difference. Lecture Notes in Computer Science, 739: 485–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwan:1993:SAD

[2060] Matthew Kwan. Simultaneous attacks in differential cryptanalysis (getting more pairs per encryption). Lecture Notes in Computer Science, 739: 489–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mjolsnes:1993:PCP

[2061] Stig Fr. Mjolsnes. Privacy, cryptographic pseudonyms, and the state of health. Lecture Notes in Computer Science, 739:493–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daemen:1993:LEM

[2062] Joan Daemen. Limitations of the evenmansour construction. Lecture Notes in Computer Science, 739:495–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bos:1993:PUS

[2063] Jurjen N. E. Bos and David Chaum. Provably unforgeable signatures. Lecture Notes in Computer Science, 740: 1–14, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400001.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/0740/07400001.
pdf.

vanHeijst:1993:NCF

[2064] Eugène van Heijst, Torben Pryds Pedersen, and Birgit Pfitzmann. New constructions of fail-stop signatures and lower bounds (extended abstract). Lecture Notes in Computer Science, 740: 15–30, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400015.
pdf.

Heijst:1993:NCF

[2065] Eugène van Heijst, Torben Pryds Pedersen, and Birgit Pfitzmann. New constructions of fail-stop signatures and lower bounds (extended abstract). Lecture Notes in Computer Science, 740: 15–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400015.
pdf.

Okamoto:1993:PSP

[2066] Tatsuaki Okamoto. Provably secure and practical identification schemes and corresponding signature schemes. Lecture Notes in Computer Science, 740:31–53, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/0740/07400031.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400031.
                                        pdf.

Okamoto:1993:EDS

[2067] Tatsuaki Okamoto, Atsushi Fujioka, and Eiichiro Fujisaki. An efficient digital signature scheme based on an elliptic curve over the ring Z_{n}. Lecture Notes in Computer Science, 740:54– 65, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400054.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400054.
                                        pdf.

Gordon:1993:DDT

[2068] Daniel M. Gordon. Designing and detecting trapdoors for discrete log cryptosystems. Lecture Notes in Computer Science, 740:66– 75, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400066.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400066.
                                        pdf.

Smid:1993:RCN

[2069] Miles E. Smid and Dennis K. Branstad. Response to comments on the NIST proposed digital signature standard (invited). Lecture Notes in Computer Science, 740:76– 88, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0740/07400076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400076.
pdf.

Chaum:1993:WDO

[2070] David Chaum and Torben Pryds Pedersen. Wallet databases with observers. Lecture Notes in Computer Science, 740:89–105, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0740/07400089.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0740/07400089.
pdf.

Hirschfeld:1993:MER

[2071] Rafael Hirschfeld. Making electronic refunds safer. Lecture Notes in Computer Science, 740:106–112, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400106.
pdf.

Micali:1993:FPK

[2072] Silvio Micali. Fair public-key cryptosystems (rough draft). Lecture Notes in Computer Science, 740:113– 138, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400113.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/0740/07400113.
                                        pdf.

Dwork:1993:PPC

[2073] Cynthia Dwork and Moni Naor. Pricing via processing or combatting junk mail. Lecture Notes in Computer Science, 740:139–147, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400139.
                                        pdf.

Blundo:1993:IRS

[2074] C. Blundo, A. De Santis, L. Gargano, and U. Vaccaro. On the information rate of secret sharing schemes. Lecture Notes in Computer Science, 740:148– 167, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400148.
                                        pdf.

Stinson:1993:NGL

[2075] Douglas R. Stinson. New general lower bounds on the information rate of secret sharing schemes. Lecture Notes in Computer Science, 740:168– 182, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400168.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/0740/07400168.
pdf.

Beimel:1993:UIS

[2076] Amos Beimel and Benny Chor. Universally ideal secret sharing schemes (preliminary version). Lecture Notes in Computer Science, 740:183–195, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400183.
pdf.

Naor:1993:PZK

[2077] Moni Naor, Rafail Ostrovsky, Ramarathnam Venkatesan, and Moti Yung. Perfect zero-knowledge arguments for np can be based on general complexity assumptions (extended abstract). Lecture Notes in Computer Science, 740:196–214, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400196.
pdf.

Dwork:1993:LCP

[2078] Cynthia Dwork, Uri Feige, Joe Kilian, Moni Naor, and Muli Safra. Low communication 2-Prover zeroknowledge proofs for NP. Lecture Notes in Computer Science, 740:215– 227, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400215.
                                        pdf.

Goldwasser:1993:ISN

[2079] Shafi Goldwasser and Rafail Ostrovsky. Invariant signatures and noninteractive zero-knowledge proofs are equivalent (extended abstract). Lecture Notes in Computer Science, 740:228– 245, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400228.
                                        pdf.

Sakurai:1993:DBS

[2080] Kouichi Sakurai and Toshiya Itoh. On the discrepancy between serial and parallel of zero-knowledge protocols (extended abstract). Lecture Notes in Computer Science, 740:246– 259, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400246.
                                        pdf.

Sivabalan:1993:DSN

[2081] M. Sivabalan, S. E. Tavares, and L. E. Peppard. On the design of SP networks from an information theoretic point of view. Lecture Notes in Computer Science, 740:260–279, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400260.
pdf.

Carlet:1993:PBF

[2082] Claude Carlet. Partially-bent functions. Lecture Notes in Computer Science, 740:280–291, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0740/07400280.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0740/07400280.pdf.

Zheng:1993:PAA

[2083] Yuliang Zheng and Jennifer Seberry. Practical approaches to attaining security against adaptively chosen ciphertext attacks (extended abstract). Lecture Notes in Computer Science, 740:292–304, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400292.
pdf.

Baritaud:1993:SPK

[2084] Thierry Baritaud, Mireille Campana, Pascal Chauvaud, and Henri Gilbert. On the security of the permuted kernel identification scheme. Lecture Notes in Computer Science, 740:305– 311, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400305.
                                        pdf.

Gordon:1993:MPC

[2085] Daniel M. Gordon and Kevin S. Mc-Curley. Massively parallel computation of discrete logarithms. Lecture Notes in Computer Science, 740:312– 323, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400312.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400312.
                                        pdf.

Peralta:1993:QSD

[2086] René Peralta. A quadratic sieve on the n-dimensional cube. Lecture Notes in Computer Science, 740:324– 332, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400324.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400324.
                                        pdf.

Meier:1993:EMC

[2087] Willi Meier and Othmar Staffelbach. Efficient multiplication on certain nonsupersingular elliptic curves. Lecture Notes in Computer Science, 740:333– 344, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0740/07400333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400333.
pdf.

Koyama:1993:SEC

[2088] Kenji Koyama and Yukio Tsuruoka. Speeding up elliptic cryptosystems by using a signed binary window method. Lecture Notes in Computer Science, 740:345–357, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400345.
pdf.

Brandt:1993:GPP

[2089] Jørgen Brandt and Ivan Bjerre Damgård. On generation of probable primes by incremental search. Lecture Notes in Computer Science, 740:358– 370, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400358.
pdf.

Fellows:1993:KKI

[2090] Michael Fellows and Neal Koblitz.

Kid Krypto (invited). Lecture Notes in Computer Science, 740:371–389, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/0740/07400371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400371.
                                        pdf.

Bellare:1993:DPK

[2091] Mihir Bellare and Oded Goldreich. On defining proofs of knowledge. Lecture Notes in Computer Science, 740:390– 420, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400390.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400390.
                                        pdf.

Herzberg:1993:PRC

[2092] Amir Herzberg and Michael Luby. Public randomness in cryptography. Lecture Notes in Computer Science, 740: 421–432, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0740/07400421.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400421.
                                        pdf.

Russell:1993:NSC

[2093] Alexander Y. Russell. Necessary and sufficient conditions for collision-free hashing. Lecture Notes in Computer Science, 740:433–441, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400433.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/0740/07400433.
pdf.

Bellare:1993:CCT

[2094] Mihir Bellare and Moti Yung. Certifying cryptographic tools: The case of trapdoor permutations. Lecture Notes in Computer Science, 740:442– 460, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400442.
pdf.

Maurer:1993:PSK

[2095] Ueli M. Maurer. Protocols for secret key agreement by public discussion based on common information. Lecture Notes in Computer Science, 740:461– 470, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400461.
pdf.

Blundo:1993:PSK

[2096] Carlo Blundo, Alfredo De Santis, Amir Herzberg, Shay Kutten, Ugo Vaccaro, and Moti Yung. Perfectly-secure key distribution for dynamic conferences. Lecture Notes in Computer Science, 740:471–486, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400471.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400471.
                                        pdf.

Biham:1993:DCF

[2097] Eli Biham and Adi Shamir. Differential cryptanalysis of the full 16-round DES. Lecture Notes in Computer Science, 740:487–496, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400487.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400487.
                                        pdf.

Knudsen:1993:ICS

[2098] L. R. Knudsen. Iterative characteristics of DES and S02 DES. Lecture Notes in Computer Science, 740:497–511, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knudsen:1993:IC

[2099] Lars R. Knudsen. Iterative characteristics of DES and s^{2}-DES. Lecture Notes in Computer Science, 740:497– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400497.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400497.
                                        pdf.

Campbell:1993:G

[2100] Keith W. Campbell and Michael J. Wiener. DES is not a group. Lecture Notes in Computer Science, 740:512– 520, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400512.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400512.
pdf.

Eberle:1993:HSI

[2101] Hans Eberle. A high-speed DES implementation for network applications. Lecture Notes in Computer Science, 740:521–539, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0740/07400521.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0740/07400521.
pdf.

Blakley:1993:TSD

[2102] Bob Blakley, G. R. Blakley, Agnes Hui Chan, and James L. Massey. Threshold schemes with disenrollment. Lecture Notes in Computer Science, 740:540– 548, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400540.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400540.
pdf.

Frankel:1993:NEH

[2103] Yair Frankel, Yvo Desmedt, and Mike V. D. Burmester. Non-existence of homomorphic general sharing schemes for some key spaces. Lecture Notes in Computer Science, 740:549–557, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400549.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400549.
                                        pdf.

Harn:1993:SGS

[2104] Lein Harn and Hung-Yu Lin. An lspan generalized secret sharing scheme. Lecture Notes in Computer Science, 740:558–565, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0740/07400558.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400558.
                                        pdf.

Nyberg:1993:PSA

[2105] Kaisa Nyberg and Lars Ramkilde Knudsen. Provable security against differential cryptanalysis. Lecture Notes in Computer Science, 740:566– 574, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0740/07400566.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0740/07400566.
                                        pdf.

Wayner:1993:CAS

[2106] Peter C. Wayner. Content-addressable search engines and DES-like systems. Lecture Notes in Computer Science, 740:575–586, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0740/07400575.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/0740/07400575.
pdf.

Vaudenay:1993:FHI

[2107] Serge Vaudenay. FFT-Hash-II is not yet collision-free. Lecture Notes in Computer Science, 740:587–593, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0740/07400587.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0740/07400587.
pdf.

DeSchutter:1993:TFA

[2108] B. De Schutter. Trends in the fight against computer-related delinquency. Lecture Notes in Computer Science, 741:3–19, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vandewalle:1993:TAT

[2109] J. Vandewalle, R. Govaerts, and B. Preneel. Technical approaches to thwart computer fraud. Lecture Notes in Computer Science, 741:20–32, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSoete:1993:PKC

[2110] M. De Soete. Public key cryptography. Lecture Notes in Computer Science, 741:33–49, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeWaleffe:1993:BLP

[2111] D. De Waleffe and J.-J. Quisquater.

Better login protocols for computer networks. Lecture Notes in Computer Science, 741:50–70, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanTilburg:1993:SKE

[2112] J. Van Tilburg. Secret-key exchange with authentication. Lecture Notes in Computer Science, 741:71–86, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preneel:1993:IAH

[2113] B. Preneel, R. Govaerts, and J. Vandewalle. Information authentication: Hash functions and digital signatures. Lecture Notes in Computer Science, 741:87–131, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Funny:1993:KM

[2114] W. Funny. Key management. Lecture Notes in Computer Science, 741:132– 150, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roberts:1993:ECI

[2115] D. W. Roberts. Evaluation criteria for IT security. Lecture Notes in Computer Science, 741:151–161, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preneel:1993:SCT

[2116] B. Preneel. Standardization of cryptographic techniques. Lecture Notes in Computer Science, 741:162–173, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaum:1993:NCB

[2117] D. Chaum. Numbers can be a better form of cash than paper. Lecture Notes in Computer Science, 741:174– 178, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verschuren:1993:IOS

[2118] J. Verschuren, R. Govaerts, and J. Vandewalle. ISO-OSI security architecture. Lecture Notes in Computer Science, 741:179–192, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vedder:1993:SAM

[2119] K. Vedder. Security aspects of mobile communications. Lecture Notes in Computer Science, 741:193–210, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fumy:1993:LAN

[2120] W. Fumy. (local area) network security. Lecture Notes in Computer Science, 741:211–226, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosselaers:1993:CWP

[2121] A. Bosselaers, R. Govaerts, and J. Vandewalle. Cryptography within phase I of the EEC-RACE programme. Lecture Notes in Computer Science, 741:227– 234, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lennox:1993:ES

[2122] G. Lennox. EDI security. Lecture Notes in Computer Science, 741:235–243, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geoffroy:1993:APF

[2123] M. Geoffroy, R. Bjones, and H. Cnudde. AXYTRANS: Physical funds transport and DES. Lecture Notes in Computer Science, 741:244–256, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeDecker:1993:USK

[2124] B. De Decker. Unix security and Kerberos. Lecture Notes in Computer Science, 741:257–274, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sargeant:1993:UFO

[2125] John Sargeant. Uniting functional and object-oriented programming. Lecture Notes in Computer Science, 742:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiczales:1993:TCM

[2126] Gregor Kiczales. Traces (A cut at the “make isn’t generic” problem). Lecture Notes in Computer Science, 742:27–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pintado:1993:GSS

[2127] Xavier Pintado. Gluons: a support for software component cooperation. Lecture Notes in Computer Science, 742:43–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamazaki:1993:TOO

[2128] Kenichi Yamazaki, Yoshiji Amagai, Masaharu Yoshida, and Ikuo Takeuchi.

TAO: An object-orientation kernel. Lecture Notes in Computer Science, 742:61–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gondow:1993:CMC

[2129] Katsuhiko Gondow, Takashi Imaizumi, Yoichi Shinoda, and Takuya Katayama. Change management and consistency maintenance in software development environments using objectoriented attribute grammars. Lecture Notes in Computer Science, 742:77–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitsui:1993:DIE

[2130] Kin’ichi Mitsui, Hiroaki Nakamura, Theodore C. Law, and Shahram Javey. Design of an integrated and extensible C++ programming environment. Lecture Notes in Computer Science, 742: 95–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okamura:1993:MDA

[2131] Hideaki Okamura, Yutaka Ishikawa, and Mario Tokoro. Metalevel decomposition in AL-1/D. Lecture Notes in Computer Science, 742:110–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mulet:1993:DRK

[2132] Philippe Mulet and Pierre Cointe. Definition of a reflective kernel for a prototype-based language. Lecture Notes in Computer Science, 742:128– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokote:1993:KSO

[2133] Yasuhiko Yokote. Kernel structuring of object-oriented operating systems: The Apertos approach. Lecture Notes in Computer Science, 742:145–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bancilhon:1993:ODS

[2134] François Bancilhon. Object database systems: Functional architecture. Lecture Notes in Computer Science, 742: 163–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergstein:1993:MBC

[2135] Paul L. Bergstein and Walter L. Hursch. Maintaining behavioral consistency during schema evolution. Lecture Notes in Computer Science, 742: 176–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1993:OCA

[2136] Ling Liu. An object-centered approach for manipulating hierarchically complex objects. Lecture Notes in Computer Science, 742:194–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Byeon:1993:TUV

[2137] Kwang June Byeon and Dennis McLeod. Towards the unification of views and versions for object databases. Lecture Notes in Computer Science, 742:220–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qiming:1993:AVO

[2138] Chen Qiming and Shan Ming-Chein.

Abstract view objects for multiple OODB integration. Lecture Notes in Computer Science, 742:237–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hai:1993:OOQ

[2139] Jin Chang Hai and Hai Jin Chang. An object-oriented query model supporting views. Lecture Notes in Computer Science, 742:251–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnson:1993:RA

[2140] Ralph E. Johnson and William F.

Opdyke. Refactoring and aggregation. Lecture Notes in Computer Science, 742:264–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kristensen:1993:TAA

[2141] Bent Bruun Kristensen. Transverse activities: Abstractions in objectoriented programming. Lecture Notes in Computer Science, 742:279–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malhotra:1993:DES

[2142] Jawahar Malhotra. Dynamic extensibility in a statically-compiled objectoriented language. Lecture Notes in Computer Science, 742:297–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkinson:1993:MCP

[2143] Malcolm P. Atkinson, D. I. K. Sjoberg, and R. Morrison. Managing change in persistent object systems. Lecture Notes in Computer Science, 742:315– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gemis:1993:OOP

[2144] Marc Gemis and Jan Paredaens. An object-oriented pattern matching language. Lecture Notes in Computer Science, 742:339–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hui:1993:CCB

[2145] Siu Cheung Hui, Angela Goh, and Jose Kolencheril Raphel. CLOG: A class-based logic language for objectoriented databases. Lecture Notes in Computer Science, 742:356–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaplan:1993:NMO

[2146] Alan Kaplan and Jack C. Wileden. Name management and object technology for advanced software. Lecture Notes in Computer Science, 742:371– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBaelen:1993:COO

[2147] Stefan Van Baelen, Johan Lewi, Eric Steegmans, and Bart Swennen. Constraints in object-oriented analysis. Lecture Notes in Computer Science, 742:393–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1993:ITA

[2148] Sai Peck Lee and Collette Rolland.

Integration of the tool (AWB) supporting the O^{∗} method in the PCTE-Based software engineering environment. Lecture Notes in Computer Science, 742:408–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lieberherr:1993:MDC

[2149] Karl J. Lieberherr and Cun Xiao. Minimizing dependency on class structures with adaptive programs. Lecture Notes in Computer Science, 742:424– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wakita:1993:FCM

[2150] Ken Wakita. First class messages as first class continuations. Lecture Notes in Computer Science, 742:442– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vasconcelos:1993:TSC

[2151] Vasco T. Vasconcelos and Mario Tokoro. A typing system for a calculus of objects. Lecture Notes in Computer Science, 742:460–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harada:1993:TMB

[2152] Yasunori Harada. A type mechanism based on restricted CCS for distributed active objects. Lecture Notes in Computer Science, 742:475–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Notkin:1993:AII

[2153] David Notkin, David Garlan, William G. Griswold, and Kevin Sullivan. Adding implicit invocation to languages: Three approaches. Lecture Notes in Computer Science, 742:489–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bandinelli:1993:REE

[2154] Sergio Bandinelli, Luciano Baresi, Alfonso Fuggetta, and Luigi Lavazza. Requirements and early experiences in the implementation of the SPADE repository using object-oriented technology. Lecture Notes in Computer Science, 742:511–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laorakpong:1993:OOF

[2155] Amarit Laorakpong and Motoshi Saeki. Object-oriented formal specification development using VDM. Lecture Notes in Computer Science, 742:529– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otsuki:1993:DLI

[2156] S. Otsuki. Discovery learning in intelligent tutoring systems. Lecture Notes in Computer Science, 743:3–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiehagen:1993:IIA

[2157] R. Wiehagen. From inductive inference to algorithmic learning theory. Lecture Notes in Computer Science, 743:13–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konagaya:1993:SAG

[2158] A. Konagaya. A stochastic approach to genetic information processing. Lecture Notes in Computer Science, 743:25–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokomori:1993:LSL

[2159] T. Yokomori. On learning systolic languages. Lecture Notes in Computer Science, 743:41–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balcazar:1993:NQC

[2160] J. L. Balcazar, J. Diaz, R. Gavalda, and O. Watanabe. A note on the query complexity of learning DFA. Lecture Notes in Computer Science, 743:53–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1993:PTM

[2161] K. Ito and A. Yamamoto. Polynomialtime MAT learning of multilinear logic programs. Lecture Notes in Computer Science, 743:63–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurita:1993:IWL

[2162] T. Kurita. Iterative weighted least squares algorithms for neural networks classifiers. Lecture Notes in Computer Science, 743:77–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niijima:1993:DAA

[2163] K. Niijima. Domains of attraction in autoassociative memory networks for character pattern recognition. Lecture Notes in Computer Science, 743:87–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akaho:1993:RLN

[2164] S. Akaho. Regularization learning of neural networks for generalization. Lecture Notes in Computer Science, 743: 99–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamimura:1993:CLE

[2165] R. Kamimura. Competitive learning by entropy minimization. Lecture Notes in Computer Science, 743:111–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukouchi:1993:IIB

[2166] Y. Mukouchi. Inductive inference with bounded mind changes. Lecture Notes in Computer Science, 743:125– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishizaka:1993:EII

[2167] H. Ishizaka, H. Arimura, and T. Shinohara. Efficient inductive inference of primitive prologs from positive data. Lecture Notes in Computer Science, 743:135–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1993:MLL

[2168] S. Kapur. Monotonic language learning. Lecture Notes in Computer Science, 743:147–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1993:PVL

[2169] S. Jain and A. Sharma. Prudence in vacillatory language identification. Lecture Notes in Computer Science, 743: 159–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ueda:1993:IHP

[2170] K. Ueda and S. Nagano. Implementation of heuristic problem solving process including analogical reasoning. Lecture Notes in Computer Science, 743:171–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okubo:1993:PAB

[2171] Y. Okubo and M. Haraguchi. Planning with abstraction based on partial predicate mappings. Lecture Notes in Computer Science, 743:183–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakai:1993:LTM

[2172] Y. Sakai and A. Maruoka. Learning k-term monotone Boolean formulae. Lecture Notes in Computer Science, 743:197–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takeuchi:1993:SIS

[2173] J. Takeuchi. Some improved sample complexity bounds in the probabilistic PAC learning model. Lecture Notes in Computer Science, 743:208–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsuoka:1993:ABP

[2174] M. Matsuoka. An application of Bernstein polynomials in PAC model. Lecture Notes in Computer Science, 743: 220–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akutsu:1993:PLF

[2175] T. Akutsu and A. Takasu. On PAC learnability of functional dependencies. Lecture Notes in Computer Science, 743:229–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mamitsuka:1993:PSS

[2176] H. Mamitsuka and K. Yamanishi. Protein secondary structure prediction based on stochastic-rule learning. Lecture Notes in Computer Science, 743: 240–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kakihara:1993:NPL

[2177] K. Kakihara and H. Imai. Notes on the PAC learning of geometric concepts with additional information. Lecture Notes in Computer Science, 743:252– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laird:1993:IUP

[2178] P. Laird. Identifying and using patterns in sequential data. Lecture Notes in Computer Science, 744:1–18, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyano:1993:LTT

[2179] S. Miyano. Learning theory toward genome informatics. Lecture Notes in Computer Science, 744:19–36, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muggleton:1993:OLL

[2180] S. Muggleton. Optimal layered learning: A PAC approach to incremental sampling. Lecture Notes in Computer Science, 744:37–44, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arima:1993:REL

[2181] J. Arima and H. Sawamura. Reformulation of explanation by linear logic — toward logic for explanation. Lecture Notes in Computer Science, 744:45–58, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barzdins:1993:TEIb

[2182] J. Barzdins, G. Barzdins, K. Apsitis, and U. Sarkans. Towards efficient inductive synthesis of expressions from input/output examples. Lecture Notes in Computer Science, 744:59–72, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagiya:1993:TLC

[2183] M. Hagiya. A typed lambda-calculus for proving-by-example and bottomup generalization procedure. Lecture Notes in Computer Science, 744:73–86, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jantke:1993:CBR

[2184] K. P. Jantke and S. Lange. Casebased representation and learning of pattern languages. Lecture Notes in Computer Science, 744:87–100, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1993:IR

[2185] T. Sato and S. Akiba. Inductive resolution. Lecture Notes in Computer Science, 744:101–110, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamamoto:1993:GUB

[2186] A. Yamamoto. Generalized unification as background knowledge in learning logic programs. Lecture Notes in Computer Science, 744:111–122, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukouchi:1993:IIM

[2187] Y. Mukouchi and S. Arikawa. Inductive inference machines that can refute hypothesis spaces. Lecture Notes in Computer Science, 744:123–136, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1993:DBM

[2188] R. Freivalds and C. H. Smith. On the duality between mechanistic learners and what it is they learn. Lecture Notes in Computer Science, 744:137– 149, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1993:ATL

[2189] S. Jain and A. Sharma. On aggregating teams of learning machines. Lecture Notes in Computer Science, 744:150– 163, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viksna:1993:LGQ

[2190] J. Viksna. Learning with growing quality. Lecture Notes in Computer Science, 744:164–172, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daley:1993:URA

[2191] R. Daley and B. Kalyanasundaram.

Use of reduction arguments in determining popperian FIN-Type learning capabilities. Lecture Notes in Computer Science, 744:173–186, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moriyama:1993:PLC

[2192] T. Moriyama and M. Sato. Properties of language classes with finite elasticity. Lecture Notes in Computer Science, 744:187–196, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1993:UCV

[2193] S. Kapur. Uniform characterizations of various kinds of language learning. Lecture Notes in Computer Science, 744:197–208, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knuutila:1993:HIC

[2194] T. Knuutila. How to invent characterizable inference methods for regular languages. Lecture Notes in Computer Science, 744:209–222, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuellar:1993:NDA

[2195] J. R. Cuellar and H. U. Simon. Neural discriminant analysis. Lecture Notes in Computer Science, 744:223–236, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwayama:1993:NAA

[2196] M. Iwayama, N. Indurkhya, and H. Motoda. A new algorithm for automatic configuration of hidden Markov models. Lecture Notes in Computer Science, 744:237–250, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakurai:1993:VDD

[2197] A. Sakurai. On the VC-Dimension of depth four threshold circuits and the complexity of Boolean-valued functions. Lecture Notes in Computer Science, 744:251–264, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takimoto:1993:SCC

[2198] E. Takimoto and A. Maruoka. On the sample complexity of consistent learning with one-sided error. Lecture Notes in Computer Science, 744:265–278, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shinohara:1993:CCV

[2199] A. Shinohara. Complexity of computing Vapnik–Chervonenkis dimension. Lecture Notes in Computer Science, 744:279–287, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1993:AKL

[2200] S. Hasegawa, H. Imai, and M. Ishiguro. xxx-approximations of kappalabel spaces. Lecture Notes in Computer Science, 744:288–299, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1993:ELL

[2201] A. Nakamura and N. Abe. Exact learning of linear combinations of monotone terms from function value queries. Lecture Notes in Computer Science, 744:300–313, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siromoney:1993:TSD

[2202] R. Siromoney, D. G. Thomas, K. G.

Subramanian, and V. R. Dare. Thue systems and DNA — A learning algorithm for a subclass. Lecture Notes in Computer Science, 744:314–327, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishigami:1993:VDF

[2203] Y. Ishigami and S. Tani. The VC-Dimensions of finite automata with n states. Lecture Notes in Computer Science, 744:328–341, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshida:1993:ULM

[2204] K. Yoshida, H. Motoda, and N. Indurkhya. Unifying learning methods by colored digraphs. Lecture Notes in Computer Science, 744:342–355, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suwa:1993:PCV

[2205] M. Suwa and H. Motoda. A perceptual criterion for visually controlling learning. Lecture Notes in Computer Science, 744:356–369, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1993:LSU

[2206] S. Kobayashi. Learning strategies using decision lists. Lecture Notes in Computer Science, 744:370–383, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1993:DBI

[2207] N. Zhong and S. Ohsuga. A decomposition based induction model for discovering concept clusters from databases. Lecture Notes in Computer Science, 744:384–397, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ganascia:1993:ASS

[2208] J.-G. Ganascia. Algebraic structure of some learning systems. Lecture Notes in Computer Science, 744:398–409, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1993:IPR

[2209] S. Tsumoto and H. Tanaka. Induction of probabilistic rules based on rough set theory. Lecture Notes in Computer Science, 744:410–423, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luccio:1993:GPC

[2210] R. Luccio. Gestalt problems in cognitive psychology: Field theory, invariance, and auto-organisation. Lecture Notes in Computer Science, 745:2–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parisi:1993:NNL

[2211] D. Parisi and S. Nolfi. Neural network learning in an ecological and evolutionary context. Lecture Notes in Computer Science, 745:20–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiGesu:1993:TPR

[2212] V. Di Gesu. Trends in pattern recognition. Lecture Notes in Computer Science, 745:41–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biancardi:1993:CVS

[2213] A. Biancardi, V. Cantoni, and L. Lombardi. Computer vision systems: Functionality and structure integration. Lecture Notes in Computer Science, 745:70–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cosi:1993:UAM

[2214] P. Cosi. On the use of auditory models in speech technology. Lecture Notes in Computer Science, 745:85–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frasconi:1993:ASR

[2215] P. Frasconi, M. Gori, and G. Soda. Automatic speech recognition with neural networks: beyond nonparametric models. Lecture Notes in Computer Science, 745:104–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chella:1993:NSS

[2216] A. Chella, S. Gaglio, and R. Pirrone. New shape from shading methods. Lecture Notes in Computer Science, 745:122–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voltolina:1993:ODE

[2217] M. Voltolina and C. Umilta. Orientationdependent effects in neural networks. Lecture Notes in Computer Science, 745:137–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santini:1993:IPD

[2218] S. Santini. Identification of pattern dimensionality by self-organization. Lecture Notes in Computer Science, 745: 142–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murino:1993:MMP

[2219] V. Murino, G. L. Foresti, and C. S. Regazzoni. A Markovian model for perceptual grouping of different shape primitives. Lecture Notes in Computer Science, 745:161–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giretti:1993:SIR

[2220] A. Giretti, P. Puliti, G. Tascini, and P. Zingaretti. Segmented image reduction. Lecture Notes in Computer Science, 745:181–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corazza:1993:DAT

[2221] A. Corazza, M. Federico, R. Gretter, and G. Lazzari. Design and acquisition of a task-oriented spontaneous speech database. Lecture Notes in Computer Science, 745:196–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tabossi:1993:LA

[2222] P. Tabossi. Language and ambiguity. Lecture Notes in Computer Science, 745:211–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelBimbo:1993:SLI

[2223] A. Del Bimbo, E. Vicario, and D. Zingoni. Spatial logic for image representation and retrieval-by-contents. Lecture Notes in Computer Science, 745: 222–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bragato:1993:CAP

[2224] P. L. Bragato and V. Roberto. Communication among perceiving agents. Lecture Notes in Computer Science, 745: 241–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragoni:1993:PMS

[2225] A. F. Dragoni and P. Puliti. Perception of mental states from communication. Lecture Notes in Computer Science, 745:263–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zoppis:1993:PGSb

[2226] B. Zoppis, O. Boissier, and J. L. Crowley. Perceptual grouping for scene interpretation in an active vision system. Lecture Notes in Computer Science, 745:276–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trucco:1993:UVC

[2227] E. Trucco and A. M. Wallace. Using viewer-centered representations in machine vision. Lecture Notes in Computer Science, 745:307–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milanese:1993:NLI

[2228] R. Milanese, T. Pun, and H. Wechsler. A non-linear integration process for the selection of visual information. Lecture Notes in Computer Science, 745:322–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bollini:1993:DPP

[2229] A. Bollini. Distributed perceptive paradigms: Preliminary statements. Lecture Notes in Computer Science, 745:337–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Floreano:1993:ATM

[2230] D. Floreano. Anticipatory tracking of a moving object. Lecture Notes in Computer Science, 745:344–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Romano:1993:TCD

[2231] M. Romano and N. Ancona. A timeto-crash detector based on area expansions: Example of opto-motor reflex. Lecture Notes in Computer Science, 745:354–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cassinis:1993:ITI

[2232] R. Cassinis and M. Terceros Rojas.

“intelligent” telepresence: Introducing virtual reality in advanced robots. Lecture Notes in Computer Science, 745: 368–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

An:1993:RFR

[2233] Z. An and M. McLeish. RES: A formalism for reasoning with relative-strength defaults. Lecture Notes in Computer Science, 747:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1993:SON

[2234] F. Baader and K. Schlechta. A semantics for open normal defaults via a modified preferential approach. Lecture Notes in Computer Science, 747:9–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benferhat:1993:PLN

[2235] S. Benferhat, D. Dubois, and H. Prade. Possibilistic logic: From nonmonotonicity to logic programming. Lecture Notes in Computer Science, 747:17–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergadano:1993:LMF

[2236] F. Bergadano and V. Cutello. Learning membership functions. Lecture Notes in Computer Science, 747:25–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bittencourt:1993:UPL

[2237] G. Bittencourt, M. Marengoni, and S. Sandri. The use of possibilistic logic PL1 in a customizable tool for the generation of production-rule based systems. Lecture Notes in Computer Science, 747:33–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouckaert:1993:PNC

[2238] R. R. Bouckaert. Probabilistic network construction using the minimum description length principle. Lecture Notes in Computer Science, 747:41–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouckaert:1993:IPM

[2239] R. R. Bouckaert. IDAGs: a perfect map for any distribution. Lecture Notes in Computer Science, 747:49–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

deCampos:1993:LNP

[2240] L. M. de Campos and J. F. Huete.

Learning non probabilistic belief networks. Lecture Notes in Computer Science, 747:57–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cravo:1993:PSD

[2241] M. R. Cravo and J. P. Martins. A practical system for defeasible reasoning and belief revision. Lecture Notes in Computer Science, 747:65–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cubero:1993:IGL

[2242] J. C. Cubero, J. M. Medina, and M. A.

Vila. Influence of granularity level in fuzzy functional dependencies. Lecture Notes in Computer Science, 747:73–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1993:LRA

[2243] S. K. Das and J. Fox. A logic for reasoning about safety in decision support systems. Lecture Notes in Computer Science, 747:79–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elvang-Goransson:1993:AAL

[2244] M. Elvang-Goransson, P. J. Krause, and J. Fox. Acceptability of arguments as ‘logical uncertainty’. Lecture Notes in Computer Science, 747:85–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1993:TMT

[2245] J. Engelfriet and J. Treur. A temporal model theory for default logic. Lecture Notes in Computer Science, 747:91–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fargier:1993:UCS

[2246] H. Fargier and J. Lang. Uncertainty in constraint satisfaction problems: a probabilistic approach. Lecture Notes in Computer Science, 747:97–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FarinasdelCerro:1993:ILC

[2247] L. Farinas del Cerro and A. Herzig. Interference logic = conditional logic + frame axiom. Lecture Notes in Computer Science, 747:105–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fehrer:1993:ULF

[2248] D. Fehrer. A unifying logical framework for reason maintenance. Lecture Notes in Computer Science, 747:113– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fouquere:1993:TLT

[2249] C. Fouquere and J. Vauzeilles. Taxonomic linear theories. Lecture Notes in Computer Science, 747:121–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbay:1993:MIR

[2250] D. Gabbay and A. Hunter. Making inconsistency respectable: Part 2 — metal-level handling of inconsistency. Lecture Notes in Computer Science, 747:129–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbay:1993:RAL

[2251] D. Gabbay and A. Hunter. Restricted access logics for inconsistent information. Lecture Notes in Computer Science, 747:137–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasquet:1993:TIW

[2252] O. Gasquet and A. Herzig. Translating inaccessible worlds logic into bimodal logic. Lecture Notes in Computer Science, 747:145–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gebhardt:1993:NAS

[2253] J. Gebhardt and R. Kruse. A new approach to semantic aspects of possibilistic reasoning. Lecture Notes in Computer Science, 747:151–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilio:1993:PCK

[2254] A. Gilio. Probabilistic consistency of knowledge bases in inference systems. Lecture Notes in Computer Science, 747:160–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guiasu:1993:WIB

[2255] S. Guiasu. Weighting independent bodies of evidence. Lecture Notes in Computer Science, 747:168–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hopkins:1993:DLO

[2256] M. Hopkins. Default logic: Orderings and extensions. Lecture Notes in Computer Science, 747:174–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huete:1993:LCP

[2257] J. F. Huete and L. M. de Campos. Learning causal polytrees. Lecture Notes in Computer Science, 747:180– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohlas:1993:SEA

[2258] J. Kohlas. Symbolic evidence, arguments, supports and valuation networks. Lecture Notes in Computer Science, 747:186–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohler:1993:DOR

[2259] A. Kohler. A dynamic ordering relation for revision. Lecture Notes in Computer Science, 747:199–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriz:1993:EMD

[2260] O. Kriz. On extensions of marginals for decision-making. Lecture Notes in Computer Science, 747:205–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laenens:1993:SNL

[2261] E. Laenens. On the semantics of negations in logic programming. Lecture Notes in Computer Science, 747:213– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larranaga:1993:SLA

[2262] P. Larranaga and Y. Yurramendi.

Structure learning approaches in causal probabilistics networks. Lecture Notes in Computer Science, 747:227–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levy:1993:WED

[2263] F. Levy. Weak extensions for default theories. Lecture Notes in Computer Science, 747:233–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1993:RIF

[2264] W. Liu, A. Bundy, and D. Robertson.

Recovering incidence functions. Lecture Notes in Computer Science, 747: 241–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1993:RBI

[2265] W. Liu, A. Bundy, and D. Robertson.

On the relations between incidence calculus and ATMS. Lecture Notes in Computer Science, 747:249–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mathieu:1993:RMN

[2266] C. Mathieu. A resolution method for a non monotonic multimodal logic. Lecture Notes in Computer Science, 747: 257–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1993:DLB

[2267] J.-J. Ch. Meyer and W. van der Hoek. A default logic based on epistemic states. Lecture Notes in Computer Science, 747:265–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moral:1993:FLC

[2268] S. Moral. A formal language for convex sets of probabilities. Lecture Notes in Computer Science, 747:274–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ngair:1993:LTA

[2269] T.-H. Ngair and G. Provan. A latticetheoretic analysis of ATMS problem solving. Lecture Notes in Computer Science, 747:282–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noormohammadian:1993:ECP

[2270] M. Noormohammadian and U. G. Oppel. Examples of causal probabilistic expert systems. Lecture Notes in Computer Science, 747:290–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papini:1993:MAR

[2271] O. Papini and A. Rauzy. A mixed approach of revision in propositional calculus. Lecture Notes in Computer Science, 747:296–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parsons:1993:IUH

[2272] S. Parsons and A. Saffiotti. Integrating uncertainty handling formalisms in distributed artificial intelligence. Lecture Notes in Computer Science, 747: 304–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaub:1993:VCD

[2273] T. Schaub. Variations of constrained default logic. Lecture Notes in Computer Science, 747:310–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shenoy:1993:ISD

[2274] P. P. Shenoy. Information sets in decision theory. Lecture Notes in Computer Science, 747:318–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shu:1993:PSM

[2275] H. Shu. The preferential semantics of A multi-modal nonmonotonic logic. Lecture Notes in Computer Science, 747:326–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smets:1993:PDB

[2276] P. Smets. Probability of deductibility and belief functions. Lecture Notes in Computer Science, 747:332–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Studeny:1993:FPC

[2277] M. Studeny. Formal properties of conditional independence in different calculi of AI. Lecture Notes in Computer Science, 747:341–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1993:PTC

[2278] Y.-H. Tan. A proof theory for constructive default logic. Lecture Notes in Computer Science, 747:349–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weydert:1993:PID

[2279] E. Weydert. Plausible inference for default conditionals. Lecture Notes in Computer Science, 747:356–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1993:DMB

[2280] N. Wilson. Decision-making with belief functions and pignistic probabilities. Lecture Notes in Computer Science, 747:364–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1993:DLD

[2281] N. Wilson. Default logic and Dempster-Shafer theory. Lecture Notes in Computer Science, 747:372–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witteveen:1993:BRE

[2282] C. Witteveen and W. van der Hoek. Belief revision by expansion. Lecture Notes in Computer Science, 747:380– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:AIc

[2283] Anonymous. Author index. Lecture Notes in Computer Science, 747:389– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halstead:1993:OPU

[2284] Robert H. Halstead and Takayasu Ito.

Overview of papers from the U.S./ Japan workshop on parallel symbolic computing. Lecture Notes in Computer Science, 748:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kale:1993:PPS

[2285] L. V. Kale, B. Ramkumar, V. Saletore, and A. B. Sinha. Prioritization in parallel symbolic computing. Lecture Notes in Computer Science, 748:12–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaneda:1993:PCS

[2286] Yukio Kaneda, Hideo Matsuda, and Shigeo Suzuka. A priority control system for OR-Parallel Prolog and its performance evaluation. Lecture Notes in Computer Science, 748:42–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Osborne:1993:EMS

[2287] Randy B. Osborne. Extending the Multilisp sponsor model to deal with intertask synchronization side effects. Lecture Notes in Computer Science, 748: 54–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohr:1993:DDP

[2288] Eric Mohr. Distillations of dynamic partitioning experience. Lecture Notes in Computer Science, 748:88–??, 1993.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feeley:1993:MPI

[2289] Marc Feeley. A message passing implementation of lazy task creation. Lecture Notes in Computer Science, 748:94–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1993:PCB

[2290] Takayasu Ito and Tomohiro Seino. Pcontinuation based implementation of PaiLisp interpreter. Lecture Notes in Computer Science, 748:108–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hagiya:1993:RHO

[2291] Masami Hagiya. Running higher-order unification in PaiLisp. Lecture Notes in Computer Science, 748:155–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1993:HMD

[2292] Hidehiko Tanaka and Jun ichi Tatemura. HyperDEBU: A multiwindow debugger for parallel logic programs. Lecture Notes in Computer Science, 748:162–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halstead:1993:MTV

[2293] Robert H. Halstead, David A. Kranz, and Patrick G. Sobalvarro. MulTVision: A tool for visualizing parallel program executions. Lecture Notes in Computer Science, 748:183–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kinson:1993:MSE

[2294] Ho Kinson and Paul N. Hilfinger. Managing side effects on shared data. Lecture Notes in Computer Science, 748: 205–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Queinnec:1993:DCD

[2295] Christian Queinnec and David De Roure. Design of a concurrent and distributed language. Lecture Notes in Computer Science, 748:234–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jagannathan:1993:TSD

[2296] Suresh Jagannathan. TS/ scheme: Distributed data structures in lisp. Lecture Notes in Computer Science, 748: 260–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuasa:1993:SET

[2297] Taiichi Yuasa. A SIMD environment TUPLE for parallel list processing. Lecture Notes in Computer Science, 748:268–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padget:1993:AIC

[2298] Julian Padget, Duncan Batey, and Simon Merrall. Architecture independence and coordination. Lecture Notes in Computer Science, 748:287– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yih:1993:PIS

[2299] Benny Yih, Mark Swanson, and Robert Kessler. Persistent immutable shared abstractions. Lecture Notes in Computer Science, 748:300–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tokoro:1993:ART

[2300] Mario Tokoro and Ichiro Satoh. Asynchrony and real-time in distributed systems. Lecture Notes in Computer Science, 748:318–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1993:ACM

[2301] Naoki Kobayashi and Akinori Yonezawa. Asynchronous communication model based on linear logic. Lecture Notes in Computer Science, 748:331– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chikayama:1993:PIS

[2302] Takashi Chikayama and Ryozo Kiyohara. Parallel inference system research in the Japanese FGCS project. Lecture Notes in Computer Science, 748:338–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waltz:1993:MPS

[2303] David L. Waltz. Massively parallel symbolic computing. Lecture Notes in Computer Science, 748:352–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agarwal:1993:SMV

[2304] Anant Agarwal, Jonathan Babb, David Chaiken, Godfrey D’Souza, Kirk Johnson, David Kranz, John Kubiatowicz, Beng-Hong Lim, Gino Maa, and Ken Mackenzie. Sparcle: A multithreaded VLSI processor for parallel computing. Lecture Notes in Computer Science, 748:359–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruggeman:1993:NAD

[2305] Carl Bruggeman and R. Kent Dybvig.

A new architecture design paradigm for parallel computing in scheme. Lecture Notes in Computer Science, 748:362– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Philbin:1993:CPM

[2306] James Philbin. Customizable policy management in the sting operating system. Lecture Notes in Computer Science, 748:380–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taura:1993:EIS

[2307] Kenjiro Taura, Satoshi Matsuoka, and Akinori Yonezawa. An efficient implementation scheme of concurrent objectoriented languages on stock multicomputers. Lecture Notes in Computer Science, 748:402–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ito:1993:MPA

[2308] Takayasu Ito, Rishiyur Nikhil, Julian Padget, Norihisa Suzuki, and Taiichi Yuasa. Massively parallel architectures and symbolic computation. Lecture Notes in Computer Science, 748:408– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halstead:1993:APS

[2309] Robert H. Halstead, Takashi Chikayama, Richard Gabriel, David Waltz, and Akinori Yonezawa. Applications for parallel symbolic computation. Lecture Notes in Computer Science, 748: 417–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:LWP

[2310] Anonymous. List of workshop participants. Lecture Notes in Computer Science, 748:419–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ducasse:1993:PSA

[2311] Mireille Ducasse. A pragmatic survey of automated debugging. Lecture Notes in Computer Science, 749:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahmehri:1993:UCA

[2312] Nahid Shahmehri, Mariam Kamkar, and Peter Fritzson. Usability criteria for automated debugging systems. Lecture Notes in Computer Science, 749:16–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrand:1993:NSE

[2313] Gerard Ferrand. The notions of symptom and error in declarative diagnosis of logic programs. Lecture Notes in Computer Science, 749:40–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1993:DDA

[2314] Luis M. Pereira, Carlos Viegas Damasio, and Jose J. Alferes. Debugging by diagnosing assumptions. Lecture Notes in Computer Science, 749:58–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuh-Jeng:1993:DLP

[2315] Lee Yuh-Jeng and Nachum Dershowitz.

Debugging logic programs using specifications. Lecture Notes in Computer Science, 749:75–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Console:1993:MBD

[2316] Luca Console, Gerhard Friedrich, and Daniele T. Dupre. Model-based diagnosis meets error diagnosis in logic programs. Lecture Notes in Computer Science, 749:85–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fromherz:1993:TDD

[2317] Markus P. J. Fromherz. Towards declarative debugging of concurrent constraint programs. Lecture Notes in Computer Science, 749:88–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alagar:1993:HTD

[2318] Sridhar Alagar and S. Venkatesan.

Hierarchy in testing distributed programs. Lecture Notes in Computer Science, 749:101–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilsson:1993:LAD

[2319] Henrik Nilsson and Peter Fritzson. Lazy algorithmic debugging: Ideas for practical implementation. Lecture Notes in Computer Science, 749:117– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hazan:1993:LEF

[2320] Jonathan E. Hazan and Richard G. Morgan. The location of errors in functional programs. Lecture Notes in Computer Science, 749:135–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirunkitti:1993:GQM

[2321] Visit Hirunkitti and C. J. Hogger. A generalized query minimization for program debugging. Lecture Notes in Computer Science, 749:153–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tobermann:1993:WTB

[2322] Gerhard Tobermann and Clemens Beckstein. What’s in a trace: The box model revisited. Lecture Notes in Computer Science, 749:171–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Binks:1993:DDA

[2323] Dominic Binks. Declarative debugging of abstract data types in godel. Lecture Notes in Computer Science, 749: 188–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ball:1993:SPA

[2324] Thomas Ball and Susan Horwitz. Slicing programs with arbitrary control flow. Lecture Notes in Computer Science, 749:206–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jingde:1993:SCP

[2325] Cheng Jingde. Slicing concurrent programs. Lecture Notes in Computer Science, 749:223–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tip:1993:AGP

[2326] Frank Tip. Animators for generated programming environments. Lecture Notes in Computer Science, 749:241– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia-Martin:1993:VDU

[2327] Julio Garcia-Martin and Juan J.

Moreno-Navarro. Visualization as debugging: Understanding/debugging the Warren Abstract Machine. Lecture Notes in Computer Science, 749: 255–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Westman:1993:GUI

[2328] Rickard Westman and Peter Fritzson.

Graphical user interfaces for algorithmic debugging. Lecture Notes in Computer Science, 749:273–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lutz:1993:TPC

[2329] Rudi Lutz. Towards a plan calculus based intelligent debugging system. Lecture Notes in Computer Science, 749:287–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reiss:1993:TBD

[2330] Steven P. Reiss. Trace-based debugging. Lecture Notes in Computer Science, 749:305–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korel:1993:IFM

[2331] Bogdan Korel. Identifying faulty modifications in software maintenance. Lecture Notes in Computer Science, 749: 315–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goyal:1993:AFS

[2332] Anoop Goyal and Sriram Sankar. The application of formal specifications to software documentation and debugging. Lecture Notes in Computer Science, 749:333–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuchcinski:1993:ADV

[2333] Krzysztof Kuchcinski, Wlodzimierz Drabent, and Jan Maluszynski. Automatic diagnosis of VLSI digital circuits using algorithmic debugging. Lecture Notes in Computer Science, 749:350– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jahne:1993:STI

[2334] Bernd Jahne. Spatio-temporal image processing. theory and scientific applications. Lecture Notes in Computer Science, 751:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Semmel:1993:DCC

[2335] Ralph D. Semmel. Discovering context in a conceptual schema. Lecture Notes in Computer Science, 752:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bach:1993:IIM

[2336] Jeffrey R. Bach, Santanu Paul, and Ramesh Jain. An interactive image management system for face information retrieval. Lecture Notes in Computer Science, 752:16–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alhajj:1993:QOO

[2337] Reda Alhajj and M. Erol Arkun.

Queries in object-oriented database systems. Lecture Notes in Computer Science, 752:36–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1993:CCO

[2338] Herve Martin, Michel Adiba, and Bruno Defude. Consistency checking in object-oriented databases: A behavioral approach. Lecture Notes in Computer Science, 752:53–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Formica:1993:ICR

[2339] Anna Formica and Michele Missikoff. Integrity constraints representation in object-oriented databases. Lecture Notes in Computer Science, 752:69–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pissinou:1993:FTO

[2340] Niki Pissinou and Kia Makki. A framework for temporal object databases. Lecture Notes in Computer Science, 752:86–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keen:1993:IDA

[2341] Debby Keen and Arcot Rajasekar. Inductive dependencies and approximate databases. Lecture Notes in Computer Science, 752:98–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Song:1993:OOD

[2342] Il-Yeol Song and E. K. Park. Objectoriented database design methodologies: A survey. Lecture Notes in Computer Science, 752:115–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodker:1993:HAC

[2343] S. Bodker. Historical analysis and conflicting perspectives — contextualizing HCI. Lecture Notes in Computer Science, 753:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brusilovsky:1993:SFH

[2344] P. Brusilovsky, I. Burmistrov, and V. Kaptelinin. Structuring the field of HCI. an empirical study of experts’ representations. Lecture Notes in Computer Science, 753:11–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kameas:1993:CIS

[2345] A. Kameas, S. Papadimitriou, and G. Pavlides. Coupling interaction specification with functionality description. Lecture Notes in Computer Science, 753:21–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacGregor:1993:EHC

[2346] R. C. MacGregor. An extension to the human-computer interaction paradigm. Lecture Notes in Computer Science, 753:29–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levin:1993:HCH

[2347] M. Sh. Levin. Hierarchical components of human-computer systems. Lecture Notes in Computer Science, 753:37–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayo:1993:SOS

[2348] K. A. Mayo and H. R. Hartson.

Synthesis-oriented situational analysis in user interface design. Lecture Notes in Computer Science, 753:53–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boeve:1993:WEW

[2349] E. Boeve, L. Barfield, and St. Pemberton. WYSIWYG editors: And what now? Lecture Notes in Computer Science, 753:68–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodart:1993:AEH

[2350] F. Bodart, A.-M. Hennebert, J.-M. Leheureux, I. Sacre, and J. Vanderdonckt. Architecture elements for highly-interactive business-oriented applications. Lecture Notes in Computer Science, 753:83–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutcliffe:1993:DMI

[2351] A. G. Sutcliffe and P. Faraday. Designing multimedia interfaces. Lecture Notes in Computer Science, 753:105– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uzilevsky:1993:ISL

[2352] G. Uzilevsky and V. Andreev. Iconic signs and languages in user interface development. Lecture Notes in Computer Science, 753:115–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leung:1993:TMG

[2353] Y. K. Leung and M. D. Apperley. E^{3}: Towards the metrication of graphical presentation techniques for large data sets. Lecture Notes in Computer Science, 753:125–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tolsby:1993:NPL

[2354] H. Tolsby. Navigating in a process landscape. Lecture Notes in Computer Science, 753:141–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chalmers:1993:VCI

[2355] M. Chalmers. Visualisation of complex information. Lecture Notes in Computer Science, 753:152–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patel:1993:TDV

[2356] U. K. Patel and A. G. Sutcliffe. Threedimensional visualisation of knowledge structures: Prototyping for design evaluation. Lecture Notes in Computer Science, 753:163–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ushakov:1993:VPR

[2357] I. Ushakov and I. Velbitskiy. Visual programming in R-technology: Concepts, systems and perspectives. Lecture Notes in Computer Science, 753: 181–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Springett:1993:ISP

[2358] M. V. Springett, A. S. Grant, and A. G. Sutcliffe. Interface semantics and procedural knowledge: A study of novice understanding of Mac-Draw. Lecture Notes in Computer Science, 753:196–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leventhal:1993:PTB

[2359] L. M. Leventhal, B. M. Teasley, D. S. Rohlman, and K. Instone. Positive test bias in software testing among professionals. Lecture Notes in Computer Science, 753:210–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salber:1993:AWO

[2360] D. Salber and J. Coutaz. Applying the Wizard of Oz technique to the study of multimodal systems. Lecture Notes in Computer Science, 753:219–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coutaz:1993:MFD

[2361] J. Coutaz, L. Nigay, and D. Salber. The MSM framework: A design space for multi-sensori-motor systems. Lecture Notes in Computer Science, 753:231–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniol:1993:RRB

[2362] G. Antoniol, R. Fiutem, R. Flor, and G. Lazzari. Radiological reporting based on voice recognition. Lecture Notes in Computer Science, 753:242– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hubbard:1993:TID

[2363] C. Hubbard and J. H. Bradford. Task interference with a discrete word recognizer. Lecture Notes in Computer Science, 753:254–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamada:1993:MUU

[2364] K. Yamada, R. Mizoguchi, N. Harada, A. Nukuzuma, K. Ishimaru, and H. Furukawa. Model of utterance and its use in cooperative response generation. Lecture Notes in Computer Science, 753:260–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karshmer:1993:SCI

[2365] A. I. Karshmer and R. L. Oliver. Special computer interfaces for the visually handicapped: F.O.B. the manufacturer. Lecture Notes in Computer Science, 753:272–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subbotin:1993:IPC

[2366] M. Subbotin and D. Subbotin. IN-TELTEXT: Producing coherent linear texts while navigating in large non-hierarchical hypertexts. Lecture Notes in Computer Science, 753:281– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Instone:1993:CEI

[2367] K. Instone, E. Brown, L. M. Leventhal, and B. M. Teasley. The challenge of effectively integrating graphics into hypertext. Lecture Notes in Computer Science, 753:290–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chelnokov:1993:CNH

[2368] V. M. Chelnokov and V. L. Zephyrova. Coherent navigation in hypertext environments: The SMIsC conception. Lecture Notes in Computer Science, 753:298–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berthome-Montoy:1993:GSA

[2369] A. Berthome-Montoy. Generating selfadaptive human-computer interfaces. Lecture Notes in Computer Science, 753:318–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cypher:1993:PUM

[2370] A. Cypher. The practical use of macro recording: A case study. Lecture Notes in Computer Science, 753:327– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conway:1993:HSO

[2371] D. Conway. HyperLecture: A selforganizing lecture presentation and revision system. Lecture Notes in Computer Science, 753:334–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brusilovsky:1993:TAH

[2372] P. Brusilovsky, L. Pesin, and M. Zyryanov. Towards an adaptive hypermedia component for an intelligent learning environment. Lecture Notes in Computer Science, 753:348–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1993:IIC

[2373] E. Wilson. An intelligent interface for computer assisted language learning. Lecture Notes in Computer Science, 753:359–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derder:1993:SMA

[2374] A. Derder and C. Garbay. A system to model, assist and control the human observation of microscopic specimen. Lecture Notes in Computer Science, 753:371–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chernin:1993:DDA

[2375] A. Chernin. The “document-driven activity” approach to modelling and designing office automation systems. Lecture Notes in Computer Science, 753: 380–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esch:1993:SCC

[2376] John W. Esch. The scope of coreference in conceptual graphs. Lecture Notes in Computer Science, 754:3–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulin:1993:UWS

[2377] Bernard Moulin and Guy W. Mineau.

Using world structures for factoring knowledge. Lecture Notes in Computer Science, 754:13–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mineau:1993:SKS

[2378] Guy W. Mineau. Sharing knowledge:

Starting with the integration of vocabularies. Lecture Notes in Computer Science, 754:34–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cambell:1993:CGA

[2379] Linda Cambell and Peter Creasy. A conceptual graphs approach to information systems design. Lecture Notes in Computer Science, 754:46–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tepfenhart:1993:RKA

[2380] William M. Tepfenhart. Representing knowledge about substances. Lecture Notes in Computer Science, 754:59–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulin:1993:RTK

[2381] Bernard Moulin. Representing temporal knowledge in discourse: an approach extending the conceptual graph theory. Lecture Notes in Computer Science, 754:72–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kocura:1993:TSI

[2382] Pavel Kocura. Towards a semantics of inchoative and causation events in conceptual graphs. Lecture Notes in Computer Science, 754:96–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oh:1993:DED

[2383] Jonathan C. Oh and Stephen Graham. Does every difference make a difference? Lecture Notes in Computer Science, 754:113–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delugach:1993:ESD

[2384] Harry S. Delugach. An exploration into semantic distance. Lecture Notes in Computer Science, 754:119–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouaud:1993:RCG

[2385] Jacques Bouaud and Pierre Zweigenbaum. A reconstruction of conceptual graphs on top of a production system. Lecture Notes in Computer Science, 754:127–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myaeng:1993:UHP

[2386] Sung H. Myaeng and Christopher Khoo. On uncertainty handling in plausible reasoning with conceptual graphs. Lecture Notes in Computer Science, 754:137–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whipple:1993:EHE

[2387] Weldon Whipple. Expert humans and expert systems: Toward a unity of uncertain reasoning. Lecture Notes in Computer Science, 754:148–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garner:1993:KF

[2388] Brian John Garner and Dickson Lukose. Knowledge fusion. Lecture Notes in Computer Science, 754:158– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berg-Cross:1993:UCS

[2389] Gary Berg-Cross. Using conceptual structures to translate data models: Concepts, context and cognitive processes. Lecture Notes in Computer Science, 754:171–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wuwongse:1993:TDO

[2390] Vilas Wuwongse and Bikash C. Gosh. Towards deductive object-oriented databases based on conceptual graphs. Lecture Notes in Computer Science, 754:188–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delugach:1993:ADI

[2391] Hary S. Delugach and Thomas H. Hinke. AERIE: Database inference modeling and detection using conceptual graphs. Lecture Notes in Computer Science, 754:206–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellis:1993:BPC

[2392] Gerard Ellis and Robert Levinson. The birth of PEIRCE: A conceptual graphs workbench. Lecture Notes in Computer Science, 754:219–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chein:1993:SWD

[2393] M. Chein and M. L. Mugnier. Specialization: Where do the difficulties occur? Lecture Notes in Computer Science, 754:229–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mugnier:1993:PAP

[2394] M. L. Mugnier and M. Chein. Polynomial algorithms for projection and matching. Lecture Notes in Computer Science, 754:239–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gi-Chul:1993:CNC

[2395] Yang Gi-Chul, Bae Choi Young, and Jonathan C. Oh. CGMA: A novel conceptual graph matching algorithm. Lecture Notes in Computer Science, 754: 252–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wermelinger:1993:XWT

[2396] Michel Wermelinger and Jose Gabriel Lopes. An X-windows toolkit for knowledge acquisition and representation based on conceptual structures. Lecture Notes in Computer Science, 754:262–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mann:1993:ACG

[2397] Graham A. Mann. Assembly of conceptual graphs from natural language by means of multiple knowledge specialists. Lecture Notes in Computer Science, 754:275–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dogru:1993:STC

[2398] Sait Dogru and James R. Slagle. A system that translates conceptual structures into english. Lecture Notes in Computer Science, 754:283–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroeder:1993:KBA

[2399] Martin Schroeder. Knowledge based analysis of radiology reports using conceptual graphs. Lecture Notes in Computer Science, 754:293–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hines:1993:OSI

[2400] Timothy R. Hines. Open systems interconnection abstract syntax notation: ASN.CG. Lecture Notes in Computer Science, 754:303–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Polovina:1993:BAB

[2401] Simon Polovina. Bridging accounting and business strategic planning using conceptual graphs. Lecture Notes in Computer Science, 754:312–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1993:SPR

[2402] Zhaohui Wu, Bernardi Ansgar, and Klauck Christoph. Skeletal plans reuse: A restricted conceptual graph approach. Lecture Notes in Computer Science, 754:322–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moller:1993:I

[2403] Bernhard Möller, Helmut Partsch, and Steve Schuman. Introduction. Lecture Notes in Computer Science, 755:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Backhouse:1993:ERT

[2404] Roland Backhouse and Paul Hoogendijk. Elements of a relational theory of datatypes. Lecture Notes in Computer Science, 755:7–42, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bird:1993:DPG

[2405] Richard Bird and Oege de Moor. From dynamic programming to greedy algorithms. Lecture Notes in Computer Science, 755:43–61, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyle:1993:PTF

[2406] James Boyle and Terence Harmer.

Practical transformation of functional programs for efficient execution: a case study. Lecture Notes in Computer Science, 755:62–88, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feather:1993:BOS

[2407] Martin Feather. Behavior-oriented specification in Gist. Lecture Notes in Computer Science, 755:89–122, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moller:1993:DGP

[2408] Bernhard Möller. Derivation of graph and pointer algorithms. Lecture Notes in Computer Science, 755:123–160, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morgan:1993:RCL

[2409] Carroll Morgan. The refinement calculus, and literate development. Lecture Notes in Computer Science, 755:161– 182, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Partsch:1993:FPS

[2410] Helmut Partsch. Formal problem specification on an algebraic basis. Lecture Notes in Computer Science, 755:183– 224, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pepper:1993:PDA

[2411] Peter Pepper. Program development in an algebraic setting. Lecture Notes in Computer Science, 755:225–262, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pettorossi:1993:RSP

[2412] Alberto Pettorossi and Maurizio Proietti. Rules and strategies for program transformation. Lecture Notes in Computer Science, 755:263–304, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sintzoff:1993:ET

[2413] Michel Sintzoff. Endomorphic typing. Lecture Notes in Computer Science, 755:305–323, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1993:ADA

[2414] Douglas Smith. Automating the design of algorithms. Lecture Notes in Computer Science, 755:324–354, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swierstra:1993:VDS

[2415] Doaitse Swierstra and Oege de Moor.

Virtual data structures. Lecture Notes in Computer Science, 755:355–371, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1993:CHP

[2416] W. Kim and G. Agha. Compilation of a highly parallel actor-based language. Lecture Notes in Computer Science, 757:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarkar:1993:CES

[2417] V. Sarkar. A concurrent execution semantics for parallel program graphs and program dependence graphs. Lecture Notes in Computer Science, 757: 16–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1993:UPI

[2418] W. Chen, R. Bringmann, S. Mahlke, S. Anik, T. Kiyohara, N. Warter, D. Lavery, W.-M. Hwu, R. Hank, and J. Gyllenhaal. Using profile information to assist advanced compiler optimization and scheduling. Lecture Notes in Computer Science, 757:31–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brownhill:1993:HPC

[2419] C. Brownhill and A. Nicolau. A hierarchical parallelizing compiler for VLIW/ MIMD machines. Lecture Notes in Computer Science, 757:49–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peterson:1993:DDA

[2420] P. Peterson and D. Padua. Dynamic dependence analysis: A novel method for data dependence evaluation. Lecture Notes in Computer Science, 757:64–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solworth:1993:FDP

[2421] J. Solworth. On the feasibility of dynamic partitioning of pointer structures. Lecture Notes in Computer Science, 757:82–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vonHanxleden:1993:CAI

[2422] R. von Hanxleden, K. Kennedy, C. Koelbel, R. Das, and J. Saltz. Compiler analysis for irregular problems in Fortran D. Lecture Notes in Computer Science, 757:97–111, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1993:DEO

[2423] C. Lin and L. Snyder. Data ensembles in Orca C. Lecture Notes in Computer Science, 757:112–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chandy:1993:CCC

[2424] K. Mani Chandy and C. Kesselman. Compositional C++: Compositional parallel programming. Lecture Notes in Computer Science, 757:124–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carriero:1993:DPL

[2425] N. Carriero and D. Gelernter. Data parallelism and linda. Lecture Notes in Computer Science, 757:145–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chien:1993:TEE

[2426] A. Chien, W. Feng, V. Karamcheti, and J. Plevyak. Techniques for efficient execution of fine-grained concurrent programs. Lecture Notes in Computer Science, 757:160–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeremiassen:1993:CPP

[2427] T. Jeremiassen and S. Eggers. Computing per-process summary side-effect information. Lecture Notes in Computer Science, 757:175–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rogers:1993:SSE

[2428] A. Rogers, J. Reppy, and L. Hendren.

Supporting SPMD execution for dynamic data structures. Lecture Notes in Computer Science, 757:192–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appelbe:1993:DTS

[2429] W. Appelbe and K. Smith. Determining transformation sequences for loop parallelization. Lecture Notes in Computer Science, 757:208–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1993:COM

[2430] M. Chen and Y. Hu. Compiler optimizations for massively parallel machines: Transformations on iterative spatial loops. Lecture Notes in Computer Science, 757:223–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:1993:HDD

[2431] B. Chapman, H. Zima, and P. Mehrotra. Handling distributed data in Vienna Fortran procedures. Lecture Notes in Computer Science, 757:248–263, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1993:SPP

[2432] S. Gupta, C.-H. Huang, P. Sadayappan, and R. Johnson. On the synthesis of parallel programs from tensor product formulas for block recursive algorithms. Lecture Notes in Computer Science, 757:264–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gao:1993:CLF

[2433] G. Gao, R. Olsen, V. Sarkar, and R. Thekkath. Collective loop fusion for array contraction. Lecture Notes in Computer Science, 757:281–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1993:PHD

[2434] Y.-F. Lee and B. Ryder. Parallel hybrid data flow algorithms. Lecture Notes in Computer Science, 757:296–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1993:CPP

[2435] H. Dietz and W. Cohen. A controlparallel programming model implemented on SIMD hardware. Lecture Notes in Computer Science, 757:311– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larus:1993:LGO

[2436] J. Larus. C^{∗∗}: A large-grain, objectoriented, data-parallel programming language. Lecture Notes in Computer Science, 757:326–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hankin:1993:CGP

[2437] C. Hankin, D. LeMetayer, and D. Sands. A calculus of gamma programs. Lecture Notes in Computer Science, 757:342–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1993:LBR

[2438] P. Ciancarini. A Linda-based runtime system for a distributed logic language. Lecture Notes in Computer Science, 757:356–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lempel:1993:PCD

[2439] O. Lempel, S. Pinter, and E. Turiel.

Parallelizing a C dialect for distributed memory MIMD machines. Lecture Notes in Computer Science, 757:369– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1993:SLT

[2440] W. Li and K. Pingali. A singular loop transformation framework based on non-singular matrices. Lecture Notes in Computer Science, 757:391–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hendren:1993:DMC

[2441] L. Hendren, C. Donawa, M. Emami, G. Gao, Justiani, and B. Sridharan. Designing the McCAT compiler based on a family of structured intermediate representations. Lecture Notes in Computer Science, 757:406–420, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolfe:1993:DJA

[2442] M. Wolfe. Doany: Not just another parallel loop. Lecture Notes in Computer Science, 757:421–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maydan:1993:DDD

[2443] D. Maydan, S. Amarsinghe, and M. Lam. Data dependence and dataflow analysis of arrays. Lecture Notes in Computer Science, 757:434–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Netzer:1993:ETR

[2444] R. Netzer and B. Miller. Experience with techniques for refining data race detection. Lecture Notes in Computer Science, 757:449–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klappholz:1993:EBW

[2445] D. Klappholz and X. Kong. Extending the Banerjee-Wolfe test to handle execution conditions. Lecture Notes in Computer Science, 757:464–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yasue:1993:FCM

[2446] T. Yasue, H. Yamana, and Y. Muraoka. A FORTRAN compiling method for dataflow machines and its prototype compiler for the parallel processing system — harray. Lecture Notes in Computer Science, 757:482–496, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duesterwald:1993:DSP

[2447] E. Duesterwald and R. Gupta.

Distributed slicing and partial reexecution for distributed programs. Lecture Notes in Computer Science, 757:497–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1993:PEV

[2448] W. Harrison and Z. Ammarguellat. A program’s eye view of Miprac. Lecture Notes in Computer Science, 757:512– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haghighat:1993:SPA

[2449] M. Haghighat and C. Polychronopoulos. Symbolic program analysis and optimization for parallelizing compilers. Lecture Notes in Computer Science, 757:538–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hinrichs:1993:UNC

[2450] S. Hinrichs and T. Gross. Utilizing new communication features in compilation for private-memory machines. Lecture Notes in Computer Science, 757:563–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teillaud:1993:TDR

[2451] Monique Teillaud. Towards dynamic randomized algorithms in computational geometry. Lecture Notes in Computer Science, 758:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehani:1993:CSC

[2452] N. Gehani, H. V. Jagadish, and O. Shmueli. COMPOSE: A system for composite specification and detection. Lecture Notes in Computer Science, 759:3–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertino:1993:ACO

[2453] E. Bertino, S. Jajodia, and P. Samarati. Access controls in OO database systems — some approaches and issues. Lecture Notes in Computer Science, 759:17–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vadaparty:1993:DPN

[2454] K. Vadaparty and S. Naqvi. The decomposition property of non-deterministic databases. Lecture Notes in Computer Science, 759:45–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sheu:1993:AOB

[2455] P. C.-Y. Sheu and L. Peterson. The architectures of an object base environment for simulation. Lecture Notes in Computer Science, 759:65–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhargava:1993:TRO

[2456] Bharat K. Bhargava, Yin he Jiang, Jagannathan Srinivasan, and Prasun Dewan. Transition from a relation to object model implementation. Lecture Notes in Computer Science, 759:87– 103, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Su:1993:OOK

[2457] S. Y. W. Su and Y. Shyy. An objectoriented knowledge model for KBMS-Supported evolutionary prototyping of software systems. Lecture Notes in Computer Science, 759:105–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Subramanian:1993:AOD

[2458] R. Subramanian and N. R. Adam. Applying OOAD in the design and implementation of an intelligent geographic information system. Lecture Notes in Computer Science, 759:127–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clifford:1993:ID

[2459] J. Clifford. Indexical databases. Lecture Notes in Computer Science, 759: 153–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elmasri:1993:TQL

[2460] R. Elmasri and V. Kouramajian. A temporal query language for a conceptual model. Lecture Notes in Computer Science, 759:175–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Segev:1993:DMT

[2461] A. Segev and R. Chandra. A data model for time-series analysis. Lecture Notes in Computer Science, 759:191– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gadia:1993:RMS

[2462] S. K. Gadia and V. Chopra. A relational model and SQL-like query language for spatial databases. Lecture Notes in Computer Science, 759:213– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1993:PQP

[2463] P. S. Yu, M.-S. Chen, J. L. Wolf, and J. Turek. Parallel query processing. Lecture Notes in Computer Science, 759:229–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flater:1993:TFD

[2464] D. W. Flater and Y. Yesha. Towards flexible distributed information retrieval. Lecture Notes in Computer Science, 759:259–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tewari:1993:EPR

[2465] R. Tewari. Efficient parallel recovery in replicated databases. Lecture Notes in Computer Science, 759:277–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siegelmann:1993:DAM

[2466] H. T. Siegelmann and O. Frieder. Document allocation in multiprocessor information retrieval systems. Lecture Notes in Computer Science, 759:289– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franchitti:1993:ATC

[2467] J.-C. Franchitti and R. King. Amalgame: A tool for creating interoperating, persistent, heterogeneous components. Lecture Notes in Computer Science, 759:313–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karabatis:1993:CEM

[2468] G. Karabatis, M. Rusinkiewicz, and A. Sheth. Correctness and enforcement of multidatabase interdependencies. Lecture Notes in Computer Science, 759:337–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersson:1993:FFM

[2469] M. Andersson, Y. Dupont, S. Spaccapietra, K. Yetongnon, M. Tresch, and H. Ye. FEMUS: A federated multilingual database system. Lecture Notes in Computer Science, 759:359–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Browne:1993:CSI

[2470] S. Browne. Communication and synchronization issues in distributed multimedia database systems. Lecture Notes in Computer Science, 759:381– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghafoor:1993:MDS

[2471] A. Ghafoor and P. B. Berra. Multimedia database systems. Lecture Notes in Computer Science, 759:397–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1993:B

[2472] Anonymous. Bibliography. Lecture Notes in Computer Science, 759:413– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maier:1993:TPO

[2473] D. Maier and J. B. Cushing. Treating programs as objects: The computational proxy experience. Lecture Notes in Computer Science, 760:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gelder:1993:FAD

[2474] A. V. Gelder. Foundations of aggregation in deductive databases. Lecture Notes in Computer Science, 760:13–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kostler:1993:DFO

[2475] G. Kostler, W. KieBling, H. Thone, and U. Guntzer. The differential fixpoint operator with subsumption. Lecture Notes in Computer Science, 760: 35–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corciulo:1993:DND

[2476] L. Corciulo, F. Giannotti, and D. Pedreschi. Datalog with nondeterministic choice computes NDB-PTIME. Lecture Notes in Computer Science, 760:49–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caseau:1993:DOO

[2477] Y. Caseau, P-Y. Guillo, and E. Levenez. A deductive and object-oriented approach to a complex scheduling problem. Lecture Notes in Computer Science, 760:67–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakshmanan:1993:LFS

[2478] L. V. S. Lakshmanan, F. Sadri, and I. N. Subramanian. On the logical foundations of schema integration and evolution in heterogeneous database systems. Lecture Notes in Computer Science, 760:81–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arora:1993:EPE

[2479] T. Arora, R. Ramakrishnan, W. G.

Roth, P. Seshadri, and D. Srivastava. Explaining program execution in deductive systems. Lecture Notes in Computer Science, 760:101–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coburn:1993:LRB

[2480] N. Coburn and G. E. Weddell. A logic for rule-based query optimization in graph-based data models. Lecture Notes in Computer Science, 760:120– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fegaras:1993:SRB

[2481] L. Fegaras, D. Maier, and T. Sheard. Specifying rule-based query optimizers in a reflective framework. Lecture Notes in Computer Science, 760:146– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoon:1993:SQO

[2482] J. P. Yoon and L. Kerschberg. Semantic query optimization in deductive objectoriented databases. Lecture Notes in Computer Science, 760:169–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramakrishnan:1993:RDO

[2483] R. Ramakrishnan. Research in deductive and object-oriented databases. Lecture Notes in Computer Science, 760:183–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramamohanarao:1993:IOA

[2484] K. Ramamohanarao. An implementation overview of the Aditi deductive database system. Lecture Notes in Computer Science, 760:184–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaniolo:1993:NAR

[2485] C. Zaniolo, N. Arni, and K. Ong.

Negation and aggregates in recursive rules: The LDL++Approach. Lecture Notes in Computer Science, 760:204– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atzeni:1993:IDL

[2486] P. Atzeni, L. Cabibbo, and G. Mecca.

ISALOG -: A deductive language with negation for complex-object databases with hierarchies. Lecture Notes in Computer Science, 760:222–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1993:ERI

[2487] X. Zhang and Z. M. Ozsoyoglu. On efficient reasoning with implication constraints. Lecture Notes in Computer Science, 760:236–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brass:1993:BQE

[2488] S. Brass and U. W. Lipeck. Bottom-up query evaluation with partially ordered defaults. Lecture Notes in Computer Science, 760:253–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandenBussche:1993:EPE

[2489] J. Van den Bussche and G. Vossen.

An extension of path expressions to simplify navigation in object-oriented queries. Lecture Notes in Computer Science, 760:267–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staudt:1993:QC

[2490] M. Staudt, M. Jarke, M. A. Jeusfeld, and H. W. Nissen. Query classes. Lecture Notes in Computer Science, 760: 283–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laurent:1993:DUR

[2491] D. Laurent, V. Phan Luong, and N. Spyratos. Database updating revisited. Lecture Notes in Computer Science, 760:296–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konomi:1993:SKC

[2492] S. Konomi, T. Furukawa, and Y. Kambayashi. Super-key classes for updating materialized derived classes in object bases. Lecture Notes in Computer Science, 760:310–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietrich:1993:ADO

[2493] S. W. Dietrich. Applications of deductive and object-oriented databases. Lecture Notes in Computer Science, 760:327–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manthey:1993:BDD

[2494] Rainer Manthey. Beyond data dictionaries: Towards a reflective architecture of intelligent database systems. Lecture Notes in Computer Science, 760:328–339, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bal:1993:DTO

[2495] R. Bal and H. Balsters. A deductive and typed object-oriented language. Lecture Notes in Computer Science, 760:340–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mumick:1993:NLD

[2496] I. S. Mumick and K. A. Ross. Noodle: A language for declarative querying in an object-oriented database. Lecture Notes in Computer Science, 760: 360–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mattox:1993:TCD

[2497] D. Mattox, K. Smith, and S. C. Y. Lu.

Tracking causal dependencies in an active object-oriented database. Lecture Notes in Computer Science, 760:379– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elmasri:1993:ACM

[2498] R. Elmasri, S. James, and V. Kouramajian. Automatic class and method generation for object-oriented databases. Lecture Notes in Computer Science, 760:395–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertino:1993:MME

[2499] E. Bertino and S. Jajodia. Modeling multilevel entities using single level objects. Lecture Notes in Computer Science, 760:415–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koh:1993:MUC

[2500] T. W. Koh, B. C. Ooi, and Y. S. Ho. A model using classes as a basic organization tool. Lecture Notes in Computer Science, 760:429–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuan:1993:KBR

[2501] L. Y. Yuan and J-H. You. Knowledge base revision using circumscription. Lecture Notes in Computer Science, 760:444–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kesim:1993:VOD

[2502] F. N. Kesim and M. Sergot. Versioning of objects in deductive databases. Lecture Notes in Computer Science, 760: 459–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobbie:1993:MSM

[2503] G. Dobbie and R. Topor. A model for sets and multiple inheritance in deductive object-oriented systems. Lecture Notes in Computer Science, 760:473– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartmanis:1993:SOA

[2504] J. Hartmanis. Some observations about the nature of computer science. Lecture Notes in Computer Science, 761:1–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanBakel:1993:EIT

[2505] S. van Bakel. Essential intersection type assignment. Lecture Notes in Computer Science, 761:13–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ait-Kaci:1993:LSL

[2506] H. Ait-Kaci and J. Garrigue. Labelselective lambda-calculus syntax and confluence. Lecture Notes in Computer Science, 761:24–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barendsen:1993:CUT

[2507] E. Barendsen and S. Smetsers. Conventional and uniqueness typing in graph rewrite systems. Lecture Notes in Computer Science, 761:41–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castagna:1993:MLT

[2508] G. Castagna. A meta-language for typed object-oriented languages. Lecture Notes in Computer Science, 761: 52–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berry:1993:PCS

[2509] G. Berry. Preemption in concurrent systems. Lecture Notes in Computer Science, 761:72–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kulkarni:1993:LVN

[2510] S. R. Kulkarni, S. K. Mitter, T. J.

Richardson, and J. N. Tsitsiklis. Local versus non-local computation of length of digitized curves. Lecture Notes in Computer Science, 761:94–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1993:DIP

[2511] K.-J. Lange and R. Niedermeier. Dataindependences of parallel random access machines. Lecture Notes in Computer Science, 761:104–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguzzi:1993:PTL

[2512] G. Aguzzi and U. Modigliani. Proving termination of logic programs by transforming them into equivalent term rewriting systems. Lecture Notes in Computer Science, 761:114–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rao:1993:CHC

[2513] M. R. K. Krishna Rao. Completeness of hierarchical combinations of term rewriting systems. Lecture Notes in Computer Science, 761:125–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1993:HOS

[2514] N. Dershowitz and S. Mitra. Higherorder and semantic unification. Lecture Notes in Computer Science, 761: 139–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1993:CEF

[2515] D. Basin and S. Matthews. A conservative extension of first-order logic and its applications to theorem proving. Lecture Notes in Computer Science, 761:151–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stuckey:1993:WFO

[2516] P. J. Stuckey and S. Sudarshan.

Well-founded ordered search. Lecture Notes in Computer Science, 761:161– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramakrishna:1993:RTI

[2517] Y. S. Ramakrishna, L. K. Dillon, L. E. Moser, P. M. Melliar-Smith, and G. Kutty. A real-time interval logic and its decision procedure. Lecture Notes in Computer Science, 761:173–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fages:1993:SOP

[2518] F. Fages. On the semantics of optimization predicates in CLP languages. Lecture Notes in Computer Science, 761: 193–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramachandran:1993:IAC

[2519] V. Ramachandran and P. V. Hentenryck. Incremental algorithms for constraint solving and entailment over rational trees. Lecture Notes in Computer Science, 761:205–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guha:1993:PPV

[2520] S. Guha and I. Suzuki. Proximity problems and the Voronoi diagram on a rectilinear plane with rectangular obstacles. Lecture Notes in Computer Science, 761:218–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asberg:1993:FDS

[2521] B. Asberg, G. Blanco, P. Bose, J. Garcia-Lopez, M. Overmars, G. Toussaint, G. Wilfong, and B. Zhu. Feasibility of design in stereolithography. Lecture Notes in Computer Science, 761:228–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radhakrishnan:1993:CLP

[2522] V. Radhakrishnan, S. O. Krumke, M. V. Marathe, D. J. Rosenkrantz, and S. S. Ravi. Compact location problems. Lecture Notes in Computer Science, 761:238–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halldorsson:1993:SCC

[2523] M. M. Halldorsson, J. Radhakrishnan, and K. V. Subrahmanyam. On some communication complexity problems related to threshold functions. Lecture Notes in Computer Science, 761:248– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cucker:1993:RCN

[2524] F. Cucker and F. Rossello. Recursiveness over the complex numbers is timebounded. Lecture Notes in Computer Science, 761:260–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mulmuley:1993:LBS

[2525] K. Mulmuley. A lower bound for solvability of polynomial equations. Lecture Notes in Computer Science, 761: 268–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reif:1993:RPS

[2526] W. Reif and K. Stenzel. Reuse of proofs in software verification. Lecture Notes in Computer Science, 761:284– ??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norvell:1993:ISI

[2527] T. S. Norvell. Induce-statements and induce-expressions: Constructs for inductive programming. Lecture Notes in Computer Science, 761:294–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoine:1993:GLB

[2528] C. Antoine, B. Le Goff, and J.-E. Pin.

A graphic language based on timing diagrams. Lecture Notes in Computer Science, 761:306–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narasimhan:1993:STI

[2529] R. Narasimhan. Software technology:

Integrating theory and practice. Lecture Notes in Computer Science, 761: 317–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bacchus:1993:GDB

[2530] F. Bacchus, A. J. Grove, J. Y. Halpern, and D. Koller. Generating degrees of belief from statistical information: An overview. Lecture Notes in Computer Science, 761:318–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1993:CRS

[2531] A. Cheng, J. Esparza, and J. Palsberg. Complexity results for 1-safe nets. Lecture Notes in Computer Science, 761:326–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janin:1993:SRA

[2532] D. Janin. Some results about logical descriptions of non-deterministic behaviours. Lecture Notes in Computer Science, 761:338–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janicki:1993:OSG

[2533] R. Janicki and M. Koutny. Order structures and generalisations of Szpilrajn’s theorem. Lecture Notes in Computer Science, 761:348–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kumar:1993:IRA

[2534] K. N. Kumar and P. K. Pandya. ICSP and its relationship with ACSP and CSP. Lecture Notes in Computer Science, 761:358–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honda:1993:RBP

[2535] K. Honda and N. Yoshida. On reduction-based process semantics. Lecture Notes in Computer Science, 761:373–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukund:1993:KTL

[2536] M. Mukund and M. Sohoni. Keeping track of the latest gossip: Bounded time-stamps suffice. Lecture Notes in Computer Science, 761:388–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1993:TOS

[2537] S. Aggarwal and S. Kutten. Time optimal self-stabilizing spanning tree algorithms. Lecture Notes in Computer Science, 761:400–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pujari:1993:EAS

[2538] A. K. Pujari. Efficient algorithm to sort linear combinations of arrays. Lecture Notes in Computer Science, 761: 411–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaishnavi:1993:SFS

[2539] V. K. Vaishnavi and D. Khanna. A simple file structure for the weighted dictionary problem. Lecture Notes in Computer Science, 761:419–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubhashi:1993:SSR

[2540] D. P. Dubhashi, K. Mehlhorn, D. Ranjan, and C. Thiel. Searching, sorting and randomised algorithms for central elements and ideal counting in posets. Lecture Notes in Computer Science, 761:436–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhattacharyya:1993:LCR

[2541] P. Bhattacharyya and G. Nagaraja. Learning classes of regular and linear languages in Valiant’s learnability framework. Lecture Notes in Computer Science, 761:444–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBerg:1993:RGD

[2542] M. De Berg, L. Guibas, D. Halperin, and M. Overmars. Reaching a goal with directional uncertainty. Lecture Notes in Computer Science, 762:1–10, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1993:CDS

[2543] G. Das and P. J. Heffernan. Constructing degree-3 spanners with other sparseness properties. Lecture Notes in Computer Science, 762:11–20, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobrindt:1993:RCH

[2544] K. Dobrindt and M. Yvinec. Remembering conflicts in history yields dynamic algorithms. Lecture Notes in Computer Science, 762:21–30, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Furer:1993:CRG

[2545] M. Furer, C. R. Subramanian, and C. E. Veni Madhavan. Coloring random graphs in polynomial expected time. Lecture Notes in Computer Science, 762:31–37, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asano:1993:GDS

[2546] T. Asano. Graphical degree sequence problems with connectivity requirements. Lecture Notes in Computer Science, 762:38–47, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dai:1993:HTD

[2547] Y. Dai, H. Imai, K. Iwano, and N. Katoh. How to treat delete requests in semi-online problems. Lecture Notes in Computer Science, 762:48–57, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carlsson:1993:FSW

[2548] S. Carlsson, H. Jonsson, and B. J. Nilsson. Finding the shortest watchman route in a simple polygon. Lecture Notes in Computer Science, 762:58–67, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1993:CSW

[2549] X. Tan and T. Hirata. Constructing shortest watchman routes by divideand-conquer. Lecture Notes in Computer Science, 762:68–77, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1993:GCR

[2550] F. Hoffmann and K. Kriegel. A graph coloring result and its consequences for some guarding problems. Lecture Notes in Computer Science, 762:78–87, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dessmark:1993:MDD

[2551] A. Dessmark, K. Jansen, and A. Lingas. The maximum k-dependent and fdependent set problem. Lecture Notes in Computer Science, 762:88–97, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahashi:1993:FSN

[2552] J.-Y. Takahashi, H. Suzuki, and T. Nishizeki. Finding shortest noncrossing rectilinear paths in plane regions. Lecture Notes in Computer Science, 762:98–107, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1993:TCG

[2553] T. Kloks. Treewidth of circle graphs. Lecture Notes in Computer Science, 762:108–117, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1993:FCH

[2554] C. Chen. A framework for constructing heap-like structures in-place. Lecture Notes in Computer Science, 762:118– 127, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khoong:1993:DEB

[2555] C. M. Khoong and H. W. Leong.

Double-ended binomial queues. Lecture Notes in Computer Science, 762:128– 137, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleischer:1993:SBS

[2556] R. Fleischer. A simple balanced search tree with O(1) worst-case update time. Lecture Notes in Computer Science, 762:138–146, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oehring:1993:MDD

[2557] S. R. Oehring and S. K. Das. Mapping dynamic data and algorithm structures into product networks. Lecture Notes in Computer Science, 762:147–156, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cogolludo:1993:PRR

[2558] J. C. Cogolludo and S. Rajasekaran.

Permutation routing on reconfigurable meshes. Lecture Notes in Computer Science, 762:157–166, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grammatikakis:1993:AOA

[2559] M. D. Grammatikakis, D. F. Hsu, and F. K. Hwang. Adaptive and oblivious algorithms for D-cube permutation routing. Lecture Notes in Computer Science, 762:167–175, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srivastav:1993:QLA

[2560] A. Srivastav and P. Stangier. On quadratic lattice approximations. Lecture Notes in Computer Science, 762: 176–184, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujito:1993:AMM

[2561] T. Fujito. A 2/3-approximation of the matroid matching problem. Lecture Notes in Computer Science, 762:185– 190, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dube:1993:UFG

[2562] S. Dube. Using fractal geometry for solving divide-and-conquer recurrences. Lecture Notes in Computer Science, 762:191–200, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruskey:1993:SCG

[2563] F. Ruskey. Simple combinatorial gray codes constructed by reversing sublists. Lecture Notes in Computer Science, 762:201–208, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borodin:1993:TST

[2564] A. Borodin. Time-space tradeoffs. Lecture Notes in Computer Science, 762: 209–220, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leung:1993:SEA

[2565] H. Leung. Separating exponentially ambiguous NFA from polynomially ambiguous NFA. Lecture Notes in Computer Science, 762:221–229, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1993:TCC

[2566] Y. Han, L. A. Hemaspaandra, and T. Thierauf. Threshold computation and cryptographic security. Lecture Notes in Computer Science, 762:230– 239, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niedermeier:1993:PRW

[2567] R. Niedermeier and P. Rossmanith.

On the power of reading and writing simultaneously in parallel computations. Lecture Notes in Computer Science, 762:240–249, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Book:1993:RCC

[2568] R. V. Book. Relativizing complexity classes with random oracles. Lecture Notes in Computer Science, 762:250– 258, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liestman:1993:PG

[2569] A. L. Liestman and D. Richards. Perpetual gossiping. Lecture Notes in Computer Science, 762:259–266, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ikeda:1993:PSN

[2570] T. Ikeda. A probabilistic selection network with butterfly networks. Lecture Notes in Computer Science, 762:267– 276, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1993:OGG

[2571] S. Fujita, M. Yamashita, and T. Ae. Optimal group gossiping in hypercubes under wormhole routing model. Lecture Notes in Computer Science, 762:277– 286, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bitan:1993:OLB

[2572] S. Bitan and S. Zaks. Optimal linear broadcast routing with capacity limitations. Lecture Notes in Computer Science, 762:287–296, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Awerbuch:1993:SRA

[2573] B. Awerbuch and T. Leighton. A survey of recent advances in multicommodity flow: Algorithms and applications. Lecture Notes in Computer Science, 762:297–302, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1993:PCC

[2574] X. He and M.-Y. Kao. Parallel construction of canonical ordering and convex drawing of triconnected planar graphs. Lecture Notes in Computer Science, 762:303–312, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asano:1993:NTH

[2575] T. Asano and N. Katoh. Number theory helps line detection in digital images. Lecture Notes in Computer Science, 762:313–322, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1993:OCS

[2576] D. Z. Chen. Optimally computing the shortest weakly visible subedge of a simple polygon. Lecture Notes in Computer Science, 762:323–332, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weihe:1993:MFE

[2577] K. Weihe. Multicommodity flows in even, planar networks. Lecture Notes in Computer Science, 762:333–342, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ripphausen-Lipa:1993:LTA

[2578] H. Ripphausen-Lipa, D. Wagner, and K. Weihe. Linear time algorithms for disjoint two-face paths. Lecture Notes in Computer Science, 762:343–352, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deng:1993:RMF

[2579] X. Deng and A. Mirzaian. Robot mapping: Foot-prints vs tokens. Lecture Notes in Computer Science, 762:353– 362, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yannakakis:1993:RDA

[2580] M. Yannakakis. Recent developments in the approximability of combinatorial problems. Lecture Notes in Computer Science, 762:363–368, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okamoto:1993:RAC

[2581] T. Okamoto. On the relationship among cryptographic physical assumptions. Lecture Notes in Computer Science, 762:369–378, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinel:1993:SCC

[2582] C. Meinel and S. Waack. Separating complexity classes related to bounded alternating w-branching programs. Lecture Notes in Computer Science, 762:379–388, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tani:1993:COV

[2583] S. Tani, K. Hamaguchi, and S. Yajima. The complexity of the optimal variable ordering problems of shared binary decision diagrams. Lecture Notes in Computer Science, 762:389–398, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kavvadias:1993:HEH

[2584] D. Kavvadias, C. H. Papadimitriou, and M. Sideri. On horn envelopes and hypergraph transversals. Lecture Notes in Computer Science, 762:399–405, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chrobak:1993:PMA

[2585] M. Chrobak, L. L. Larmore, N. Reingold, and J. Westbrook. Page migration algorithms using work functions. Lecture Notes in Computer Science, 762:406–415, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feuerstein:1993:MPC

[2586] E. Feuerstein and A. Marchetti-Spaccamela. Memory paging for connectivity and path problems in graphs. Lecture Notes in Computer Science, 762:416–425, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ChiKwongHui:1993:RCA

[2587] L. Chi Kwong Hui and C. U. Martel. Randomized competitive algorithms for successful and unsuccessful search on self-adjusting linear lists. Lecture Notes in Computer Science, 762:426–435, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koga:1993:RLA

[2588] H. Koga. Randomized on-line algorithms for the page replication problem. Lecture Notes in Computer Science, 762:436–445, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hambrusch:1993:NAM

[2589] S. E. Hambrusch and H.-Y. T. Tu.

New algorithms for minimizing the longest wire length during circuit compaction. Lecture Notes in Computer Science, 762:446–455, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenberg:1993:PAS

[2590] R. I. Greenberg, S.-C. Hung, and J.- D. Shih. Parallel algorithms for singlelayer channel routing. Lecture Notes in Computer Science, 762:456–465, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1993:CIQ

[2591] A. Aggarwal and T. Tokuyama. Consecutive interval query and dynamic programming on intervals. Lecture Notes in Computer Science, 762:466– 475, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1993:IAT

[2592] A. Aggarwal and T. Tokuyama. An improved algorithm for the Traveler’s Problem. Lecture Notes in Computer Science, 762:476–485, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karuno:1993:VST

[2593] Y. Karuno, H. Nagamochi, and T. Ibaraki. Vehicle scheduling on a tree with release and handling times. Lecture Notes in Computer Science, 762:486–495, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lok:1993:SAC

[2594] Chan Chi Lok and G. Young. Scheduling algorithms for a chain-like task system. Lecture Notes in Computer Science, 762:496–505, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1993:WIP

[2595] G. J. Chang, C. Pandu Rangan, and S. R. Coorg. Weighted independence perfect domination on cocomparability graphs. Lecture Notes in Computer Science, 762:506–514, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1993:PSA

[2596] D. P. Wang, N. F. Huang, H. S.

Chao, and R. C. T. Lee. Plane sweep algorithms for the polygonal approximation problems with applications. Lecture Notes in Computer Science, 762:515–522, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1993:ORS

[2597] S.-W. Cheng, A. Lim, and C.-T. Wu.

Optimal rectilinear Steiner tree for extremal point sets. Lecture Notes in Computer Science, 762:523–532, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1993:FAA

[2598] U. Foessmeier, M. Kaufmann, and A. Zelikovsky. Faster approximation algorithms for the rectilinear Steiner tree problem. Lecture Notes in Computer Science, 762:533–542, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nitzberg:1993:FSD

[2599] M. (Mark) Nitzberg, David Mumford, and Takahiro Shiota. Filtering, segmentation, and depth, volume 662 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-56484-5. ISSN 0302-9743 (print), 1611-3349 (electronic). vi + 143 pp. LCCN TA1634 .N58 1993.

Tanguiane:1993:APM

[2600] Andranick S. Tanguiane. Artificial perception and music recognition, volume 746 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CO-DEN LNCSD9. ISBN 3-540-57394-1 (Berlin), 0-387-57394-1 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 210 pp. LCCN ML74 .T36 1993.

Pieprzyk:1993:DHA

[2601] Josef Pieprzyk and Babak Sadeghiyan. Design of hashing algorithms, volume 756 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1993. CODEN LNCSD9. ISBN 3-540-57500-6 (Berlin), 0-387-57500-6 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). xiii + 194 pp. LCCN QA76.9.H36 P53 1993.